A Complete Bibliography of Publications in the Journal of Mathematical Physics: 1985–1989

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

 E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
            WWW URL: https://www.math.utah.edu/~beebe/

08 November 2023 Version 2.07

Title word cross-reference

#1023 [1442].

(2 + 1) [1201, 1824, 2001, 853, 1702]. (2J + 1) [730]. (3 + 2) [1694]. (4 + 1) [18]. (4, 1) [19]. (D + 1) [1940]. (d > 1) [681]. (N) [397]. (N − 1) [397, 1858, 1133]. (N = 1) [62]. (N = 3) [815]. (T, T_{z}) [570]. ∗ [883, 1420, 1596]. 0 [675, 1360]. 0 + 1 [820]. 1 [851, 330, 1816, 1641]. 1 + 1 [1145, 29]. 1/2 [412, 1136, 675, 1360, 229]. 1/N [1176]. 1080 [456]. 11 [411, 938]. 1421 [695]. 2 [1314, 1641]. 2 + 1 [1712, 2111, 368, 2048, 1394, 2026]. 230 [694]. 29 [1623]. 2k_{F} [718]. 2ν [677]. 2ω [1479]. 2 × 2 [1353]. 2 × N [772, 376, 580, 910, 139]. 3 [82, 330, 930, 245, 620, 1314]. 3 + 1 [904, 285]. 3/2 [2008]. 32 [694]. 38 [695]. 4 + K [871]. 6 [851, 82, 330, 930, 999, 245, 620, 852, 902, 1328]. 9 [1856]. ^{∗} [1418]. ^{2} [2018]. ^{4} [425, 822, 917]. ^{7} [1381]. _{0} [588]. _{8} [1096]. a [1580]. A_{2}^{(2)} [1160]. A_{n−1} [1434]. B [1822, 1580]. B(0, n) [1420, 883]. B_{l}^{(1)} [1380]. b_{Kk}(LM|l) [486]. C [128, 1403]. c < 1 [1708]. c = 1 [1575]. C^{∗} [2003, 777, 214]. C^{+} [128]. C^{∞}(M^{3}, R^{2})/C^{∞}(M^{3}, GL(2, R)) [1227]. C_{7} [228].

H [646]. HH [36]. CP^{2} [562]. CP^{N−1} [283]. CP_{2}^{n−1} [1905]. D [1365, 275, 1969, 1468, 795, 524, 1221, 718, 88, 1470, 1622, 373, 100, 1337, 635]. D(2, 1; α) [156]. d = 4 [411]. d > 1 [1493]. d > 4 [1094]. q̈ + α(t)q̇ + q^{n} = 0 [389]. δ [2105, 1650, 1433]. d ≥ 4 [1750]. Diff S^{1}/SL(2, R) [1610]. E(3) [1313]. E_{l} [913]. E_{n} [1868]. F_{4}(−20) [386]. FS × W [298]. G [1426, 1224]. G(2) ⊃ SU(2) × SU(2) [154]. g_{2} [1446, 1445, 797, 988, 1633, 328, 14, 852]. G_{3} [1681]. Γ [481]. GH^{∞} [1552]. gl(2)gl(2) [1790]. GL(2, C) [957]. gl(2/2) [1790]. GL(4, R) [759, 273, 619, 1563]. gl(m + n) [1915]. gl(m/n) [1915]. GL(n, R) [52]. gl(n/1) [1916]. h [2067]. H^{(0)} + λ⟨r^{p}⟩r^{q} [1464]. h_{n,2n−i}(LM|l) [486]. i [486]. ℑ(ψ) [1730]. j [851, 82, 330, 930, 999, 126, 245, 620, 852, 902, 1328, 1314, 1856]. J_{0}(x) [121]. jm [245]. K [413]. L [1524]. L = (κ/2)ẍ^{2} + (m/2)ẋ^{2} + (k/2)x^{2}j(τ)x(τ) [908]. λ [1999]. L × M × N [387, 1304]. m [1496]. N [120, 1810, 2032, 1801, 1722, 1183, 1930, 2139, 807, 1263, 634, 1144, 2142, 1045, 356, 704, 790, 392, 2103, 1805, 1948, 982, 1881, 108, 898, 715, 1519, 731, 305, 596, 103, 128, 1434, 709, 22, 1224, 486]. N n [1498]. N = 1 [1263, 1734]. N = 2 [2121, 981, 65, 1478, 1550, 1604, 1023]. N = 3 [2122]. N = 4 [1509]. n^{k} [1906]. ∇^{2n}Y_{l_{1}}^{m_{1}}(∇)F_{l_{2}}^{m_{2}}(r) [572]. O(3, 1) [945]. o(4, 2) × o(4, 2) [2107]. O(5) ⊃ O(3) [1053]. O(6) ⊃ O(5) [1053]. O(N) [820, 763]. OSp(1, 2) [1695, 617]. OSP(1/2) [1754]. Osp(1/2N, R) [1559]. OSp(2, 1) [1678]. OSp(2/2N, R) [1752]. OSp(2n/2m, R) [1667]. OSP(2|2) [496]. OSP(2|2p) [884]. Osp(4/2, R) [2087]. OSP(m|4) [1200]. Osp(N, 4) [1023]. \overline{∂} [1243, 1027]. \overline{S}(a, b, c/m) [1532]. P [1090, 2016, 2104, 1331]. p(A) [1511]. p4m [249]. Φ [1625]. Φ^{4} [1625, 1741, 1841, 595]. ϕ^{6} [1225]. ψ_{xx} + uψ = σψ_{y} [1027]. R [1061, 748, 1874]. R^{2} [1774, 1756, 435]. R^{3} [2065]. R^{3,1} [507]. R^{n} [731]. D [607]. ℜ(ψ) [1730]. H [781]. HH [712]. R × S^{3} [558]. Y_{l_{1}}^{m_{1}}(∇)F_{l_{2}}^{m_{2}}(r) [572]. S [1257, 1690, 1524, 1431, 651, 1539]. S^{1} [153]. S^{3} × S^{1} [1282]. S_{2} [1681]. S_{f} [620]. S_{f} ⊃ S_{f_{1}} × S_{f_{2}} [620]. S_{n} [731]. σ [1306, 1567, 1425, 1537, 1187, 1264, 913, 1640, 2024, 1125]. sl(1, 3) [267, 766, 950]. sl(1, n) [1235, 1701]. SL(2, R) [15]. SL(2, C) [1769, 1233, 2120]. SL(2, R) [987, 1718]. sl(3, C) [2093]. SL(3, R) [1345, 1574, 1584, 17, 16]. SL(4, R) [759, 619, 381]. SO(2N) [999]. SO(2N + 1) [999]. SO(2Ω + 1) [1745]. SO(3) [1190, 31, 473, 26, 32, 1271, 1449, 1844]. SO(4) [584, 1445]. SO(4, 2) [998]. SO(5) ⊃ SO(3) [333]. SO(6) [1271, 1449]. SO(9) [551]. SO(N, 1) [2142, 1144]. SO(n − 2, 2) [659]. SO(p + q, p + q) [1485]. SO(p, q) [1485]. SO^{∗}(8) [1098]. SO_{0}(1, 2) [153]. SO_{0}(4, 1) [68]. SO_{N} ⊃ SO_{N−1} [1669]. SP(2N) [999, 243, 1058]. sp(2n, R) [69]. sp(2N − 2) × sp(2) [1058]. Sp(3, R) [492]. Sp(4) [765, 332]. sp(4, R) [622]. Sp(6) ⊃ Sp(2) × O(3) [458]. sp^{∗}(3, R) [492]. spl(1, 2) [617]. SU(1, 1) [1000, 1700, 1852, 1693, 1985, 1793]. SU(1|2) [496]. SU(2) [1233, 920, 1000, 1700, 423, 1026, 1259, 1604, 1876, 1693, 1793]. SU(2) × SU(2) [1605]. SU(2) × SU(2) × U(1) [1605]. SU(2, 2) [51]. SU(2, 2/1) [289]. SU(3) [1199, 1666, 244, 473, 456, 1981, 1489, 1372, 457, 2092, 1446]. SU(3/1)

[1814, 2137]. SU(5) [1605, 1844, 949]. SU(6) [302]. SU(8) [1138]. SU(m/n) [768]. SU(N) [1104, 584, 583, 553, 1978, 1215, 300, 420]. SU(n|m) [295]. SU_{N} ⊃ SU_{N−1} ⊃ SO_{N−1} [1669]. su_{q}(2) [2046]. T [10, 461, 785]. t = const [475]. t = constant [135]. τ [1517, 1160]. θ_{3} [1990]. P̃(1, n) [1127]. U(1) [157, 1659]. U(4) ⊃ U(3) [1098]. u(6/2j + 1) [689]. U(N) [2024, 762, 385, 299]. U(n − 1) [1519]. U(N/M) ⊃ OSp(N/M) ⊃ O(N) × Sp(M) [288]. U_{mn} ⊃ U_{m} × U_{n} [245]. U_{n} [245]. u_{t} = u_{xx}/(1 + u_{x}^{2}) [683, 990]. u_{t} = u_{xx}/1 + u_{x}^{2} [1339]. u_{t} = u_{xxx} + 3(u_{xx}u^{2} + 3u_{x}^{2}u) + 3u_{x}u^{4} [991]. VI_{0} [448, 1408, 1007]. W^{∗} [1251]. ℘m [249]. x^{2} + λx^{2}/(1 + gx^{2}) [674]. Z [1603]. Z′(s) [202]. z − ct = 0 [1644, 1578]. Z_{n} [2091]. ζ [322].

-adic [2104]. -algebraic [2003, 1251, 214]. -algebras [1418, 1596]. -arbitrary [790]. -body [1496, 1519, 128, 709, 1930, 2139]. -boson [1224]. -boundary [1403]. -branes [2016]. -component [730]. -D [1816]. -de [19, 1694]. -dependent [785]. -dimensional [1948, 1881, 108, 898, 681, 1824, 938, 715, 731, 373, 1702, 635, 1479, 2032, 1969, 1940]. -dimensions [100]. -dynamical [777]. -electron [397]. -element [456]. -equation [646]. -field [1225]. -force [1133]. -function [322]. -functions [1160]. -gauge [157]. -gluon [982]. -graded [481]. -independent [785]. -level [1498, 392]. -matrix [1690]. -mer [748]. -modified [1999]. -particle [704]. -positive [2067]. -potential [595]. -representations [883, 1420]. -root [1623]. -separation [1061]. -shell [2105]. -soliton [1045]. -space [731]. -spaces [781]. -sphere [2142, 731, 1144]. -spheres [22]. -type [1259, 913, 1090]. -wave [651, 1539, 1801, 1722, 1805].

İnönü [1545]. İnönü-Wigner [1545].

24 [47, 546, 240, 697]. 25 [151, 327, 547, 509, 48, 152, 461]. 26 [488, 793, 654, 958, 2138, 328, 794, 548, 1025, 1304, 1442]. 27 [1379, 2128, 758, 759, 1085, 1021, 832, 870, 1886, 823, 1118, 1095, 795, 1789, 698, 796, 984, 1513]. 28 [1544, 1416, 2132, 2138, 1231, 1465, 1624, 2142, 1339, 2141, 1512, 1232]. 29 [1983, 1660, 1661, 2053, 1947, 1747, 2127, 1691].

30 [2139, 2135, 2126, 2131, 2137, 2133, 2140, 2136, 2134].

4 [1442].

58204a [2131].

83024a [870]. 83046a [2132]. 84g [240]. 84j [47]. 84m [697]. 85d [48]. 85e [546, 547]. 85i [327]. 85m [461]. 86e [509]. 86h [958]. 86j [793, 794, 1025]. 86m [1304]. 87c [1886]. 87f [2128, 1021]. 87g [984, 1513]. 87h [1085]. 87i [1118]. 87j [1789]. 87k [1095]. 88a [1379, 2138, 870]. 88d [1232]. 88e [1231, 1465, 1339]. 88g [1624]. 88i [1512]. 88k [1691]. 89a [1747]. 89c [2141]. 89i [1947]. 89k [1983]. 8XX [832].

90a [2137, 2127]. 90b [2053]. 90d [2140]. 90e [2132]. 90h [2139]. 90i [2135]. 90j [2134]. 90m [2131].

abelian [1687, 1598, 1163, 463, 263, 819, 1213, 708]. aberrating [691, 692, 1268]. aberration [2054, 2055, 691, 692]. aberrations [1268]. Absence [1493, 1263, 1374, 2076]. absolute [681]. Absolutely [1325]. absorbing [1083]. abstract [1377]. abundant [1939]. accelerated [1250, 2141]. accelerating [1838, 86]. acceleration [2029, 133]. accidental [1213]. accuracy [437, 77]. acoustic [728, 202, 1312, 1570, 452, 238]. acting [1681, 1264, 1376]. Action [633, 828, 534, 845, 1010, 41, 2138, 1642, 153, 1014, 1943, 1439, 29, 632, 605]. action-angle [29]. Action-at-a-distance [633, 632]. actions [958, 65, 979, 75]. adaptation [191]. Adapted [658, 539]. Addendum [2133, 2127, 2140, 1512]. Addition [1037, 1822, 1397, 115]. additional [522]. Additive [1990, 1767, 2092]. Adiabatic [1703, 920, 1422]. adic [2104]. admissible [1380]. admit [1107]. admitting [259, 1387, 1681, 1386, 317, 363, 134, 713, 1405, 1534, 2007]. adsorption [376]. Affine [1682, 1205, 1800, 811, 2082, 484, 1608, 928, 2123]. affinizations [1843]. against [1871]. Aharonov [2108, 1802, 349, 1497, 708]. aided [1043]. AKNS [952, 1517]. Algebra [1035, 381, 1223, 1574, 1285, 1127, 228, 40, 986, 1346, 68, 64, 1791, 1058, 1739, 1380, 241, 662, 473, 614, 2089, 715, 1659, 296, 1236, 1411, 1745, 975, 1820, 1421, 1098, 1584, 1099, 1152, 1630, 250, 1417, 664, 1449, 1461, 1749, 1664, 1508, 1448]. Algebraic [993, 245, 852, 1817, 1708, 2003, 1666, 1251, 1692, 507, 862, 620, 1960, 1583, 876, 1095, 1363, 1550, 591, 1766, 1036, 1779, 519, 214, 1383, 1361]. algebraical [1787]. algebraically [364, 143, 712, 1777, 2008]. Algebras [1460, 2091, 1461, 549, 489, 411, 1537, 182, 1575, 1353, 952, 1792, 621, 1096, 1572, 1984, 1381, 1418, 1847, 168, 1128, 1979, 1843, 841, 918, 1264, 504, 1272, 378, 1748, 1612, 352, 1277, 655, 590, 2094, 70, 334, 887, 1702, 1434, 1417, 1950, 1662, 1986, 1059, 1596, 1632, 1720, 393, 1859, 1626]. algorithm [38]. Aligned [691]. allowed [373]. almost [1652, 1906, 218]. almost-tangent [1906]. almost-thermodynamic [218]. along [713]. alpha [159, 486, 1025]. already [869]. alternate [750, 248]. Alternative [1549, 265]. Ambiguities [1257, 1829, 2030, 46]. ambiguity [1480, 972]. among [2032, 1940, 557]. amplification [186, 652]. amplitude [595, 964, 1680, 46, 1253, 625]. amplitudes [1073, 1500, 1166, 1616, 1291]. analog [1027]. analogs [193, 2108]. analogy [1347, 2009]. Analysis [569, 800, 1696, 1354, 1199, 1744, 83, 1700, 759, 619, 254, 626, 853, 24, 1630, 1611, 856, 246, 247, 117, 342, 1022, 825, 481, 255, 1835, 1124]. Analytic [1763, 413, 1372, 503, 422, 1718, 1413, 1500, 1463, 765, 284, 550, 618, 77, 455, 858, 1721]. Analytical [1003, 1542, 1055, 1666, 2078, 2077, 1981, 1080]. analyticity [1073]. Anderson [730]. angle [235, 2080, 324, 29]. angles [1097]. Angular [2098, 1587, 967, 1139, 126, 1099, 1523, 1360, 365, 1854, 3, 697, 355, 1213, 1022, 1057, 1815, 1286]. Anharmonic [1619, 743, 201, 1169, 1002, 1463, 165, 1772, 1047, 635, 1929]. anharmonicities [1721]. Anisotropic [476, 317, 1259, 947, 954, 836, 441, 363, 905, 427, 1080, 1714, 25]. annihilating [649]. annihilation [1039, 1950, 618]. Annihilation-operator [618]. Anomalies [913, 648, 1187, 2023, 2018, 1419, 223, 1536, 1583]. Anomalous [1688, 234, 396]. Anomaly [1443, 1137, 802, 2042]. ansatz [1960, 1868, 1267]. Anstey [631]. antenna [2088]. anti [323, 751, 878, 1064, 474, 1475, 80, 574, 687]. anti-de [323, 751, 878, 1064, 474, 574, 687]. anti-self-dual [1475, 80]. antisoliton [1606]. antisymmetric [302]. Antone [1025]. Any [1133, 223]. aperture [1108]. appearing [606, 572]. Application [1905, 2101, 73, 155, 2100, 2136, 1556, 894, 1174, 1175, 1229, 1144, 2142, 373, 198, 1382, 272, 38]. applications [617, 1920, 410, 1494, 185, 418, 371, 2067, 107, 1377, 669, 32]. applied [1690, 1518, 388, 256]. Approach [829, 201, 648, 1711, 1918, 372, 2073, 1494, 1161, 1970, 31, 674, 1124, 1236, 1960, 747, 1234, 919, 1527, 1328, 1772, 937, 1882, 1363, 882, 979, 2095, 356, 591, 881, 248, 519, 303, 1273, 1368, 1924, 542, 32, 44, 1114, 1705, 803]. approaches [1666]. approximants [321]. Approximate [1359, 920, 893, 2103, 279, 1705]. Approximation [2071, 1266, 595, 413, 202, 992, 1465, 634, 235, 171, 167, 1193, 1651, 200, 1047, 1422, 650, 77, 145, 1105, 956, 405, 312]. approximations [121, 1093, 1459, 1585]. arbitrary [749, 1072, 976, 721, 1931, 1263, 1103, 788, 30, 1082, 2115, 305, 755, 790, 36, 1108, 266]. arbitrary-helicity [976]. area [947, 451]. arise [604]. arising [2118, 1842, 1027]. Ashtekar [2005]. aspect [509, 823]. aspects [1245, 2096, 316, 1154, 425, 822, 917, 1102, 2042]. assemblies [695]. assisted [1209]. associated [489, 1173, 1201, 1718, 383, 753, 384, 1039, 1444, 1584, 2094, 388, 899]. associative [1380]. associativity [1665]. assumptions [1735]. astrophysical [1818]. asymmetric [1268]. asymmetry [1717]. Asymptotic [118, 2034, 1388, 1889, 63, 2097, 869, 1949, 1963, 1024, 146, 449]. Asymptotically [921, 399, 1581, 474, 1435]. Asymptotics [703, 367, 130, 1432]. atom [1185, 997, 101, 1178, 1467]. atomic [397, 1092, 1603, 186, 652, 1414, 1679, 625]. atoms [1498, 198]. attractive [582]. atypical [884]. Auto [855, 1855]. Auto-Bäcklund [855, 1855]. autocorrelation [404]. Automorphisms [876, 1095, 1791, 1979, 1914]. average [166, 570, 702]. averaging [630]. avoid [1971]. axial [1137, 1309, 284, 455]. Axially [780, 1044, 1045, 1705]. axis [1142, 692]. axis-symmetric [1142, 692]. axisymmetric [358, 2072, 2027, 1219, 969, 1435, 359, 401].

B [2128, 870, 2132]. background [169, 1367, 1875, 859]. backgrounds [941, 1412, 275, 1503, 1470, 1622]. Bäcklund [746, 1855, 1999, 827, 140, 281, 649, 203, 1475, 855, 1375, 73, 283, 1562, 53, 668, 857, 1202, 1697]. backscatter [690]. Baker [1995, 1754, 688, 1265]. ballooning [1441, 791]. bang [217]. bare [1971]. Bargmann [1591, 1568]. barrier [315]. barriers [556, 1255]. base [1585]. based [228, 1381, 213, 243, 1106, 1438, 539, 626, 913]. bases [1692, 621, 467, 1668, 590, 1605, 1985]. basic [2040, 2123, 1018, 1065, 1726]. Basis [847, 987, 1591, 985, 325, 1519, 1790, 297, 1446, 1445, 1099, 1987, 458, 3, 697, 267, 766, 950, 1916, 344, 610, 1224, 2033, 1271, 1844, 551, 1286]. Bassichis [644]. Batalin [974, 973]. bath [1993]. Baxter [2091, 1895, 876, 1095, 2046, 1874]. be [1404, 1617, 1111, 1529]. beam [812]. beams [148]. Behavior [1393, 1894, 996, 164, 1362, 1542, 755, 1963, 1585, 449]. Belinskiĭ [1045]. Belinskiĭ-Zakharov [1045]. Bell [1261, 1262]. belonging [1431]. Benjamin [1161, 1427]. Berezin [598]. Bergman [994]. Berry [1618, 1958, 1803, 1170, 1213]. Bessel [869, 112, 111, 592, 593, 9, 327, 113, 121, 610]. best [2003]. Bethe [614, 371, 1960, 651, 1539, 1284]. between [394, 900, 947, 397, 1730, 541, 1642, 176, 496, 1347, 705, 1884, 271, 1047, 1316]. bi [1707, 1555, 1427]. bi-Hamiltonian [1707, 1555, 1427]. Bianchi [796, 460, 523, 2073, 221, 714, 870, 1118, 448, 588, 1408, 1819, 545, 1007, 1156]. bicommutants [918, 189]. bidirectional [1891]. bifurcation [552, 409, 1269]. Bifurcations [207, 1165, 1088]. big [217]. big-bang [217]. Bilinear [313, 1167, 1212, 1279, 1430, 1068]. Billiard [1825]. billiards [1548]. bimetric [320]. bimodular [912]. binary [1474, 1611]. binding [675]. binomial [1770]. biorthogonal [1666, 1799]. Birkhoff [743, 1888, 1531, 1684, 446, 641]. Birkhoff-type [641]. bispinor [241]. bistable [1354]. black [725, 1289, 1139, 361, 1290, 969, 1898]. Bleuler [1931, 291]. Bloch [1000, 1615, 375]. Bloch-type [1000]. blowup [252]. body [1728, 1769, 725, 1008, 116, 1930, 2139, 1073, 1496, 847, 1623, 1109, 1054, 1519, 265, 33, 1680, 128, 709, 1705]. Bogoliubov [644, 246, 247]. Bohm [2108, 1802, 349, 1497, 708]. Boltzmann [63, 581, 1903, 205, 681, 2052, 449, 1083]. Bondi [1296, 574]. Bonnet [1734, 1301]. boost [149]. boosts [1353, 510, 984, 1513]. Borel [1490, 1853]. Born [2072, 1904, 413, 171, 145]. Bose [698, 90, 415, 909, 518, 478, 532, 1053]. Boson [1285, 332, 622, 418, 1793, 325, 321, 568, 814, 1084, 1110, 1655, 1438, 765, 1745, 1154, 1981, 1224]. boson-quasifermion [1745]. bosonic [1905, 1931, 1601, 1291, 1398]. bosons [1894, 1786, 1505, 438, 1473, 1186]. Bound [234, 2068, 356, 26, 1994, 1728, 1769, 1863, 2013, 349, 296, 1542, 1360, 704, 1055, 1586, 1530, 1773]. bound-state [1863]. boundaries [1359]. boundary [1757, 391, 4, 546, 1647, 371, 1644, 1578, 114, 1886, 1403, 505, 73, 1831, 1758, 1083]. boundary-value [1831]. bounded [1982]. boundedness [858]. Bounds [2003]. Boussinesq [1457, 2021, 1606, 53, 787]. Boutros [1118]. brachistochrone [565, 1652]. bracket [1392]. brackets [561]. Branching [1575, 1058, 2056, 334, 1914]. branes [2016]. Brans [448, 588]. Bravais [373]. breakdown [1266, 1267, 1342, 1780]. breaking [689, 537, 1784, 1001, 978]. brief [671]. broadened [293]. Broadwell [1824]. broken [1191]. Brownian [2009, 1370, 405]. BRST [1476, 1781]. BS [1403]. bubbles [1685]. Bundle [1042, 57, 1378, 1371, 1709, 767, 1961, 2130]. bundles [258, 2023, 1887, 1609, 290, 271, 196, 197]. Burgers [160, 700, 1060, 1671].

C [2133]. Cabibbo [324]. Calculating [172, 548]. Calculation [873, 1443, 101, 1623, 1170, 1024, 388, 455, 486, 1504]. calculations [530, 1120, 1121, 1169, 847, 1052, 1720]. calculus [350, 529, 1029, 904, 762, 763, 1637, 1515, 1370]. Campbell [1995, 1754, 688, 1265]. Canonical [1305, 1544, 309, 2099, 939, 1276, 833, 1134, 1288, 303, 1273, 1466, 942, 1505, 438, 244, 1391, 2, 240, 1330, 1964, 2061, 1489, 2071, 834, 1851, 1547, 36, 578]. canonically [1466]. Canonoid [1638, 1248, 1998, 2028]. capacity [1455]. capture [413, 145]. CAR [813]. Carleman [497]. carrying [1511]. Cartan [178, 1245, 746, 1208, 1811, 1634, 1979, 1840, 1560, 1503, 1564]. Cartan-preserving [1979]. Carter [1900]. Cartesian [495]. case [1483, 1302, 1228, 1849, 1931, 582, 1871, 730, 451, 247, 455]. cases [1019]. Casimir [1873, 2129, 1421, 1090]. Cauchy [63, 54, 1597, 1423, 174, 2114, 158, 1963, 516, 1954, 2134, 25]. Caudrey [1999]. causal [27, 864]. causality [640]. caustics [236]. cavities [1945]. cavity [1354]. CCR [1013, 215]. celestial [1705]. cell [1124]. censorship [1294, 933]. center [663, 903, 1771]. Central [891, 1263, 282, 65, 264]. Central-limit [891]. centrifugal [701]. certain [68, 1801, 1624, 1239, 276, 203, 604, 148, 1155, 1501]. chain [468, 1058, 954, 288]. chains [1948, 1977, 191, 467]. Chandrasekhar [646]. channeling [303, 1273]. chaos [1648]. chaotic [1809, 198, 1686]. Character [456, 382, 1720, 874, 320]. Characteristic [1887, 1491, 470, 471, 736, 1996, 201, 734]. characteristics [4, 546, 600]. characterization [1620, 1804, 2117]. characterizing [1949]. characters [483]. Charge [1220, 840, 945, 1603, 2009, 1956, 1138]. charge-monopole [945]. Charged [277, 522, 679, 2069, 64, 1110, 169, 86, 363, 2007, 1787, 1081, 1469, 1502, 625, 1297, 1815, 1742, 1186, 1705]. charges [1587, 1483, 1071, 1980, 1263, 65, 1522]. Chebyshev [1233, 901]. Chemical [1969]. Cheng [801]. Chern [1734, 535]. Chiral [1266, 1267, 1342, 223, 95, 2001, 1152, 1150, 2024, 2042, 1656, 1394]. Choosing [408]. Choubey [2128]. chromodynamics [1016]. chronoprojective [56]. Circle [1456]. circular [1590, 1108]. class [118, 222, 1718, 1645, 2125, 1300, 1492, 999, 1629, 1794, 2031, 9, 327, 1103, 177, 890, 1094, 276, 492, 1026, 426, 1074, 1126, 604, 210, 1278, 535, 1043, 1714, 1607, 1237, 1317, 733, 1159, 1297, 1426, 809, 624]. Classes [1100, 264, 1660, 1450, 1887, 1118, 535, 2024, 1204]. Classical [1306, 894, 1030, 925, 838, 20, 423, 1895, 652, 1760, 1780, 1593, 193, 1966, 62, 1169, 89, 561, 1010, 141, 874, 1216, 938, 667, 169, 868, 1655, 28, 1889, 528, 600, 1925, 1123, 470, 471, 736, 1996, 79, 683, 1339, 990, 919, 2061, 590, 1976, 1523, 1764, 1993, 979, 1651, 270, 774, 1588, 1997, 1766, 1179, 1957, 1076, 559, 1225, 701, 1996]. Classification [824, 988, 1387, 1200, 1572, 938, 298, 2057, 2135, 1399, 1407, 355, 1604, 1268]. classified [288]. classifying [655]. Clebsch [1666, 987, 822, 917, 2046, 571, 949]. Clifford [549, 489, 182, 1285, 228, 1353, 1236, 1748, 352, 655, 1986, 1749, 1665, 1664]. Clifford- [352]. close [2052]. Closed [765, 103, 1772, 1258, 1348, 1066, 651, 2045, 624]. closed-form [1772]. closure [261]. clouds [1706]. Cluster [2014, 1500]. cocycles [2023, 1234, 822, 917]. coefficient [2063, 855, 856, 1856, 1697]. Coefficients [1524, 1666, 987, 1443, 851, 656, 82, 330, 930, 1002, 999, 191, 762, 763, 297, 902, 1328, 1314, 2050, 570, 2046, 1669, 355, 571, 733, 1053, 949]. Coexistence [1048, 1795, 586]. coherence [1973]. Coherent [1559, 1752, 967, 1913, 1690, 2106, 550, 618, 897, 804, 1942, 1591, 1959, 1326, 868, 1803, 492, 2047, 433, 1382, 478, 532, 676, 1077, 348]. Coherent-state [1690]. cohomological [648]. Cohomology [1247, 1582, 1983, 1163, 1600, 842, 1308, 1419, 995]. Cole [826]. Collapse [1537, 750]. collective [418, 2049, 270, 1414, 1679]. Collectivity [1228, 1849, 458]. Colliding [1218, 1334, 1436, 1532, 1837, 2038, 1807, 783]. collineation [811]. collineations [1608, 836, 1298, 1116, 1682, 1205, 1222]. Collision [444, 837, 1706, 404, 1126, 1806, 1437, 404]. Collision-free [444, 837]. Collisions [1733, 1565]. Color [481, 982, 629]. coma [2054, 2055]. combinations [539]. combinatoric [420]. combining [1014]. commensurate [1715]. Comment [1379, 958, 2138, 706, 1624, 832, 1339, 1886, 509, 823, 1118, 2131, 1025, 1304]. Comments [968, 1465, 347, 1075, 870]. commutant [187]. Commutants [918, 1204, 1672]. Commutation [1039, 378, 1295, 438, 437, 1107, 939, 1547, 1768]. commutativity [1321, 840]. commuting [333, 154, 1399]. Compact [159, 390, 1025, 1306, 1326, 1198, 1545, 1884, 1244, 354]. compactification [771, 684, 1710]. compactness [129]. comparison [1812, 1316, 1169, 1047]. Compatibility [1171, 637, 1811, 1321]. Complete [1431, 851, 597, 1177, 556, 2057, 2135, 1399]. completely [210, 1278]. Completeness [232]. completion [1008]. Complex [797, 2116, 1484, 1582, 1983, 160, 924, 958, 2138, 667, 854, 1162, 1172, 1630, 657, 550, 618, 1604, 1516, 1397, 401, 75, 147, 1430]. Component [1324, 1329, 798, 1012, 730]. composed [1270]. Composite [1451, 1323, 1716, 1149, 296, 1324, 1909, 1223]. composites [1270]. composition [1662]. compressed [1178, 1467]. compressible [756]. Computation [1755, 385, 1930, 2139]. Computer [1043, 1209].

Computer-aided [1043]. computer-assisted [1209]. computing [1995]. concentration [71]. concentric [1433]. concept [2077, 1926, 1343]. Concepts [307]. concerning [1428, 757, 354, 1547]. concomitants [1181, 1796]. condensates [99]. condition [439, 556, 1167, 1212, 1279, 1430, 395, 261, 1798, 858, 972]. conditional [307]. conditionally [972]. Conditions [1880, 1617, 1529, 391, 1647, 537, 1085, 1644, 1578, 114, 1106, 579, 752, 769, 1997, 1111, 1758, 1083, 1023]. conductive [690]. conductivities [427]. conductivity [1270]. conductors [1270]. cone [367, 925, 1110, 1431, 1821, 585, 1479]. cones [1501]. configuration [198]. Configurational [1831]. configurations [1081, 1469, 1502, 136]. Confined [441]. Confinement [1071, 180, 1016, 632, 453]. Confining [573, 623]. confocal [388]. Conformal [808, 836, 817, 183, 1914, 1681, 636, 2075, 1296, 1899, 1298, 1116, 948, 1659, 921, 317, 363, 134, 1411, 701, 1684, 914, 905, 1839, 445, 1222, 2007, 1674, 1708, 633]. Conformally [779, 642, 1535, 1568, 442, 2006, 1119, 2025, 1297]. congruence [1405]. congruences [1688, 1443, 445]. conjecture [531, 1390]. conjectured [2118]. conjectures [1785]. conjugation [479]. connected [233, 541, 1231, 965, 2091]. connecting [944]. connection [394, 1163, 1865, 705, 1503, 1284, 161, 1564]. Connections [196, 828, 1188, 273, 781, 506, 197, 1693]. Conservation [124, 1481, 1815, 1742, 50, 627, 1069, 1451, 1161, 1917, 2122, 1845, 210, 1557, 2037, 633]. conserved [904, 527, 1411, 785, 879, 1702, 916]. Consistent [72]. consisting [128]. constant [1266, 259, 1387, 459, 1139, 41, 1716, 2099, 79, 745, 451, 278, 861, 1622, 1156]. constants [1890, 1211, 183, 509, 882, 1248, 1997, 302, 11, 1641, 1448]. constants/symmetries [1890]. constituent [675, 1360]. constrained [672, 900, 1922, 1554]. Constraint [1782, 946]. Constraints [1794, 775, 1495, 904, 568, 1444, 1371, 834, 1879, 382, 1951]. construct [310]. constructing [539, 1761, 1619]. Construction [1848, 536, 79, 1315, 1962, 1326, 2085, 1380, 2122, 1843, 1886, 1817, 1676, 505, 1130, 1708, 1530, 1361, 1773, 554]. constructions [1403]. Constructive [929, 1625]. contain [1010]. containing [888]. context [1784, 1027]. continua [911]. continual [350]. continuation [1003, 346, 709, 858]. Continued [514, 1577, 644, 1361, 1721]. Continuity [961, 777, 374]. Continuous [1753, 986, 1325, 101, 1735, 1823, 603, 1546, 1985]. continuum [890, 1493, 511]. contour [598, 1994]. contours [1778]. contraction [549, 1545]. convective [294]. Convergence [459, 1287, 114, 312]. Convergent [368, 285, 55, 972]. convolution [1617]. cooperative [748]. Coordinate [722, 1107, 495]. Coordinate-independent [722]. coordinates [1838, 1097, 1759, 678, 1021, 265, 1192, 1193, 270, 966, 36, 760]. Core [349, 651, 1539]. corepresentations [43, 799]. corrected [304]. Correction [2132]. corrections [1049, 704]. Correlated [1414, 1679, 348]. Correlation [1303, 1409, 947, 874, 1236, 1566]. correspondence [258, 1335, 539]. corresponding [1178, 1467]. cosine [519]. cosmic [1294, 933].

Cosmological [1192, 1193, 476, 460, 5, 47, 544, 1364, 696, 714, 174, 1367, 1775, 933, 745, 1118, 1621, 1963, 1776, 278, 1819, 416, 1622, 217, 1156]. Cosmologies [710, 938]. cosmology [870, 448, 588, 1007, 641]. cotangent [1371]. Couette [1311]. Coulomb [272, 1678, 2032, 2096, 898, 1802, 282, 1340, 1017, 1602, 113, 33, 1176, 966, 80, 1055, 519, 542, 778, 1835, 2107, 10, 461]. Coulomb-like [33]. Coulomb-modified [778]. countable [104, 793]. counterterms [751, 1082]. counting [1627]. Coupled [1842, 234, 1457, 1642, 1893, 2013, 1599, 843, 1210, 1527, 1615, 1975, 516, 128, 368, 285, 872]. coupling [466, 1104, 28, 167, 1524, 542, 1619, 1053, 2099, 1641]. couplings [97]. covariance [438, 437, 1388, 28, 2078]. Covariant [1782, 1783, 287, 906, 1307, 1088, 1943, 1439, 1153, 2045]. covering [759, 619]. crack [1570]. Creating [649]. creation [1838, 1039, 1950, 641]. creation-annihilation [1039, 1950]. criterion [122]. Critical [1094, 1729, 1149, 1055]. cross [410, 413, 171, 1663, 720]. cross-section [720]. Crossing [495]. crystal [1624, 971, 1155]. crystallographic [49]. crystallography [419]. crystals [225, 303, 1273]. cubic [1648, 379]. Cumulant [212]. Current [1152, 64, 1659, 296, 1511, 975, 2088, 1954, 2134]. current-carrying [1511]. currents [1071, 322]. Curvature [1401, 1551, 1740, 666, 1517, 259, 1387, 1610, 1675, 1294, 1796, 1733, 402, 934, 861]. Curvatures [1734]. curved [636, 941, 490, 169, 2068, 1082, 2116, 774, 1260, 1902]. curves [1258, 1550, 1758, 2025]. curvilinear [1229]. cutoff [1010]. cycles [425]. cyclic [67, 736, 1099, 1295]. cylinder [1560, 146]. Cylindrical [1089, 326, 1677, 1458, 502]. cylindrically [1933, 745].

D [1886, 1816]. d’Alembert [1922]. damped [1727, 2100, 2136]. damping [517]. data [586, 961, 2052]. Davey [1344, 1712, 1702]. Debye [1257, 1902]. decay [982, 252, 1388, 1592, 130, 405]. declining [1822]. Decomposition [1249, 1743, 1098, 107, 715, 318, 1292, 1990, 625, 2017, 208]. decompositions [386]. decoupling [181, 752]. deduced [941]. Deficiency [391]. definite [1834]. Definition [1126]. Deformation [549, 1277, 2101]. Deformations [2089, 221, 2035]. degeneracies [772, 1487]. degeneracy [118, 2121, 1302, 1254, 1213]. Degenerate [1157, 1649, 489, 993, 65, 659, 136]. degenerated [883, 1420]. degree [1444]. degrees [153, 270]. delay [435]. Delbourgo [1342]. delta [2012]. Demazure [1447]. denominator [244, 1489]. dense [498]. densities [410, 1576, 241, 854, 916, 1579]. density [725, 472, 1735, 1103, 1406, 909, 1737, 1404, 1337, 580, 525, 709, 212, 1138, 370]. dependence [1266, 1603, 631]. dependences [1943]. dependent [1359, 1980, 230, 2028, 1890, 517, 1553, 673, 469, 1327, 1928, 1335, 232, 1254, 1018, 1330, 2099, 235, 423, 1061, 785, 1526, 1725, 264, 1019, 1528, 899, 198, 1703, 718, 348, 605, 1738, 470]. depolarization [690]. dequantization [58]. Derivation [1730, 31, 1438, 1854, 643, 366, 32, 1656, 115]. derivations [1596]. derivative [1143, 959, 908]. derivatives [287, 2032, 1872, 2140]. derived [426]. describing [1505, 1801, 1722, 296, 1579]. description [1966, 1967, 1968, 2010, 1378, 652, 1868, 1685, 392, 1220, 1787, 687]. desingularization [1598]. despite [1293]. Detailed [1585]. determinant [846]. determinants [1530, 1773]. Determination [586, 851, 1451, 420, 760]. determinations [46]. determines [1133]. determining [1772, 1336]. determinism [600]. develop [1733]. development [1946, 1528]. deviation [1965, 780]. deviations [1857]. DeWitt [821]. diagonal [526]. Diagonalization [1953]. diagram [541]. diagrams [944, 12, 151, 1711, 386, 574, 1232, 951]. diamond [1624, 1155]. diatomic [971]. Dichotomies [1919]. Dicke [1230, 448, 588]. Dielectric [687, 1716]. dielectrics [1056]. diffeomorphism [1187]. Diffeomorphisms [1424, 296, 1964, 818, 914, 1378]. difference [952, 606, 119, 311, 1027]. different [944]. differentiable [561]. Differential [952, 1245, 483, 253, 684, 848, 1763, 1920, 1660, 1450, 661, 421, 991, 1634, 606, 251, 491, 1577, 1877, 254, 1428, 1488, 1148, 499, 567, 470, 471, 1996, 626, 158, 1696, 426, 299, 2094, 1352, 1747, 1919, 388, 1246, 834, 1179, 1389, 733, 2025, 255]. Differential-difference [952]. Differential-Stäckel [253]. Differentiation [555, 630, 529, 1066, 453]. diffraction [388, 1, 1442, 803]. diffusion [2118, 1946, 923, 94, 114, 199, 1846, 1736, 1579, 877, 1504]. dilatation [1452]. dilation [1500]. dimension [223, 1493, 1363, 59, 8, 152, 820]. Dimensional [506, 488, 239, 1354, 106, 1345, 1574, 1888, 120, 118, 201, 1942, 824, 814, 1084, 2039, 1948, 2032, 1881, 1866, 967, 108, 1429, 480, 1373, 1228, 1849, 1344, 45, 749, 83, 129, 898, 185, 1935, 681, 1824, 968, 912, 938, 667, 1146, 1147, 715, 308, 892, 875, 1018, 1349, 38, 1088, 1357, 701, 1264, 731, 784, 1790, 79, 373, 1570, 1255, 609, 1421, 821, 1044, 1045, 1969, 1019, 1817, 1940, 375, 587, 1911, 1741, 1336, 1238, 1352, 1747, 91, 1814, 2137, 7, 48, 684, 1751, 1246, 3, 697, 887, 1702, 267, 766, 950, 1235, 1701, 1916, 2006, 750]. dimensional [436, 1521, 1879, 635, 1586, 737, 844, 1479, 1389, 1632, 885, 1398, 1749, 810, 809, 1972, 1830, 1951, 776, 1459, 2070]. dimensionality [1082]. dimensions [1201, 648, 842, 1197, 1107, 595, 2012, 1161, 1145, 430, 1424, 673, 1750, 90, 415, 677, 373, 1094, 2036, 1743, 871, 2001, 853, 1712, 1839, 100, 1337, 2111, 562, 535, 1434, 2024, 2048, 1614, 162, 163, 560, 718, 29, 336, 1929, 1394, 2026, 1070, 1907, 1508]. dimer [38, 279]. dimers [387, 645, 1304]. Diophantine [656, 330]. Dirac [577, 1135, 1810, 234, 396, 438, 64, 124, 480, 483, 241, 2031, 178, 2105, 1040, 977, 1340, 1941, 1357, 701, 177, 1117, 754, 784, 1399, 909, 584, 1518, 1630, 103, 657, 1371, 834, 685, 1260, 1392, 723, 2011, 2043, 2102, 1377, 272, 1907, 1835]. Dirac-type [1518]. Direct [1170, 726, 727, 962, 202, 1047, 1027]. directed [961]. directing [1858]. directions [57]. Dirichlet [394]. discontinuities [27, 555, 1917]. discontinuous [1459, 2070]. Discrete [739, 1670, 932, 1011, 1329, 1687, 1145, 1565, 1103, 373, 1813, 46, 1027, 1924, 2087, 449]. discrete-time [1103]. discussion [138]. disks [937, 91]. dislocations [374]. disordered [1065, 882]. dispersion [719, 202, 660, 1093, 2058, 1540]. dispersive [60, 1134]. Displacement [1762]. Disproof [2118]. Dissipation [1808, 485, 727]. dissipative [1305, 1544, 23, 406, 1330, 726, 727, 962, 602, 1288, 1643]. dissipativity [983]. dissociation [1717]. distance [1010, 1627, 633, 632]. distinct [12, 151]. distorted [171, 1370]. distribution [313, 403, 417, 919, 1136, 2009, 1617, 2088, 280]. Distributions [2067, 367, 1385, 2030, 1338, 1460, 1617, 348, 1461]. divergence [555]. divergency [1024]. divergent [1062]. division [1580]. DLR [892]. Do [1107, 1733]. Dodd [1999]. d’Olival [870]. domain [1982, 726, 727, 962]. dominant [626]. Donder [439]. Double [1079, 1876, 1888, 695, 694, 1503, 482, 401]. double-dual [1503]. Double-solution [1876]. double-valuedness [482]. double-well [1888]. doubling [482]. down [1738]. Drinfel’d [699]. Driven [1413, 343, 186, 652, 2071, 1047]. drop [1269]. droplet [335]. Dual [173, 781, 50, 1866, 1412, 957, 1087, 1026, 583, 639, 640, 1217, 1475, 1713, 1376, 1503, 1102, 1375, 457, 80, 1068, 1812, 1876]. Duality [1373, 1899, 1221, 611, 724, 841, 1277, 1658, 1947, 369, 2084, 382]. Duals [1391]. ducts [940]. due [209]. dumbbells [139]. dust [1300, 443, 1706, 710]. Dynamic [674, 343, 1422, 1970, 905, 1393]. Dynamic-group [674]. Dynamical [467, 273, 998, 531, 1444, 153, 470, 471, 736, 1996, 1696, 1352, 1747, 1276, 777, 1832, 1861, 122, 1872, 2140, 1048]. Dynamics [532, 1305, 1544, 1645, 185, 1921, 1554, 1823, 452, 1976, 1222, 1191, 1001, 1910, 1400, 935, 1318, 916, 513, 405]. Dynkin [1995, 1232, 951]. dyonic [871]. dyonium [351]. Dyons [2082]. Dyson [298, 1128, 2124, 835]. Dyson-like [835].

early [1685]. Easier [486]. eating [554]. Eddington [574]. edh [1724]. effect [2108, 1146, 1147, 631, 1343, 1497, 2051]. Effective [41, 600, 1270, 1336, 1716, 1263, 1599, 38, 1943, 427, 1439, 1361, 1773]. effects [1862, 349, 522, 452, 748, 708]. Efficient [1586]. eigenfunction [30, 1521, 1607]. eigenfunctions [421, 1035, 311, 2071, 557]. eigenspectrum [22]. eigenspeed [1390]. eigenstate [513]. eigenstates [1360, 1962]. Eigenvalue [1847, 520, 1758, 1818, 118, 388, 1111, 1529, 610, 1924, 370, 1955]. Eigenvalues [1028, 1359, 1173, 6, 547, 2071, 301, 791, 635]. eigenvectors [1028]. eight [45, 1715]. eight-dimensional [45]. eight-vertex [1715]. Einstein [1332, 475, 222, 1333, 746, 1006, 50, 1866, 2072, 1208, 221, 178, 1934, 1116, 873, 1366, 1653, 909, 276, 1219, 871, 87, 61, 576, 1807, 2111, 1503, 1776, 577, 1685, 1221, 1944, 1368, 1562, 574, 1301, 88, 872, 2008, 810, 401, 1705]. elastic [653, 1257, 2124, 882, 625, 637]. elasticity [1385, 123]. Elasticoviscous [209, 1311]. elastoplastic [2030]. electric [326, 1703]. electrodynamic [213]. Electrodynamics [45, 1229, 227, 794, 1925, 1523, 1764, 1360, 1310, 2084]. Electrohydrodynamic [1311, 326, 2051]. electrolyte [1343, 1717]. electrolytes [24]. Electromagnetic [1587, 1746, 675, 1714, 1108, 137, 60, 89, 1072, 911, 318, 310, 540, 2112, 690, 929, 278, 1733, 1432, 1787, 1221, 348, 803]. electromagnetism [1980, 1014, 1220, 402]. Electron [1337, 1341, 397, 487, 413, 1236, 1017, 1602, 1788, 718, 145, 1738]. electronic [247]. electrostatics [1291]. electrovac [1807]. electroweak [324, 98, 1936]. element [456]. Elementary [51, 742, 759, 619, 84, 1223]. elements [1340, 1833, 2098, 1090, 659, 1605, 1844, 1186]. elimination [972]. ellipsoid [1319, 1455]. elliptic [1755, 2091, 956]. Elliptical [1548]. embedded [638, 1422]. embedding [1786, 497, 1454]. Emden [389]. encoding [81]. end [2088]. endospectral [850]. energies [1073, 1603, 790, 1530]. Energy [201, 211, 635, 1341, 120, 1878, 479, 1794, 178, 1873, 2129, 854, 173, 1178, 232, 1735, 875, 873, 1180, 1473, 1406, 1017, 1602, 1362, 335, 1099, 1542, 755, 100, 816, 867, 2081, 782, 1534, 1467, 1909]. energy-cyclic [1099]. energy-dependent [232]. Energy-momentum [211, 178, 1180, 1534]. energy-time [479]. Enlargeable [1059]. ensemble [1146, 1147]. ensembles [874, 224]. entailing [215]. entire [1106]. entropic [2002]. entropy [1757, 1920, 1062, 531, 199, 1123, 1217, 893, 777, 559]. entropy-minimum [1920]. envelopes [282, 634]. enveloping [473, 1098, 1449]. environment [1737]. equal [330]. equation [1763, 1135, 488, 239, 932, 1904, 1810, 1016, 1707, 722, 396, 63, 1891, 953, 54, 1429, 1329, 1855, 56, 991, 1999, 2063, 807, 581, 749, 1730, 1903, 83, 1177, 1161, 2021, 512, 1133, 205, 606, 662, 1003, 500, 1395, 735, 614, 1037, 371, 254, 892, 1113, 843, 1038, 1039, 1441, 699, 1340, 1941, 2091, 773, 1816, 177, 1117, 199, 756, 754, 275, 403, 417, 1351, 1639, 683, 1339, 990, 341, 2124, 726, 727, 962, 389, 1696, 1772, 305, 663, 1277, 432, 1280, 1314, 1956, 1875, 940, 1606, 1630, 1689, 520, 369, 501, 388, 1868, 835]. equation [229, 1648, 855, 856, 1702, 651, 1539, 246, 247, 2052, 347, 791, 888, 591, 1027, 248, 1932, 1237, 1284, 2011, 2043, 1831, 1586, 1256, 2102, 1131, 1929, 1318, 1141, 36, 712, 449, 1083, 25, 1076, 857, 312, 738, 1070, 1691, 1643, 1350, 2065, 1657, 646, 877, 1046, 255, 1390, 889, 1907, 1697, 1384, 502, 147, 787, 1469]. Equations [1759, 1332, 475, 1173, 1201, 222, 1920, 1729, 1101, 797, 988, 1589, 104, 793, 1333, 952, 1517, 746, 1660, 1355, 1450, 1006, 358, 2015, 1863, 661, 1866, 2120, 656, 2072, 1329, 330, 865, 1811, 1391, 1477, 484, 1939, 195, 1890, 1481, 1344, 1457, 1087, 960, 1982, 221, 827, 597, 1634, 976, 681, 361, 2031, 1000, 383, 251, 491, 923, 1590, 178, 1577, 886, 721, 1934, 1100, 1847, 119, 1116, 1762, 1427, 1877, 753, 826, 1633, 1428, 1488, 926, 1210, 1075, 1870, 815, 1366, 788, 1568, 1167, 1212, 1279, 1430, 1644, 1578, 1148, 2083, 499, 1313, 1895]. equations [1061, 470, 471, 1996, 57, 626, 276, 158, 2064, 1612, 1219, 203, 663, 871, 584, 1712, 293, 87, 61, 576, 1807, 426, 1613, 2044, 1940, 1518, 2094, 876, 1095, 1658, 1947, 1472, 1374, 1625, 1741, 1841, 1475, 1238, 1352, 1747, 1032, 893, 1919, 2111, 1990, 516, 1713, 562, 899, 1842, 2046, 1033, 2133, 210, 1278, 1874, 1372, 887, 685, 888, 1685, 206, 610, 102, 436, 2062, 700, 1060, 1671, 1068, 1221, 1237, 1368, 465, 1924, 1736, 1540, 825, 1562, 1954, 2134, 1389, 946, 1812, 733, 801, 734, 88, 872, 1470, 2074, 2008, 481, 161]. equations [44, 340, 53, 1579, 720, 1225, 810, 1383, 671, 401, 1705, 1835, 989, 1738]. equilateral [432]. Equilibrium [179, 2125, 136, 1908, 874, 1970, 1511, 682, 829]. Equivalence [900, 1840, 440, 1654, 1932, 1181, 833, 18, 657, 1079]. equivalent [508, 832]. equivalents [398, 1329]. equivariant [2044, 1033, 2133]. era [696]. Ergodic [417, 403]. Ermakov [602]. Ernst [305, 596, 1868]. Erratum [488, 1544, 1983, 151, 793, 2128, 1660, 1416, 758, 2139, 47, 654, 546, 2138, 759, 2129, 1231, 240, 327, 2135, 1661, 1021, 547, 328, 2142, 2053, 794, 823, 1095, 1947, 2126, 795, 1747, 1789, 2137, 48, 2141, 697, 698, 152, 548, 796, 984, 1513, 2136, 2134, 2130, 1232, 1691, 1442, 461, 2132]. errors [503]. Espindola [832, 832]. essential [1646]. estimates [1209]. estimation [1055]. Euclidean [1567, 1537, 1881, 898, 731, 1026, 583, 1238, 117, 342, 1938, 533, 671]. Euler [1097, 700, 1060, 1671]. Eulerian [798]. evaluating [486]. Evaluation [1711, 335, 604, 322, 145]. Evaporation [1289, 1290]. even [919]. Evolution [1700, 805, 1789, 991, 1722, 1481, 960, 383, 1395, 1100, 1327, 1928, 94, 1834, 167, 203, 888, 1400, 1237, 481, 1579]. Exact [1355, 1946, 1802, 472, 205, 1145, 1565, 101, 1463, 338, 94, 1038, 1366, 1117, 1362, 1561, 937, 364, 644, 1238, 679, 984, 1513, 844, 1929, 1225, 147, 475, 1353, 1333, 1891, 360, 681, 1824, 2031, 968, 1085, 714, 2083, 1725, 579, 1118, 856, 577, 1288, 2043, 370, 175, 510]. exactly [915, 1379, 1100]. example [811, 2030, 483, 173, 1027]. examples [1562]. Exceptional [1950, 15, 797, 988, 1792, 1633, 174, 657, 386, 156]. exchange [294]. excitation [1603]. excitations [582]. excited [2013, 368]. excitonic [1057]. excitonic-polaron [1057]. exhibiting [1780]. Existence [1878, 40, 1073, 1511, 587, 1115, 1105, 63, 372, 1293, 230, 1106, 1581, 1945, 424, 511, 449, 1083, 1626, 1853, 814]. existing [869]. exomorphic [589]. expanding [1746, 849]. Expansion [99, 2012, 1821, 169, 1257, 35, 1405, 1176, 1737, 2126, 1991, 1885, 2131, 1274, 835, 966, 146, 1669, 2014, 1665]. expansions [466, 665, 2097, 30, 113, 2099, 250, 368, 285, 572, 102, 1521, 212, 1619, 1464, 55]. expectations [307]. explanation [515]. Explicit [617, 299, 1868, 1640, 283, 1504, 2122, 1112, 1489]. Explicitly [1890, 214]. explosive [1579]. exponent [1094]. Exponential [1327, 114, 1590, 519, 107]. exponentially [1822]. exponentials [920, 1104, 1877, 107]. exponentiation [1537]. exponents [1809]. expression [159, 571, 1112, 1025, 702, 2107]. Expressions [1873, 2129, 852, 301, 572]. Extended [1729, 2001, 1510, 1127, 2122, 84, 1411, 1154, 1206, 364, 1393, 1301, 1232, 1023, 951]. Extension [621, 127, 28, 1526, 477, 1689]. Extensions [1136, 411, 2018]. Exterior [529, 71, 1705]. external [1135, 1071, 1101, 802, 41, 963, 1017, 1602, 1012, 1993, 1372, 1432, 2011, 2043, 1954, 2134]. extinction [2050]. extra [1487]. extracting [172, 548]. extrapolation [503, 422, 466]. extremely [1046]. extremum [422].

F [1304]. Factor [2020, 982, 1782, 1783]. Factorization [1675, 499, 1070, 1691, 1643, 1350, 241]. Factorizations [1286]. Factors [21, 66, 852, 378, 1841, 931, 1669]. Faddeev [1003, 846]. failures [1293]. fall [1923, 20]. families [1204]. family [406, 1163, 1807, 1952, 1776, 844, 1876]. fast [413]. Federbush [649, 785, 879, 1482]. Fermi [1015, 321, 90, 415, 1092, 1440, 543, 909, 1969, 1053, 256]. Fermi-type [256]. Fermion [582, 915, 1379, 325, 322, 282, 2042, 1566]. Fermionic [97, 120, 1829, 1950, 1398]. Fermions [558, 1015, 228, 1505, 438, 188, 1226, 912, 2110, 963, 95]. ferromagnets [1094, 1409]. few [1085, 33]. few-body [33]. Feynman [151, 944, 12, 1711, 1128, 585, 730, 1892, 455, 2000, 708, 292]. fiber [1378, 290, 271, 196, 692]. fibers [1842]. Fibonacci [377]. Field [1593, 1187, 2016, 161, 1135, 1973, 1483, 1305, 1544, 814, 1084, 1015, 1908, 1505, 2039, 104, 793, 1711, 286, 534, 845, 179, 1006, 1645, 2125, 64, 39, 561, 2040, 1029, 2082, 1412, 1072, 41, 1388, 323, 219, 326, 2051, 1601, 242, 211, 125, 582, 86, 1507, 926, 318, 28, 282, 450, 1366, 788, 841, 30, 1627, 164, 600, 1438, 701, 1864, 184, 1123, 677, 638, 1172, 1712, 238, 1541, 1374, 181, 752, 1839, 1012, 652, 1181, 2116, 1788, 1867, 2081, 1371, 1372, 774, 1588, 834, 457, 1369, 80, 1260, 1779, 750, 1007, 1787, 1938, 197, 220]. field [1562, 1261, 1262, 1708, 180, 453, 647, 1398, 214, 1225, 1186, 810, 2045, 1906, 1193]. Field-diffeomorphism [1187]. fields [1305, 1544, 1071, 942, 612, 813, 1013, 157, 89, 2085, 636, 233, 957, 1087, 1858, 838, 1931, 19, 1601, 406, 1149, 921, 1694, 236, 963, 1869, 1330, 831, 441, 1211, 1746, 1064, 1594, 264, 305, 596, 1026, 583, 2112, 755, 1975, 346, 1563, 42, 601, 1310, 978, 1246, 506, 1503, 85, 929, 1150, 1432, 577, 81, 392, 172, 548, 2104, 1703, 2011, 2043, 26, 844, 1318, 143, 872, 1051, 1902, 1777, 170, 1398, 1628, 1626, 1876]. Fierz [182, 721, 1751, 881]. fifth [1351]. fifth-order [1351]. filled [1629]. filling [748]. filter [631]. filtering [1174, 1175]. final [670]. finding [248]. finest [1434]. Finite [1088, 1790, 403, 404, 417, 1472, 1814, 2137, 267, 766, 950, 1235, 1701, 1986, 1303, 1908, 468, 437, 1169, 883, 1420, 108, 749, 1146, 1147, 1507, 167, 919, 1833, 1969, 1987, 1352, 1747, 1028, 1916, 1027, 1988, 643, 463, 659, 2014, 1252, 1937, 1951]. finite-degenerated [883, 1420]. Finite-dimensional [1088, 1790, 1814, 2137, 267, 766, 950, 1235, 1701, 1352, 1747, 1916, 1951]. finite-set-of-moments [919]. finite-temperature [1507, 643]. Finitely [1647, 1433]. Finslerian [164]. First [389, 916, 894, 1494, 1634, 732, 471, 264, 1584, 1086]. First-order [916, 894, 1494, 1634, 732, 471]. fitting [484]. five [1778, 885]. five-dimensional [885]. five-simplex [1778]. fixed [1354, 234, 570, 857, 1202]. fixed-isospin [570]. fixing [291]. flat [399, 921, 1652, 569, 509, 771, 1435, 1119, 11, 2025, 642, 1535, 1297, 810, 809]. flights [996]. Flow [294, 2128, 1429, 209, 1311, 338, 756, 149, 693]. flows [1845]. Fluctuation [485, 207]. Fluctuation-dissipation [485]. fluctuations [1699, 365]. Fluid [523, 525, 460, 1836, 1101, 237, 2072, 360, 1069, 277, 544, 2113, 1933, 1921, 1365, 132, 678, 1021, 134, 1406, 177, 1219, 1259, 440, 1468, 795, 524, 91, 1534, 588, 1408, 1007, 2062, 150, 1562, 1081, 1469, 1502, 136, 809]. fluid-saturated [237]. fluids [1333, 358, 1183, 808, 316, 357, 836, 1298, 1901, 441, 1347, 1142, 1560, 425, 822, 917, 905, 679, 970, 1566, 359, 642, 1535, 1046]. flux [2118, 252, 440]. Fock [246, 247, 1828, 2010, 288, 21, 34]. focus [1168]. focusing [1293, 1356]. Fokker [118, 1920, 429, 310, 682, 465, 2000, 1383]. Fokker-type [429]. Foldy [1941, 644]. folk [1157]. followed [1706]. force [1133, 522, 335, 264]. forced [517, 564, 605]. Forces [511, 337]. form [398, 743, 1888, 1245, 1855, 1730, 1100, 1840, 566, 1527, 1772, 1489, 1956, 299, 1468, 795, 524, 1292, 369, 1796, 834, 651, 643, 1740, 1464, 1657, 624]. Formal [815, 1528, 2048]. formalism [1305, 1544, 894, 900, 1857, 974, 1391, 636, 1364, 2079, 1690, 1870, 1330, 227, 794, 819, 2099, 1925, 784, 290, 973, 127, 1872, 2140, 1041]. forms [394, 797, 1429, 664, 424, 283, 1090, 1377]. Formula [300, 581, 1627, 1772, 385, 103, 1854, 1274, 650, 1516, 1686, 1462]. formulas [797, 988, 1307, 304, 869, 1633, 245, 620, 1438, 705, 107, 989]. Formulation [429, 738, 1707, 722, 981, 1495, 332, 622, 1427, 1441, 566, 1613, 1976, 940, 414, 1066, 816, 1322, 2002, 559, 1132, 1704, 1832]. Four [814, 1084, 1523, 842, 1623, 701, 675, 1360, 1680, 1238, 562, 2092, 2006, 671, 1907, 1508, 624]. four-body [1623, 1680]. Four-dimensional [814, 1084, 701, 1238, 2006]. Four-momentum [1523, 675, 1360]. four-space [671]. Fourier [1974, 623, 1452, 1028, 370, 55]. fourth [1763, 251, 1705]. fourth-order [1763, 1705]. Fractal [996]. Fractals [1826]. Fractional [595, 121, 1736, 956, 2032, 202, 1524, 877]. fractions [1577, 644, 514, 1361, 1721]. Fradkin [973, 974]. frame [713, 1181, 1250, 2141]. frames [271, 1798]. framework [2015]. Fredholm [1292, 893]. Free [1601, 1735, 2033, 1183, 2085, 316, 487, 1388, 1931, 1923, 125, 875, 1677, 762, 763, 17, 1620, 841, 20, 852, 1775, 1969, 444, 837, 571, 718, 381, 1081, 1469, 1502, 16]. free-fall [20]. freedom [1568, 153, 270, 645]. frequencies [172, 548]. frequency [1732, 83, 404]. Front [566]. fronted [278]. frustration [261]. fulfilling [813, 1013]. function [503, 159, 1978, 201, 468, 987, 244, 1490, 1137, 719, 907, 923, 1873, 2129, 242, 322, 9, 327, 1103, 288, 1340, 113, 403, 404, 417, 747, 305, 1489, 121, 971, 882, 610, 2062, 155, 2033, 1585, 486, 1025, 778, 370, 1240, 1241, 1242, 1656, 1426, 272, 401]. Functional [698, 518, 120, 630, 21, 470, 471, 736, 1996, 354]. functionals [785, 1767, 1189, 533]. Functions [67, 1894, 1303, 725, 1718, 612, 1678, 1326, 1517, 947, 2032, 1774, 1491, 1226, 586, 874, 898, 555, 1755, 1590, 607, 724, 854, 1151, 869, 1178, 1759, 112, 111, 592, 593, 1018, 113, 1106, 1724, 1506, 313, 1834, 705, 1362, 1884, 909, 1019, 1990, 395, 966, 1160, 110, 572, 1771, 660, 1093, 2058, 77, 1566, 1708, 956, 256, 1822, 1186, 1467, 1853, 624, 1431]. functorial [258]. fundamental [105]. further [1700]. future [1581].

G [509, 823, 2133]. gain [603]. Galilean [216, 2089, 1396]. gap [1994]. gas [1966, 1967, 1968, 756, 1969, 1472, 1337, 718, 916]. gases [444, 837]. Gauge [1600, 89, 2018, 216, 291, 818, 914, 1378, 1943, 880, 1068, 143, 1961, 2130, 1697, 1384, 828, 1538, 1342, 40, 1905, 157, 319, 286, 534, 845, 480, 141, 1412, 464, 4, 546, 958, 2138, 2122, 1931, 1040, 2013, 1309, 963, 798, 1568, 629, 819, 1163, 1054, 184, 784, 142, 1026, 284, 821, 585, 583, 1765, 271, 781, 231, 1152, 1012, 1784, 1782, 1783, 1563, 42, 1191, 1798, 2081, 368, 285, 457, 34, 881, 197, 455, 1315, 2014, 872, 1051, 180, 453, 647, 995, 1564, 75, 1876, 1266, 1005]. gauges [1307]. Gauss’ [1922, 1734, 1301]. Gaussian [1576, 1047]. Gegenbauer [624]. Gel’fand [15, 987, 1804, 267, 766, 950, 1916]. Gell [1516]. Gell-Mann [1516]. General [1333, 1406, 1956, 669, 894, 1836, 137, 439, 534, 845, 1293, 758, 575, 744, 1732, 1183, 808, 316, 1145, 1933, 31, 1716, 1364, 836, 779, 1309, 400, 1254, 1401, 1402, 695, 694, 1094, 663, 293, 1259, 655, 1404, 745, 329, 1299, 748, 1126, 1014, 1868, 301, 782, 8, 152, 2024, 1635, 525, 1080, 1551, 1112, 96, 220, 32, 1297, 1962, 384, 138, 818, 246, 1261]. Generalization [1078, 1263, 199, 142, 266, 1006, 37, 2077, 1358, 1698, 1723, 329, 639, 730, 1770]. Generalizations [1417, 1900]. Generalized [1942, 319, 2032, 1087, 1590, 1974, 854, 1097, 298, 1803, 1091, 339, 1162, 425, 752, 1250, 2141, 1521, 700, 1060, 1671, 994, 1053, 256, 1830, 67, 1753, 998, 1959, 985, 27, 54, 1531, 1139, 581, 129, 1821, 219, 242, 1335, 1554, 884, 2057, 2135, 1438, 297, 378, 800, 1684, 428, 596, 663, 1615, 1468, 795, 524, 2035, 395, 1595, 200, 1503, 1497, 1998, 1397, 393, 1768, 646, 1044]. Generating [799, 437, 494, 1196, 1719, 987, 622, 714, 655, 1118, 1771, 155, 1812]. Generation [401, 508, 832, 1813]. generator [695, 694]. generators [561, 2037, 1605, 1844]. generic [1551, 1932]. Geodesic [1965, 1857, 1293, 1858, 527, 1405, 1192, 1193]. Geodesically [2005]. geodesics [1857, 713, 1827]. Geometric [897, 672, 1494, 1004, 1005, 979, 1253, 1936, 1245, 1638, 211, 2016, 1109, 1519, 231, 2095, 1260, 1852, 1516, 994, 1114]. Geometrical [507, 1572, 1620, 341, 123, 157, 454, 846, 1328, 2054]. geometries [361, 599, 1986]. Geometrostatis [975]. Geometry [1923, 789, 390, 319, 56, 1228, 1849, 1496, 2041, 1827, 1206, 1784, 7, 48, 458, 684, 1635, 320, 2025]. geophysical [859]. Geroch [2004, 1194, 1512]. ghost [1782, 1783]. Gibbon [1999]. Gibbs [581, 200, 1462]. Ginzberg [1982]. GKP [1403]. glass [889]. glasses [1347]. Glauber [676, 625]. Global [316, 406, 1583, 640, 1348, 2052, 1766, 449, 1156, 63, 1774, 1208, 230, 629, 1709, 424, 1502, 1083]. globally [1423, 1192, 1193]. glueball [368, 285]. gluon [725, 982]. gluons [180]. Gödel [1143, 133, 1258, 680]. Gödel-type [1143, 680]. Goldstone [651, 1539]. good [1606]. googly [830, 831]. Gordan [1666, 987, 2046, 571, 949]. Gordon [54, 807, 1151, 892, 1279, 341, 1238, 347, 206, 750, 1393, 1835]. governed [1831]. Graded [258, 1206, 1059, 481]. gradient [552, 1725]. gradient-type [1725]. gradings [1434, 2093]. graph [2045]. graphs [377, 850, 501, 829, 292]. Grassmann [1306, 1567, 193, 1100, 1921, 352, 1515, 702, 1448]. Grassmann-algebra-valued [1448]. Grassmann-like [352]. Grassmann-valued [1921]. Grassmannian [1425, 1537, 1640, 723, 283]. gravitating [1300, 937]. gravitation [476, 2114, 1865, 1563, 970, 402]. Gravitational [648, 2108, 443, 1621, 812, 1369, 849, 712, 273, 1008, 89, 907, 223, 219, 1218, 1334, 1436, 86, 873, 1837, 2038, 441, 780, 164, 1864, 638, 1594, 1217, 260, 1867, 278, 1733, 1936, 1706, 1437, 711, 1009, 783]. graviton [948]. gravitons [781]. Gravity [1564, 1143, 438, 802, 1300, 484, 1597, 943, 1599, 1044, 1045, 684, 320, 935, 578]. Green [468, 2032, 898, 854, 1151, 1340, 1724, 971, 882, 2033, 778, 1240, 1241, 1242, 272]. Greenberg [2039]. Gribov [1480]. groove [1946]. ground [2121, 397, 1794, 1473, 1542, 212]. ground-state [1473, 1542]. Group [804, 985, 2023, 1154, 853, 246, 247, 760, 828, 488, 239, 1345, 1574, 1199, 1978, 1600, 1127, 998, 1591, 797, 988, 439, 1197, 1881, 1681, 1791, 1866, 1226, 2018, 2085, 1344, 1214, 1215, 464, 379, 654, 418, 938, 2056, 19, 674, 1754, 456, 948, 1633, 493, 467, 296, 1236, 384, 620, 328, 14, 317, 363, 134, 1144, 2142, 1452, 153, 567, 1089, 1668, 142, 1265, 492, 1415, 1882, 818, 914, 1378, 1447, 18, 1376, 691, 2047, 550, 618, 380, 2048, 117, 342, 248, 1422, 43, 510, 984, 1513, 192, 1800, 386, 1852, 1271, 2017, 995, 692, 1268]. group [1114, 1508, 2107]. Group-related [804]. Group-theoretic [760]. group-theoretical [1199, 938, 691, 692, 1268]. groups [489, 1011, 193, 824, 1753, 1198, 108, 1477, 1848, 331, 1419, 999, 491, 295, 958, 2138, 216, 759, 619, 1486, 762, 763, 1883, 884, 1200, 1402, 798, 1471, 2053, 20, 658, 1949, 52, 1834, 297, 695, 694, 786, 1884, 1234, 914, 891, 913, 1515, 249, 1382, 463, 1194, 1512, 767, 49, 1516, 2014, 453, 647, 578, 75]. growth [934]. guides [1356]. Gupta [1931, 291]. Gustavson [743, 1888].

H [1339]. Haar [299]. Hadamard [941, 304]. hadron [1215]. hadron-type [1215]. hadrons [687]. Hahn [986, 1546, 1799]. Hajj [1118]. Hajj-Boutros [1118]. half [1231, 965, 1570]. half-line [1231, 965]. half-space [1570]. Hamilton [1857, 735, 1061, 1613, 414, 501, 1556, 1657]. Hamiltonian [547, 1904, 1707, 900, 372, 89, 160, 1495, 2028, 204, 827, 487, 490, 469, 1912, 1427, 904, 263, 1759, 1254, 6, 153, 1351, 563, 1520, 1975, 2111, 1953, 1278, 1910, 1371, 834, 1557, 2119, 1555, 1614, 542, 935, 1761, 170, 1076, 787]. Hamiltonians [832, 1888, 1646, 109, 353, 508, 242, 1928, 310, 1676, 2103, 1464, 1361]. Hamiltonization [1031, 1906]. Hankel [1723, 919]. Hansen [2005]. hard [403, 404, 417, 91, 1539]. Harmonic [117, 342, 1894, 1786, 1559, 1569, 1948, 1959, 1525, 517, 1553, 1455, 1642, 1862, 1327, 776, 896, 305, 1478, 903, 1066, 3, 697, 1771, 2033, 320, 1768, 1286]. harmonics [131, 379, 654, 176, 607, 1144, 2142, 380, 74, 115]. Hartree [1528, 246, 247]. Hausdorff [1995, 1754, 688, 1265]. Hauser [1534]. head [1166]. head-wave [1166]. Heat [1821, 1388, 440, 2126, 1991, 1885, 2131, 1993, 294]. heavy [100]. Heisenberg [1591, 1569, 986, 952, 954, 493, 1820, 1172]. helically [170]. helicities [1170]. Helicity [1360, 611, 976, 1431, 1931, 639]. Helmholtz [2055, 1037, 1997, 1831, 115]. help [388]. Hermite [1274]. Hermitian [2064]. heterogeneous [1429]. hidden [1036, 1779, 1927]. hierarchies [1939, 1609, 785, 879]. hierarchy [50, 160, 886, 1243, 1376, 1832, 1861, 888, 1112]. Higgs [1598, 423, 978, 1244]. High [77, 1732, 83, 755, 1909]. High-accuracy [77]. high-energy [755]. high-frequency [1732, 83]. Higher [1935, 7, 48, 790, 1143, 974, 1491, 195, 1890, 1184, 185, 2122, 1840, 626, 2036, 1044, 1045, 684, 535, 1872, 2140, 935, 1070, 1972, 1830, 1697, 1384, 1906]. higher-derivative [1143]. Higher-dimensional [1935, 7, 48, 185, 1044, 1045, 684, 1972, 1830]. Higher-order [790, 974, 195, 1890, 1184, 1840, 626, 535, 935, 1697, 1906]. Highest [1915, 1514]. Highly [2013]. Hilbert [1718, 187, 1930, 2139, 1739, 383, 1984, 125, 1882, 880, 1237, 1195, 1484, 1767, 946]. Hill [1773, 1607, 1530]. Hirota [1167, 1212, 1279, 1430, 769, 1068, 340]. Hölder [257]. hole [1289, 1139, 361, 1290, 1745, 969, 1898]. Holonomy [1402, 880]. Homogeneous [1143, 1201, 952, 1492, 1913, 938, 19, 66, 2057, 2135, 362, 444, 601, 1819, 150, 1389]. homogeneously [293]. Hopf [826, 409]. horizon [1963]. horizons [174, 446]. horizontal [2049]. Hubbard [415]. Hurwitz [1662]. Hydrodynamic [416, 5, 47, 783]. Hydrodynamics [863, 1385, 366]. hydrogen [1287, 997, 705, 625]. hydrogen-oscillator [705]. hyper [927, 1927]. hyper-Kähler [1927]. hyper-relativistic [927]. hyperbolic [987, 536, 357, 41, 958, 2138, 1590, 1423, 628, 1280, 789, 1953, 1557, 75]. hyperbolicity [1348]. Hyperfunction [2040]. Hyperfunctions [613]. hypergeometric [985, 1203]. hyperheavens [1079]. hyperoctahedral [2056]. hyperplane [1644, 1578]. hyperspherical [265, 74]. hyperspin [1471, 2053]. hypersurface [362]. hypersurface-homogeneous [362]. hypersurfaces [475, 810, 809]. Hypervirial [1066]. hypoelliptic [1756]. hypothesis [933, 346]. hysteresis [1611].

icosahedral [1573, 1239, 1415, 380]. Ideal [1188, 746, 1441, 1255, 757, 791]. Ideal-preserving [1188]. idempotents [1748]. identical [377, 1473, 931]. identification [1174, 1175]. identities [182, 985, 2115, 284, 1478, 1074, 1751, 881, 1607, 1622]. identity [221, 557, 1761]. II [654, 796, 984, 1513, 106, 1574, 799, 1769, 2055, 1267, 1030, 814, 1967, 1120, 988, 845, 1420, 930, 1477, 622, 2005, 1838, 361, 556, 353, 295, 838, 1334, 1147, 1124, 415, 112, 593, 763, 714, 528, 2038, 831, 1212, 2101, 1602, 1840, 870, 442, 403, 417, 471, 1339, 990, 727, 822, 1446, 1045, 1193, 914, 1118, 1489, 1005, 1741, 181, 1468, 652, 1783, 280, 59, 856, 1796, 1278, 1861, 247, 950, 1701, 279, 618, 1679, 1529, 380, 545, 342, 1060, 463, 1273, 510, 163, 359, 1262, 1469, 486, 1535, 32, 1461, 1641, 647]. II [1241, 692, 1467, 408, 1773]. III [1342, 1084, 1968, 1121, 523, 1196, 1436, 1633, 1778, 1279, 404, 736, 962, 917, 1378, 1841, 451, 1819, 1671, 560, 1502, 1242, 1268]. Illner [1145]. images [260, 49]. Imaging [27]. imbedding [1356]. Immiscible [149]. Imparting [870, 796, 545]. Imperfect [970]. Implementability [963, 1040]. implications [2083]. imply [1800]. Improved [1016]. improvement [979]. impulses [1919]. impulsive [1570, 1706, 1437]. impurity [1912, 59, 1501]. incident [1146, 1147]. including [812, 1237]. incoming [2076]. increasing [609]. Indecomposable [948, 1099, 2120, 1848]. indefinite [993, 412, 504]. independence [1082]. independent [722, 345, 1356, 1996, 785, 1336, 1586]. independent-particle [1336]. index [1137, 999, 38, 1163, 1952, 1842, 334]. Indices [1158, 391]. induced [1769, 567, 539, 18]. Inelastic [1909]. Inequalities [257, 450, 1409, 1261, 1262, 10, 461]. inequality [2118, 1172]. inequivalent [233, 215]. inertial [1133]. Infeld [1904, 2072]. Infinite [1411, 1264, 785, 1702, 1389, 1354, 1537, 67, 466, 50, 1866, 484, 1344, 1970, 1636, 235, 879, 1570, 1911, 940, 1337, 887, 651, 1515, 1766, 1614, 1879, 1715, 1927, 829, 1159, 1749, 1530]. Infinite-dimensional [1264, 1702, 1389, 1354, 1866, 1344, 887, 1879, 1749]. infinite-order [235]. infinite-particle [1970]. infinitely [1878, 1786, 1106, 210]. Infinitesimal [106, 1668, 76, 206, 437, 1739, 470, 471, 736, 1996, 196]. infinities [866]. infinitude [704]. infinity [2004, 2005, 1869, 1594, 1628]. inflationary [696, 1661, 1543]. influence [326, 1993]. information [1321, 603]. infraparticle [1153]. infrared [194, 284]. inheritance [2075]. inhomogeneities [653]. inhomogeneity [237]. inhomogeneous [475, 286, 179, 1848, 205, 1677, 2050, 1776, 1521, 692]. Initial [1837, 2038, 2052, 1237, 1009, 783, 670]. injection [293]. inner [504, 553, 1672]. instabilities [744]. instability [2051, 406, 1621]. instanton [1972]. instantons [1958, 1977]. integer [1103]. Integrability [2022, 1320, 1641, 68, 661, 1722, 960, 597, 1177, 122, 825, 502]. Integrable [1201, 1429, 469, 1832, 1861, 1801, 886, 1496, 263, 1759, 384, 310, 1949, 1895, 1960, 2001, 1352, 1747, 1281, 210, 1278, 269, 1555, 1317, 481, 1761, 1394, 1951]. Integral [1694, 1518, 1718, 2015, 1802, 512, 1850, 351, 869, 1992, 742, 1690, 1723, 908, 158, 2064, 1584, 1952, 940, 893, 388, 424, 598, 723, 1237, 1466, 1585, 25, 1579, 848, 533, 889, 803]. Integral-spin [1694]. Integrals [112, 111, 1052, 555, 353, 928, 263, 1309, 2110, 2, 240, 592, 593, 9, 327, 264, 389, 585, 474, 903, 1066, 1024, 901, 604, 535, 698, 518, 1771, 1093, 1892, 455, 2000, 1479, 55, 1426]. Integration [1778, 1198, 1881, 1634, 826, 354, 72, 162, 163, 560, 1586, 702, 605, 647]. integrational [2096]. integrations [1833, 300]. integrity [467]. integro [1352, 1747]. integro-differential [1352, 1747]. interacting [1894, 1786, 612, 438, 101, 2083, 776, 1154, 335, 1981, 675, 1363, 392, 2103, 139]. interaction [597, 490, 2105, 945, 454, 226, 866, 790, 1464, 1398, 407, 408, 1805, 1853]. interactions [1801, 1722, 1078, 430, 1647, 961, 543, 33, 1606, 269, 324, 98, 1936, 1433, 2000, 1957, 1320, 1641]. interelectronic [2107]. interface [1570, 238]. interfaces [911, 326, 1725]. Interference [1522, 1831]. Interior [917, 1219]. intermediate [1646, 1871]. Internal [188, 894, 1030, 390, 164, 1109, 1519, 1275]. interpretation [157, 360, 173, 1200, 226, 846, 341, 274, 1260]. interpreted [1532]. intertwining [51]. interval [1642, 2109, 939]. Intervals [2051]. Intrinsic [753, 1097, 164, 135]. Introducing [1580]. invariance [953, 1029, 56, 1764, 1005, 1181, 1709]. Invariant [1681, 715, 318, 1090, 1190, 190, 1356, 226, 1103, 1568, 1313, 1239, 300, 899, 1246, 506, 301, 2084, 1130, 511, 1466, 1189, 533]. Invariants [1553, 23, 1850, 1573, 673, 1018, 1019, 1288, 1179, 43]. Inverse [1650, 59, 859, 1256, 60, 27, 1490, 653, 83, 129, 1319, 1455, 954, 732, 1428, 1816, 1725, 726, 727, 609, 962, 2044, 1129, 1033, 2133, 1027, 436, 419, 650, 1318, 738, 1643, 848, 729, 1132, 1805]. inversion [27, 241, 587]. investigation [1243, 1179]. inviscid [2062]. involution [1874]. involutional [329]. involutions [1162]. involving [959, 592, 593, 1103, 181]. Ion [202, 1312]. Ion-acoustic [202]. ionic [1624, 1155]. ions [413, 100]. Irreducible [1966, 1967, 1968, 883, 1420, 1916, 156, 813, 1013, 1505, 187, 298, 17, 1509, 850, 1814, 2137, 18, 250, 766, 1271, 716, 551]. irregular [115]. irreps [16]. irreversible [1903]. isentropic [2062]. Ising [1149]. Ising-type [1149]. Islands [924]. isogroup [1102]. isolated [1293, 780, 1331]. isometries [1275]. isometry [1935]. isoscalar [852, 1669]. isospectral [1112]. isospin [570]. Isotropic [363, 150, 1598, 678, 1021, 1286]. isotropy [824, 76, 1800]. isovectors [627]. iteration [1625]. iterative [1810, 1955]. itself [1617]. IV [1016, 1430, 1996, 1719, 458]. Iwasawa [386].

J [488, 1544, 1983, 1379, 151, 793, 2128, 1660, 1416, 2132, 758, 2139, 47, 654, 546, 958, 2138, 759, 1231, 1465, 240, 327, 2135, 1661, 1085, 1021, 1624, 547, 328, 2142, 832, 2053, 794, 870, 1339, 1886, 509, 823, 1118, 1095, 1947, 2126, 2131, 795, 1747, 1789, 2137, 48, 2141, 2133, 697, 698, 152, 2127, 548, 796, 2140, 984, 1513, 1512, 2136, 1025, 2134, 1232, 1691, 1304, 1442, 461]. J.-M [1085]. Jacobi [1857, 735, 1636, 1061, 414, 501, 966, 1556, 208]. Jacobi-matrix [966]. Jacobian [1656]. Jacobson [69]. Jauch [1251]. Jet [28]. Joint [744]. Jordan [1364, 1272]. Josephson [643]. Jost [1362]. Julia [2009]. jump [500, 78].

Kac [1739, 205, 1979, 70, 1914]. Kadomtsev [662, 886, 1427, 340, 1657]. Kähler [1634, 685, 1927]. Kahn [909]. Kaluza [95, 185, 938, 1364, 97, 364, 969, 816, 306, 1709, 1220]. KAM [1209]. Kantowski [696, 217]. Kasner [2102]. KdV [1167, 1243, 1639, 1189, 1390]. KdV-invariant [1189]. KdV-type [1167]. Kepler [895, 30, 701]. kernel [622, 1388, 1821, 1518, 2126, 1991, 1885, 2131, 1292, 130]. kernels [941, 49, 994]. Kerr [758, 795, 575, 1139, 173, 37, 2041, 1367, 638, 1468, 524, 1595, 274, 1042, 1079, 1898]. Killing [2132, 1139, 2075, 628, 921, 1386, 2041, 713, 905, 445, 1683, 2127, 2007, 2037, 1295, 861, 711, 2074]. kind [742]. Kinematic [905, 928]. kinematical [824, 1620, 1444]. Kinematics [235, 1222, 896, 1396, 1227]. kinetic [1794, 1565, 1880, 2060]. kink [1151, 452]. kinks [1731, 1159]. Kirkpatrick [889]. Klein [1364, 807, 185, 938, 97, 1151, 364, 1238, 969, 95, 816, 306, 1709, 1220, 1393, 1835]. knots [2014]. known [503]. Komar [474]. Korteweg [1855, 2063, 843, 1351, 1689, 855, 856, 591, 206, 1112, 857, 1697, 1384, 502].

Kotera [1999]. KP [195, 1890]. Kramer [1044]. Krawtchouk [2090]. Kreın [1506, 34, 1137, 1307]. Krichever [1847]. Kronecker [1196]. Kronig [1912].

L [832]. label [333]. labeling [761, 379, 654, 243, 380]. laced [786]. Ladder [171]. Lagrangian [900, 1494, 1495, 185, 976, 1931, 1263, 2079, 1925, 2114, 1064, 908, 1765, 1872, 2140, 1318]. Lagrangians [1491, 1184, 1549, 1840, 2099, 1864]. Lanczos [1896]. Landau [1707, 1190, 1982, 34]. Langevin [722, 899]. language [328, 14]. Laplace [500, 71, 149, 250, 78]. Laplace-type [250]. Laplacian [22]. Large [1810, 807, 1164, 120, 634, 92, 684, 1432, 514]. Large- [1810, 807, 120, 634]. laser [487, 293, 652]. latent [224]. lateral [114]. lattice [459, 1970, 2013, 772, 321, 1699, 450, 376, 1074, 1911, 1012, 580, 910, 979, 368, 285, 2080, 247, 279, 387, 139, 1304, 972]. lattices [308, 1085, 38, 373, 428, 579, 1911, 1910, 645, 1555, 1937]. law [906]. laws [124, 50, 1481, 627, 1069, 1451, 1161, 1917, 2122, 1845, 210, 1557, 2037, 633]. Lax [1855, 1161, 311, 855, 1390]. Lax-pair [311]. Layer [729, 728, 1570]. Layer-stripping [729]. layered [1355, 1714]. leading [1729]. Lee [1690, 1086]. Legendre [112, 111]. Leibbrandt [1309]. Lemaître [574, 574]. lemma [69, 1897]. length [280]. lens [260, 691]. less [1254, 1046]. level [1498, 1359, 101, 434, 186, 652, 392]. levels [201, 1178, 1467]. Levinson [1078, 314]. Lévy [996]. Lewis [1738]. Lie [488, 239, 1574, 1127, 1101, 1948, 1346, 883, 1420, 68, 1791, 1848, 621, 1572, 331, 464, 1380, 491, 168, 1486, 762, 763, 1264, 1949, 52, 1790, 140, 281, 649, 1272, 378, 492, 1234, 1612, 203, 590, 2094, 1447, 301, 334, 267, 766, 950, 1235, 1701, 1916, 1434, 1417, 591, 664, 248, 465, 1090, 1194, 1512, 1950, 767, 386, 155, 1562, 1516, 107, 1059, 1632, 156, 885, 1232, 951, 1720, 1383, 1462, 1695]. Lie-admissible [1380]. Lienard [81]. life [521]. Lifshiftz [1707]. lifts [1371]. light [925, 1110, 1431, 585, 1867, 812, 790, 1479]. light-cone [585, 1479]. like [330, 456, 1113, 33, 352, 835, 1635]. Limit [1687, 1593, 120, 1034, 466, 1732, 1104, 403, 404, 417, 1065, 92, 1981, 1012, 280, 891, 269, 198, 1400, 1194, 1512, 1076, 559]. limits [920, 644, 1648]. line [2023, 1231, 965, 1496, 961, 269]. Linear [943, 1067, 1082, 2086, 839, 2060, 2076, 1345, 1574, 468, 958, 2138, 838, 1762, 1488, 1864, 539, 902, 2061, 1919, 506, 301, 1235, 1701, 1851, 1932, 1556, 633, 632, 1383, 75, 82]. Linearization [735, 158, 801, 594, 744, 1182, 1482, 1068]. linearized [286, 204, 1875]. Linearly [19]. lineshape [101]. linked [73]. links [1939]. Liouville [1947, 314, 1658, 1238, 1129, 835, 1635, 801]. Liouville-like [1635]. liquid [1269]. liquids [1311]. little [798, 2017]. Lobachevskian [677]. Local [1774, 1182, 717, 2025, 1240, 1241, 1242, 1483, 942, 1308, 1003, 958, 2138, 296, 841, 177, 1117, 59, 2052, 1036, 198, 393, 1626, 75, 840, 1853]. Localizability [2078]. localization [2019]. localized [1912, 1259, 844, 1579, 1147]. Locally [331, 1486]. Location [1646]. locations [147]. Lochak [1085]. log [1846]. log-normal [1846]. Logarithmic [704, 1711, 212]. logics [13]. Long [1894, 1612, 1191, 268, 597, 1395, 521, 1001]. Long-range [1191, 268, 1001].

Long-time [1894]. long-wave-short-wave [597]. Loop [1381, 995, 662, 1309, 1599, 751, 1052, 585, 1291, 1439]. loops [805, 1789]. López [2133]. Lorentz [984, 1513, 66, 1249, 1088, 297, 2086, 1181, 1250, 2141, 510, 2017, 669, 670]. Lorentzian [1188, 1797]. Lorenz [1148]. losses [388]. Low [1737, 867, 1632, 1968, 1866, 1073, 1362, 1994]. Low-density [1737]. Low-dimensional [1632, 1866]. Low-energy [867]. low-temperature [1968]. Löwdin [159, 486, 1025]. lower [2003, 109, 535]. lowest [1667]. lumps [141]. Lund [888]. Lyapunov [531, 1809].

M [1465, 1085, 832, 870, 2133]. Madelung [1101, 459]. Madivanane [329]. magic [1624, 1155]. Magnetic [37, 1900, 234, 396, 372, 2118, 1980, 447, 895, 30, 695, 1541, 1788, 1439, 1007, 359, 170, 1186, 1937]. Magnetized [526, 758, 575, 86]. magnetogas [2128, 693]. magnetohydrodynamic [440]. magnetohydrodynamics [1818, 1441]. magnetoplasma [1312]. magnetotellurics [587]. Maillard [1085]. Malhiot [1534]. Mandelstam [1309]. manifold [1388, 1630, 2116]. manifolds [1306, 1567, 1011, 258, 1326, 1797, 628, 2057, 2135, 1698, 731, 975, 771, 2006, 1389, 1830]. Mann [1516]. many [1878, 1786, 1647, 1073, 1236, 1106, 1976, 1012, 270, 210, 162, 163, 560, 1433]. many-body [1073]. many-component [1012]. many-electron [1236]. many-time [1976]. map [492, 564, 303, 1273, 1795]. Mapping [197, 320]. mappings [470, 471, 736, 1996, 471]. maps [1354, 2028, 1103, 1067, 767, 1456]. Marchenko [83, 129]. Markov [959, 52, 343, 1047]. Markov-type [52]. Markovian [1034, 1126]. maser [1230]. Mass [1049, 30, 1217, 1289, 1431, 1290, 86, 1211, 412, 683, 1339, 990, 781, 639, 640, 1432, 849, 1994]. masses [181, 752, 368, 285, 1220]. Massive [1628, 1483, 1226, 1064, 140, 281, 1975, 577, 750, 1806, 686]. Massless [636, 1869, 64, 721, 1931, 788, 1482, 1089, 1052, 1421, 455, 29]. Mastersymmetries [2080]. matching [377]. material [218]. Math [488, 1544, 1983, 1379, 151, 793, 2128, 1660, 1416, 2132, 758, 2139, 47, 654, 546, 958, 2138, 759, 1231, 1465, 240, 327, 2135, 1661, 1085, 1021, 1624, 547, 328, 2142, 832, 2053, 794, 870, 1339, 1886, 509, 823, 1118, 1095, 1947, 2126, 2131, 795, 1747, 1789, 2137, 48, 2141, 2133, 697, 698, 152, 2127, 548, 796, 2140, 984, 1513, 1512, 2136, 1025, 2134, 1232, 1691, 1304, 1442, 461]. Mathematical [425, 822, 917, 1001, 1700, 1203, 2101]. matrices [799, 67, 1353, 1164, 472, 494, 1196, 1418, 253, 339, 768, 1719, 385, 103, 1654, 1434, 212, 1487, 720]. Matrix [1609, 754, 193, 332, 1257, 1690, 1340, 1833, 1446, 1445, 2098, 966, 269, 659, 1186, 989, 10, 461, 1605, 1844]. matter [1078, 1259]. Maximal [1173, 1584, 797, 1959, 1321, 399, 6, 547, 1813]. maximally [1262]. Maximum [1062, 1757, 1920, 703, 556, 199, 893]. maximum-entropy [893]. Maxwell [222, 1139, 2031, 1116, 926, 1644, 57, 276, 1032, 516, 1033, 577, 1944, 1470, 2074, 2008, 671, 1705]. Maxwellian [2052]. Maxwellians [681]. may [1971]. McLaughlin [1945]. McVittie [936]. Mean [737, 179, 1645, 2125, 450, 238]. mean-field [179, 1645, 2125, 238]. meaning [287, 826]. means [158, 127, 1077, 1393]. measure [1809, 353, 1599, 299, 1907]. measurements [350, 603, 2002, 741]. measures [109, 928, 1171]. mechanical [1525, 1922, 353, 1463, 296]. mechanics [1727, 515, 1251, 2121, 1878, 1908, 350, 1174, 1175, 1107, 1499, 89, 1137, 391, 1647, 490, 925, 1110, 673, 1862, 1031, 537, 84, 1254, 431, 706, 1460, 2078, 2077, 739, 1889, 566, 442, 747, 1527, 409, 1172, 1283, 663, 2061, 1765, 91, 270, 1322, 1323, 1997, 1766, 1036, 1179, 102, 2104, 1284, 1558, 58, 511, 1466, 107, 1998, 1461, 393, 559, 741, 608, 806]. mechanism [978, 1648]. media [137, 60, 947, 286, 1355, 1429, 512, 345, 1356, 1677, 540, 2050, 427, 1714, 1521, 1459, 2070, 2060]. medium [237, 653, 2109, 149, 839, 1621, 867, 1932, 737, 1691, 1350]. medium-range [867]. Meixner [1820, 2090]. Mellin [2126, 1991, 1885, 2131]. Membranes [1785, 2124, 123]. memory [1303]. memory-time [1303]. mer [748]. metaplectic [493]. metastable [405]. method [1757, 1920, 466, 1995, 1410, 500, 1100, 165, 263, 322, 714, 1438, 423, 539, 821, 655, 1118, 1625, 893, 270, 966, 1932, 1130, 78, 1556, 1656, 401, 1805]. methods [1863, 541, 923, 345, 1356, 2116, 2119, 1047, 123, 1179, 1316, 848]. metric [811, 758, 575, 484, 1296, 838, 1901, 37, 1661, 1543, 993, 1180, 863, 1366, 936, 1367, 1886, 1181, 505, 1796, 1859, 1705, 1192]. metric-affine [484]. metrics [2132, 1139, 1532, 1900, 713, 76, 1684, 2112, 477, 1709, 1776, 2020, 526, 1927, 809]. metrizable [1612]. Meyer [1886]. MHD [757, 791]. microeffects [1826]. Microscopic [2049, 643]. Mills [423, 584, 1244, 1600, 842, 1308, 1477, 1226, 957, 1087, 1124, 817, 815, 831, 1347, 194, 409, 2044, 1658, 1947, 1875, 1475, 973, 601, 1713, 1376, 562, 2133, 1503, 1372, 1102, 1375, 1954, 2134, 1812, 1282, 1631]. Mills-type [194]. Milne [25]. Minimal [1659, 133, 1173, 1734, 28, 694, 1360, 1708]. minimally [675]. Minimization [120, 1190]. minimizing [747]. minimum [1920, 534, 845]. minisuperspace [138, 528, 1778]. Minkowski [1399, 57, 1748, 181, 752, 80, 81, 1800]. Minkowskian [1211]. misleading [400]. mismatch [962]. missing [333, 1349]. Mixed [295, 1593, 168, 1650]. mixture [2049, 91, 710]. mixtures [1474]. mKdV [1212]. mKdV-type [1212]. Mobility [741]. Möbius [1291]. Mode [1898, 1124, 478, 542]. model [1878, 1786, 1942, 1165, 617, 1187, 476, 460, 1905, 1946, 1413, 1598, 472, 205, 1145, 1824, 1970, 101, 418, 1655, 1178, 325, 415, 892, 1623, 875, 1230, 696, 1690, 1778, 1411, 751, 140, 281, 649, 785, 879, 1154, 452, 2001, 1259, 1981, 1012, 644, 730, 1717, 1945, 1191, 294, 324, 98, 1936, 1932, 1715, 366, 1138, 1780, 1957, 1994, 1686, 1398, 1394, 1830, 1467, 1710, 1962, 1693, 1985, 686, 1504]. modeling [293, 1459, 2070, 1738]. models [1538, 1306, 1567, 1425, 1537, 689, 2073, 141, 1829, 5, 47, 1629, 544, 1794, 1565, 1364, 2049, 1149, 1659, 1698, 1599, 1085, 714, 1620, 1264, 1949, 1960, 1482, 890, 1493, 975, 1775, 866, 1813, 1050, 1192, 1193, 579, 1118, 1533, 1191, 913, 602, 2119, 1640, 2024, 1080, 588, 1819, 416, 283, 820, 1740, 449, 215, 217, 949, 1125]. modes [1168]. Modified [340, 438, 1999, 1971, 954, 592, 593, 1092, 1351, 1639, 389, 591, 448, 588, 206, 778, 115, 1697]. Modular [1987, 897, 1978, 1990]. modules [1949, 1099, 1726]. molecular [663, 645, 760]. molecules [790]. Molien [242]. Möller [1216]. Moloney [1945]. Moment [512, 1441, 234, 396, 1349, 1106, 1472, 1439]. Moment-equation [512]. momenta [264]. moments [2027, 919, 1322, 256]. momentum [1587, 1863, 967, 1139, 178, 126, 211, 904, 1018, 706, 1180, 1099, 1019, 1523, 675, 1360, 365, 1854, 2081, 782, 3, 697, 1534, 355, 1213, 1022, 1479, 738, 1815, 1742, 1286, 10, 461]. monomer [279]. monomer-dimer [279]. monomode [1842]. monoparametric [406]. monopole [1600, 234, 1133, 483, 895, 176, 607, 945, 701, 1875, 26, 1972]. Monopoles [1958, 447, 536, 789]. monotonic [1103, 1603]. Montroll [971]. Moody [1739, 1979, 70, 1914]. Morozov [69]. Morphisms [49]. Morse [998, 1326, 1035, 1526]. most [758, 575, 659, 1962]. motion [1133, 1231, 965, 1759, 527, 183, 79, 563, 509, 2009, 1956, 1625, 1741, 1841, 1299, 1248, 1370, 1997, 2062, 778, 1086, 11]. motions [1681, 1069, 1296, 317, 363, 134, 1109]. moving [271, 8, 152, 1787]. Moyal [569, 1461]. MR [1983, 1379, 793, 2128, 2132, 2139, 47, 546, 958, 2138, 1231, 1465, 240, 327, 2135, 1085, 1021, 1624, 547, 832, 2053, 794, 870, 1339, 1886, 509, 1118, 1095, 1947, 2131, 1747, 1789, 2137, 48, 2141, 697, 2127, 2140, 984, 1513, 1512, 1025, 2134, 1232, 1691, 1304, 461, 1442]. Mullins [1946]. Multi [1904, 1278]. Multi-Hamiltonian [1904, 1278]. multicomponent [1409, 1621]. Multidimensional [1132, 960, 1629, 732, 506, 1555, 682, 1225, 848, 729]. Multigroup [720, 252]. multilevel [1414, 1679]. multilinear [1090]. Multiparameter [1004]. multiparticle [2098]. Multiple [1934, 254, 255, 35, 87, 1320, 1641]. Multiple-soliton [1934]. multiplets [1734]. Multiplication [1438, 2030, 1397]. Multiplications [1385]. Multiplicative [959]. multiplicities [420, 553]. Multiplicity [762, 763, 17, 999, 852, 381, 16]. Multiplicity-free [762, 763, 17, 852, 381, 16]. Multiply [233]. Multipole [2027, 1594]. Multiseries [1949]. Multispinor [721]. multitype [93, 280, 451]. Multivariable [1546, 1799, 2090, 1989, 747]. must [1404]. mutual [777].

N [1379, 2128, 1624, 832]. Nahm [825, 44]. natural [172, 548]. nature [1540]. Navier [1762]. near [236, 1963, 2088, 1193]. Near-field [1193]. Nearest [1911, 772, 580]. nearly [2064, 563]. Necessary [1085, 579, 395, 261, 556]. neighbor [772, 1911, 580]. nets [1626]. Neugebauer [1044]. Neumann [1358, 1937]. neutral [234, 1081, 1469, 1502]. Neutrino [1653, 369, 577, 872]. neutron [586, 252]. neutron-flux [252]. Newell [1945]. Newlander [1860]. Newman [758, 575, 1139, 173, 37, 1211, 1367]. Newton [83]. Newtonian [1008, 1300]. Nijenhuis [481]. nilpotent [184]. Nirenberg [1860]. no [1983, 1379, 793, 2128, 2132, 2139, 47, 546, 490, 958, 2138, 1231, 1465, 921, 240, 327, 2135, 1085, 1021, 1624, 547, 832, 2053, 794, 870, 404, 1339, 890, 1886, 509, 1118, 1095, 1947, 2131, 1747, 1789, 2137, 48, 2141, 2133, 697, 2127, 2140, 984, 1513, 1512, 1025, 2134, 1232, 1691, 1304, 1331, 461]. no-collision [404]. no-interaction [490]. nodal [337]. Noether [1905, 1069, 429, 1637]. Noether-type [1069]. noise [1993, 1047]. Non [1213, 708, 787, 819, 563, 1630, 95, 1126, 602]. Non-Abelian [1213, 708, 819]. non-Ermakov [602]. Non-Hamiltonian [787, 563]. non-Markovian [1126]. non-Riemannian [1630, 95]. Nonabelian [144, 141, 4, 546, 2081, 2014]. nonadditive [1179]. nonassociative [1843]. nonbarotropic [1845]. Noncanonical [1419, 1668]. noncentral [1038, 337]. noncollinear [1218, 1334, 1436]. noncommutative [428]. noncommutativity [1283]. noncompact [1306, 1567, 1752, 1198, 1514, 1667, 2047, 2087, 1125]. nonconnected [1486]. nonconservative [1730, 1556]. nonconstant [1504]. Noncoordinated [344]. nondiagonal [576, 720]. Nonequilibrium [1123, 92]. Nonexistence [1119, 173, 1907]. nonfractal [996]. nonintegrable [1393]. noninteracting [710]. Noninteraction [566, 567]. nonisospectral [2063]. Nonlinear [797, 988, 1197, 252, 383, 462, 1633, 1280, 774, 1301, 1464, 488, 239, 1818, 106, 932, 1306, 1173, 1187, 1101, 63, 1918, 661, 50, 1946, 930, 865, 1391, 1801, 1722, 1939, 1481, 960, 1598, 1903, 1177, 205, 491, 923, 1590, 1974, 1577, 1100, 23, 1151, 753, 1097, 254, 582, 94, 1623, 1210, 1039, 1075, 114, 2114, 1960, 683, 1339, 990, 975, 158, 167, 203, 1696, 1712, 1808, 564, 1940, 781, 1374, 1238, 1352, 1747, 1281, 1963, 913, 899, 1842, 210, 1278, 1588, 887, 2119, 888, 1635, 2048, 73, 1221, 2084, 1237, 737, 844, 1389, 1740, 1083]. nonlinear [1105, 1076, 1955, 741, 255, 889, 1962, 1805]. nonlinearities [865]. Nonlinearly [1527, 19]. Nonlocal [281, 1482, 1375, 430, 778]. Nonnegative [919]. Nonorientable [1291]. nonorthogonal [665]. nonperiodic [1415]. Nonperturbative [1016, 1577]. nonplanar [1691, 1350]. Nonpolynomial [1308]. Nonpropagating [2063]. nonquadratic [936]. nonrelativistic [84, 866, 2058, 1780, 608]. nonresonant [1801]. Nonrotating [822]. nonseparable [1929]. nonseparated [1647]. nonsingular [1031, 948]. Nonstatic [2007]. nonstationary [874, 756]. Nonsymmetric [1982, 1865, 1563]. Nontrivial [190, 82, 330, 930, 656, 245, 863, 1366]. Nontriviality [1084]. Nontypical [950, 1701]. nonuniformities [1855]. nonunitary [2085, 17]. nonzero [1289, 1290, 1406, 639]. Nordström [1369]. Nordtvedt [476]. norm [1920]. Normal [1377, 398, 743, 1888, 1100, 1775, 1527, 1846, 542, 760]. normalized [1178, 1467]. note [1341, 287, 611, 1828, 1258, 860, 1945, 355, 1856, 680, 16]. Notes [1492, 1663]. notion [724]. novel [9, 327, 1426]. Novikov [1847]. nuclear [617, 418, 2049, 1603]. nuclei [1053]. nucleon [1022]. Null [1957, 1149, 1869, 57, 76, 1594, 1405, 1827, 1295, 2025, 1628]. number [182, 12, 151, 703, 1411, 879, 260, 1702, 704, 1580, 1715, 1927]. numbers [1785, 1624, 2092, 1155, 1456, 1397]. Numerical [162, 163, 560]. NUT [2035].

O [832]. object [638]. objects [105, 1629]. observability [1115]. observables [1321, 1171, 1783, 603, 1766, 393, 608, 1626]. Observation [1834]. observations [500]. observer [925, 1110]. observers [1838]. obtained [1939, 576, 1435, 1225]. obtaining [38]. occupation [910]. occupied [580]. Occurrence [497]. occurring [1024]. octonion [7, 48]. Octonions [1792, 477]. odd [821, 260, 1053]. odd-dimensional [821]. odd-odd [1053]. O’Doherty [631]. offset [631]. old [1734, 352]. old-minimal [1734]. ondelettes [2019]. One [2066, 1398, 488, 239, 106, 1345, 1574, 1888, 120, 118, 1429, 999, 430, 968, 1146, 1147, 1309, 308, 1018, 1349, 317, 363, 134, 1255, 609, 1019, 1817, 375, 587, 1741, 1336, 1363, 59, 8, 152, 709, 1291, 1586, 737, 538, 266, 1439, 1459]. one- [308, 1741, 709, 266]. One-dimensional [1398, 488, 239, 106, 1345, 1574, 1888, 120, 118, 1429, 968, 1146, 1147, 1018, 1349, 1255, 609, 1019, 1817, 375, 587, 1336, 1586, 737, 1459]. one-loop [1309, 1291, 1439]. one-parameter [317, 363, 134, 538]. Ono [1161, 1427]. open [1797, 458, 1291]. operating [331, 1486]. operation [350, 1034]. operation-valued [1034]. Operator [250, 1571, 1346, 129, 101, 31, 1793, 1040, 1893, 71, 1327, 1928, 2123, 1103, 1843, 38, 1017, 1602, 1926, 754, 1399, 33, 1171, 224, 880, 127, 166, 128, 2080, 1555, 1111, 1529, 643, 1204, 1377, 32, 2065, 1695, 1835, 550, 618]. operator-valued [1171]. Operators [3, 697, 118, 244, 1756, 421, 187, 1325, 1215, 1380, 1216, 333, 2105, 51, 1418, 715, 1128, 1039, 6, 547, 257, 311, 918, 1438, 1724, 504, 194, 1399, 1668, 726, 962, 167, 1421, 1489, 299, 1246, 301, 834, 1414, 1679, 1607, 1950, 1672, 1195, 1673, 1767, 107, 498, 481, 1223, 266, 840, 1286]. Oppenheimer [1141]. optical [1354, 983, 573, 1541, 1945, 1842]. optics [2054, 2055, 606, 1960, 692]. options [1992]. orbifold [1710]. orbit [270]. orbital [1360]. orbitals [1937]. orbits [1681, 2022, 977, 174]. order [1763, 894, 437, 974, 1491, 1494, 195, 1890, 1184, 1634, 251, 304, 732, 1488, 873, 1870, 1340, 1941, 1840, 567, 235, 171, 470, 471, 1996, 1351, 626, 92, 1696, 2094, 882, 691, 535, 790, 1988, 935, 733, 1318, 1449, 1252, 916, 514, 692, 1697, 1384, 1705, 1906, 1030]. Ordered [1877, 920, 1104, 2110, 1128]. ordering [606, 1000, 1992, 1782, 1783]. orderings [1571]. orders [2035]. ordinary [661, 251, 254, 1428, 1488, 310, 567, 2094, 733, 255, 494]. orientation [1424]. Origin [1495, 781]. orthochronous [1800]. orthogonal [1373, 615, 1545, 1438, 1203, 1382]. orthogonalization [1132]. orthonormal [1666, 1431, 1798]. orthosymplectic [1559, 1569, 1948, 1754, 884, 764, 1200, 301]. Oscillating [681, 209]. oscillations [1750, 1699, 375]. oscillator [1727, 743, 1894, 1786, 201, 1942, 1559, 998, 1569, 1948, 1959, 1169, 1525, 967, 517, 1553, 1862, 1463, 1327, 165, 705, 896, 903, 1066, 3, 697, 1047, 1771, 635, 2033, 2100, 2136, 1929, 1619, 1768, 605, 1286]. oscillators [1642, 1893, 167, 1527, 1772, 564, 1704]. other [1096, 1822]. outer [553]. overlap [2003]. Ovsiannikov [1130].

P [2128, 1085]. packet [2109]. packets [1926, 375]. Padé [1704, 200]. Painless [665]. Painlevé [1855, 1457, 308, 626, 1712, 855, 1435, 122, 700, 1060, 1671, 801, 340, 53, 668]. pair [1161, 1893, 154, 311, 403, 417]. paired [1022]. pairs [1855, 580, 855]. paper [1379, 1465, 1085, 1624, 832, 509, 823, 1304]. papers [870]. para [478, 532]. para-Bose [478, 532]. parabolic [1038, 297, 966, 146]. parallax [1620]. parallax-free [1620]. Parallel [713, 209]. Parallel-propagated [713]. parameter [1359, 1174, 1175, 1103, 317, 363, 134, 1603, 1066, 1055, 538]. parameter-dependent [1359]. parameters [1971, 1146, 1147, 1766, 348]. Parametrization [902, 1410, 798, 550, 618]. parametrized [2059, 305]. Parastatistics [1236, 1839, 1346, 271]. Parent [1159]. parentage [1524]. Partial [1418, 1875, 625, 1660, 1450, 991, 1634, 1984, 1148, 1672, 1389]. partially [1581, 1130, 1083]. particle [234, 719, 1078, 1838, 1970, 490, 1216, 667, 169, 1880, 28, 1840, 1357, 20, 1745, 432, 1976, 1336, 280, 1993, 704, 1580, 2103, 1284, 1558, 829, 1781, 1909, 266]. particle-hole [1745]. particle-matter [1078]. Particlelike [1588]. particles [894, 1030, 2069, 1289, 1678, 703, 1290, 1146, 1147, 84, 847, 522, 1113, 296, 961, 376, 412, 776, 1089, 909, 1136, 181, 752, 1363, 910, 931, 8, 356, 152, 1760, 1787, 1806, 455, 625, 1815, 1742, 1223]. particular [305, 596]. Partition [725, 907, 1708]. passage [705]. passing [1167, 1212, 1279, 1430]. past [925, 1110, 1431]. Patel [1661, 1543]. Path [1198, 1881, 1980, 1307, 351, 908, 605, 2015, 2096, 1802, 512, 353, 928, 2110, 1992, 2, 240, 742, 1690, 2099, 1952, 901, 604, 723, 1466, 1957, 1907, 803]. Path-dependent [1980, 2099]. path-integral [512, 1690]. path-integral-Riemann-space [803]. Path-space [1307, 1907]. paths [708]. pattern [762, 763]. Pauli [721, 1434]. PCT [755, 43]. PDEs [1801, 1722, 1635]. Pedestrian [1234]. Pell [1314]. pencil [1867]. penetrable [1, 1442]. Penney [1912]. Penrose [1211, 639, 574]. Percus [749]. Perfect [134, 588, 1046, 1326, 1333, 358, 1183, 1069, 277, 2113, 316, 357, 1933, 1365, 132, 678, 1021, 177, 1219, 1468, 795, 524, 1534, 1408, 359, 1562, 1081, 1469, 1502, 642, 1535, 809]. perfect-fluid [1468, 795, 524]. perfectly [690]. Periodic [1713, 857, 1202, 1049, 1412, 968, 326, 2051, 1270, 563, 1919, 270, 59, 1048, 1795]. periodically [2071]. periodicity [225]. perspective [1638]. pertaining [422, 682]. Perturbation [109, 452, 1062, 1287, 1151, 400, 522, 213, 203, 563, 1004, 1005, 2035, 2071, 1047, 102, 212, 2100, 2136, 514, 1464, 1704]. perturbations [1395, 1004, 712, 711, 2008, 2076]. Perturbative [1973, 1176, 1918, 466]. perturbed [812, 1648, 347]. Petrov [795, 1183, 1468, 524]. Petviashvili [662, 886, 1427, 340, 1657]. Pfaffian [1665]. Phares [1304]. Phase [2019, 869, 1806, 1744, 1958, 749, 997, 1124, 2, 240, 1803, 1460, 1438, 313, 1170, 149, 890, 1493, 727, 962, 589, 1322, 261, 1213, 1585, 336, 1461, 559, 1937]. Phase-integral [869, 1585]. phase-plane [313]. phase-space [2, 240, 1460, 1438, 1322, 1461, 559]. phases [1618, 1715]. phenomena [1729, 1880]. Phenomenological [1229, 1932]. phenomenon [1598, 200]. phonon [1968, 452]. photoionization [1287]. photon [948, 369]. photon-Weyl [948]. Phys [488, 1544, 1983, 1379, 151, 793, 2128, 1660, 1416, 2132, 758, 2139, 47, 654, 546, 958, 2138, 759, 1231, 240, 327, 2135, 1661, 1021, 1624, 547, 328, 2142, 832, 2053, 794, 870, 1339, 1886, 509, 823, 1118, 1095, 1947, 2126, 2131, 795, 1747, 1789, 2137, 48, 2141, 2133, 697, 698, 152, 2127, 548, 796, 2140, 984, 1513, 1512, 2136, 1025, 2134, 1232, 1691, 1304, 1442, 461]. Phys. [1465, 1085]. Physical [1214, 537, 972, 1133, 173, 2083, 936, 585, 274, 1224, 215]. physics [617, 479, 104, 793, 1987, 1203, 1024, 1580, 1422, 381, 107]. picture [487, 407]. Pin [1424]. pinning [1542]. Piron [1251]. Planar [66, 1054, 1519, 1269, 1437, 292]. Planck [118, 1920, 310, 682, 465, 1383]. Plane [278, 1891, 357, 41, 2031, 1862, 1311, 854, 1218, 1334, 1436, 1231, 965, 1837, 2038, 540, 313, 373, 1570, 2112, 645, 1393, 55, 2076, 1108]. Plane-fronted [278]. plane-symmetric [2112]. plasma [485, 1511, 426, 660, 1093, 2058, 436, 366, 542]. plate [209]. plus [1802]. Pohlmeyer [888]. Poincaré [1764, 1153, 1538, 1127, 439, 1197, 2085, 715, 567, 1265, 1421, 1099, 564, 18, 2048, 1800]. point [1763, 1587, 1894, 612, 379, 654, 430, 1338, 456, 1020, 543, 566, 1696, 2061, 1956, 8, 152, 269, 380, 825, 498]. point-group-like [456]. pointed [1862]. pointlike [26]. points [1354, 955, 1048, 1795, 857, 1202]. Poisson [561, 2028, 2101]. polarizable [1056]. polarization [941, 1218, 1334, 1436]. polarized [1440, 2076]. polaron [1542, 1541, 1057]. Pole [1292]. poles [1892]. Pollaczek [1820, 2090]. polyacetylene [452]. polyacetylene-type [452]. Polyakov [2015, 846, 1480]. Polynomial [1573, 509, 249, 11, 1199, 865, 421, 2022, 1858, 353, 251, 1489, 1189]. polynomial-type [1199]. Polynomials [1571, 1233, 594, 986, 615, 377, 1636, 38, 1438, 1820, 1203, 901, 1546, 1799, 2090, 1989]. polytopes [1750]. Poncelet [1548, 1825]. Popov [846]. porous [237, 947, 1429, 149]. Pöschl [168]. posedness [62, 1738]. Positive [1824, 2087, 114, 1834, 1171, 100, 816, 2067]. positive-energy [816]. Positivity [503, 313, 395]. possess [264]. possessing [377, 658, 1584]. possible [515, 1483, 2077, 1987]. post [1008, 216]. post-Galilean [216]. post-Newtonian [1008]. Potential [634, 1896, 1744, 1101, 1678, 2096, 595, 749, 1802, 1794, 556, 413, 315, 1035, 674, 1146, 1147, 168, 282, 1017, 1602, 701, 756, 776, 1526, 609, 1969, 1336, 520, 59, 1539, 519, 145, 1835, 2107]. potentials [1190, 1002, 2022, 1549, 1003, 1500, 469, 232, 1018, 573, 1038, 623, 1176, 264, 1019, 1650, 1310, 867, 269, 1055, 1317, 778, 1929, 1902, 1159, 633, 1951]. Potts [1012, 971]. Power [906, 911, 815, 1627, 737, 1530, 1773]. power-series [1530, 1773]. powerful [1761]. powers [330, 813]. PPN [1041]. Practical [407, 408, 500]. precession [666, 510, 984, 1513, 208]. predictable [1581]. prediction [324]. predictive [566]. predictor [500, 78]. prescription [1309]. prescriptions [1892]. presence [1135, 209, 962, 1156]. preserving [1188, 1979, 2035]. Press [1072]. pressure [1511]. presymplectic [309]. pretheories [407, 408]. Principal [885, 1271]. principle [1254, 328, 14, 199, 1765, 1845, 1014, 704]. principles [661, 50, 1922, 615, 1451, 491, 28, 429, 818, 1879]. probabilities [515, 2034]. Probability [1287, 909, 370, 2003, 404, 707, 740]. problem [1818, 422, 1008, 915, 1379, 594, 60, 63, 27, 54, 1410, 1597, 472, 761, 2030, 895, 556, 1319, 954, 383, 945, 732, 325, 349, 1428, 243, 1236, 1441, 1349, 1837, 2038, 629, 30, 1816, 1106, 701, 1054, 2114, 609, 663, 2044, 587, 520, 1129, 833, 1024, 388, 516, 1322, 880, 1033, 2133, 1207, 1454, 279, 1027, 1036, 1779, 709, 419, 650, 1237, 1924, 336, 1954, 2134, 946, 1318, 78, 738, 783, 803]. problems [1757, 1918, 862, 1496, 114, 859, 2048, 73, 1932, 1831, 1105, 1955, 848, 729]. procedure [1971, 486]. process [959, 1034, 1846, 93, 280, 451]. processes [959, 350, 1335, 94, 1834, 343, 92, 1474, 1126, 1592, 829, 2117, 806]. produced [441]. produces [260]. product [1498, 1353, 504, 569, 1239, 1804, 1990, 103, 128, 1663, 13, 510, 984, 1513, 1558, 1672, 1462]. products [1196, 2097, 1636, 110, 1022]. profile [1511]. profiles [727]. programme [58]. progressive [144]. projection [243]. projection-based [243]. Projective [43, 1537, 1133, 770, 463]. Projector [1692]. Prolongation [887, 664, 746, 1161, 954]. Proof [1952, 314, 181, 1531, 841, 933, 1988, 750]. propagated [713]. Propagation [540, 570, 512, 345, 1356, 371, 414, 166, 1586, 1301, 734, 348]. propagator [286, 517, 1623, 1497, 723, 676]. Propagators [1064, 1391, 323, 262, 878, 929]. Proper [1608, 1674, 66, 821]. proper-time [821]. properties [1818, 1498, 396, 953, 108, 1029, 507, 827, 868, 1912, 742, 450, 403, 417, 936, 717, 1541, 905, 587, 100, 189, 1126, 1709, 8, 152, 1253, 1679, 1295, 74, 407]. property [131, 531, 1855, 552, 308, 683, 1339, 990, 1712, 855, 122, 801, 734, 340, 53, 668]. proportional [1405]. proving [437]. pseudomechanics [981]. pulsating [1727]. Pure [591, 1593, 1146]. purely [2076].

QCD [286, 485, 1052]. QED [286, 291, 1943, 1439, 407, 408]. Quadratic [1328, 1928, 264, 901, 1614, 1961, 2130, 1859, 605]. quadratic-Hamiltonian [1614]. Quadratures [957]. quadric [1825]. quadrupole [190, 1854, 1009]. qualitative [1156]. quantities [1199, 527, 1411, 879, 1702, 1766]. Quantization [1568, 270, 1316, 1593, 1538, 897, 672, 942, 1015, 480, 1931, 351, 1554, 775, 963, 1220, 1516, 1227, 1476, 1781, 994]. quantized [439]. Quantum [1727, 350, 1321, 353, 1110, 922, 1231, 965, 296, 213, 412, 2083, 896, 365, 1432, 2104, 366, 864, 1227, 1805, 1125, 515, 398, 1251, 1483, 983, 980, 2121, 1878, 1894, 1786, 1728, 1769, 1016, 479, 942, 1908, 530, 1302, 1120, 1121, 390, 915, 1379, 228, 1013, 1505, 2039, 1948, 1711, 1175, 1107, 1169, 179, 1525, 1499, 1137, 1645, 2125, 39, 2040, 391, 2085, 1930, 2139, 1647, 606, 673, 838, 1862, 1463, 868, 2068, 1545, 537, 84, 1254, 1699, 1020, 1460, 739, 1039, 817, 841, 30, 227, 794, 1627, 568, 20, 1109, 1054, 1519, 1895, 1960, 184, 776, 1123, 1834]. quantum [677, 1396, 33, 1527, 425, 822, 917, 1172, 1283, 717, 663, 2061, 927, 414, 752, 2116, 1651, 1024, 270, 1310, 835, 1322, 2071, 1323, 602, 929, 307, 603, 2092, 2095, 707, 1766, 13, 1036, 1779, 1179, 1400, 102, 2103, 1284, 1558, 58, 2002, 1938, 1466, 1566, 1261, 1262, 107, 1461, 180, 453, 647, 393, 559, 1398, 214, 608, 578, 1626, 1114, 1962]. Quantum-mechanical [353, 296]. quark [725]. quark-gluon [725]. quarkonium [982]. quarks [180]. quartic [2022, 635]. quasi [1596]. quasi- [1596]. quasifermion [1745]. quasilattice [1415]. quasilinear [1280]. quasineutral [1511]. quasiparticle [1967]. quasiperiodic [932, 1325, 1817]. Quasiprimary [1149]. Quasiresonance [521]. quasispin [1091]. quaternion [7, 48, 978, 1417, 1726]. quaternionic [477, 1324, 1323, 1195, 1767]. Quotient [1011].

R [2128, 2133]. Racah [656, 190, 82, 930, 1328, 250, 2046, 1695]. radial [572]. Radially [141]. radiated [1587, 873]. radiating [779]. radiation [101, 125, 780, 1522, 443, 812, 392, 26, 710, 1009]. Radiative [2050]. radius [1455]. Radon [27]. Rainich [2113]. Ramanujan [1785]. Random [262, 179, 1164, 512, 345, 2109, 257, 540, 846, 1737, 839, 2050, 93, 280, 451, 690, 1458, 198, 1932, 737, 829]. range [268, 749, 1146, 1147, 522, 35, 748, 1191, 1001, 867, 707]. range- [748]. Rarita [721, 275, 1902, 1777]. rate [2066, 792, 1592, 405]. Rational [1018, 1019, 340, 1853]. ratios [1641]. ray [1234]. Rayleigh [2051, 165, 346, 514, 1464, 272]. reaction [792, 1579]. reaction-diffusion [1579]. reactive [235]. Real [1536, 182, 1582, 1983, 797, 187, 1730, 418, 1588, 269, 550, 618, 664, 1484, 1487, 1632]. Realization [567, 224, 15, 987, 1107, 332, 622, 325, 568, 1745, 492, 289, 34]. realizations [1346, 331, 1793, 1486, 765, 2086, 18, 1574]. Realizing [598]. Reboucas [870]. Reciprocal [427, 1939, 1456]. Reconstruction [513, 727]. rectangular [1164, 38, 1911, 580, 910, 387, 1304, 989]. Recurrence [903, 1771, 455, 1833]. recurrent [2006]. Recursion [1555, 2080]. redistribution [1071]. reduced [1034, 472, 312, 738]. reduces [701]. reducible [552]. reducing [799, 494, 1196, 1719]. Reduction [2111, 897, 617, 1101, 1855, 1373, 1058, 662, 263, 775, 764, 1941, 153, 1421, 584, 1050, 2098, 1975, 1713, 1246, 506, 128, 1225, 358]. reductions [2021]. Reeh [709]. reference [1250, 2141]. refinement [329]. Reflection [728, 1281, 238, 737, 2065]. reflectionless [1159]. reflectivity [1453]. refraction [728]. regeneration [2066, 1592]. Regge [529, 904, 888]. regions [1809, 750]. Regular [399, 1232, 951, 208, 2005, 922, 1750, 1581]. regularization [1829, 284]. regularized [1873, 2129, 1656]. Reissner [1369]. related [804, 1838, 1148, 1960, 1352, 1747, 570, 1874]. Relation [541, 496, 104, 793, 397, 330, 1495, 1730, 202, 869, 232, 271, 128, 643, 1194, 1512]. Relations [1884, 2046, 799, 438, 437, 1107, 2032, 410, 1214, 1072, 31, 924, 494, 1196, 1754, 431, 1039, 688, 1265, 1719, 1833, 939, 1940, 1680, 903, 427, 1771, 1547, 455, 32, 1768]. Relationship [947, 176]. Relativistic [1744, 228, 1589, 358, 1298, 1901, 565, 1340, 175, 660, 1093, 2058, 359, 1186, 1728, 1769, 1942, 2054, 2055, 390, 915, 1379, 1013, 1525, 89, 2120, 1811, 357, 1216, 847, 1113, 282, 1406, 1522, 30, 566, 1560, 927, 675, 1360, 1472, 2080, 1779, 1760, 2103, 1284, 1558, 29, 608, 272, 1046]. relativity [1836, 137, 1353, 439, 1293, 744, 1732, 1183, 808, 316, 1933, 1364, 836, 779, 2027, 400, 1401, 1402, 1407, 663, 1259, 1404, 745, 1222, 1299, 1014, 1868, 782, 525, 1080, 1551, 1297]. Relaxation [1230, 24]. relevant [1685]. remark [1547, 1476]. remarkable [131]. Remarks [552, 2, 240, 184, 1203, 562, 354, 980, 813, 1085, 306, 1944, 889]. Renewal [1474]. Renormalization [1226, 1971, 1330, 1082, 866, 613]. renormalized [1395]. replacement [621]. Representation [1499, 1883, 1978, 1769, 1678, 1891, 1576, 912, 1716, 1485, 213, 1882, 385, 1550, 18, 433, 1422, 723, 1938, 767, 1852, 659, 292, 170]. Representations [1718, 64, 1096, 1313, 70, 1023, 1575, 1559, 530, 1120, 1121, 15, 1197, 883, 1420, 68, 920, 2120, 1848, 2085, 552, 1215, 418, 494, 51, 759, 619, 1675, 1035, 2123, 462, 753, 948, 298, 1157, 296, 384, 17, 1514, 1667, 262, 1790, 539, 1272, 1668, 1719, 492, 1234, 822, 917, 1421, 1446, 1445, 1915, 1987, 590, 1814, 2137, 1158, 18, 929, 267, 766, 950, 1235, 1701, 1916, 2093, 2048, 463, 1604, 2087, 386, 381, 156, 1271, 1605, 1844, 453, 647, 215, 533, 551, 1446, 1445]. represented [1171]. representing [406]. repulsion [1369]. repulsive [703, 961, 520, 651, 970]. requirements [133]. rescaling [23, 1696, 1782]. reservoir [1303]. resistance [92]. resolution [2030]. resolvent [194, 1292, 25, 702]. Resolving [599, 46]. resonance [1019, 1280, 1805]. resonances [234, 1106, 172, 548]. Resonant [1642, 792, 853, 1648]. resonator [71]. resonators [388]. respect [925, 1110, 522]. response [1808]. rest [1289, 1290, 1211]. restricted [1646, 1320, 1641]. restrictions [192]. result [1988]. results [2040, 384, 2101, 1176, 757, 2126, 1885, 2131, 709, 510, 984, 1513, 1083]. retarded [555, 318, 1925]. retrieval [336]. reversal [294]. reversible [1048]. revisited [1157, 1392, 1540]. ribbon [1, 1442]. Riccati [1087, 102, 989]. Ricci [1610, 1116, 771, 642, 1535]. Ricci-flat [771, 642, 1535]. Riemann [1811, 1739, 383, 1609, 1882, 880, 1244, 946, 1670, 803]. Riemannian [1388, 599, 731, 1630, 7, 95, 48, 2006, 1043, 1830]. rigged [125]. rigid [277, 1299, 2052, 175]. rigidity [218, 637, 1041]. rigidly [1365]. Rigorous [1209, 1092, 757]. Rindler [1229]. Ring [1804, 1354, 493, 1945]. rings [937]. Robertson [1661, 811, 1608, 1543, 76, 1897, 1205]. Robinson [2039]. rod [403, 404, 417]. Rodrigues [1772]. role [1917]. root [1623]. Rosen [1035, 1776, 320]. rotated [1528]. Rotating [1299, 1365, 86, 1661, 1543, 1560, 1561, 1269, 1788, 1250, 2141]. Rotation [870, 1198, 277, 1833, 796, 545]. rotational [177, 1117]. rotations [798, 373, 1221, 2084, 659]. rotator [2071]. rough [541, 371, 238]. rule [12, 151, 931, 435]. rules [1575, 39, 1058, 2056, 126, 495, 730, 570, 334, 1914]. runaway [1880]. Rytov [312].

S [2128, 870, 2018, 1381]. Sachs [696, 217]. Salpeter [614, 371, 1284]. Satake [386]. satisfying [1046]. Sato [2020]. Satsuma [340]. saturated [237]. saturation [280]. Sawada [1999]. Sawi [1465]. Scalar [323, 227, 794, 819, 1310, 1796, 942, 476, 1891, 1110, 838, 219, 715, 1869, 1509, 1561, 1804, 1421, 1743, 2094, 774, 1588, 446, 1558, 96, 220, 1119, 1449]. scalar-tensor [476, 96, 220, 1119]. scalar-tetradic [446]. scalars [473, 154]. scale [906]. scales [254, 255]. Scaling [100, 269]. scatterers [348]. Scattering [2069, 237, 268, 2109, 84, 690, 1458, 146, 1501, 980, 1744, 27, 410, 653, 1930, 2139, 83, 129, 1319, 101, 1216, 541, 383, 1073, 1500, 838, 168, 371, 349, 961, 35, 1106, 1570, 1725, 726, 727, 609, 962, 33, 1650, 2050, 1680, 59, 867, 859, 1179, 1714, 436, 1521, 2070, 1932, 650, 1256, 1377, 625, 25, 214, 1643, 272, 1114, 848, 729, 1132, 1909, 1805, 1108, 726, 962]. Scattering-into-cones [1501]. scenario [1661, 1543]. scheme [998, 1572, 1330, 931, 1524, 1709]. schemes [1226, 218, 637]. Schifini [2133]. Schild [795, 1468, 524, 1595, 344, 1079]. Schlieder [709]. Schlieder-type [709]. Schoenmaker [1016]. Schrödinger [85, 1938, 932, 394, 1753, 1729, 1589, 109, 1329, 865, 56, 1457, 1730, 1325, 1177, 165, 582, 1113, 1038, 1039, 1075, 1061, 1772, 432, 520, 516, 2071, 899, 1842, 1648, 1111, 1529, 709, 1256, 1929, 1076, 514, 1464, 2117, 1962, 1773, 1835]. Schrödinger-Langevin [899]. Schrödinger-like [1113]. Schücking [1207]. Schwartz [464]. Schwarz [257]. Schwarzschild [1965, 838, 1869]. Schwinger [721, 1040, 1506, 275, 1272, 821, 1191, 1902, 1777, 533, 1853]. Screen [608]. screened [1176, 1055, 519]. screws [1852]. search [1167, 1212, 1279, 1430]. Second [304, 1870, 1444, 1318, 1763, 437, 1488, 873, 1340, 1941, 567, 171, 470, 1996, 79, 908, 92, 1696, 2094, 733, 1030]. Second-degree [1444]. second-derivative [908]. Second-order [304, 1870, 1318, 1763, 1488, 873, 1340, 1941, 567, 171, 470, 1996, 92, 1696, 2094, 733, 1030]. second-order-infinitesimal [437]. section [720]. sections [410, 413]. sector [1600]. sectors [1153]. secular [497]. seed [576]. Selberg [1686]. selection [1648]. Self [1017, 1602, 1341, 611, 50, 1866, 1300, 1412, 957, 1087, 5, 47, 1140, 338, 826, 626, 335, 1026, 583, 937, 1277, 781, 1533, 1658, 1947, 1475, 1713, 1376, 1102, 1375, 457, 80, 416, 641, 1812, 1464, 1876]. self-creation [641]. self-dominant [626]. self-dual [50, 1866, 1412, 957, 1087, 1026, 583, 781, 1713, 1376, 1102, 1375, 457, 1812, 1876]. self-duality [611, 1277, 1658, 1947]. Self-energy [1017, 1602, 1341, 335]. self-gravitating [1300, 937]. self-interaction [1464]. self-similar [5, 47, 1140, 338, 826, 1533, 416]. selfadjoint [498]. selfadjointness [109]. semi [169, 1570, 2112, 1911, 940, 1337]. semi-classical [169]. semi-infinite [1570, 1911, 940, 1337]. semi-plane-symmetric [2112]. semiboundedness [109]. semicircle [370]. Semiclassical [91, 1077, 1410, 213, 392, 1400, 676]. semigroups [85, 1462]. Semilinear [1673]. semisimple [621, 762, 763, 1720]. seniority [931]. sense [393]. Sensitive [1992]. Separability [1139, 177, 1470, 2074]. Separable [1838, 232, 1520, 778]. separated [141]. separates [1117]. Separation [731, 275, 685, 754, 1061, 2041, 284, 2011]. sequence [53]. Sequences [2061]. sequential [93, 280, 451]. Series [530, 1120, 1121, 1062, 985, 1995, 955, 165, 1624, 815, 1856, 1093, 212, 2087, 1155, 1530, 1773]. set [187, 333, 919, 1672]. sets [1359, 104, 793, 1431, 1399, 695, 694, 2009, 2092]. setting [454, 1109, 1519, 442, 1261]. sextic [1463, 2100, 2136]. shallow [1134]. shape [1056]. shaped [910]. shapes [1228, 1849]. sheaf [1582, 1983]. Shear [1183, 1775, 316, 1298, 1405, 1081, 1469, 1502]. Shear-free [1183, 1775, 316, 1081, 1469, 1502]. sheet [1964]. shell [413, 2105, 325, 1108]. shell-model [325]. shells [458]. shift [1744, 1137, 1774, 1791, 31, 1479, 32]. shifts [1806]. shock [1332, 2030, 1824]. Short [874, 597, 2066, 522, 35, 1627, 1834]. short-range [522, 35]. Short-time [874, 2066]. sidebands [186, 652]. sided [2006]. Siegert [1649]. sigma [1599, 975, 2119, 283, 820]. Signals [2084]. similar [5, 47, 1140, 338, 826, 870, 1533, 416]. similarity [1101, 2021, 658, 423, 856]. Simon [1407]. Simons [1734]. Simple [491, 1545, 756, 572, 486, 215, 2010, 528, 841, 376, 2083, 1050, 590, 1472, 1717, 1447, 1158, 334, 1434, 1988, 723, 1516, 88, 115, 554]. simplex [1778]. Simplicial [138, 528, 1778]. simplified [617, 1441, 791, 2117, 1710]. simply [658, 786]. Simulation [667]. simultaneity [1923]. Simultaneous [727]. sine [54, 892, 341, 206, 750, 1237, 1279, 347]. sine-Gordon [54, 892, 341, 206, 750, 347]. Sine-Gordon-type [1279]. sine-Hilbert [1237]. Singh [2128, 2128]. single [370]. Singular [2120, 955, 1721, 865, 391, 315, 1031, 1823, 940, 1872, 2140, 825]. Singularities [346, 906, 1293, 1500, 599, 1627, 626, 1294, 284, 1733, 446, 1585, 1469]. singularity [194, 718, 934]. Sites [830, 831, 179]. Sitter [291, 2142, 1538, 19, 323, 1731, 943, 1694, 1144, 751, 878, 1064, 914, 474, 18, 574, 687]. six [1683, 2127]. sixfold [1165]. size [1809, 349, 280]. Skeleton [450]. Skyrme [1830]. slab [971]. Slater [1576, 909, 1337, 1336]. Slater-type [1576]. slicings [399, 658]. slightly [690]. Sloane [1624]. slow [207]. slow-fluctuation [207]. slowing [1738]. Small [1750, 1699, 1374, 181, 752, 348]. small-spaced [348]. smooth [1355]. smoothing [541]. Smoothness [828]. soft [1319]. sojourn [1926]. Sokolov [699]. solid [968, 414, 175, 115]. solids [375, 882]. Solitary [1354, 1210, 61, 375]. Soliton [699, 1044, 1045, 583, 576, 1606, 457, 1394, 1934, 1167, 1212, 1279, 1430, 87, 206, 1924, 1393]. Solitonic [1364, 937]. solitons [1165, 1015, 1049, 2063, 1312, 1615, 1944]. solubility [1085, 579]. soluble [915, 1379, 1623]. Solution [422, 1757, 54, 1816, 2124, 432, 388, 2102, 646, 1810, 807, 581, 1903, 1802, 761, 205, 606, 1933, 243, 1661, 1543, 1038, 1349, 788, 2091, 1264, 683, 1339, 990, 936, 1725, 1696, 305, 596, 293, 238, 576, 1277, 940, 893, 1945, 501, 856, 651, 1539, 2052, 344, 279, 274, 1812, 1141, 1876]. Solutions [1567, 1801, 361, 178, 1644, 1578, 1219, 663, 1685, 1579, 1763, 488, 239, 932, 1332, 475, 222, 1101, 1333, 1891, 1355, 1006, 2072, 82, 930, 865, 1300, 230, 360, 1413, 1598, 681, 1145, 1565, 1824, 2031, 968, 251, 1364, 1577, 1463, 1934, 1037, 1762, 1532, 1365, 843, 1210, 132, 678, 1021, 815, 1366, 528, 773, 2083, 1117, 423, 341, 276, 1407, 1561, 2036, 1405, 203, 1772, 1026, 871, 2001, 87, 61, 583, 576, 1807, 1118, 1940, 675, 1360, 1374, 364, 1741, 1468, 795, 524, 1238, 812, 835, 535, 1372, 1588, 457, 175, 679, 577, 1640]. solutions [2024, 1635, 448, 206, 1771, 1221, 1368, 1130, 2043, 96, 1119, 1256, 1780, 844, 1740, 1393, 1929, 88, 872, 2008, 1105, 1282, 1955, 1297, 1394, 340, 1225, 810, 1383, 1972, 729, 401, 1705, 147, 554, 1861, 359]. solvability [1946]. solvable [1477, 1100, 415, 1352, 1747, 733, 214]. Solving [1920, 1863, 500, 1949, 78]. Some [1332, 980, 475, 862, 500, 592, 593, 504, 757, 337, 2126, 1885, 2131, 306, 1709, 1669, 1102, 577, 1435, 2024, 1944, 1672, 96, 88, 74, 889, 1818, 985, 360, 2085, 94, 112, 245, 620, 194, 852, 1561, 1940, 1203, 8, 152, 1635, 571, 465, 1204, 107, 1105, 1955, 637, 106, 930, 528]. source [1408, 1562, 147]. sources [523, 780, 1975, 1372, 1369, 1221, 88, 1051]. Space [1415, 135, 1143, 824, 942, 1718, 2019, 523, 1965, 666, 1828, 1107, 1525, 1063, 1416, 2132, 259, 1387, 1681, 1863, 1229, 2010, 616, 1492, 399, 636, 187, 1307, 45, 898, 997, 1140, 2075, 1072, 1608, 490, 667, 1423, 323, 1750, 169, 599, 125, 90, 415, 921, 1694, 2, 240, 1460, 1698, 1869, 362, 1386, 1682, 376, 1211, 1163, 1357, 1438, 177, 1117, 751, 1082, 1264, 504, 878, 1064, 658, 870, 731, 1652, 569, 1399, 776, 57, 695, 694, 1570, 1886, 1142, 933, 1581, 509, 1026, 1897, 583, 1827, 444, 1205, 474, 639, 837, 1004, 1005, 133, 1258]. space [181, 752, 1348, 1453, 1238, 580, 910, 2116, 280, 505, 601, 2111, 2035, 1322, 657, 506, 166, 774, 685, 1435, 249, 80, 81, 1042, 1683, 2127, 1408, 1260, 796, 545, 680, 1043, 1551, 117, 342, 1368, 1194, 1512, 1800, 558, 1672, 1195, 1484, 1674, 1767, 1315, 2102, 320, 11, 143, 712, 1902, 1777, 2025, 1461, 139, 559, 1749, 1859, 738, 1937, 671, 1907, 803, 10, 461]. Space-group [1415]. Space-times [135]. spaced [348]. Spacelike [445, 1681, 1386, 600]. spaces [1201, 894, 1030, 50, 1811, 1307, 1930, 2139, 185, 770, 1913, 19, 964, 788, 830, 831, 1748, 1804, 781, 135, 7, 48, 18, 278, 1534, 354, 740, 1253, 861]. Sparling [773]. spatial [947, 2004, 2005, 1869, 1648]. Spatially [1819, 205, 444, 1348]. Special [1353, 865, 2075, 773, 1940, 1235, 1701, 571, 486, 143, 712, 1777, 2008]. specific [947]. Spectra [820, 1756, 1325, 473, 168, 1228, 1849]. Spectral [396, 923, 1912, 71, 1615, 1541, 435, 1195, 1137, 1774, 410, 854, 2101, 1243, 757, 130]. spectroscopic [570, 931]. Spectrum [482, 1818, 1728, 1646, 101, 1653, 434, 1650, 1518, 8, 152, 498, 1449]. Spencer [1738]. sphere [1647, 1144, 2142, 731, 1297]. spheres [1836, 770, 779, 628, 317, 363, 134, 1406, 2052, 525, 1080, 2007, 22, 1433, 136]. spherical [728, 131, 703, 1333, 237, 357, 176, 607, 1037, 1762, 1759, 1677, 780, 1144, 2142, 1884, 238, 1404, 250, 790, 1108]. Spherically [678, 1021, 2036, 871, 220, 1063, 1416, 2132, 1549, 2105, 132, 685, 1683, 2127, 1081, 1469, 1502]. spheroidal [663]. Spin [1440, 1142, 1503, 2008, 489, 1008, 179, 1797, 1687, 954, 721, 176, 607, 119, 1655, 847, 1694, 450, 788, 1347, 412, 2115, 1560, 1136, 675, 1360, 229, 1515, 698, 518, 1077, 1138]. Spin- [2008, 412, 675, 1360, 229]. spin-Bose [698, 518]. spin-boson [1655]. Spin-connection [1503].

Spin-polarized [1440]. spin-weighted [176, 607]. spinning [1867, 1760, 1781]. Spinor [770, 1168, 878, 1563, 192, 447, 862, 462, 753, 977, 328, 14, 262, 1561, 1115, 1181, 844]. Spinorial [484, 1748]. Spinors [105, 616, 1412, 1692, 108, 2010, 507, 628, 1097, 798, 1206, 657, 711, 2074]. spins [1170]. Splitting [345, 1356, 1275, 2065, 266]. Spontaneous [771, 1275, 978, 1780]. spreading [1926]. spurious [1892]. square [384, 279, 1585]. squared [311]. squared-difference [311]. squared-sum [311]. squares [557]. Stability [1269, 564, 63, 531, 1182, 924, 1311, 326, 1067, 1765, 2047, 750, 303, 1273, 1131, 498, 1898]. stabilizers [503]. Stable [1122, 2013, 858]. Stäckel [253, 339]. standard [1855, 495, 1539, 1795]. standing [1256]. Starobinsky [2115, 1622]. stars [1404]. starts [209]. state [2121, 1728, 1769, 1591, 1863, 1413, 761, 537, 243, 1473, 1690, 403, 417, 1737, 1542, 414, 2106, 127, 1669, 433, 1382, 212, 1932, 1469, 1393, 1141, 405, 578, 1046]. States [1596, 1593, 897, 1498, 804, 1251, 1878, 1942, 725, 1559, 1752, 1908, 234, 1959, 1678, 1326, 179, 2125, 39, 967, 397, 252, 1122, 1794, 1700, 1913, 868, 2068, 1146, 1147, 2013, 90, 415, 847, 349, 1803, 1649, 1519, 1065, 492, 458, 1324, 2047, 356, 704, 550, 618, 1414, 1679, 1055, 478, 532, 1022, 1586, 676, 1077, 1530, 1773, 1783]. Static [357, 219, 87, 2112, 1404, 1683, 2127, 1333, 1063, 1416, 2132, 1933, 30, 442, 745, 837, 96, 1119]. Stationary [2072, 969, 534, 845, 358, 209, 1142, 1594, 1219, 87, 1435, 359, 401, 1738]. Statistical [371, 442, 747, 631, 1908, 1970, 1146, 1147, 1085, 92, 224, 1474, 579, 91, 1024, 682, 366, 107]. statistics [530, 1302, 1120, 1121, 1020, 38, 1871, 426, 1911, 910, 139]. steady [2128, 252, 693]. step [70, 1842]. step-index [1842]. Stewartson [1344, 1712, 1702]. Stochastic [1015, 1908, 630, 119, 1507, 1370, 1185, 959, 350, 1034, 1174, 1169, 920, 1335, 84, 1992, 431, 706, 92, 1765, 1276, 1637, 682, 2060, 58, 829, 1994, 806]. Stochasticity [1631]. Stokes [1762]. strategy [1930, 2139]. stratified [1691, 1350]. streamfunctions [1343]. strength [1032]. stress [211]. stresses [511]. String [2015, 1905, 1601, 1785, 1480, 1964, 1291, 29, 2045]. strings [1425, 2083, 745, 1227, 292]. stringy [1597]. strip [1291]. stripping [729]. Strong [167, 1294, 186, 652, 413, 1104, 1787, 145, 1619]. strong-potential [413]. strongly [2013, 626, 186, 652, 368, 285, 198]. Structural [56, 759, 619, 2040]. Structure [489, 1368, 273, 1904, 894, 1030, 1582, 1983, 802, 160, 188, 447, 1939, 2004, 2005, 1133, 954, 1899, 406, 1401, 699, 1471, 2053, 1163, 470, 471, 736, 1996, 1351, 1594, 1259, 1765, 1115, 231, 1563, 601, 1278, 2020, 1555, 1422, 1787, 302, 1800, 1484, 1761, 716, 1664, 1448]. structureless [625]. structures [862, 1797, 2005, 770, 1424, 993, 1624, 1506, 1065, 428, 1001, 1134, 1953, 1874, 1557, 887, 1042, 664, 1155, 864, 1076, 1579, 1906]. Studies [1386]. study [1302, 1842, 1043, 1156, 671]. Stueckelberg [1064]. Sturm [314, 1129]. Sturmian [1340, 610]. Sturmians [966]. subalgebra [1584]. subalgebras [1127, 797, 1792, 885, 1232, 951]. Subduced [1719]. subgroup [456, 1447, 1271, 1605]. subgroups [1753, 1866, 491, 538]. subspaces [1646].

Substitutional [1487]. substitutions [944]. sudden [235]. Sufficient [1111, 1085, 579, 395]. suitable [1460, 1461]. sum [1636, 311, 1337, 1336, 570, 931, 334, 435, 1856]. Summable [1977, 1490]. Summation [571, 743, 1062, 126, 165]. Summation-free [571]. summations [1833]. Sums [110, 459, 330, 972, 624]. super [1575, 1828, 1887, 715, 1639, 1265, 2086, 1421, 1670, 481, 1686]. super-Lorentz [2086]. super-Poincaré [715, 1265, 1421]. Super-Selberg [1686]. super-Virasoro [1575]. superalgebra [1790, 1814, 2137, 267, 766, 950, 1235, 1701, 1916, 2087, 1695]. Superalgebras [1448, 1569, 1948, 883, 1420, 1381, 2123, 1158, 301, 1090, 155, 156, 885, 1232, 951]. supercharges [182]. superconductivity [1138]. superconformal [1499]. superevaluation [827]. Superfield [65, 716, 715, 1421, 1743]. superfields [1509, 1162, 1448]. superfluid [175]. supergravities [290]. supergravity [62, 1734, 1023]. supergroup [1559, 1667, 288, 975]. supergroups [1752, 496, 1514, 1091, 688, 1670, 538]. superheavy [1220]. Superinvariant [1150]. supermanifolds [1247, 1582, 1983, 1029, 1552, 116, 860, 1860, 424]. supermatrices [1962]. Supermultiplet [1678]. superposed [1311]. Superposition [989, 797, 988, 661, 50, 491, 1633]. superpotentials [981]. superprojection [768]. superrotation [2086]. superselection [39]. Superspace [981, 1634, 1029, 1162, 1478, 1751, 72, 598]. superspaces [800]. superstability [1735]. superstring [1813, 1275, 1784]. Supersymmetric [1786, 1689, 980, 2121, 1878, 689, 1137, 1177, 1850, 2122, 1263, 1358, 815, 1050, 1478, 913, 1604, 820, 1090, 946]. supersymmetries [1206]. Supersymmetry [1346, 411, 287, 617, 537, 1507, 1059]. supertableaux [295, 884, 496, 764, 1200]. supertori [1550]. Supertwistor [290, 289]. Superunitary [764, 295]. supplementary [1601]. support [367]. supports [1433]. Surface [535, 1071, 703, 947, 541, 543, 690, 1244, 1458, 1048]. surfaces [372, 1825, 1423, 1609, 846, 1670]. susceptibility [1094, 718]. swarms [1880]. sweeping [1841]. switching [1703]. symbols [245, 620, 852]. Symmetric [1144, 2142, 22, 1201, 68, 1063, 1416, 2132, 1549, 1933, 2105, 779, 620, 132, 678, 1021, 780, 1142, 2036, 1772, 871, 1044, 1045, 2112, 745, 1637, 685, 1683, 2127, 881, 302, 1674, 1022, 1487, 220, 1081, 1469, 1502, 1795, 170, 1271, 1605, 1844, 551, 1705, 692]. Symmetric-tensor [22]. symmetrical [1002, 1518]. Symmetries [1063, 1416, 2132, 1866, 195, 1488, 2079, 527, 174, 1639, 42, 1562, 502, 686, 475, 234, 1660, 1450, 1977, 1890, 487, 721, 1901, 338, 1428, 1444, 184, 754, 1482, 2086, 1684, 203, 1813, 135, 1191, 1375, 1068, 465, 1674, 1487, 1295, 1053, 787]. Symmetry [1498, 689, 627, 662, 191, 1050, 2035, 578, 106, 1345, 1574, 1537, 1266, 1267, 1342, 1165, 1646, 1959, 1333, 1245, 1866, 1939, 1687, 1344, 552, 2075, 357, 958, 2138, 945, 2077, 780, 183, 20, 177, 1117, 470, 471, 736, 1996, 1239, 539, 1784, 1001, 978, 1702, 247, 294, 1780, 1389, 453, 647, 1225, 1114, 75, 687, 1679]. symmetry-adapted [539]. Symplectic [1496, 1285, 1913, 418, 1883, 2057, 2135, 1067, 492, 822, 917, 550, 618, 1484]. synchronization [259, 1387]. system [1498, 120, 358, 1139, 1457, 1730, 556, 119, 1655, 948, 263, 1759, 1210, 1473, 2101, 1148, 1109, 1519, 567, 423, 167, 434, 1584, 596, 853, 1712, 1518, 1520, 186, 652, 2111, 1611, 1324, 247, 1102, 1375, 790, 1872, 2140, 1224, 1077, 25, 1768, 1909, 1805, 1223]. systematic [31, 1924, 32]. Systematics [976, 626]. systems [106, 1345, 1574, 983, 980, 1359, 672, 390, 900, 531, 179, 268, 661, 1922, 309, 1494, 2028, 204, 1687, 1825, 252, 1974, 469, 495, 1545, 23, 2027, 732, 168, 321, 90, 415, 1735, 1254, 1699, 1092, 1488, 993, 1823, 2059, 450, 499, 153, 1895, 776, 403, 404, 417, 470, 471, 736, 1996, 79, 768, 265, 343, 2064, 717, 1615, 927, 1352, 1747, 1782, 1783, 1281, 1276, 270, 1323, 602, 691, 1953, 1371, 777, 1832, 1861, 1557, 929, 698, 518, 1432, 1288, 704, 1766, 392, 1414, 1679, 122, 682, 2103, 1048, 1393, 1761, 1815, 1742, 692, 1268]. systems [1951, 736]. Szafron [1006].

table [456]. tables [1791]. Tadpole [2045]. tangent [1906]. tangential [326, 2051]. target [451]. Taub [1684, 2035]. Taylor [2051, 955]. Technical [1863]. technique [1342, 207, 38, 1991, 145, 1393]. techniques [422, 1980, 512, 606, 1388, 640, 77, 155, 1812]. Teichmüller [2111]. Teixeira [832]. telegraph [1131]. Teller [168]. temperature [1968, 1908, 1169, 1507, 643, 1994, 1504]. temperatures [1969]. temporal [94]. ten [824, 912, 1421, 1743, 1751]. ten-dimensional [824, 912, 1421, 1751]. Tensor [1613, 13, 1966, 1967, 1968, 476, 244, 1183, 1215, 178, 219, 173, 873, 1144, 2142, 713, 1407, 1489, 506, 2081, 1534, 790, 2037, 22, 1787, 96, 220, 1119, 359]. Tensors [1239, 211, 1180, 2041, 250, 249, 861]. tentative [89]. term [1040, 1292]. terms [1995, 108, 1380, 497, 1992, 798, 171, 1669, 1997, 2062, 1950, 2014, 292, 670]. Ternary [1662, 1565]. test [1457]. tetrad [161]. tetradic [446]. tetrahedron [922]. Teukolsky [1072, 2115, 663, 1622]. their [1173, 157, 797, 108, 1443, 1977, 1329, 1495, 2085, 418, 958, 2138, 759, 619, 869, 1843, 1144, 2142, 2083, 403, 404, 417, 373, 626, 203, 1550, 18, 1771, 2037, 1464, 55, 75]. them [172, 548]. Theorem [204, 40, 1905, 2019, 813, 2039, 531, 1531, 1208, 1184, 485, 1548, 1825, 1078, 1634, 1850, 555, 490, 2066, 1157, 1358, 429, 1163, 567, 1834, 539, 1065, 1952, 329, 1510, 755, 314, 639, 1658, 1947, 181, 752, 1860, 1129, 1637, 1066, 1770, 816, 1592, 127, 1726, 198, 446, 1614, 1662, 641, 1009, 208, 1501, 266]. Theorems [1272, 682, 1687, 241, 1000, 1037, 1762, 566, 1684, 337, 891, 1874, 934, 115, 1822]. theoretic [760]. theoretical [1199, 985, 938, 242, 226, 1154, 691, 246, 247, 692, 1268, 2107]. theories [273, 1600, 1505, 2039, 1645, 2125, 802, 974, 1029, 1010, 2082, 1226, 185, 2122, 211, 2013, 97, 629, 1865, 184, 271, 231, 1152, 1839, 816, 880, 2081, 368, 285, 1588, 196, 197, 1708, 935, 2014, 214, 1906, 181]. Theory [1699, 1566, 1973, 897, 1483, 980, 1303, 814, 1084, 1908, 476, 228, 906, 104, 793, 1711, 1678, 1245, 1174, 1175, 534, 845, 1499, 89, 268, 2118, 39, 124, 719, 410, 561, 2040, 1208, 480, 1287, 309, 1209, 595, 1930, 2139, 2113, 4, 546, 216, 838, 922, 219, 1601, 1124, 1151, 1880, 371, 84, 400, 522, 1507, 1018, 1092, 1440, 213, 1883, 454, 1785, 1616, 841, 2016, 35, 30, 412, 819, 1627, 600, 701, 1054, 2114, 846, 784, 1123, 677, 1415, 919, 2064, 563, 757, 452, 821, 631, 839, 1474, 818, 914, 1378, 414, 364, 2050, 181, 752]. theory [1336, 973, 2116, 95, 1563, 2035, 2071, 1244, 1371, 774, 834, 246, 307, 457, 123, 790, 1779, 750, 709, 1179, 433, 1422, 1382, 446, 2060, 1938, 1195, 1852, 96, 220, 1119, 2100, 2136, 320, 1057, 1261, 1262, 1377, 1301, 481, 538, 1631, 180, 453, 647, 995, 514, 1556, 214, 889, 1564, 2045, 1909, 1704, 31]. There [1331, 632]. thermal [414, 2106, 212]. thermally [1330]. thermo [1305, 1544]. Thermodynamic [2113, 581, 1903, 1122, 403, 404, 417, 218]. Thermodynamics [387, 645, 1304, 137, 1903, 911, 92, 1968]. thermonuclear [792]. Theta [1978, 1624, 1155]. thin [728]. third [882, 691, 1449, 692]. third-order [882, 1449]. Thirring [1411, 140, 281, 1482, 686]. Thomas [984, 1513, 666, 1092, 1440, 510]. Thompson [823, 509]. Three [813, 83, 784, 1570, 1813, 1363, 436, 1942, 1197, 1107, 480, 1413, 1849, 749, 1161, 430, 90, 415, 847, 112, 111, 592, 593, 875, 1167, 1212, 1279, 1430, 38, 701, 1109, 1054, 373, 265, 853, 1019, 135, 186, 652, 1798, 535, 8, 152, 117, 342, 1521, 1320, 1641, 1398, 810, 809, 2070]. three-body [847, 1109, 1054, 265]. Three-dimensional [83, 784, 1570, 436, 1942, 480, 1849, 749, 875, 38, 701, 1521, 810, 809, 2070]. three-frames [1798]. three-level [186, 652]. three-resonance [1019]. three-soliton [1167, 1212, 1279, 1430]. three-space [117, 342]. three-spaces [135]. three-state [1413]. three-wave [853, 1320, 1641, 1398]. Threshold [1500, 268]. tightly [296]. Tiling [225]. tilings [1817]. Time [606, 673, 1128, 1330, 1725, 1528, 899, 130, 1894, 1303, 959, 479, 942, 1965, 1525, 230, 2028, 1890, 636, 874, 517, 1140, 1553, 1000, 1423, 323, 345, 1356, 2066, 1327, 1928, 1335, 1254, 1018, 1103, 1698, 1869, 1682, 1211, 1357, 878, 1064, 870, 1926, 423, 1061, 1399, 1834, 1526, 1886, 726, 727, 962, 167, 264, 821, 1019, 1976, 1205, 1004, 1005, 133, 1258, 1348, 2116, 505, 657, 506, 774, 435, 80, 81, 1042, 796, 545, 1800, 1703, 558, 1586, 1315, 2102, 2025, 1907, 605, 1738, 1400]. Time-dependent [673, 1330, 1725, 1528, 899, 230, 2028, 1890, 517, 1553, 1327, 1928, 1335, 1254, 1018, 423, 1061, 1526, 264, 1019, 1703, 605, 1738]. time-evolution [1327]. time-independent [345, 1356]. time-of-sojourn [1926]. Time-ordered [1128]. Time-ordering [606, 1000]. timelike [1298, 1258]. times [1143, 523, 1063, 1416, 2132, 259, 1387, 1681, 616, 1492, 399, 2075, 1072, 1608, 490, 921, 362, 1386, 177, 1117, 1082, 658, 1142, 933, 1581, 1897, 444, 474, 639, 837, 135, 1453, 2035, 685, 1435, 1683, 2127, 1408, 680, 1043, 1551, 1194, 1512, 1674, 143, 712, 1902, 1777]. Titchmarsh [410]. Tits [1447]. Toda [1517, 308, 428, 2080]. Tolman [1533, 1141]. Tomimatsu [2020]. tool [1243, 1761]. Topological [189, 261, 2042, 4, 546, 1862, 38, 781, 1179]. topologically [12, 151, 863, 1366]. Topologies [1338, 1461]. Topology [2059, 1583, 860, 1194, 1512, 708]. Torsion [148, 771, 306]. torus [38]. Total [782, 12, 151, 1794, 1283, 100]. Totally [1453, 1271]. trace [128, 1761, 1686]. trajectories [667, 1651]. transcendent [596]. transcendents [1435, 700, 1060, 1671]. transfer [586, 1642, 2050, 1654]. transform [1591, 27, 1739, 623, 1723, 149, 919, 1615, 2126, 1991, 1885, 2131, 1028, 370, 1853]. Transformation [1376, 828, 488, 239, 1855, 331, 827, 958, 2138, 826, 1486, 1882, 1250, 2141, 855, 1988, 1252, 668, 671, 1384, 75]. transformations [106, 1345, 1574, 1905, 1101, 437, 746, 309, 1638, 1419, 2028, 808, 627, 1999, 464, 66, 1040, 2, 240, 963, 798, 1249, 1618, 140, 281, 649, 917, 1044, 1045, 2061, 329, 1475, 833, 1595, 1248, 1375, 1288, 1851, 206, 73, 1112, 283, 1466, 1998, 1961, 2130, 2017, 53, 857, 1202, 1697, 669, 670]. transforming [19]. Transforms [384, 1718, 1974, 1452]. Transient [137, 172, 548]. Transition [696, 964, 269, 740, 2003, 2034, 1355, 749, 1254, 1253, 1414, 1679, 889]. transitions [890, 1493, 589, 1542, 261]. transitive [658]. transitivity [1373]. Translating [1655]. translational [1389]. translations [1315]. Transmission [1002, 556]. transparency [556]. transport [719, 683, 1339, 990, 839, 1474, 940, 1932, 720]. Transverse [772, 2128, 926, 693]. trapped [1788]. travel [133]. Traveling [843, 1891]. Traveling-wave [843]. treatment [1016, 2096, 607, 691, 553, 1077, 692, 1268]. Tree [501]. trees [1785]. Triality [328, 14]. triangle [432]. triangular [645]. triangulations [528]. triclinic [247]. trilinear [1107]. triple [1856]. triples [1804]. triplets [291]. Trotter [1462]. truncated [166]. tubular [2088]. Tunneling [315, 1255, 2064]. turning [1763]. twenty [624]. twenty-four [624]. Twist [786, 554]. twist-eating [554]. twisted [2097, 1979]. twistor [724, 830, 831, 81, 44]. Twistors [2029, 2026]. Two [129, 544, 1075, 751, 585, 2006, 478, 1587, 1894, 201, 1728, 1769, 648, 1008, 915, 1379, 612, 2039, 1353, 1355, 358, 967, 1329, 586, 1228, 1930, 2139, 1145, 430, 1424, 101, 1216, 1500, 667, 673, 1311, 326, 308, 892, 1386, 798, 1522, 38, 1357, 153, 235, 79, 1270, 149, 1052, 1234, 425, 822, 917, 434, 1737, 663, 1019, 1118, 1911, 1741, 1839, 91, 1990, 903, 1617, 3, 697, 2024, 1851, 1771, 750, 1932, 510, 984, 1513, 1284, 1558, 336, 844, 1929, 1398, 605, 1909, 266, 1838, 776]. two- [38, 1019]. two-angle-dependent [235]. two-body [1728, 1769, 1008]. two-center [663, 903, 1771]. two-cluster-two-cluster [1500]. two-cocycles [1234, 822, 917]. two-component [1329, 798]. two-composite-particle [1909]. two-cycles [425]. Two-dimensional [129, 201, 2039, 967, 1228, 667, 308, 892, 1357, 79, 1911, 1741, 91, 3, 697, 750, 844, 776]. two-fermion [915, 1379]. Two-fluid [544]. two-layered [1355]. two-level [101, 434]. Two-loop [751, 585, 1052]. Two-mode [478]. two-neutron-transfer [586]. two-particle [1216, 1284, 1558, 266]. two-phase [149]. two-point [1587, 1894, 612]. Two-sided [2006]. two-state [1737, 1932]. two-time [605]. two-velocity [1145]. Type [1365, 1143, 1199, 1306, 460, 319, 1517, 1576, 1183, 1184, 1069, 1215, 1000, 1149, 714, 257, 429, 1167, 1212, 1279, 1941, 870, 194, 52, 275, 1725, 425, 452, 1259, 1118, 1940, 1518, 795, 524, 913, 2067, 250, 1434, 448, 588, 1408, 1819, 796, 545, 680, 709, 1221, 465, 1090, 641, 88, 1470, 1622, 256]. Type- [1365, 448, 1470]. type-VI [588]. types [589, 1468, 637]. Typical [267, 1790, 1235].

ultrarelativistic [1141]. ultrasmooth [1826]. ultraspherical [594, 110]. ultraviolet [1267, 284, 750]. Umehara [1339]. Ünal [1379]. Unbounded [1984, 453, 647, 918, 504, 189, 1204, 1672]. unboundedness [907]. Uncertainty [431, 704, 2002]. unconventional [942]. underlying [1139, 1206]. unequal [1270]. unfolding [406]. unidimensional [338]. unification [89, 1935, 306, 1138, 949]. unified [607, 553, 196]. uniform [2029, 1404, 146, 525, 858]. uniform-density [1404]. uniformly [1056]. unifying [402]. unimodality [503]. Unique [1056, 1730, 709]. uniquely [1133]. Uniqueness [1180, 2048, 63, 1511, 2044, 587, 1129, 2081, 2133, 1892, 1083]. unitarity [1680, 1023]. unitarization [1793]. Unitary [1040, 1514, 1667, 18, 1575, 487, 1893, 1545, 1104, 1236, 17, 1471, 2053, 1618, 1834, 492, 1234, 1851, 1195, 381, 16, 550]. Unitary-operator [550]. universal [761, 2088]. Universality [1722, 924, 1864]. universe [1629, 1746, 1685, 849, 150]. universes [233, 1653, 1561, 574, 1156]. unsaturated [1429]. untwisted [1979]. unusual [126, 1020]. upper [2003]. Use [265, 2085, 1971, 55]. using [1757, 1174, 2032, 662, 1577, 23, 873, 1064, 199, 1725, 1630, 1782, 1783, 1771, 248, 370, 702, 1656]. Usual [126]. utilization [149]. Utiyama [1208, 454, 226].

V [2073, 1228, 1930, 2139]. vacant [280]. vacua [1838, 1275]. Vacuum [39, 941, 1405, 883, 1420, 1410, 595, 221, 1406, 275, 1407, 871, 576, 220, 1262, 1622, 1777, 810, 830]. vacuum-vacuum [595]. Vaidya [1661, 1543]. validity [346]. value [1757, 1837, 2038, 73, 1237, 1831, 1009, 783]. valued [350, 1034, 1921, 1171, 1448]. valuedness [482]. values [1055, 486]. vanishing [178, 359, 486]. vanishingly [181, 752]. variable [2063, 164, 2041, 855, 856, 859, 1779, 436, 733, 1697]. variables [1525, 928, 1100, 754, 731, 1061, 275, 736, 1782, 1783, 2080, 1851, 1036, 198, 1368, 2011, 29, 1057, 1927]. Variational [1029, 615, 1845, 1540, 806, 466, 1451, 429, 1765, 356, 1879, 2117]. variations [1637]. varieties [876, 1095]. various [1264, 1911]. variously [910]. Vector [236, 433, 1382, 258, 1887, 2075, 1037, 406, 1609, 262, 504, 1064, 2064, 369, 1310, 85, 1150, 2007, 324, 98, 718, 2011, 2043, 1286, 1693, 1985]. Vectorial [1361]. Vectorlike [2047]. vectors [2132, 1681, 1791, 1149, 1386, 905, 445, 657, 1663, 1683, 2127, 578, 670]. velocities [1889]. velocity [666, 1858, 1145, 469, 404, 1996, 727, 962, 436, 449, 470]. velocity-dependent [469, 470]. velocity-independent [1996]. velocity-space [666]. Verma [590]. version [160, 415]. versions [806]. versus [1321, 298, 1085, 1834, 579].

Vertex [2123, 1843, 1266, 1267, 1380, 1715]. vertical [653]. very [1270]. VI [1849, 588]. via [1388, 776, 335, 1952, 564, 712]. vibration [1556, 760]. vicious [1453]. view [1696, 1502]. Vilkovisky [974, 973]. violated [1262]. violation [640]. Virasoro [1575, 1882, 1508]. Virtual [911]. Viscous [1836, 460, 360, 808, 440, 679, 1007, 150]. Vlasov [488, 239, 2101, 248]. Volkov [1141]. Volterra [614]. vortex [230, 1910]. vortices [1982, 1020, 2095]. Vries [1855, 2063, 843, 1351, 1689, 855, 856, 591, 206, 1112, 857, 1697, 1384, 502].

Wake [1677]. Wake-free [1677]. walk [262, 1737]. Walker [1661, 811, 1608, 1543, 76, 1897, 1205]. walks [829]. Ward [284, 1478]. warm [1312]. water [1134]. Wave [718, 734, 2065, 1135, 2055, 728, 1678, 1891, 953, 1355, 2015, 2120, 1329, 1811, 1801, 1722, 586, 597, 512, 976, 361, 2031, 1590, 721, 1166, 1395, 345, 1356, 2109, 1037, 1178, 753, 1759, 1532, 843, 1210, 288, 1168, 1578, 1816, 113, 1926, 705, 726, 727, 962, 33, 2064, 1612, 663, 2001, 853, 238, 61, 1807, 375, 414, 229, 210, 1278, 651, 1539, 436, 2062, 1736, 1586, 1256, 625, 1301, 1320, 1641, 1398, 741, 55, 312, 738, 1070, 1691, 1643, 1350, 1331, 1467, 1805, 147, 1108, 1186]. wavefunctions [1981]. waveguides [2068]. Wavelength [348]. Wavelength-dependent [348]. Wavelike [872]. Waves [1165, 1354, 1332, 1824, 1218, 1334, 1436, 1677, 1837, 2038, 540, 756, 1570, 1875, 1281, 1963, 690, 1134, 1458, 144, 278, 1733, 1138, 1706, 1437, 783, 2076, 1, 1442]. way [78]. Weak [218, 1041, 1343, 1717, 122, 637]. weak-Painlevé [122]. Weakly [55]. wedge [803]. Weight [1314, 851, 330, 1514, 1667, 768, 1915, 420]. Weight- [1314, 851]. weighted [176, 607]. weights [1572]. Weil [1785]. Weinberg [1510]. Weisskopf [1185]. Well [1738, 1888, 62, 35]. Well-posedness [1738, 62]. Wess [2121, 751]. Weyl [1591, 319, 1791, 410, 1183, 1184, 1214, 1215, 948, 1335, 493, 2110, 1180, 142, 526, 359]. Weyl-ordered [2110]. Weyl-type [319, 1184]. Wheeler [2000]. which [1010, 1971, 209, 1117, 264]. Wick [127]. widths [1911]. Wiechert [81]. Wien [1343, 1717]. Wiener [353, 928, 1024, 1274, 1611]. Wightman [612, 1506]. Wigner [1185, 82, 930, 607, 1545, 762, 763, 798, 1568, 776, 297, 1136, 395, 1617, 1988, 1252, 2017, 1695]. Wilson [699, 1989]. winding [1456]. within [493]. without [1525, 1601, 1735, 1651, 1220, 402, 2000, 840]. Witten [1137, 1510]. WKB [2064, 1651]. WKBJ [992, 1465]. World [1964, 1620]. Wouthuysen [1941]. Wouthuysen-type [1941]. wreath [1239]. Wronskian [1243]. Wronskians [1635]. Wunderlich [1304].

xxxxx [832].

Yang [1600, 842, 1308, 1477, 1226, 957, 1087, 1124, 817, 815, 831, 1347, 2091, 194, 423, 1895, 409, 584, 2044, 876, 1095, 1658, 1947, 1875, 1475, 973, 601, 1713, 1376, 562, 1244, 2046, 2133, 1503, 1874, 1372, 1102, 1375, 1954, 2134, 1812, 1086, 1282, 1631]. Yano [2041, 713]. Yeh [1442]. Yevick [749]. yield [2003]. yielded [2112]. Yukawa [335].

Z [958, 2138]. Zakharov [1045]. Zamolodchikov [875]. Zeeman [1897]. zero [1878, 1517, 595, 2012, 1431, 1675, 1211, 1407, 675, 1360, 1741, 1111, 1529, 1086]. zero- [1741]. zero-energy [1878]. zero-rest-mass [1211]. zeros [719, 851, 190, 82, 330, 930, 421, 902, 1328, 1314, 355, 1585, 1530]. zeta [1873, 2129, 1656]. zeta-function [1873, 2129]. Zetlin [267, 766, 950, 1916]. Zhong [2138, 958]. Zitterbewegung [1357]. Zumino [2121, 751].

References

Yeh:1963:DWP

[1] C. Yeh. The diffraction of waves by a penetrable ribbon. Journal of Mathematical Physics, 4(1):65–71, February 1963. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [1442].

Gerry:1983:RCT

[2] Christopher C. Gerry. Remarks on canonical transformations in phasespace path integrals. Journal of Mathematical Physics, 24(4):874–877, April 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [240].

Odundun:1983:OTD

[3] Olúwolé Òdúndún. Operators for the two-dimensional harmonic oscillator in an angular momentum basis. Journal of Mathematical Physics, 24(9):2340–2344, September 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [697].

Cronstrom:1983:TBC

[4] C. Cronström and J. Mickelsson. On topological boundary characteristics in nonabelian gauge theory. Journal of Mathematical Physics, 24(10): 2528–2531, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [546].

Chi:1983:HSS

[5] L. K. Chi. On the hydrodynamic self-similar cosmological models. Journal of Mathematical Physics, 24(10):2532–2533, October 1983. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [47].

Harrell:1984:HOM

[6] Evans M. Harrell, II. Hamiltonian operators with maximal eigenvalues. Journal of Mathematical Physics, 25(1):48–51, January 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [547].

Moffat:1984:HDR

[7] J. W. Moffat. Higher-dimensional Riemannian geometry and quaternion and octonion spaces. Journal of Mathematical Physics, 25(2):347–350, February 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [48].

Perez:1984:SGP

[8] J. Fernando Perez, C. P. Malta, and F. A. B. Coutinho. On some general properties of the point spectrum of three particles moving in one dimension. Journal of Mathematical Physics, 25(9):2589–2592, September 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [152].

Glasser:1984:NCB

[9] M. L. Glasser. A novel class of Bessel function integrals. Journal of Mathematical Physics, 25(10):2933–2934, October 1984. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [327].

vanHaeringen:1984:NIC

[10] H. van Haeringen and L. P. Kok. New inequalities for the Coulomb T matrix in momentum space. Journal of Mathematical Physics, 25(10): 3033–3038, October 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [461].

Thompson:1984:PCM

[11] Gerard Thompson. Polynomial constants of motion in flat space. Journal of Mathematical Physics, 25(12):3474–3478, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [509, 823].

Battaglia:1984:RTN

[12] Franco Battaglia and Thomas F. George. A rule for the total number of topologically distinct Feynman diagrams. Journal of Mathematical Physics, 25(12):3489–3491, December 1984. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See erratum [151].

Pulmannova:1985:TPQ

[13] Sylvia Pulmannová. Tensor product of quantum logics. Journal of Mathematical Physics, 26(1):1–5, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i1/p1_s1.

Hasiewicz:1985:TPG

[14] Z. Hasiewicz and A. K. Kwaśniewski. Triality principle and G_{2} group in spinor language. Journal of Mathematical Physics, 26(1):6–11, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v26/i1/p6_s1. See erratum [328].

Basu:1985:GRE

[15] Debabrata Basu and T. Bhattacharya. The Gel’fand realization and the exceptional representations of SL(2, R). Journal of Mathematical Physics, 26(1):12–17, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i1/p12_s1.

Weaver:1985:NMF

[16] L. Weaver and L. C. Biedenharn. A note on the multiplicity-free unitary irreps of \overline{SL(3, R)}. Journal of Mathematical Physics, 26(1):18–20, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p18_s1.

Guler:1985:MFU

[17] Y. Güler. Multiplicity-free, unitary and nonunitary irreducible representations of \overline{SL(3, R)}. Journal of Mathematical Physics, 26(1):21–28, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p21_s1.

Moylan:1985:URS

[18] P. Moylan. Unitary representations of the (4 + 1) de Sitter group on unitary irreducible representation spaces of the Poincaré group: equivalence with their realizations as induced representations. Journal of Mathematical Physics, 26(1):29–40, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i1/p29_s1.

Drechsler:1985:LNT

[19] W. Drechsler. Linearly and nonlinearly transforming fields on homogeneous spaces of the (4, 1)-de Sitter group. Journal of Mathematical Physics, 26(1):41–54, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i1/p41_s1.

Iwai:1985:CQS

[20] Toshihiro Iwai and See-Gew Rew. Classical and quantum symmetry groups of a free-fall particle. Journal of Mathematical Physics, 26(1): 55–61, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p55_s1.

Hegerfeldt:1985:FFF

[21] Gerhard C. Hegerfeldt. Factors of the Fock functional. Journal of Mathematical Physics, 26(1):62–64, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i1/p62_s1.

Rubin:1985:STE

[22] Mark A. Rubin and Carlos R. Ordóñez. Symmetric-tensor eigenspectrum of the Laplacian on n-spheres. Journal of Mathematical Physics, 26(1): 65–67, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p65_s1.

Feix:1985:IDN

[23] Marc R. Feix and H. Ralph Lewis. Invariants for dissipative nonlinear systems by using rescaling. Journal of Mathematical Physics, 26(1):68– 73, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p68_s1.

Levine:1985:ARE

[24] Harold Levine. On the analysis of relaxation in electrolytes. Journal of Mathematical Physics, 26(1):74–84, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i1/p74_s1.

Ueno:1985:CSR

[25] S. Ueno. Cauchy system for the resolvent of Milne’s integral equation with anisotropic scattering. Journal of Mathematical Physics, 26(1):85– 88, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p85_s1.

Sorg:1985:BRF

[26] M. Sorg. Bound and radiation fields of a pointlike SO(3) monopole. Journal of Mathematical Physics, 26(1):89–98, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p89_s1.

Beylkin:1985:IDI

[27] G. Beylkin. Imaging of discontinuities in the inverse scattering problem by inversion of a causal generalized Radon transform. Journal of Mathematical Physics, 26(1):99–108, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i1/p99_s1.

Grassini:1985:JEC

[28] Renato Grassini. Jet extension of a classical particle field: principles of covariance and minimal coupling. Journal of Mathematical Physics, 26(1):109–111, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i1/p109_s1.

Soderberg:1985:AAV

[29] B. Söderberg, B. Andersson, and G. Gustafson. The action-angle variables for the massless relativistic string in 1 + 1 dimensions. Journal of Mathematical Physics, 26(1):112–123, January 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i1/p112_s1.

Hostler:1985:MEE

[30] Levere C. Hostler. Mass eigenfunction expansions for the relativistic Kepler problem and arbitrary static magnetic field in relativistic quantum theory. Journal of Mathematical Physics, 26(1):124–127, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p124_s1.

DeMeyer:1985:GAS

[31] H. De Meyer and G. Vanden Berghe. A general approach to the systematic derivation of SO(3) shift operator relations. I. Theory. Journal of Mathematical Physics, 26(1):128–132, January 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i1/p128_s1.

VandenBerghe:1985:GAS

[32] G. Vanden Berghe and H. De Meyer. A general approach to the systematic derivation of SO(3) shift operator relations. II. Applications. Journal of Mathematical Physics, 26(1):133–138, January 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i1/p133_s1.

Kroger:1985:WOF

[33] Helmut Kröger. On the wave operator in few-body quantum scattering with Coulomb-like interactions. Journal of Mathematical Physics, 26(1):139–142, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i1/p139_s1.

Pierotti:1985:FKR

[34] Dario Pierotti. A Fock–Krein realization of the Landau gauge. Journal of Mathematical Physics, 26(1):143–144, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i1/p143_s1.

Holden:1985:SRE

[35] H. Holden, R. Høegh-Krohn, and M. Mebkhout. The short-range expansion for multiple well scattering theory. Journal of Mathematical Physics, 26(1):145–151, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i1/p145_s1.

TorresdelCastillo:1985:EAC

[36] G. F. Torres del Castillo. The HH equation in arbitrary canonical coordinates. Journal of Mathematical Physics, 26(1):152–154, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p152_s1.

GarciaDiaz:1985:MGK

[37] Alberto García Díaz. Magnetic generalization of the Kerr–Newman metric. Journal of Mathematical Physics, 26(1):155–156, January 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p155_
s1.

Hosoya:1985:EAO

[38] Haruo Hosoya and Akiko Motoyama. An effective algorithm for obtaining polynomials for dimer statistics. Application of operator technique on the topological index to two- and three-dimensional rectangular and torus lattices. Journal of Mathematical Physics, 26(1):157–167, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p157_s1.

Borek:1985:VSS

[39] R. Borek. Vacuum states and superselection rules in quantum field theory. Journal of Mathematical Physics, 26(1):168–171, January 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p168_
s1.

Batalin:1985:ETG

[40] I. A. Batalin and G. A. Vilkovisky. Existence theorem for gauge algebra. Journal of Mathematical Physics, 26(1):172–184, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p172_s1.

Comtet:1985:EAH

[41] A. Comtet and P. J. Houston. Effective action on the hyperbolic plane in a constant external field. Journal of Mathematical Physics, 26(1):185– 191, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p185_s1.

Molelekoa:1985:SGF

[42] Moshe Molelekoa. Symmetries of gauge fields. Journal of Mathematical Physics, 26(1):192–197, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i1/p192_s1.

Rudra:1985:PCI

[43] P. Rudra. Projective corepresentations and invariants of PCT group. Journal of Mathematical Physics, 26(1):198–201, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p198_s1.

Wainwright:1985:TAN

[44] P. R. Wainwright. A twistor approach to Nahm’s equations. Journal of Mathematical Physics, 26(1):202–207, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i1/p202_s1.

Chandola:1985:EED

[45] H. C. Chandola and B. S. Rajput. Electrodynamics in eight-dimensional space. Journal of Mathematical Physics, 26(1):208–210, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/
p208_s1.

Moravcsik:1985:RDA

[46] Michael J. Moravcsik. Resolving the discrete ambiguities in amplitude determinations. Journal of Mathematical Physics, 26(1):211–213, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i1/p211_s1.

Chi:1985:EHS

[47] L. K. Chi. Erratum: “On the hydrodynamic self-similar cosmological models”, [J. Math. Phys. 24 (1983), no. 10, 2532–2533; MR 84j:83047]. Journal of Mathematical Physics, 26(1):214, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i1/p214_s1. See

[5].

Moffat:1985:EHD

[48] J. W. Moffat. Erratum: “Higher-dimensional Riemannian geometry and quaternion and octonion spaces”, [J. Math. Phys. 25 (1984), no. 2, 347–350; MR 85d:83072]. Journal of Mathematical Physics, 26(1): 214, January 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i1/p214_s2. See [7].

Senechal:1985:MCG

[49] Marjorie Senechal. Morphisms of crystallographic groups: kernels and images. Journal of Mathematical Physics, 26(2):219–228, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i2/
p219_s1.

Boyer:1985:IHC

[50] Charles P. Boyer and Jerzy F. Plebański. An infinite hierarchy of conservation laws and nonlinear superposition principles for self-dual Einstein spaces. Journal of Mathematical Physics, 26(2):229–234, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i2/p229_s1.

Dobrev:1985:ERI

[51] V. K. Dobrev. Elementary representations and intertwining operators for SU(2, 2). I. Journal of Mathematical Physics, 26(2):235–251, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i2/p235_s1.

Johnson:1985:MTL

[52] Joseph E. Johnson. Markov-type Lie groups in GL(n, R). Journal of Mathematical Physics, 26(2):252–257, February 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i2/p252_s1.

Weiss:1985:PPB

[53] John Weiss. The Painlevé property and Bäcklund transformations for the sequence of Boussinesq equations. Journal of Mathematical Physics, 26(2):258–269, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i2/p258_s1.

Boiti:1985:SCP

[54] Marco Boiti, Jérôme J. P. Leon, and Flora Pempinelli. Solution of the Cauchy problem for a generalized sine-Gordon equation. Journal of Mathematical Physics, 26(2):270–275, February 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i2/p270_s1.

Weniger:1985:WCE

[55] E. Joachim Weniger. Weakly convergent expansions of a plane wave and their use in Fourier integrals. Journal of Mathematical Physics, 26(2):276–291, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i2/p276_s1.

Burdet:1985:SIS

[56] G. Burdet and M. Perrin. Structural invariance of the Schrödinger equation and chronoprojective geometry. Journal of Mathematical Physics, 26(2):292–299, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i2/p292_s1.

Kent:1985:MEB

[57] Steven L. Kent, Carlos N. Kozameh, and E. T. Newman. Maxwell’s equations and the bundle of null directions on Minkowski space. Journal of Mathematical Physics, 26(2):300–305, February 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i2/p300_s1.

Schroeck:1985:DPS

[58] F. E. Schroeck, Jr. The dequantization programme for stochastic quantum mechanics. Journal of Mathematical Physics, 26(2):306–310, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i2/p306_s1.

Newton:1985:ISL

[59] Roger G. Newton. Inverse scattering by a local impurity in a periodic potential in one dimension. II. Journal of Mathematical Physics, 26(2):311– 316, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i2/p311_s1.

Beezley:1985:EIP

[60] R. S. Beezley and R. J. Krueger. An electromagnetic inverse problem for dispersive media. Journal of Mathematical Physics, 26(2):317– 325, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i2/p317_s1.

Letelier:1985:SWS

[61] P. S. Letelier. Solitary wave solutions to the Einstein equations. Journal of Mathematical Physics, 26(2):326–328, February 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i2/p326_s1.

Bao:1985:WPC

[62] David Bao, Yvonne Choquet-Bruhat, James Isenberg, and Philip B. Yasskin. The well-posedness of (N = 1) classical supergravity. Journal of Mathematical Physics, 26(2):329–333, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i2/p329_s1.

Bellomo:1985:CPN

[63] N. Bellomo and G. Toscani. On the Cauchy problem for the nonlinear Boltzmann equation: global existence, uniqueness and asymptotic stability. Journal of Mathematical Physics, 26(2):334–338, February 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i2/p334_
s1.

Borek:1985:RCA

[64] R. Borek. Representations of the current algebra of a charged massless Dirac field. Journal of Mathematical Physics, 26(2):339–344, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i2/p339_s1.

Hassoun:1985:SAD

[65] J. Hassoun, A. Restuccia, and J. G. Taylor. Superfield actions for N = 2 degenerate central charges. Journal of Mathematical Physics, 26(2):345– 351, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i2/p345_s1.

Fahnline:1985:PFP

[66] Donald E. Fahnline. Planar factors of proper homogeneous Lorentz transformations. Journal of Mathematical Physics, 26(2):352–358, February 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i2/p352_s1.

Audit:1985:FIG

[67] Philippe Audit. Functions of infinite generalized cyclic matrices. Journal of Mathematical Physics, 26(3):361–364, March 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i3/p361_s1.

Bohm:1985:ICS

[68] A. Bohm and P. Moylan. On the integrability of certain symmetric representations of the Lie algebra of \overline{SO_{0}(4, 1)}. Journal of Mathematical Physics, 26(3):365–374, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i3/p365_s1.

Kocak:1985:JML

[69] Hüseyin Koçak. A Jacobson–Morozov lemma for sp(2n, R). Journal of Mathematical Physics, 26(3):375–376, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i3/p375_s1.

Mickelsson:1985:RKM

[70] Jouko Mickelsson. Representations of Kac–Moody algebras by step algebras. Journal of Mathematical Physics, 26(3):377–382, March 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p377_
s1.

Fernandez:1985:SCL

[71] Claudio Fernández. Spectral concentration for the Laplace operator in the exterior of a resonator. Journal of Mathematical Physics, 26(3):383– 384, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p383_s1.

Rogers:1985:CSI

[72] Alice Rogers. Consistent superspace integration. Journal of Mathematical Physics, 26(3):385–392, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i3/p385_s1.

Rogers:1985:ALB

[73] C. Rogers. Application of linked Bäcklund transformations to nonlinear boundary value problems. Journal of Mathematical Physics, 26(3):393– 395, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p393_s1.

Wen:1985:SPH

[74] Zhen Yi Wen and John Avery. Some properties of hyperspherical harmonics. Journal of Mathematical Physics, 26(3):396–403, March 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p396_
s1.

Zhong:1985:HCL

[75] Zai Zhe Zhong. On the hyperbolic complex linear symmetry groups and their local gauge transformation actions. Journal of Mathematical Physics, 26(3):404–406, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i3/p404_s1. See comment [958]..

Koch-Sen:1985:INI

[76] Lisa Koch-Sen. Infinitesimal null isotropy and Robertson–Walker metrics. Journal of Mathematical Physics, 26(3):407–410, March 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p407_
s1.

Schwartz:1985:HAA

[77] Charles Schwartz. High-accuracy approximation techniques for analytic functions. Journal of Mathematical Physics, 26(3):411–415, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/
p411_s1.

Vega-Fernandez:1985:NWS

[78] J. M. Vega-Fernández, J. F. Duque-Carrillo, and J. J. Peña-Bernal. A new way for solving Laplace’s problem (the predictor jump method). Journal of Mathematical Physics, 26(3):416–419, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p416_s1.

Kaushal:1985:CSC

[79] R. S. Kaushal, S. C. Mishra, and K. C. Tripathy. Construction of the second constant of motion for two-dimensional classical systems. Journal of Mathematical Physics, 26(3):420–427, March 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i3/p420_s1.

Porter:1985:ASD

[80] John R. Porter. The anti-self-dual Coulomb field in Minkowski space– time. Journal of Mathematical Physics, 26(3):428–430, March 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p428_
s1.

Porter:1985:TEL

[81] John R. Porter. A twistor encoding of Lienard–Wiechert fields in Minkowski space–time. Journal of Mathematical Physics, 26(3):431– 433, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p431_s1.

Brudno:1985:NZWa

[82] Simcha Brudno. Nontrivial zeros of the Wigner (3-j) and Racah (6-j) coefficients. I. Linear solutions. Journal of Mathematical Physics, 26(3): 434–435, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p434_s1.

Cheney:1985:TDIa

[83] Margaret Cheney and James H. Rose. Three-dimensional inverse scattering: high-frequency analysis of Newton’s Marchenko equation. Journal of Mathematical Physics, 26(3):436–439, March 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i3/p436_s1.

Gagnon:1985:STE

[84] R. Gagnon. Scattering theory for extended elementary particles in nonrelativistic stochastic quantum mechanics. Journal of Mathematical Physics, 26(3):440–452, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i3/p440_s1.

Osborn:1985:SSV

[85] T. A. Osborn, R. A. Corns, and Y. Fujiwara. Schrödinger semigroups for vector fields. Journal of Mathematical Physics, 26(3):453–464, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i3/p453_s1.

GarciaDiaz:1985:GFC

[86] Alberto García Díaz and Nora Bretón Baez. The gravitational field of a charged, magnetized, accelerating, and rotating mass. Journal of Mathematical Physics, 26(3):465–466, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i3/p465_s1.

Letelier:1985:SSM

[87] Patricio S. Letelier. Static and stationary multiple soliton solutions to the Einstein equations. Journal of Mathematical Physics, 26(3):467– 476, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p467_s1.

TorresdelCastillo:1985:SST

[88] G. F. Torres del Castillo and Jerzy F. Plebański. Some simple type D solutions to the Einstein equations with sources. Journal of Mathematical Physics, 26(3):477–481, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i3/p477_s1.

Boisseau:1985:GTR

[89] Bruno Boisseau. Gauge theory in relativistic Hamiltonian classical mechanics: tentative unification of electromagnetic and gravitational fields. Journal of Mathematical Physics, 26(3):482–489, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p482_s1.

Garbaczewski:1985:FSBa

[90] Piotr Garbaczewski. Fermi states of Bose systems in three space dimensions. Journal of Mathematical Physics, 26(3):490–494, March 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p490_
s1.

Mishra:1985:SSM

[91] B. M. Mishra and S. K. Sinha. Semiclassical statistical mechanics of two-dimensional fluid mixture of hard disks. Journal of Mathematical Physics, 26(3):495–504, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i3/p495_s1.

Lavenda:1985:NST

[92] B. H. Lavenda and R. Serra. Nonequilibrium statistical thermodynamics of second-order stochastic processes in the limit of large resistance. Journal of Mathematical Physics, 26(3):505–513, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p505_s1.

Mellein:1985:MRSa

[93] B. Mellein and E. E. Mola. A multitype random sequential process. Journal of Mathematical Physics, 26(3):514–521, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p514_s1.

Garrido:1985:ETE

[94] L. Garrido and J. Masoliver. Exact temporal evolution for some nonlinear diffusion processes. Journal of Mathematical Physics, 26(3):522– 527, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p522_s1.

Moffat:1985:CFN

[95] J. W. Moffat. Chiral fermions in non-Riemannian Kaluza–Klein theory. Journal of Mathematical Physics, 26(3):528–531, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/p528_s1.

Singh:1985:SSS

[96] T. Singh and Tarkeshwar Singh. Some static solutions in the general scalar-tensor theory. Journal of Mathematical Physics, 26(3):532– 535, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i3/p532_s1.

Finkelstein:1985:FCK

[97] R. Finkelstein and M. Villasante. Fermionic couplings in Kaluza–Klein theories. Journal of Mathematical Physics, 26(3):536–541, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i3/
p536_s1.

Reifler:1985:VME

[98] Frank Reifler. A vector model for electroweak interactions. Journal of Mathematical Physics, 26(3):542–550, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i3/p542_s1.

Chakrabarti:1985:EC

[99] J. Chakrabarti and M. Sajjad Zahir. Expansion in condensates. Journal of Mathematical Physics, 26(3):551–553, March 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i3/p551_s1.

March:1985:SPT

[100] N. H. March. Scaling properties of total energy of heavy positive ions in d-dimensions. Journal of Mathematical Physics, 26(3):554–555, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i3/p554_s1.

Davidson:1985:ECL

[101] Russell Davidson and John J. Kozak. Exact calculation of the lineshape and the scattering operator for a model of a two-level atom interacting with a continuous spectrum of radiation. Journal of Mathematical Physics, 26(3):556–565, March 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i3/p556_s1.

Rogers:1985:REP

[102] George W. Rogers. Riccati equations and perturbation expansions in quantum mechanics. Journal of Mathematical Physics, 26(4):567–575, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p567_s1.

Moreno:1985:CFP

[103] Matías Moreno. Closed formula for the product of n Dirac matrices. Journal of Mathematical Physics, 26(4):576–584, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p576_s1.

BenDaniel:1985:RTC

[104] D. J. BenDaniel. The relation of a theory of countable sets to the field equations of physics. Journal of Mathematical Physics, 26(4):585–587, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v26/i4/p585_s1. See erratum [793].

Budinich:1985:SFO

[105] Paolo Budinich and Krystyna Bugajska. Spinors as fundamental objects. Journal of Mathematical Physics, 26(4):588–592, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p588_s1.

Aguirre:1985:IST

[106] M. Aguirre and J. Krause. Infinitesimal symmetry transformations. II. Some one-dimensional nonlinear systems. Journal of Mathematical Physics, 26(4):593–600, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p593_s1.

Suzuki:1985:DFE

[107] Masuo Suzuki. Decomposition formulas of exponential operators and Lie exponentials with some applications to quantum mechanics and statistical physics. Journal of Mathematical Physics, 26(4):601–612, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p601_s1.

Braden:1985:DST

[108] H. W. Braden. n-dimensional spinors: their properties in terms of finite groups. Journal of Mathematical Physics, 26(4):613–620, April 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p613_
s1.

Brasche:1985:PSH

[109] Johannes Brasche. Perturbation of Schrödinger Hamiltonians by measures—selfadjointness and lower semiboundedness. Journal of Mathematical Physics, 26(4):621–626, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p621_s1.

Rahman:1985:SPU

[110] Mizan Rahman and Mihr J. Shah. Sums of products of ultraspherical functions. Journal of Mathematical Physics, 26(4):627–632, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p627_s1.

Gervois:1985:ITB

[111] A. Gervois and H. Navelet. Integrals of three Bessel functions and Legendre functions. I. Journal of Mathematical Physics, 26(4):633–644, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p633_s1.

Gervois:1985:IST

[112] A. Gervois and H. Navelet. Integrals of some three Bessel functions and Legendre functions. II. Journal of Mathematical Physics, 26(4): 645–655, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p645_s1.

Humblet:1985:BFE

[113] J. Humblet. Bessel function expansions of Coulomb wave functions. Journal of Mathematical Physics, 26(4):656–659, April 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i4/p656_s1.

Holland:1985:ECN

[114] Charles J. Holland and James G. Berryman. Exponential convergence for nonlinear diffusion problems with positive lateral boundary conditions. Journal of Mathematical Physics, 26(4):660–663, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p660_s1.

Weniger:1985:SDA

[115] E. Joachim Weniger and E. Otto Steinborn. A simple derivation of the addition theorems of the irregular solid harmonics, the Helmholtz harmonics, and the modified Helmholtz harmonics. Journal of Mathematical Physics, 26(4):664–670, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p664_s1.

Catenacci:1985:BS

[116] Roberto Catenacci, Cesare Reina, and Paolo Teofilatto. On the body of supermanifolds. Journal of Mathematical Physics, 26(4):671–674, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p671_s1.

Rno:1985:HAEa

[117] Jung Sik Rno. Harmonic analysis on the Euclidean group in three-space. Journal of Mathematical Physics, 26(4):675–677, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p675_s1.

Angeletti:1985:AED

[118] Angelo Angeletti, Cinzia Castagnari, and Francesco Zirilli. Asymptotic eigenvalue degeneracy for a class of one-dimensional Fokker–Planck operators. Journal of Mathematical Physics, 26(4):678–691, April 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p678_
s1.

Falk:1985:SDE

[119] H. Falk. Stochastic difference equations for a spin system. Journal of Mathematical Physics, 26(4):692–697, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i4/p692_s1.

Andric:1985:MEF

[120] I. Andrić and A. Mikelić. Minimization of the energy functional of a one-dimensional fermionic system in the large-N limit. Journal of Mathematical Physics, 26(4):698–704, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p698_s1.

Martin:1985:FAB

[121] Pablo Martín and Antonio L. Guerrero. Fractional approximations to the Bessel function J_{0}(x). Journal of Mathematical Physics, 26(4):705– 707, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p705_s1.

Ranada:1985:WPP

[122] A. F. Rañada, A. Ramani, B. Dorizzi, and B. Grammaticos. The weak-Painlevé property as a criterion for the integrability of dynamical systems. Journal of Mathematical Physics, 26(4):708–710, April 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p708_
s1.

Peterson:1985:GME

[123] Mark A. Peterson. Geometrical methods for the elasticity theory of membranes. Journal of Mathematical Physics, 26(4):711–717, April 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p711_
s1.

Boudet:1985:CLD

[124] Roger Boudet. Conservation laws in the Dirac theory. Journal of Mathematical Physics, 26(4):718–724, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p718_s1.

Gadella:1985:RHS

[125] M. Gadella. A rigged Hilbert space for the free radiation field. Journal of Mathematical Physics, 26(4):725–727, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i4/p725_s1.

Elbaz:1985:UUS

[126] E. Elbaz. Usual and unusual summation rules over j angular momentum. Journal of Mathematical Physics, 26(4):728–731, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p728_s1.

Nomura:1985:EWT

[127] Masao Nomura. Extension of Wick’s theorem by means of the state operator formalism. Journal of Mathematical Physics, 26(4):732–737, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p732_s1.

Nomura:1985:RRB

[128] Masao Nomura. A reduction relation for an n-body trace of an operator product consisting of coupled C^{+}’s and C’s. Journal of Mathematical Physics, 26(4):738–742, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p738_s1.

Cheney:1985:TDIb

[129] Margaret Cheney. Two-dimensional inverse scattering: compactness of the generalized Marchenko operator. Journal of Mathematical Physics, 26(4):743–752, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p743_s1.

Osborn:1985:TDS

[130] T. A. Osborn and R. Wong. Time decay and spectral kernel asymptotics. Journal of Mathematical Physics, 26(4):753–768, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p753_s1.

Beig:1985:RPS

[131] R. Beig. A remarkable property of spherical harmonics. Journal of Mathematical Physics, 26(4):769–770, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i4/p769_s1.

Hajj-Boutros:1985:SSP

[132] J. Hajj-Boutros. On spherically symmetric perfect fluid solutions. Journal of Mathematical Physics, 26(4):771–773, April 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i4/p771_s1.

Malament:1985:MAR

[133] David B. Malament. Minimal acceleration requirements for “time travel” in Gödel space–time. Journal of Mathematical Physics, 26(4):774–777, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p774_s1.

Herrera:1985:PFS

[134] L. Herrera and J. Ponce de León. Perfect fluid spheres admitting a one-parameter group of conformal motions. Journal of Mathematical Physics, 26(4):778–784, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p778_s1.

Martinez:1985:STI

[135] E. Martinez and J. L. Sanz. Space-times with intrinsic symmetries on the three-spaces t = constant. Journal of Mathematical Physics, 26(4): 785–791, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p785_s1.

Whitman:1985:ECD

[136] Patrick G. Whitman. Equilibrium configurations of degenerate fluid spheres. Journal of Mathematical Physics, 26(4):792–797, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/
p792_s1.

Barrabes:1985:TTE

[137] Claude Barrabès. Transient thermodynamics of electromagnetic media in general relativity. Journal of Mathematical Physics, 26(4):798–803, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p798_s1.

Hartle:1985:SMG

[138] James B. Hartle. Simplicial minisuperspace. I. General discussion. Journal of Mathematical Physics, 26(4):804–814, April 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i4/p804_s1.

Walikainen:1985:SID

[139] D. Walikainen and R. B. McQuistan. The statistics of interacting dumbbells on a 2 × N lattice space. Journal of Mathematical Physics, 26(4): 815–821, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p815_s1.

Kersten:1985:LBT

[140] Paul H. M. Kersten and Ruud Martini. Lie–Bäcklund transformations for the massive Thirring model. Journal of Mathematical Physics, 26(4): 822–825, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p822_s1.

Burzlaff:1985:RSC

[141] Jürgen Burzlaff. Radially separated classical lumps in nonabelian gauge models. Journal of Mathematical Physics, 26(4):826–830, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/
p826_s1.

Komar:1985:GWG

[142] Arthur Komar. Generalization of Weyl’s gauge group. Journal of Mathematical Physics, 26(4):831–835, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p831_s1.

TorresdelCastillo:1985:GFA

[143] G. F. Torres del Castillo. Gauge fields in algebraically special space– times. Journal of Mathematical Physics, 26(4):836–840, April 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/p836_
s1.

Oh:1985:NPW

[144] C. H. Oh and Rosy Teh. Nonabelian progressive waves. Journal of Mathematical Physics, 26(4):841–844, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i4/p841_s1.

Sil:1985:TES

[145] N. C. Sil and J. H. McGuire. A technique for evaluation of the strong potential Born approximation for electron capture. Journal of Mathematical Physics, 26(4):845–853, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p845_s1.

Ott:1985:SPC

[146] R. H. Ott. Scattering by a parabolic cylinder—a uniform asymptotic expansion. Journal of Mathematical Physics, 26(4):854–860, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i4/
p854_s1.

Ziolkowski:1985:ESW

[147] Richard W. Ziolkowski. Exact solutions of the wave equation with complex source locations. Journal of Mathematical Physics, 26(4):861–863, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p861_s1.

Obaid:1985:TCB

[148] Evelyn E. Obaid and S. A. Obaid. Torsion of certain beams. Journal of Mathematical Physics, 26(4):864–869, April 1985. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i4/p864_s1.

King:1985:ITP

[149] Michael J. King. Immiscible two-phase flow in a porous medium: utilization of a Laplace transform boost. Journal of Mathematical Physics, 26(4):870–877, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i4/p870_s1.

Santos:1985:IHU

[150] N. O. Santos, R. S. Dias, and A. Banerjee. Isotropic homogeneous universe with viscous fluid. Journal of Mathematical Physics, 26(4):878– 881, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i4/p878_s1.

Battaglia:1985:ERT

[151] Franco Battaglia and Thomas F. George. Erratum: “A rule for the total number of topologically distinct Feynman diagrams”, [J. Math. Phys. 25, 3489 (1984)]. Journal of Mathematical Physics, 26(4):882, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i4/p882_s2. See [12].

Perez:1985:ESG

[152] J. Fernando Perez, C. P. Malta, and F. A. B. Coutinho. Erratum: “On some general properties of the point spectrum of three particles moving in one dimension”, [J. Math. Phys. 25, 2589 (1984)]. Journal of Mathematical Physics, 26(4):882, April 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i4/p882_s1. See [8].

Iwai:1985:RTD

[153] Toshihiro Iwai. On reduction of two degrees of freedom Hamiltonian systems by an S^{1} action, and SO_{0}(1, 2) as a dynamical group. Journal of Mathematical Physics, 26(5):885–893, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p885_s1.

Hughes:1985:PCS

[154] J. W. B. Hughes and J. Van der Jeugt. A pair of commuting scalars for G(2) ⊃ SU(2) × SU(2). Journal of Mathematical Physics, 26(5):

894–900, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p894_s1.

Sharp:1985:AGF

[155] R. T. Sharp, J. Van der Jeugt, and J. W. B. Hughes. Application of generating function techniques to Lie superalgebras. Journal of Mathematical Physics, 26(5):901–912, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i5/p901_s1.

VanderJeugt:1985:IRE

[156] Joris Van der Jeugt. Irreducible representations of the exceptional Lie superalgebras D(2, 1; α). Journal of Mathematical Physics, 26(5):913–924, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i5/p913_s1.

Beckers:1985:GFT

[157] J. Beckers and V. Hussin. On U(1)-gauge fields and their geometrical interpretation. Journal of Mathematical Physics, 26(5):925–928, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i5/p925_s1.

Kroger:1985:LND

[158] Helmut Kröger. Linearization of nonlinear differential equations by means of Cauchy’s integral. Journal of Mathematical Physics, 26(5): 929–939, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p929_s1.

Antone:1985:CEL

[159] A. A. Antone. Compact expression for Löwdin’s alpha function. Journal of Mathematical Physics, 26(5):940–942, May 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i5/p940_s1. See com-

ment [1025].

Bruschi:1985:HSC

[160] M. Bruschi and O. Ragnisco. The Hamiltonian structure of a complex version of the Burgers hierarchy. Journal of Mathematical Physics, 26(5):

943–945, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p943_s1.

Urani:1985:FET

[161] John R. Urani and Frank J. Kutchko. Field equations and the tetrad connection. Journal of Mathematical Physics, 26(5):946–950, May 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i5/p946_
s1.

Schwartz:1985:NIMa

[162] Charles Schwartz. Numerical integration in many dimensions. I. Journal of Mathematical Physics, 26(5):951–954, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p951_s1.

Schwartz:1985:NIMb

[163] Charles Schwartz. Numerical integration in many dimensions. II. Journal of Mathematical Physics, 26(5):955–957, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p955_s1.

Ikeda:1985:IBI

[164] Satoshi Ikeda. On the intrinsic behavior of the internal variable in the Finslerian gravitational field. Journal of Mathematical Physics, 26(5): 958–960, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p958_s1.

Friedlander:1985:SMR

[165] Leonid Friedlander. A summation method for the Rayleigh–Schrödinger series for the anharmonic oscillator. Journal of Mathematical Physics, 26(5):961–964, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i5/p961_s1.

Nomura:1985:POA

[166] Masao Nomura. On the propagation of the operator average in truncated space. Journal of Mathematical Physics, 26(5):965–969, May 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i5/p965_
s1.

Kroger:1985:SAT

[167] Helmut Kröger. Strong approximation of time evolution operators for a finite system of oscillators with nonlinear coupling. Journal of Mathematical Physics, 26(5):970–972, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i5/p970_s1.

Frank:1985:LAS

[168] Alejandro Frank and Kurt Bernardo Wolf. Lie algebras for systems with mixed spectra. I. The scattering Pöschl–Teller potential. Journal of Mathematical Physics, 26(5):973–983, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p973_s1.

Elworthy:1985:SCE

[169] David Elworthy, Aubrey Truman, and Keith Watling. The semi-classical expansion for a charged particle on a curved space background. Journal of Mathematical Physics, 26(5):984–990, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p984_s1.

Turner:1985:HRH

[170] Leaf Turner. Hamiltonian representation for helically symmetric magnetic fields. Journal of Mathematical Physics, 26(5):991–993, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i5/
p991_s1.

Jin:1985:LCT

[171] Y. Q. Jin and J. A. Kong. Ladder and cross terms in second-order distorted Born approximation. Journal of Mathematical Physics, 26(5): 994–1011, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p994_s1.

Ramm:1985:CRN

[172] A. G. Ramm. Calculating resonances (natural frequencies) and extracting them from transient fields. Journal of Mathematical Physics, 26(5): 1012–1020, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p1012_s1. See erratum [548].

Ferdon:1985:DPI

[173] D. W. Ferdon and J. L. Safko. Dual physical interpretation of the energy tensor: nonexistence of the Kerr–Newman example. Journal of Mathematical Physics, 26(5):1021–1023, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i5/p1021_s1.

Isenberg:1985:SCC

[174] James Isenberg and Vincent Moncrief. Symmetries of cosmological Cauchy horizons with exceptional orbits. Journal of Mathematical Physics, 26(5):1024–1027, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i5/p1024_s1.

Papini:1985:RSR

[175] G. Papini and M. Weiss. Relativistic superfluid in a rigid solid. Exact solutions. Journal of Mathematical Physics, 26(5):1028–1029, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i5/p1028_s1.

Dray:1985:RBM

[176] Tevian Dray. The relationship between monopole harmonics and spinweighted spherical harmonics. Journal of Mathematical Physics, 26(5): 1030–1033, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p1030_s1.

Iyer:1985:SDE

[177] B. R. Iyer and C. V. Vishveshwara. Separability of the Dirac equation in a class of perfect fluid space–times with local rotational symmetry. Journal of Mathematical Physics, 26(5):1034–1039, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i5/p1034_s1.

Dimakis:1985:SEC

[178] A. Dimakis and F. Müller-Hoissen. Solutions of the Einstein–Cartan– Dirac equations with vanishing energy-momentum tensor. Journal of Mathematical Physics, 26(5):1040–1048, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i5/p1040_s1.

Blobel:1985:ESI

[179] A. Blobel and J. Messer. Equilibrium states of inhomogeneous meanfield quantum spin systems with random sites. Journal of Mathematical Physics, 26(5):1049–1056, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i5/p1049_s1.

Volkel:1985:CQG

[180] A. H. Völkel. Confinement of quarks and gluons in gauge quantum field theory. Journal of Mathematical Physics, 26(5):1057–1064, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i5/p1057_s1.

Manoukian:1985:PDT

[181] Edward B. Manoukian. Proof of the decoupling theorem of field theory in Minkowski space. II. Theories involving particles with vanishingly small masses. Journal of Mathematical Physics, 26(5):1065–1069, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i5/p1065_s1.

Baake:1985:FIR

[182] M. Baake, P. Reinicke, and V. Rittenberg. Fierz identities for real Clifford algebras and the number of supercharges. Journal of Mathematical Physics, 26(5):1070–1071, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i5/p1070_s1.

Hussin:1985:CSC

[183] V. Hussin and S. Sinzinkayo. Conformal symmetry and constants of motion. Journal of Mathematical Physics, 26(5):1072–1076, May 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i5/p1072_
s1.

Kaempffer:1985:RAN

[184] F. A. Kaempffer. Remarks about the nilpotent symmetries of quantum gauge field theories. Journal of Mathematical Physics, 26(5):1077– 1082, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i5/p1077_s1.

Chiang:1985:LDH

[185] C. C. Chiang, S. C. Lee, and G. Marmo. Lagrangian dynamics on higherdimensional spaces with applications to Kaluza–Klein theories. Journal of Mathematical Physics, 26(5):1083–1092, May 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i5/p1083_s1.

Mavroyannis:1985:SAS

[186] Constantine Mavroyannis. Strong amplification of sidebands in a strongly driven three-level atomic system. Journal of Mathematical Physics, 26(5):1093–1101, May 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i5/p1093_s1.

Castrigiano:1985:CIS

[187] D. P. L. Castrigiano and U. Mutze. On the commutant of an irreducible set of operators in real Hilbert space. Journal of Mathematical Physics, 26(6):1107–1110, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i6/p1107_s1.

Bugajska:1985:ISF

[188] Krystyna Bugajska. Internal structure of fermions. Journal of Mathematical Physics, 26(6):1111–1117, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i6/p1111_s1.

Mathot:1985:TPU

[189] Françoise Mathot. Topological properties of unbounded bicommutants. Journal of Mathematical Physics, 26(6):1118–1124, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1118_s1.

Brudno:1985:NZR

[190] Simcha Brudno and James D. Louck. Nontrivial zeros of the Racah quadrupole invariant. Journal of Mathematical Physics, 26(6):1125– 1127, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1125_s1.

Fritzer:1985:SCA

[191] Harald P. Fritzer and Bruno Gruber. Symmetry chains and adaptation coefficients. Journal of Mathematical Physics, 26(6):1128–1139, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1128_s1.

Sarma:1985:SGR

[192] C. R. Sarma and J. Paldus. Spinor group and its restrictions. Journal of Mathematical Physics, 26(6):1140–1145, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i6/p1140_s1.

Backhouse:1985:GAC

[193] N. B. Backhouse and A. G. Fellouris. Grassmann analogs of classical matrix groups. Journal of Mathematical Physics, 26(6):1146–1151, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1146_s1.

Jensen:1985:ISR

[194] Arne Jensen and Wolfgang Krüger. On the infrared singularity of the resolvent of some Yang–Mills-type operators. Journal of Mathematical Physics, 26(6):1152–1157, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1152_s1.

Case:1985:SHO

[195] K. M. Case. Symmetries of the higher-order KP equations. Journal of Mathematical Physics, 26(6):1158–1159, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i6/p1158_s1.

Sen:1985:CIF

[196] D. K. Sen. Connections on infinitesimal fiber bundles and unified theories. Journal of Mathematical Physics, 26(6):1160–1161, June 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1160_
s1.

Sen:1985:MCB

[197] D. K. Sen. Mapping of connections on bundles and gauge field theories. Journal of Mathematical Physics, 26(6):1162–1164, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1162_s1.

Reichert:1985:LLT

[198] P. Reichert and R. Schilling. A local limit theorem for strongly dependent random variables and its application to a chaotic configuration of atoms. Journal of Mathematical Physics, 26(6):1165–1172, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1165_s1.

Jumarie:1985:GDE

[199] Guy Jumarie. Generalization of the diffusion equation by using the maximum entropy principle. Journal of Mathematical Physics, 26(6):1173– 1174, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1173_s1.

Nemeth:1985:GPG

[200] G. Németh and G. Páris. The Gibbs phenomenon in generalized Padé approximation. Journal of Mathematical Physics, 26(6):1175–1178, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1175_s1.

Ari:1985:ELT

[201] Naşit Arı and Metin Demiralp. Energy levels of a two-dimensional anharmonic oscillator: characteristic function approach. Journal of Mathematical Physics, 26(6):1179–1185, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i6/p1179_s1.

Donoso:1985:IAD

[202] Guillermo Donoso and Pablo Martín. Ion-acoustic dispersion relation with direct fractional approximation for Z′(s). Journal of Mathematical Physics, 26(6):1186–1188, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1186_s1.

Lakshmanan:1985:LBS

[203] M. Lakshmanan and K. M. Tamizhmani. Lie–Bäcklund symmetries of certain nonlinear evolution equations under perturbation around their solutions. Journal of Mathematical Physics, 26(6):1189–1200, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/
p1189_s1.

Case:1985:TLH

[204] K. M. Case. Theorem on linearized Hamiltonian systems. Journal of Mathematical Physics, 26(6):1201–1202, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i6/p1201_s1.

Cornille:1985:ESS

[205] Henri Cornille. Exact solution for the spatially inhomogeneous nonlinear Kac model of the Boltzmann equation. Journal of Mathematical Physics, 26(6):1203–1214, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1203_s1.

Rao:1985:ITA

[206] Jyoti A. Rao and Abbas A. Rangwala. Infinitesimal transformations about soliton solutions of sine-Gordon and modified Korteweg–de Vries equations. Journal of Mathematical Physics, 26(6):1215–1218, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/
p1215_s1.

Augusteijn:1985:BSF

[207] Marijke F. Augusteijn and Ernst Breitenberger. Bifurcations in the slowfluctuation technique. Journal of Mathematical Physics, 26(6):1219– 1227, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1219_s1.

Yamada:1985:RPJ

[208] Katsuya Yamada. Regular precession in Jacobi’s decomposition theorem. Journal of Mathematical Physics, 26(6):1228–1233, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1228_s1.

Choubey:1985:EFD

[209] Krishna Raj Choubey. Elasticoviscous flow due to a plate which starts oscillating in the presence of a parallel stationary plate. Journal of Mathematical Physics, 26(6):1234–1236, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i6/p1234_s1.

Nutku:1985:NCC

[210] Y. Nutku. On a new class of completely integrable nonlinear wave equations. I. Infinitely many conservation laws. Journal of Mathematical Physics, 26(6):1237–1242, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1237_s1.

Ferraris:1985:EMT

[211] M. Ferraris and M. Francaviglia. Energy-momentum tensors and stress tensors in geometric field theories. Journal of Mathematical Physics, 26(6):1243–1252, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i6/p1243_s1.

Royer:1985:CET

[212] Antoine Royer. Cumulant expansions for thermal density matrices and ground state logarithmic perturbation series. Journal of Mathematical Physics, 26(6):1253–1260, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1253_s1.

Gordov:1985:QEP

[213] E. P. Gordov and A. Z. Fazliev. Quantum electrodynamic perturbation theory based on semiclassical representation. Journal of Mathematical Physics, 26(6):1261–1263, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1261_s1.

Warchall:1985:AST

[214] Henry A. Warchall. C^{∗}-algebraic scattering theory and explicitly solvable quantum field theories. Journal of Mathematical Physics, 26(6):1264– 1279, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1264_s1.

Warchall:1985:SPM

[215] Henry A. Warchall. Simple physical models entailing inequivalent representations of the CCR. Journal of Mathematical Physics, 26(6):1280– 1297, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1280_s1.

Dimakis:1985:GTP

[216] Aristophanes Dimakis. Gauge theory of the post-Galilean groups. Journal of Mathematical Physics, 26(6):1298–1307, June 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1298_s1.

Weber:1985:KSC

[217] Edmond Weber. Kantowski–Sachs cosmological models as big-bang models. Journal of Mathematical Physics, 26(6):1308–1310, June 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1308_
s1.

delOlmo:1985:WRA

[218] V. del Olmo and J. Olivert. Weak rigidity in almost-thermodynamic material schemes. Journal of Mathematical Physics, 26(6):1311–1314, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1311_s1.

Duttachoudhury:1985:SGF

[219] S. B. Duttachoudhury and N. Banerjee. Static gravitational field in generalized scalar tensor theory. Journal of Mathematical Physics, 26(6): 1315–1319, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1315_s1.

Singh:1985:SSV

[220] Tarkeshwar Singh. Spherically symmetric vacuum field in the general scalar-tensor theory. Journal of Mathematical Physics, 26(6):1320–1322, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1320_s1.

Chinea:1985:VEE

[221] F. J. Chinea and F. Guil Guerrero. The vacuum Einstein equations and deformations of the Bianchi identity. Journal of Mathematical Physics, 26(6):1323–1329, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1323_s1.

Bajer:1985:CSE

[222] Konrad Bajer and Jerzy Klemens Kowalczyński. A class of solutions of the Einstein–Maxwell equations. Journal of Mathematical Physics, 26(6): 1330–1333, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1330_s1.

Delbourgo:1985:CGA

[223] R. Delbourgo and T. Matsuki. Chiral and gravitational anomalies in any dimension. Journal of Mathematical Physics, 26(6):1334–1336, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1334_s1.

Leaf:1985:RLE

[224] Boris Leaf. Realization of latent ensembles of the statistical operator. Journal of Mathematical Physics, 26(6):1337–1341, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1337_s1.

Radin:1985:TPC

[225] Charles Radin. Tiling, periodicity, and crystals. Journal of Mathematical Physics, 26(6):1342–1344, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1342_s1.

Grassini:1985:UIT

[226] Renato Grassini. On Utiyama’s invariant theoretical interpretation of interaction. Journal of Mathematical Physics, 26(6):1345–1347, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1345_s1.

Hostler:1985:SFQ

[227] Levere C. Hostler. Scalar formalism for quantum electrodynamics. Journal of Mathematical Physics, 26(6):1348–1354, June 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i6/p1348_s1. See er-

ratum [794].

Basri:1985:RQT

[228] S. A. Basri and A. O. Barut. Relativistic quantum theory of fermions based on the Clifford algebra C_{7}. Journal of Mathematical Physics, 26(6):

1355–1362, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1355_s1.

Nash:1985:NSW

[229] Patrick L. Nash. A new spin-1/2 wave equation. Journal of Mathematical Physics, 26(6):1363–1367, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1363_s1.

Burzlaff:1985:GET

[230] Jürgen Burzlaff and Vincent Moncrief. The global existence of timedependent vortex solutions. Journal of Mathematical Physics, 26(6): 1368–1372, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1368_s1.

Mangiarotti:1985:GSG

[231] Luigi Mangiarotti and Marco Modugno. On the geometric structure of gauge theories. Journal of Mathematical Physics, 26(6):1373–1379, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1373_s1.

Garcilazo:1985:CRE

[232] Humberto Garcilazo. Completeness relation for energy-dependent separable potentials. Journal of Mathematical Physics, 26(6):1380–1386, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i6/p1380_s1.

Catenacci:1985:MCU

[233] Roberto Catenacci and Franco Salmistraro. Multiply connected universes and inequivalent fields. Journal of Mathematical Physics, 26(6):1387– 1389, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1387_s1.

Barut:1985:BSR

[234] A. O. Barut and A. J. Bracken. Bound states, resonances, and symmetries of a neutral Dirac particle with anomalous magnetic moment, coupled to a fixed monopole. Journal of Mathematical Physics, 26(6):

1390–1396, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1390_s1.

Jellinek:1985:KTA

[235] Julius Jellinek. Kinematics of the two-angle-dependent reactive infiniteorder sudden approximation. Journal of Mathematical Physics, 26(6): 1397–1403, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1397_s1.

Gorman:1985:VFN

[236] Arthur D. Gorman. Vector fields near caustics. Journal of Mathematical Physics, 26(6):1404–1407, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1404_s1.

Berryman:1985:SSI

[237] James G. Berryman. Scattering by a spherical inhomogeneity in a fluidsaturated porous medium. Journal of Mathematical Physics, 26(6):1408– 1419, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1408_s1.

Lerche:1985:MFS

[238] I. Lerche and N. R. Hill. A mean-field solution of the reflection of a spherical acoustic wave from a rough interface. Journal of Mathematical Physics, 26(6):1420–1427, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i6/p1420_s1.

Abraham-Shrauner:1985:LTG

[239] B. Abraham-Shrauner. Lie transformation group solutions of the nonlinear one-dimensional Vlasov equation. Journal of Mathematical Physics, 26(6):1428–1435, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i6/p1428_s1. See erratum [488].

Gerry:1985:ERC

[240] Christopher C. Gerry. Erratum: “Remarks on canonical transformations in phase-space path integrals”, [J. Math. Phys. 24 (1983), no. 4, 874–877; MR 84g:81016]. Journal of Mathematical Physics, 26(6):

1436, June 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i6/p1436_s1. See [2].

Crawford:1985:ADB

[241] J. P. Crawford. On the algebra of Dirac bispinor densities: factorization and inversion theorems. Journal of Mathematical Physics, 26(7):1439– 1441, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1439_s1.

Felix:1985:GMF

[242] Jeffrey W. Felix and Dorian M. Hatch. A generalized Molien function for field theoretical Hamiltonians. Journal of Mathematical Physics, 26(7): 1442–1445, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1442_s1.

Gould:1985:PBS

[243] M. D. Gould and E. G. Kalnins. A projection-based solution to the Sp(2n) state labeling problem. Journal of Mathematical Physics, 26(7): 1446–1457, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1446_s1.

Biedenharn:1985:DFC

[244] L. C. Biedenharn, M. A. Lohe, and James D. Louck. On the denominator function for canonical SU(3) tensor operators. Journal of Mathematical Physics, 26(7):1458–1492, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1458_s1.

Haase:1985:AFS

[245] Richard W. Haase and Philip H. Butler. Algebraic formulas for some nontrivial U_{n} 6j symbols and U_{mn} ⊃ U_{m} × U_{n} 3jm symbols. Journal of Mathematical Physics, 26(7):1493–1513, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1493_s1.

Ozaki:1985:GTAa

[246] Masa aki Ozaki. Group theoretical analysis of the Hartree–Fock– Bogoliubov equation. I. General theory. Journal of Mathematical Physics, 26(7):1514–1520, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i7/p1514_s1.

Ozaki:1985:GTAb

[247] Masa aki Ozaki. Group theoretical analysis of the Hartree–Fock– Bogoliubov equation. II. the case of the electronic system with triclinic lattice symmetry. Journal of Mathematical Physics, 26(7):1521–1528, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1521_s1.

Roberts:1985:AAF

[248] Dana Roberts. An alternate approach to finding and using the Lie group of the Vlasov equation. Journal of Mathematical Physics, 26(7):1529– 1533, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1529_s1.

Phaneuf:1985:PTS

[249] D. Phaneuf and R. T. Sharp. Polynomial tensors for the space groups ℘m and p4m. Journal of Mathematical Physics, 26(7):1534–1539, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1534_s1.

Niukkanen:1985:ORA

[250] A. W. Niukkanen. Operator Racah algebra and Laplace-type expansions of irreducible spherical tensors. Journal of Mathematical Physics, 26(7): 1540–1546, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1540_s1.

Dehesa:1985:PSO

[251] J. S. Dehesa, E. Buendía, and M. A. Sanchez-Buendia. On the polynomial solutions of ordinary differential equations of the fourth order. Journal of Mathematical Physics, 26(7):1547–1552, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1547_s1.

Chen:1985:NMN

[252] Gen Shun Chen and Anthony W. Leung. Nonlinear multigroup neutronflux systems: blowup, decay, and steady states. Journal of Mathematical Physics, 26(7):1553–1559, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1553_s1.

Kalnins:1985:DSM

[253] E. G. Kalnins and Willard Miller, Jr. Differential-Stäckel matrices. Journal of Mathematical Physics, 26(7):1560–1565, July 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1560_s1.

Gang:1985:MSA

[254] Xu Gang, Shu Changqing, and Lin Lei. Multiple scales analysis of a nonlinear ordinary differential equation. Journal of Mathematical Physics, 26(7):1566–1569, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1566_s1.

Xu:1985:MSA

[255] Gang Xu, Chang Qing Shu, and Lei Lin. Multiple scales analysis of a nonlinear ordinary differential equation. Journal of Mathematical Physics, 26(7):1566–1569, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Wendel:1985:GMA

[256] M. H. Wendel and E. R. Hilf. Generalized moments applied to Fermitype functions. Journal of Mathematical Physics, 26(7):1570–1575, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1570_s1.

Hegerfeldt:1985:ISH

[257] Gerhard C. Hegerfeldt. Inequalities of Schwarz and Hölder type for random operators. Journal of Mathematical Physics, 26(7):1576–1577, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1576_s1.

Batchelor:1985:GMV

[258] M. Batchelor. Graded manifolds and vector bundles: a functorial correspondence. Journal of Mathematical Physics, 26(7):1578–1582, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1578_s1.

Bona:1985:STA

[259] Carlos Bona and Bartolomé Coll. On the space–times admitting a synchronization of constant curvature. Journal of Mathematical Physics, 26(7):1583–1591, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i7/p1583_s1.

McKenzie:1985:GLP

[260] Ross H. McKenzie. A gravitational lens produces an odd number of images. Journal of Mathematical Physics, 26(7):1592–1596, July 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1592_
s1.

Nencka-Ficek:1985:TCN

[261] H. Nencka-Ficek. Topological closure as the necessary condition for frustration or phase transitions. Journal of Mathematical Physics, 26(7): 1597–1599, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1597_s1.

Jacobson:1985:RWR

[262] Ted Jacobson. Random walk representations for spinor and vector propagators. Journal of Mathematical Physics, 26(7):1600–1604, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/
p1600_s1.

Gagnon:1985:AIR

[263] L. Gagnon, J. Harnad, P. Winternitz, and J. Hurtubise. Abelian integrals and the reduction method for an integrable Hamiltonian system. Journal of Mathematical Physics, 26(7):1605–1612, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1605_s1.

Leach:1985:CPT

[264] P. G. L. Leach. Classes of potentials of time-dependent central force fields which possess first integrals quadratic in the momenta. Journal of Mathematical Physics, 26(7):1613–1620, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1613_s1.

Klar:1985:UAH

[265] H. Klar. Use of alternative hyperspherical coordinates for three-body systems. Journal of Mathematical Physics, 26(7):1621–1625, July 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1621_
s1.

Zivkovic:1985:GST

[266] Tomislav P. Živković. Generalization of the splitting theorem to arbitrary one- and two-particle operators. Journal of Mathematical Physics, 26(7): 1626–1639, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1626_s1.

Palev:1985:FDR

[267] Tchavdar D. Palev. Finite-dimensional representations of the Lie superalgebra sl(1, 3) in a Gel’fand–Zetlin basis. I. typical representations. Journal of Mathematical Physics, 26(7):1640–1660, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1640_s1.

Bolle:1985:STL

[268] D. Bollé, F. Gesztesy, and W. Schweiger. Scattering theory for longrange systems at threshold. Journal of Mathematical Physics, 26(7): 1661–1674, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1661_s1.

Pittner:1985:TMP

[269] L. Pittner and M. Valjavec. Transition matrix of point interactions as the scaling limit of integrable potentials on the real line. Journal of Mathematical Physics, 26(7):1675–1679, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1675_s1.

Muller-Kirsten:1985:QSM

[270] H. J. W. Müller-Kirsten and A. Wiedemann. Quantization of systems with many degrees of freedom by the method of collective coordinates: quantum mechanics around a classical periodic orbit. Journal of Mathematical Physics, 26(7):1680–1694, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i7/p1680_s1.

Macrae:1985:RBG

[271] Kenneth I. Macrae. On the relation between gauge theories, moving frames, fiber bundles, and parastatistics. Journal of Mathematical Physics, 26(7):1695–1700, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1695_s1.

Wong:1985:DCG

[272] M. K. F. Wong and E. H. Y. Yeh. The Dirac Coulomb Green’s function and its application to relativistic Rayleigh scattering. Journal of Mathematical Physics, 26(7):1701–1710, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i7/p1701_s1.

Antonowicz:1985:DSG

[273] Marek Antonowicz and Wiktor Szczyrba. The dynamical structure of gravitational theories with GL(4, R) connections. Journal of Mathematical Physics, 26(7):1711–1727, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i7/p1711_s1.

Punsly:1985:PIK

[274] Brian Punsly. A physical interpretation of the Kerr solution. Journal of Mathematical Physics, 26(7):1728–1739, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1728_s1.

Kamran:1985:SVR

[275] N. Kamran. Separation of variables for the Rarita–Schwinger equation on all type D vacuum backgrounds. Journal of Mathematical Physics, 26(7): 1740–1742, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1740_s1.

Kowalczynski:1985:CCS

[276] Jerzy Klemens Kowalczyński. On a certain class of solutions of the Einstein–Maxwell equations. Journal of Mathematical Physics, 26(7): 1743–1751, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1743_s1.

Chakraborty:1985:CPF

[277] S. K. Chakraborty and N. Bandyopadhyaya. Charged perfect fluid in rigid rotation. Journal of Mathematical Physics, 26(7):1752–1754, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1752_s1.

Ozsvath:1985:PFG

[278] István Ozsváth, Ivor Robinson, and Krzysztof Rózga. Plane-fronted gravitational and electromagnetic waves in spaces with cosmological constant. Journal of Mathematical Physics, 26(7):1755–1761, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1755_s1.

Phares:1985:ASM

[279] Alain J. Phares, Donald E. Shaw, and Francis J. Wunderlich. An approximate solution of the monomer-dimer problem on a square lattice. II. Journal of Mathematical Physics, 26(7):1762–1768, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1762_s1.

Mellein:1985:MRSb

[280] B. Mellein. A multitype random sequential process. II. Distribution of particle size and vacant space length in the saturation limit. Journal of Mathematical Physics, 26(7):1769–1774, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1769_s1.

Kersten:1985:NLB

[281] Paul H. M. Kersten and Ruud Martini. Nonlocal Lie–Bäcklund transformations of the massive Thirring model. Journal of Mathematical Physics, 26(7):1775–1778, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1775_s1.

Hall:1985:CPE

[282] Richard L. Hall. Coulomb potential envelopes for a relativistic fermion in a central field. Journal of Mathematical Physics, 26(7):1779–1785, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1779_s1.

Sasaki:1985:EFB

[283] R. Sasaki. Explicit forms of Bäcklund transformations for the Grassmannian and CP^{N−1} sigma models. Journal of Mathematical Physics, 26(7): 1786–1792, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1786_s1.

Lee:1985:AGW

[284] Hoong Chien Lee and Michael S. Milgram. On the axial gauge: Ward identities and the separation of infrared and ultraviolet singularities by analytic regularization. Journal of Mathematical Physics, 26(7):1793– 1804, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1793_s1.

OCarroll:1985:CEG

[285] Michael O’Carroll and Wilson Dantas Barbosa. Convergent expansions for glueball masses in strongly coupled 3+1 lattice gauge theories. Journal of Mathematical Physics, 26(7):1805–1809, July 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1805_s1.

Bickeboller:1985:GFP

[286] M. Bickeböller, R. Goldflam, and L. Wilets. The gauge field propagator for QED and linearized QCD in inhomogeneous media. Journal of Mathematical Physics, 26(7):1810–1817, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i7/p1810_s1.

Aldaya:1985:NMC

[287] Víctor Aldaya and José A. de Azcárraga. A note on the meaning of covariant derivatives in supersymmetry. Journal of Mathematical Physics, 26(7):1818–1821, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i7/p1818_s1.

Han:1985:FWF

[288] Qi Zhi Han, Hong Zhou Sun, Mei Zhang, and Da hsuan Feng. The Fock wave function as classified by the supergroup chain U(N/M) ⊃ OSp(N/M) ⊃ O(N) × Sp(M). Journal of Mathematical Physics, 26(7): 1822–1827, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1822_s1.

Luehr:1985:SR

[289] C. P. Luehr and M. Rosenbaum. Supertwistor realization of SU(2, 2/1). Journal of Mathematical Physics, 26(7):1828–1833, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1828_s1.

Luehr:1985:SFB

[290] C. P. Luehr and M. Rosenbaum. Supertwistor fiber bundles as a formalism for supergravities. Journal of Mathematical Physics, 26(7):1834– 1846, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1834_s1.

Gazeau:1985:GFG

[291] J. P. Gazeau. Gauge fixing and Gupta–Bleuler triplets in de Sitter QED. Journal of Mathematical Physics, 26(7):1847–1854, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/p1847_s1.

Tuan:1985:NRP

[292] R. Hong Tuan. A new representation for planar Feynman graphs in terms of strings. Journal of Mathematical Physics, 26(7):1855–1857, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i7/p1855_s1.

Lerche:1985:GSE

[293] I. Lerche. On the general solution to equations modeling a homogeneously broadened injection laser. Journal of Mathematical Physics, 26(7):1858– 1859, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i7/p1858_s1.

Pignotti:1985:FRS

[294] Alberto Pignotti. Flow reversal symmetry in a convective heat exchange model. Journal of Mathematical Physics, 26(7):1860–1862, July 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i7/
p1860_s1.

Delduc:1985:MSS

[295] F. Delduc and M. Gourdin. Mixed supertableaux of the superunitary groups. II. SU(n|m). Journal of Mathematical Physics, 26(8):1865– 1879, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i8/p1865_s1.

Goldin:1985:QMR

[296] Gerald A. Goldin and Ralph Menikoff. Quantum-mechanical representations of the group of diffeomorphisms and local current algebra describing tightly bound composite particles. Journal of Mathematical Physics, 26(8):1880–1884, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p1880_s1.

Kerimov:1985:WCG

[297] G. A. Kerimov. On the Wigner coefficients of the generalized Lorentz groups in the parabolic basis. Journal of Mathematical Physics, 26 (8):1885–1888, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p1885_s1.

Garola:1985:GVD

[298] Claudio Garola and Luigi Solombrino. Generalized FS × W versus Dyson’s classification of irreducible representations. Journal of Mathematical Physics, 26(8):1889–1901, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i8/p1889_s1.

Maekawa:1985:EFH

[299] Takayoshi Maekawa. Explicit form of the Haar measure of U(n) and differential operators. Journal of Mathematical Physics, 26(8):1902– 1909, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i8/p1902_s1.

Maekawa:1985:FII

[300] Takayoshi Maekawa. Formula for invariant integrations on SU(n). Journal of Mathematical Physics, 26(8):1910–1913, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p1910_s1.

Nwachuku:1985:NEE

[301] C. O. Nwachuku and M. A. Rashid. New expressions for the eigenvalues of the invariant operators of the general linear and the orthosymplectic Lie superalgebras. Journal of Mathematical Physics, 26(8):1914– 1920, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i8/p1914_s1.

Sarid:1985:SAS

[302] Uri Sarid. The symmetric and antisymmetric structure constants for SU(6). Journal of Mathematical Physics, 26(8):1921–1924, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/
p1921_s1.

Saenz:1985:CMA

[303] A. W. Sáenz. Canonical map approach to channeling stability in crystals. Journal of Mathematical Physics, 26(8):1925–1935, August 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p1925_
s1.

Epele:1985:SOC

[304] L. N. Epele, H. Fanchiotti, and C. A. García Canal. Second-order corrected Hadamard formulas. Journal of Mathematical Physics, 26 (8):1936–1937, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p1936_s1.

Leaute:1985:PSE

[305] B. Léauté and G. Marcilhacy. On a particular solution of the equation of Ernst with n fields, parametrized by an arbitrary harmonic function. Journal of Mathematical Physics, 26(8):1938–1941, August 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p1938_
s1.

Muzinich:1985:SRT

[306] I. J. Muzinich. Some remarks on torsion in Kaluza–Klein unification. Journal of Mathematical Physics, 26(8):1942–1947, August 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p1942_
s1.

Ozawa:1985:CCE

[307] Masanao Ozawa. Concepts of conditional expectations in quantum theory. Journal of Mathematical Physics, 26(8):1948–1955, August 1985.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/
p1948_s1.

Gibbon:1985:OTD

[308] J. D. Gibbon and M. Tabor. On the one- and two-dimensional Toda lattices and the Painlevé property. Journal of Mathematical Physics, 26(8):1956–1960, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p1956_s1.

Carinena:1985:CTT

[309] J. F. Cariñena, J. Gomis, L. A. Ibort, and N. Román. Canonical transformations theory for presymplectic systems. Journal of Mathematical Physics, 26(8):1961–1969, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i8/p1961_s1.

Hietarinta:1985:HCI

[310] J. Hietarinta. How to construct integrable Fokker–Planck and electromagnetic Hamiltonians from ordinary integrable Hamiltonians. Journal of Mathematical Physics, 26(8):1970–1975, August 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i8/p1970_s1.

Ichikawa:1985:LPO

[311] Yoshi Hiko Ichikawa and Kazu hiro Ino. Lax-pair operators for squaredsum and squared-difference eigenfunctions. Journal of Mathematical Physics, 26(8):1976–1978, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i8/p1976_s1.

Weston:1985:CRA

[312] V. H. Weston. On the convergence of the Rytov approximation for the reduced wave equation. Journal of Mathematical Physics, 26(8):1979– 1985, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i8/p1979_s1.

Janssen:1985:BPP

[313] A. J. E. M. Janssen. Bilinear phase-plane distribution functions and positivity. Journal of Mathematical Physics, 26(8):1986–1994, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i8/p1986_s1.

Ma:1985:PLT

[314] Zhong Qi Ma. Proof of the Levinson theorem by the Sturm–Liouville theorem. Journal of Mathematical Physics, 26(8):1995–1999, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i8/p1995_s1.

Dittrich:1985:TTS

[315] J. Dittrich and P. Exner. Tunneling through a singular potential barrier. Journal of Mathematical Physics, 26(8):2000–2008, August 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p2000_
s1.

Collins:1985:GAS

[316] C. B. Collins. Global aspects of shear-free perfect fluids in general relativity. Journal of Mathematical Physics, 26(8):2009–2017, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i8/p2009_s1.

Herrera:1985:ASA

[317] L. Herrera and J. Ponce de León. Anisotropic spheres admitting a oneparameter group of conformal motions. Journal of Mathematical Physics, 26(8):2018–2023, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p2018_s1.

Graells:1985:IDR

[318] J. Graells, C. Martín, and J. M. Codina. Invariant decomposition of the retarded electromagnetic field. Journal of Mathematical Physics, 26(8):2024–2029, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i8/p2024_s1.

Bergmann:1985:GWT

[319] Peter G. Bergmann and Arthur B. Komar. Generalized Weyl-type gauge geometry. Journal of Mathematical Physics, 26(8):2030–2031, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i8/p2030_s1.

Stoeger:1985:HMC

[320] W. R. Stoeger, A. P. Whitman, and R. J. Knill. The harmonic mapping character of Rosen’s bimetric theory of gravity and the geometry of its harmonic mapping space. Journal of Mathematical Physics, 26(8):2032– 2038, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i8/p2032_s1.

Garbaczewski:1985:BAL

[321] Piotr Garbaczewski. Boson approximants for lattice Fermi systems. Journal of Mathematical Physics, 26(8):2039–2044, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p2039_s1.

GamboaSaravi:1985:FME

[322] R. E. Gamboa Saraví, M. A. Muschietti, F. A. Schaposnik, and J. E. Solomin. ζ-function method and the evaluation of fermion currents. Journal of Mathematical Physics, 26(8):2045–2049, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p2045_s1.

Dullemond:1985:SFP

[323] C. Dullemond and E. van Beveren. Scalar field propagators in anti-de Sitter space–time. Journal of Mathematical Physics, 26(8):2050–2058, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i8/p2050_s1.

Reifler:1985:PCA

[324] Frank Reifler and Randall Morris. A prediction of the Cabibbo angle in the vector model for electroweak interactions. Journal of Mathematical Physics, 26(8):2059–2066, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i8/p2059_s1.

Gambhir:1985:BBR

[325] Y. K. Gambhir, R. S. Nikam, C. R. Sarma, and J. A. Sheikh. Boson basis realization for shell-model fermion problem. Journal of Mathematical Physics, 26(8):2067–2071, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v26/i8/p2067_s1.

ElDabe:1985:EST

[326] Nabil T. El Dabe, El Sayed F. El Shehawey, Galal M. Moatimid, and Abou El Magd A. Mohamed. Electrohydrodynamic stability of two cylindrical interfaces under the influence of a tangential periodic electric field. Journal of Mathematical Physics, 26(8):2072–2081, August 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i8/p2072_
s1.

Glasser:1985:ENC

[327] M. L. Glasser. Erratum: “A novel class of Bessel function integrals”, [J. Math. Phys. 25 (1984), no. 10, 2933–2934; MR 85i:33007]. Journal of Mathematical Physics, 26(8):2082, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i8/p2082_s1. See [9].

Hasiewicz:1985:ETP

[328] Z. Hasiewicz and A. K. Kwaśniewski. Erratum: “Triality principle and G_{2} group in spinor language”, [J. Math. Phys. 26, 6 (1985)]. Journal of Mathematical Physics, 26(8):2082, August 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i8/p2082_s2. See [14].

Lomont:1985:GRM

[329] J. S. Lomont. A generalization and refinement of Madivanane’s theorem on general involutional transformations. Journal of Mathematical Physics, 26(9):2087–2091, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2087_s1.

Brudno:1985:NZWb

[330] Simcha Brudno and James D. Louck. Nontrivial zeros of weight 1 3-j and 6-j coefficients: relation to Diophantine equations of equal sums of like powers. Journal of Mathematical Physics, 26(9):2092–2095, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2092_s1.

Carinena:1985:LOR

[331] José F. Cariñena, Mariano A. del Olmo, and Mariano Santander. Locally operating realizations of transformation Lie groups. Journal of Mathematical Physics, 26(9):2096–2106, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i9/p2096_s1.

Castanos:1985:BRM

[332] O. Castaños, E. Chacón, M. Moshinsky, and C. Quesne. Boson realization of sp(4). I. The matrix formulation. Journal of Mathematical Physics, 26(9):2107–2123, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2107_s1.

DeMeyer:1985:SCM

[333] H. De Meyer, G. Vanden Berghe, J. Van der Jeugt, and P. De Wilde. A set of commuting missing label operators for SO(5) ⊃ SO(3). Journal of Mathematical Physics, 26(9):2124–2126, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2124_s1.

Okubo:1985:BIS

[334] Susumu Okubo. Branching index sum rules for simple Lie algebras. Journal of Mathematical Physics, 26(9):2127–2137, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2127_s1.

Krappe:1985:ESE

[335] Hans J. Krappe. Evaluation of the self-energy of a droplet interacting via a Yukawa force. Journal of Mathematical Physics, 26(9):2138–2140, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2138_s1.

Stefanescu:1985:PRP

[336] I. Sabba Stefanescu. On the phase retrieval problem in two dimensions. Journal of Mathematical Physics, 26(9):2141–2160, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2141_
s1.

Levinger:1985:SNT

[337] J. S. Levinger. Some nodal theorems for noncentral forces. Journal of Mathematical Physics, 26(9):2161–2164, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2161_s1.

Gaffet:1985:ESU

[338] B. Gaffet. Exact symmetries of unidimensional self-similar flow. Journal of Mathematical Physics, 26(9):2165–2167, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2165_s1.

Kalnins:1985:GSM

[339] E. G. Kalnins and W. Miller, Jr. Generalized Stäckel matrices. Journal of Mathematical Physics, 26(9):2168–2173, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2168_s1.

Weiss:1985:MER

[340] John Weiss. Modified equations, rational solutions, and the Painlevé property for the Kadomtsev–Petviashvili and Hirota–Satsuma equations. Journal of Mathematical Physics, 26(9):2174–2180, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2174_
s1.

Klein:1985:GIS

[341] J. J. Klein. Geometrical interpretation of the solutions of the sine-Gordon equation. Journal of Mathematical Physics, 26(9):2181–2185, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2181_s1.

Rno:1985:HAEb

[342] Jung Sik Rno. Harmonic analysis of the Euclidean group in threespace. II. Journal of Mathematical Physics, 26(9):2186–2188, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2186_s1.

Korzeniowski:1985:DSD

[343] Andrzej Korzeniowski. Dynamic systems driven by Markov processes. Journal of Mathematical Physics, 26(9):2189–2191, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2189_
s1.

Palou:1985:NBS

[344] P. Palou, J. Stela, and L. Mas. Noncoordinated basis and Schild’s solution. Journal of Mathematical Physics, 26(9):2192–2195, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2192_s1.

Evans:1985:SMT

[345] J. W. Evans. Splitting methods for time-independent wave propagation in random media. Journal of Mathematical Physics, 26(9):2196–2200, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2196_s1.

Maystre:1985:SCF

[346] D. Maystre and M. Cadilhac. Singularities of the continuation of fields and validity of Rayleigh’s hypothesis. Journal of Mathematical Physics, 26(9):2201–2204, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2201_s1.

Papastamatiou:1985:CPS

[347] N. J. Papastamatiou, H. Matsumoto, and H. Umezawa. Comments on the perturbed sine-Gordon equation. Journal of Mathematical Physics, 26(9):2205–2207, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2205_s1.

Twersky:1985:WDE

[348] Victor Twersky. Wavelength-dependent electromagnetic parameters for coherent propagation in correlated distributions of small-spaced scatterers. Journal of Mathematical Physics, 26(9):2208–2217, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2208_s1.

Gauthier:1985:CSE

[349] N. Gauthier and P. Rochon. Core size effects, bound states, and scattering states of the Aharonov–Bohm problem. Journal of Mathematical Physics, 26(9):2218–2221, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2218_s1.

Barchielli:1985:QSC

[350] A. Barchielli and G. Lupieri. Quantum stochastic calculus, operation valued stochastic processes, and continual measurements in quantum mechanics. Journal of Mathematical Physics, 26(9):2222–2230, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2222_s1.

Durr:1985:PIQ

[351] H. Dürr and A. Inomata. Path integral quantization of the dyonium. Journal of Mathematical Physics, 26(9):2231–2233, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2231_
s1.

Kwasniewski:1985:CGL

[352] A. K. Kwaśniewski. Clifford- and Grassmann-like algebras—old and new. Journal of Mathematical Physics, 26(9):2234–2238, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2234_
s1.

Daubechies:1985:QMP

[353] Ingrid Daubechies and John R. Klauder. Quantum-mechanical path integrals with Wiener measure for all polynomial Hamiltonians. II. Journal of Mathematical Physics, 26(9):2239–2256, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2239_s1.

Patrascioiu:1985:RCF

[354] Adrian Patrascioiu and Jean-Louis Richard. Remarks concerning functional integration on compact spaces. Journal of Mathematical Physics, 26(9):2257–2259, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2257_s1.

Rao:1985:NCZ

[355] K. Srinivasa Rao. A note on the classification of the zeros of angular momentum coefficients. Journal of Mathematical Physics, 26(9):2260– 2261, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2260_s1.

Perez:1985:BSP

[356] J. Fernando Perez, C. P. Malta, and F. A. B. Coutinho. Bound states of N particles: a variational approach. Journal of Mathematical Physics, 26(9):2262–2267, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2262_s1.

Collins:1985:SRP

[357] C. B. Collins. Static relativistic perfect fluids with spherical, plane, or hyperbolic symmetry. Journal of Mathematical Physics, 26(9):2268–2274, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2268_s1.

Bonanos:1985:RSA

[358] S. Bonanos and D. Sklavenites. Relativistic, stationary, axisymmetric perfect fluids. I. Reduction to a system of two equations. Journal of Mathematical Physics, 26(9):2275–2278, September 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2275_s1.

Sklavenites:1985:RSA

[359] D. Sklavenites. Relativistic, stationary, axisymmetric perfect fluids. II. Solutions with vanishing magnetic Weyl tensor. Journal of Mathematical Physics, 26(9):2279–2281, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2279_s1.

Carot:1985:VFI

[360] J. Carot and J. Iba nez. On the viscous fluid interpretation of some exact solutions. Journal of Mathematical Physics, 26(9):2282–2285, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2282_s1.

Couch:1985:SWE

[361] W. E. Couch. Solutions to wave equations on black hole geometries. II. Journal of Mathematical Physics, 26(9):2286–2296, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2286_
s1.

Hajj-Boutros:1985:HHS

[362] J. Hajj-Boutros. On hypersurface-homogeneous space–times. Journal of Mathematical Physics, 26(9):2297–2301, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2297_s1.

Herrera:1985:IAC

[363] L. Herrera and J. Ponce de León. Isotropic and anisotropic charged spheres admitting a one-parameter group of conformal motions. Journal of Mathematical Physics, 26(9):2302–2307, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2302_s1.

Mann:1985:ESA

[364] R. B. Mann. Exact solutions of an algebraically extended Kaluza–Klein theory. Journal of Mathematical Physics, 26(9):2308–2316, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i9/p2308_s1.

Nachtergaele:1985:QAM

[365] Bruno Nachtergaele. Quantum angular momentum fluctuations. Journal of Mathematical Physics, 26(9):2317–2323, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2317_s1.

Sewell:1985:QSD

[366] Geoffrey L. Sewell. Quantum statistical derivation of the hydrodynamics of a plasma model. Journal of Mathematical Physics, 26(9):2324–2334, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2324_s1.

Berceanu:1985:ADS

[367] S. Berceanu and A. Gheorghe. On the asymptotics of distributions with support in cone. Journal of Mathematical Physics, 26(9):2335–2341, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2335_s1.

OCarroll:1985:CEE

[368] Michael O’Carroll. Convergent expansions for excited glueball masses in 2 + 1 strongly coupled lattice gauge theories. Journal of Mathematical Physics, 26(9):2342–2345, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2342_s1.

Mickelsson:1985:VFN

[369] Jouko Mickelsson. The vector form of the neutrino equation and the photon neutrino duality. Journal of Mathematical Physics, 26(9):2346– 2349, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2346_s1.

Ullah:1985:PDF

[370] Nazakat Ullah. Probability density function of the single eigenvalue outside the semicircle using the exact Fourier transform. Journal of Mathematical Physics, 26(9):2350–2351, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i9/p2350_s1.

Furutsu:1985:STS

[371] K. Furutsu. Statistical theory of scattering and propagation through a rough boundary—the Bethe–Salpeter equation and applications. Journal of Mathematical Physics, 26(9):2352–2369, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2352_s1.

Bernardin:1985:HAE

[372] M. P. Bernardin and J. A. Tataronis. Hamiltonian approach to the existence of magnetic surfaces. Journal of Mathematical Physics, 26 (9):2370–2380, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i9/p2370_s1.

Kim:1985:PRA

[373] Shoon K. Kim. The plane rotations allowed for d-dimensional discrete lattices and their application to the Bravais lattices in three dimensions. Journal of Mathematical Physics, 26(9):2381–2382, September 1985. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2381_
s1.

Bloomer:1985:CD

[374] Iris Bloomer and John M. Charap. The continuity of dislocations. Journal of Mathematical Physics, 26(9):2383–2385, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i9/p2383_s1.

Luban:1985:BOO

[375] Marshall Luban. Bloch oscillations in one-dimensional solids and solitary wave packets. Journal of Mathematical Physics, 26(9):2386–2391, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2386_s1.

Hock:1985:ASP

[376] J. L. Hock and R. B. McQuistan. The adsorption of simple particles on a 2 × N lattice space. Journal of Mathematical Physics, 26(9):2392–2395, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i9/p2392_s1.

El-Basil:1985:FGP

[377] Sherif El-Basil, Pavel Křivka, and Nenad Trinajstić. Fibonacci graphs possessing identical matching polynomials. Journal of Mathematical Physics, 26(9):2396–2398, September 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i9/p2396_s1.

Kleeman:1985:CFG

[378] R. Kleeman. Commutation factors on generalized Lie algebras. Journal of Mathematical Physics, 26(10):2405–2412, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2405_s1.

Conte:1985:LPG

[379] Robert Conte and Jacques Raynal. On a labeling for point group harmonics. I. Cubic group. Journal of Mathematical Physics, 26(10):2413– 2440, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2413_s1. See errata [654, 380].

Raynal:1985:LPG

[380] Jacques Raynal. On a labeling for point group harmonics. II. Icosahedral group. Journal of Mathematical Physics, 26(10):2441–2456, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v26/i10/p2441_s1. See errata [654, 379].

Sijacki:1985:APU

[381] Dj. Šijački and Y. Ne’eman. Algebra and physics of the unitary multiplicity-free representations of \overline{SL}(4, R). Journal of Mathematical Physics, 26(10):2457–2464, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i10/p2457_s1.

Sullivan:1985:CCD

[382] John J. Sullivan. Character constraints on duality. Journal of Mathematical Physics, 26(10):2465–2468, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i10/p2465_s1.

Degasperis:1985:NEE

[383] A. Degasperis, P. M. Santini, and M. J. Ablowitz. Nonlinear evolution equations associated with a Riemann–Hilbert scattering problem. Journal of Mathematical Physics, 26(10):2469–2472, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2469_s1.

Grossmann:1985:TAS

[384] A. Grossmann, J. Morlet, and T. Paul. Transforms associated to square integrable group representations. I. General results. Journal of Mathematical Physics, 26(10):2473–2479, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i10/p2473_s1.

Maekawa:1985:CFR

[385] Takayoshi Maekawa. On a computation formula for the representation matrices of U(n). Journal of Mathematical Physics, 26(10):2480– 2484, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2480_s1.

Sharma:1985:SDI

[386] Veena Sharma and K. C. Tripathy. Satake diagrams, Iwasawa decompositions, and representations of the exceptional Lie group F_{4}(−20). Journal of Mathematical Physics, 26(10):2485–2490, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2485_s1.

Phares:1985:TDR

[387] Alain J. Phares and Francis J. Wunderlich. Thermodynamics of dimers on a rectangular L × M × N lattice. Journal of Mathematical Physics, 26(10):2491–2499, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2491_s1. See comment [1304].

Muller-Kirsten:1985:SEP

[388] H. J. W. Müller-Kirsten and E. Sattel. Solution of the eigenvalue problem of an integral equation with the help of its associated differential equation applied to the calculation of diffraction losses in confocal resonators. Journal of Mathematical Physics, 26(10):2500–2509, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2500_s1.

Leach:1985:FIM

[389] P. G. L. Leach. First integrals for the modified Emden equation q̈ + α(t)q̇ + q^{n} = 0. Journal of Mathematical Physics, 26(10):2510– 2514, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2510_s1.

Barut:1985:CQS

[390] A. O. Barut and A. J. Bracken. Compact quantum systems: internal geometry of relativistic systems. Journal of Mathematical Physics, 26 (10):2515–2519, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2515_s1.

Bulla:1985:DIS

[391] W. Bulla and F. Gesztesy. Deficiency indices and singular boundary conditions in quantum mechanics. Journal of Mathematical Physics, 26(10):2520–2528, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2520_s1.

Raggio:1985:SDL

[392] G. A. Raggio and H. S. Zivi. On the semiclassical description of Nlevel systems interacting with radiation fields. Journal of Mathematical Physics, 26(10):2529–2539, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i10/p2529_s1.

Wan:1985:ALO

[393] K. Kong Wan and R. G. D. McLean. On the algebras of local observables in the generalized sense in quantum mechanics. Journal of Mathematical Physics, 26(10):2540–2545, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i10/p2540_s1.

Albeverio:1985:CBS

[394] S. Albeverio, F. Gesztesy, W. Karwowski, and L. Streit. On the connection between Schrödinger and Dirichlet forms. Journal of Mathematical Physics, 26(10):2546–2553, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i10/p2546_s1.

Mourgues:1985:NSC

[395] G. Mourgues, M. R. Feix, J. C. Andrieux, and P. Bertrand. Not necessary but sufficient condition for the positivity of generalized Wigner functions. Journal of Mathematical Physics, 26(10):2554–2555, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2554_s1.

Behncke:1985:SPD

[396] Horst Behncke. Spectral properties of the Dirac equation with anomalous magnetic moment. Journal of Mathematical Physics, 26(10):2556– 2559, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2556_s1.

Briet:1985:RBE

[397] P. H. Briet. The relation between the (N)- and (N − 1)-electron atomic ground states. Journal of Mathematical Physics, 26(10):2560–2564, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2560_s1.

Ali:1985:QNF

[398] M. K. Ali. The quantum normal form and its equivalents. Journal of Mathematical Physics, 26(10):2565–2572, October 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2565_s1.

Carfora:1985:RMS

[399] Mauro Carfora. Regular maximal slicings in asymptotically flat space– times. Journal of Mathematical Physics, 26(10):2573–2580, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2573_s1.

Geroch:1985:PTM

[400] Robert Geroch and Lee Lindblom. Is perturbation theory misleading in general relativity. Journal of Mathematical Physics, 26(10):2581– 2588, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2581_s1.

Zhong:1985:GNS

[401] Zai Zhe Zhong. Generation of new solutions of the stationary axisymmetric Einstein equations by a double complex function method. Journal of Mathematical Physics, 26(10):2589–2595, October 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2589_s1.

Schutze:1985:UEG

[402] Dieter Schütze. On unifying electromagnetism and gravitation without curvature. Journal of Mathematical Physics, 26(10):2596–2600, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2596_s1.

Kasperkovitz:1985:FHRa

[403] P. Kasperkovitz and J. Reisenberger. Finite hard rod systems and their thermodynamic limit. II. Ergodic properties, pair distribution function, and equation of state. Journal of Mathematical Physics, 26(10):2601– 2616, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2601_s1.

Kasperkovitz:1985:FHRb

[404] P. Kasperkovitz and J. Reisenberger. Finite hard rod systems and their thermodynamic limit. III. Collision frequency, no-collision probability, and velocity autocorrelation function. Journal of Mathematical Physics, 26(10):2617–2626, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2617_s1.

Weaver:1985:BDA

[405] D. L. Weaver. Brownian dynamics approximation to the decay rate of a metastable state. Journal of Mathematical Physics, 26(10):2627– 2631, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i10/p2627_s1.

Fernandez:1985:GIM

[406] Ariel Fernández. Global instability of a monoparametric family of vector fields representing the unfolding of a dissipative structure. Journal of Mathematical Physics, 26(10):2632–2633, October 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2632_s1.

Yoakam:1985:PPQa

[407] M. C. Yoakam. Practical pretheories of QED. I. Properties of the interaction picture. Journal of Mathematical Physics, 26(10):2634–2638, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2634_s1.

Yoakam:1985:PPQb

[408] M. C. Yoakam. Practical pretheories of QED. II. Choosing the interaction. Journal of Mathematical Physics, 26(10):2639–2642, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2639_s1.

Kursawe:1985:HBY

[409] Ulrich Kursawe and Edward Malec. Hopf bifurcation in Yang–Mills mechanics. Journal of Mathematical Physics, 26(10):2643–2647, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2643_s1.

Brandas:1985:TWT

[410] Erkki Brändas, Magnus Rittby, and Nils Elander. Titchmarsh–Weyl theory and its relations to scattering theory: spectral densities and cross sections; theory and applications. Journal of Mathematical Physics, 26 (10):2648–2658, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2648_s1.

Ahmed:1985:ESA

[411] E. Ahmed, B. L. Rands, and J. G. Taylor. On extensions of supersymmetry algebras in d = 4 and 11. Journal of Mathematical Physics, 26(10):2659–2665, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2659_s1.

Hostler:1985:QTP

[412] Levere Hostler. Quantum theory of particles of indefinite mass: spin-1/2. Journal of Mathematical Physics, 26(10):2666–2669, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i10/
p2666_s1.

Deb:1985:ASP

[413] N. C. Deb, N. C. Sil, and J. H. McGuire. Analytic strong-potential Born approximation cross sections for electron capture from the K shell by fast ions. Journal of Mathematical Physics, 26(10):2670–2675, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i10/p2670_s1.

Mandelis:1985:HJF

[414] Andreas Mandelis. Hamilton–Jacobi formulation and quantum theory of thermal wave propagation in the solid state. Journal of Mathematical Physics, 26(10):2676–2683, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i10/p2676_s1.

Garbaczewski:1985:FSBb

[415] Piotr Garbaczewski. Fermi states of Bose systems in three space dimensions. II. The solvable version of the Hubbard model. Journal of Mathematical Physics, 26(10):2684–2692, October 1985. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v26/i10/p2684_s1.

Sanyal:1985:HSS

[416] Abhik Kumar Sanyal, A. Banerjee, and Dipankar Ray. Hydrodynamic self-similar cosmological models. Journal of Mathematical Physics, 26 (10):2693–2699, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i10/p2693_s1.

Kasperkovitz:1985:FHRc

[417] P. Kasperkovitz and J. Reisenberger. Finite hard rod systems and their thermodynamic limit. II. ergodic properties, pair distribution function, and equation of state. Journal of Mathematical Physics, 26(10):26010– 26016, October 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Deenen:1985:BRR

[418] J. Deenen and C. Quesne. Boson representations of the real symplectic group and their applications to the nuclear collective model. Journal of Mathematical Physics, 26(11):2705–2716, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/p2705_s1.

Rubinstein:1985:IPC

[419] R. Rubinstein and J. E. Fitzgerald. An inverse problem in crystallography. Journal of Mathematical Physics, 26(11):2717–2720, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2717_s1.

Schmidt:1985:CDW

[420] Jeffrey R. Schmidt. On combinatoric determination of SU(n) weight multiplicities. Journal of Mathematical Physics, 26(11):2721–2728, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2721_s1.

Buendia:1985:ZEP

[421] E. Buendía, J. S. Dehesa, and M. A. Sanchez-Buendia. On the zeros of eigenfunctions of polynomial differential operators. Journal of Mathematical Physics, 26(11):2729–2736, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i11/p2729_s1.

Auberson:1985:SEP

[422] G. Auberson and S. Ciulli. Solution of an extremum problem pertaining to analytic extrapolation techniques. Journal of Mathematical Physics, 26(11):2737–2745, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2737_s1.

Joseph:1985:CYM

[423] K. Babu Joseph and B. V. Baby. Classical SU(2) Yang–Mills–Higgs system: time-dependent solutions by similarity method. Journal of Mathematical Physics, 26(11):2746–2748, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i11/p2746_s1.

Rogers:1985:EGI

[424] Alice Rogers. On the existence of global integral forms on supermanifolds. Journal of Mathematical Physics, 26(11):2749–2753, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2749_s1.

Kupershmidt:1985:MAQ

[425] B. A. Kupershmidt. Mathematical aspects of quantum fluids. I. generalized two-cycles of ^{4}He type. Journal of Mathematical Physics, 26 (11):2754–2758, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2754_s1.

Liboff:1985:CDE

[426] Richard L. Liboff. On a class of differential equations derived from plasma statistics. Journal of Mathematical Physics, 26(11):2759–2760, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2759_s1.

Nevard:1985:RRE

[427] John Nevard and Joseph B. Keller. Reciprocal relations for effective conductivities of anisotropic media. Journal of Mathematical Physics, 26 (11):2761–2765, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2761_s1.

Kupershmidt:1985:SGN

[428] B. A. Kupershmidt. The structures of generalized noncommutative Toda lattices. Journal of Mathematical Physics, 26(11):2766–2768, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2766_s1.

Herman:1985:FNT

[429] Warren N. Herman. Formulation of Noether’s theorem for Fokker-type variational principles. Journal of Mathematical Physics, 26(11):2769– 2776, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2769_s1.

Dabrowski:1985:NPI

[430] L. Dabrowski and H. Grosse. On nonlocal point interactions in one, two, and three dimensions. Journal of Mathematical Physics, 26(11):2777– 2780, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2777_s1.

Golin:1985:URS

[431] S. Golin. Uncertainty relations in stochastic mechanics. Journal of Mathematical Physics, 26(11):2781–2783, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i11/p2781_s1.

Li:1985:SSE

[432] Wai Kee Li and S. M. Blinder. Solution of the Schrödinger equation for a particle in an equilateral triangle. Journal of Mathematical Physics, 26(11):2784–2786, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2784_s1.

Rowe:1985:VCS

[433] D. J. Rowe, G. Rosensteel, and R. Gilmore. Vector coherent state representation theory. Journal of Mathematical Physics, 26(11):2787–2791, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2787_s1.

Kus:1985:STL

[434] Marek Kuś. On the spectrum of a two-level system. Journal of Mathematical Physics, 26(11):2792–2795, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i11/p2792_s1.

Osborn:1985:SSR

[435] T. A. Osborn, K. B. Sinha, D. Bollé, and C. Danneels. Spectral sum rule for time delay in R^{2}. Journal of Mathematical Physics, 26(11):2796– 2802, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2796_s1.

Rose:1985:TDI

[436] James H. Rose, Margaret Cheney, and Brian DeFacio. Three-dimensional inverse scattering: plasma and variable velocity wave equations. Journal of Mathematical Physics, 26(11):2803–2813, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/p2803_s1.

Belinfante:1985:GIT

[437] Frederik J. Belinfante. Generating infinitesimal transformations with second-order-infinitesimal accuracy for proving covariance of commutation relations under finite transformations. Journal of Mathematical Physics, 26(11):2814–2826, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i11/p2814_s1.

Belinfante:1985:CDM

[438] Frederik J. Belinfante. On the covariance of Dirac’s modified canonical commutation relations for fermions interacting with gravity and bosons. Journal of Mathematical Physics, 26(11):2827–2835, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2827_s1.

Belinfante:1985:DCP

[439] Frederik J. Belinfante. The De Donder condition and the Poincaré group in quantized general relativity. Journal of Mathematical Physics, 26 (11):2836–2839, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2836_s1.

Maiti:1985:EMF

[440] S. R. Maiti and S. N. Das. Equivalence of a magnetohydrodynamic fluid and a viscous fluid with heat flux. Journal of Mathematical Physics, 26 (11):2840–2846, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2840_s1.

Herrera:1985:CGF

[441] L. Herrera and J. Ponce de León. Confined gravitational fields produced by anisotropic fluids. Journal of Mathematical Physics, 26(11):2847– 2849, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2847_s1.

Kandrup:1985:SMC

[442] Henry E. Kandrup. Statistical mechanics in a conformally static setting. II. Journal of Mathematical Physics, 26(11):2850–2858, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2850_s1.

Isaacson:1985:GRD

[443] R. A. Isaacson, J. S. Welling, and J. Winicour. Gravitational radiation from dust. Journal of Mathematical Physics, 26(11):2859–2868, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2859_s1.

Maartens:1985:CFG

[444] R. Maartens and S. D. Maharaj. Collision-free gases in spatially homogeneous space–times. Journal of Mathematical Physics, 26(11):2869– 2880, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2869_s1.

Mason:1985:SCK

[445] D. P. Mason and M. Tsamparlis. Spacelike conformal Killing vectors and spacelike congruences. Journal of Mathematical Physics, 26(11):2881– 2901, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2881_s1.

Saez:1985:STT

[446] D. Sáez and J. V. Arnau. On the scalar-tetradic theory A—the Birkhoff theorem: horizons and singularities. Journal of Mathematical Physics, 26(11):2902–2908, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p2902_s1.

Bugajska:1985:SSM

[447] Krystyna Bugajska. From spinor structure to magnetic monopoles. Journal of Mathematical Physics, 26(11):2909–2914, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/
p2909_s1.

Ram:1985:BTS

[448] Shri Ram. Bianchi type-VI_{0} solutions in modified Brans–Dicke cosmology. Journal of Mathematical Physics, 26(11):2915–2917, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2915_s1.

Toscani:1985:GEA

[449] G. Toscani. Global existence and asymptotic behavior for the discrete velocity models of the Boltzmann equation. Journal of Mathematical Physics, 26(11):2918–2921, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i11/p2918_s1.

Hara:1985:SIM

[450] Takashi Hara, Tetsuya Hattori, and Hal Tasaki. Skeleton inequalities and mean field properties for lattice spin systems. Journal of Mathematical Physics, 26(11):2922–2929, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i11/p2922_s1.

Mellein:1985:MRSc

[451] B. Mellein. A multitype random sequential process. III. The case of constant target area. Journal of Mathematical Physics, 26(11):2930– 2939, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2930_s1.

Leblanc:1985:PTP

[452] Y. Leblanc, H. Matsumoto, and H. Umezawa. Perturbation theory for polyacetylene-type kink dynamics model with acoustic phonon effects. Journal of Mathematical Physics, 26(11):2940–2955, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/
p2940_s1.

Volkel:1985:URS

[453] A. H. Völkel. Unbounded representations of symmetry groups in gauge quantum field theory. I. Confinement and differentiation. Journal of Mathematical Physics, 26(11):2956–2973, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/p2956_s1.

Grassini:1985:GSU

[454] Renato Grassini. The geometrical setting of Utiyama’s interaction theory. Journal of Mathematical Physics, 26(11):2974–2976, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/
p2974_s1.

Slim:1985:RRA

[455] H. A. Slim. Recurrence relations for the analytic calculation of Feynman integrals in the axial gauge: the case of massless particles. Journal of Mathematical Physics, 26(11):2977–2984, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/p2977_s1.

Flyvbjerg:1985:CTE

[456] Henrik Flyvbjerg. Character table for the 1080-element point-group-like subgroup of SU(3). Journal of Mathematical Physics, 26(11):2985– 2989, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i11/p2985_s1.

Papadopoulos:1985:SSS

[457] Demetrios Papadopoulos. Soliton solutions for self-dual SU(3) gauge field theory. Journal of Mathematical Physics, 26(11):2990–2994, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i11/p2990_s1.

Moshinsky:1985:CGI

[458] M. Moshinsky, M. Nicolescu, and R. T. Sharp. Collectivity and geometry. IV. Sp(6) ⊃ Sp(2) × O(3) basis states for open shells. Journal of Mathematical Physics, 26(11):2995–2998, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i11/p2995_s1.

Borwein:1985:CLS

[459] David Borwein, Jonathan M. Borwein, and Keith F. Taylor. Convergence of lattice sums and Madelung’s constant. Journal of Mathematical Physics, 26(11):2999–3009, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i11/p2999_s1.

Banerjee:1985:BTC

[460] A. Banerjee, S. B. Dutta Choudhury, and Abhik Kumar Sanyal. Bianchi type I cosmological model with a viscous fluid. Journal of Mathematical Physics, 26(11):3010–3015, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i11/p3010_s1.

vanHaeringen:1985:ENI

[461] H. van Haeringen and L. P. Kok. Erratum: “New inequalities for the Coulomb T matrix in momentum space”, [J. Math. Phys. 25 (1984), no. 10, 3033–3038; MR 85m:81129]. Journal of Mathematical Physics, 26(11):3016, November 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i11/p3016_s1. See [10].

Furlan:1985:NSR

[462] P. Furlan and R. Raczka. Nonlinear spinor representations. Journal of Mathematical Physics, 26(12):3021–3032, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3021_s1.

Saeed-ul-Islam:1985:PRF

[463] M. Saeed ul Islam. On the projective representations of finite abelian groups. II. Journal of Mathematical Physics, 26(12):3033–3035, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3033_s1.

Cirelli:1985:GGT

[464] Renzo Cirelli and Alessandro Manià. The group of gauge transformations as a Schwartz–Lie group. Journal of Mathematical Physics, 26(12):3036– 3041, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3036_s1.

Sastri:1985:LSS

[465] C. C. A. Sastri and K. A. Dunn. Lie symmetries of some equations of the Fokker–Planck type. Journal of Mathematical Physics, 26(12):3042– 3047, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3042_s1.

Bonnier:1985:ICL

[466] B. Bonnier, M. Hontebeyrie, and E. H. Ticembal. The infinite coupling limit of perturbative expansions from a variational extrapolation method. Journal of Mathematical Physics, 26(12):3048–3052, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3048_s1.

Gilmore:1985:DGC

[467] R. Gilmore and J. P. Draayer. Dynamical group chains and integrity bases. Journal of Mathematical Physics, 26(12):3053–3067, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3053_s1.

Bass:1985:GFF

[468] Ronald Bass. The Green’s function for a finite linear chain. Journal of Mathematical Physics, 26(12):3068–3069, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3068_s1.

Dorizzi:1985:IHS

[469] B. Dorizzi, B. Grammaticos, A. Ramani, and P. Winternitz. Integrable Hamiltonian systems with velocity-dependent potentials. Journal of Mathematical Physics, 26(12):3070–3079, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3070_s1.

Katzin:1985:CFSa

[470] Gerald H. Katzin and Jack Levine. Characteristic functional structure of infinitesimal symmetry mappings of classical dynamical systems. I. Velocity-dependent mappings of second-order differential equations. Journal of Mathematical Physics, 26(12):3080–3099, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/
p3080_s1.

Katzin:1985:CFSb

[471] Gerald H. Katzin and Jack Levine. Characteristic functional structure of infinitesimal symmetry mappings of classical dynamical systems. II. Mappings of first-order differential equations. Journal of Mathematical Physics, 26(12):3100–3104, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i12/p3100_s1.

Cohen:1985:ERD

[472] Leon Cohen and Chongmoon Lee. Exact reduced density matrices for a model problem. Journal of Mathematical Physics, 26(12):3105–3108, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3105_s1.

DeMeyer:1985:SSE

[473] H. De Meyer, G. Vanden Berghe, and J. Van der Jeugt. On the spectra of SO(3) scalars in the enveloping algebra of SU(3). Journal of Mathematical Physics, 26(12):3109–3111, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v26/i12/p3109_s1.

Magnon:1985:KIA

[474] Anne Magnon. On Komar integrals in asymptotically anti-de Sitter space–times. Journal of Mathematical Physics, 26(12):3112–3117, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3112_s1.

Argueso:1985:SEI

[475] F. Argüeso and J. L. Sanz. Some exact inhomogeneous solutions of Einstein’s equations with symmetries on the hypersurfaces t = const. Journal of Mathematical Physics, 26(12):3118–3124, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3118_s1.

Banerjee:1985:ACM

[476] A. Banerjee, N. Banerjee, and N. O. Santos. Anisotropic cosmological model in Nordtvedt’s scalar-tensor theory of gravitation. Journal of Mathematical Physics, 26(12):3125–3130, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3125_s1.

Marques:1985:EQM

[477] S. Marques and C. G. Oliveira. An extension of quaternionic metrics to octonions. Journal of Mathematical Physics, 26(12):3131–3139, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3131_s1.

Saxena:1985:TMP

[478] G. M. Saxena and C. L. Mehta. Two-mode para-Bose coherent states. Journal of Mathematical Physics, 26(12):3140–3145, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/
p3140_s1.

Bailly:1985:ETC

[479] F. Bailly. On the energy-time conjugation in quantum physics. Journal of Mathematical Physics, 26(12):3146–3154, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3146_s1.

Burnel:1985:DQT

[480] A. Burnel and M. Van Der Rest-Jaspers. Dirac quantization of a three-dimensional gauge theory. Journal of Mathematical Physics, 26 (12):3155–3159, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v26/i12/p3155_s1.

Trostel:1985:CAT

[481] Robert Trostel. Color analysis, theory of Γ-graded integrable evolution equations, and super Nijenhuis operators. Journal of Mathematical Physics, 26(12):3160–3171, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v26/i12/p3160_s1.

Stacey:1985:SDD

[482] Richard Stacey. Spectrum doubling and double-valuedness. Journal of Mathematical Physics, 26(12):3172–3175, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3172_s1.

Coquereaux:1985:DCD

[483] Robert Coquereaux. Differential characters: the Dirac monopole as an example. Journal of Mathematical Physics, 26(12):3176–3179, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3176_s1.

Cant:1985:SIE

[484] A. Cant and Y. Ne’eman. Spinorial infinite equations fitting metric-affine gravity. Journal of Mathematical Physics, 26(12):3180–3189, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3180_s1.

Chakrabarti:1985:FDT

[485] J. Chakrabarti. Fluctuation-dissipation theorem for QCD plasma. Journal of Mathematical Physics, 26(12):3190–3192, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v26/i12/p3190_s1.

Suzuki:1985:SCL

[486] Noboru Suzuki. Simple calculation of Löwdin’s alpha function. II. Easier procedure for evaluating b_{Kk}(LM|l), and vanishing of h_{n,2n−i}(LM|l) for special values of i and n. Journal of Mathematical Physics, 26(12):3193– 3199, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3193_s1.

Dattoli:1985:HPF

[487] G. Dattoli and R. Mignani. Hamiltonian picture of the free electron laser and unitary symmetries. Journal of Mathematical Physics, 26(12):3200– 3203, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v26/i12/p3200_s1.

Abraham-Shrauner:1985:ELT

[488] B. Abraham-Shrauner. Erratum: “Lie transformation group solutions of the nonlinear one-dimensional Vlasov equation”, [J. Math. Phys. 26, 1428 (1985)]. Journal of Mathematical Physics, 26(12):3204, December 1985. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v26/i12/p3204_s1. See [239].

Ablamowicz:1986:SSG

[489] Rafal Ablamowicz. Structure of spin groups associated with degenerate Clifford algebras. Journal of Mathematical Physics, 27(1):1–6, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p1_s1.

DeBievre:1986:HPM

[490] Stephan De Bièvre. Hamiltonian particle mechanics on curved space– times—a no-interaction theorem. Journal of Mathematical Physics, 27(1):7–13, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p7_s1.

DelOlmo:1986:SSS

[491] M. A. Del Olmo, M. A. Rodríguez, and P. Winternitz. Simple subgroups of simple Lie groups and nonlinear differential equations with superposition principles. Journal of Mathematical Physics, 27(1):14– 23, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p14_s1.

Kramer:1986:CUR

[492] P. Kramer, Z. Papadopolos, and W. Schweizer. A class of unitary representations of the Lie group Sp(3, R), its coherent states, and its map to a symplectic realization on sp^{∗}(3, R). Journal of Mathematical Physics, 27(1):24–28, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p24_s1.

Garcia-Bulle:1986:MGW

[493] Mauricio García-Bullé, Wolfgang Lassner, and Kurt Bernardo Wolf. The metaplectic group within the Heisenberg–Weyl ring. Journal of Mathematical Physics, 27(1):29–36, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p29_s1.

Dirl:1986:GRR

[494] R. Dirl, P. Kasperkovitz, M. I. Aroyo, J. N. Kotzev, and M. Angelova-Tjurkedjieva. Generating relations for reducing matrices. I. Ordinary representations. Journal of Mathematical Physics, 27(1):37–52, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p37_s1.

Elbaz:1986:CRC

[495] E. Elbaz. Crossing rules in Cartesian and standard coordinate systems. Journal of Mathematical Physics, 27(1):53–60, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/p53_s1.

Gourdin:1986:RBS

[496] Michel Gourdin. Relation between the supertableaux of the supergroups OSP(2|2) and SU(1|2). Journal of Mathematical Physics, 27(1):61– 65, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p61_s1.

Esperidiao:1986:OST

[497] A. S. C. Esperidião and R. F. S. Andrade. Occurrence of secular terms in the Carleman embedding. Journal of Mathematical Physics, 27(1): 66–70, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p66_s1.

Thomas:1986:SDP

[498] Lawrence E. Thomas and C. Eugene Wayne. On the stability of dense point spectrum for selfadjoint operators. Journal of Mathematical Physics, 27(1):71–75, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p71_s1.

Humi:1986:FSD

[499] Mayer Humi. Factorization of systems of differential equations. Journal of Mathematical Physics, 27(1):76–81, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p76_s1.

Duque-Carrillo:1986:SPO

[500] J. F. Duque-Carrillo, J. M. Vega-Fernández, J. J. Peña-Bernal, and M. A. Rossell-Bueno. Some practical observations on the predictor jump method for solving the Laplace equation. Journal of Mathematical Physics, 27(1):82–87, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p82_s1.

Molzahn:1986:TGS

[501] F. H. Molzahn and T. A. Osborn. Tree graphs and the solution to the Hamilton–Jacobi equation. Journal of Mathematical Physics, 27(1):88– 99, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p88_s1.

Zhu:1986:SIC

[502] Guo Cheng Zhu and H. H. Chen. Symmetries and integrability of the cylindrical Korteweg–de Vries equation. Journal of Mathematical Physics, 27(1):100–103, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p100_s1.

Antolin:1986:PUS

[503] J. Antolín and A. Cruz. Positivity and unimodality as stabilizers of the analytic extrapolation of a function known with errors. Journal of Mathematical Physics, 27(1):104–115, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p104_s1.

Jakobczyk:1986:SAU

[504] Lech Jakóbczyk. Some algebras of unbounded operators in an indefinite inner product vector space. Journal of Mathematical Physics, 27(1):116– 123, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p116_s1.

Meyer:1986:MSC

[505] David A. Meyer. A metric space construction for the boundary of space– time. Journal of Mathematical Physics, 27(1):124–131, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/

p124_s1. See comment [1886].

Nikolova:1986:DRI

[506] Lora Nikolova and V. A. Rizov. Dimensional reduction of invariant linear connections and tensor fields on multidimensional space–time. Journal of Mathematical Physics, 27(1):132–142, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p132_s1.

Bugajska:1986:GPA

[507] Krystyna Bugajska. Geometrical properties of the algebraic spinors for R^{3,1}. Journal of Mathematical Physics, 27(1):143–150, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/
p143_s1.

Espindola:1986:GEH

[508] Oslim Espindola, Nelson L. Teixeira, and Maria L. Espindola. On the generation of equivalent Hamiltonians. Journal of Mathematical Physics, 27(1):151–152, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p151_s1. See comment [832].

Leach:1986:CAP

[509] P. G. L. Leach. Comment on an aspect of a paper: “Polynomial constants of motion in flat space”, [J. Math. Phys. 25 (1984), no. 12, 3474–3478; MR 86e:70007 ] by G. Thompson. Journal of Mathematical Physics, 27(1):153–156, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p153_s1. See [11] and erratum

[823].

Salingaros:1986:LGT

[510] Nikos Salingaros. The Lorentz group and the Thomas precession. II. Exact results for the product of two boosts. Journal of Mathematical Physics, 27(1):157–162, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p157_s1. See errata [984, 1513].

Segev:1986:FES

[511] Reuven Segev. Forces and the existence of stresses in invariant continuum mechanics. Journal of Mathematical Physics, 27(1):163–170, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p163_s1.

Codona:1986:MEP

[512] Johanan L. Codona, Dennis B. Creamer, Stanley M. Flatté, R. G. Frehlich, and Frank S. Henyey. Moment-equation and path-integral techniques for wave propagation in random media. Journal of Mathematical Physics, 27(1):171–177, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p171_s1.

VilelaMendes:1986:RDE

[513] R. Vilela Mendes. Reconstruction of dynamics from an eigenstate. Journal of Mathematical Physics, 27(1):178–184, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/p178_s1.

Vrscay:1986:CFR

[514] E. R. Vrscay and J. Cizek. Continued fractions and Rayleigh–Schrödinger perturbation theory at large order. Journal of Mathematical Physics, 27(1):185–201, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p185_s1.

Aerts:1986:PEP

[515] Dirk Aerts. A possible explanation for the probabilities of quantum mechanics. Journal of Mathematical Physics, 27(1):202–210, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p202_s1.

Nakamitsu:1986:CPC

[516] Kuniaki Nakamitsu and Masayoshi Tsutsumi. The Cauchy problem for the coupled Maxwell–Schrödinger equations. Journal of Mathematical Physics, 27(1):211–216, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p211_s1.

Cheng:1986:PTD

[517] Bin Kang Cheng. The propagator of the time-dependent forced harmonic oscillator with time-dependent damping. Journal of Mathematical Physics, 27(1):217–220, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p217_s1.

Papadopoulos:1986:FIS

[518] G. J. Papadopoulos. Functional integrals for spin-Bose systems. Journal of Mathematical Physics, 27(1):221–228, January 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i1/p221_s1. See erra-

tum [698].

Roychoudhury:1986:AAE

[519] R. K. Roychoudhury and Barnana Roy. An algebraic approach to the exponential cosine screened Coulomb potential. Journal of Mathematical Physics, 27(1):229–231, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p229_s1.

Matsumoto:1986:EPS

[520] S. Matsumoto, K. Kakazu, and T. Nagamine. Eigenvalue problem for Schrödinger’s equation with repulsive potential. Journal of Mathematical Physics, 27(1):232–237, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p232_s1.

Meyer:1986:QLL

[521] R. E. Meyer. Quasiresonance of long life. Journal of Mathematical Physics, 27(1):238–248, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p238_s1.

Gesztesy:1986:CPS

[522] F. Gesztesy, G. Karner, and L. Streit. Charged particles with a shortrange force: perturbation theory with respect to the range and to additional effects. Journal of Mathematical Physics, 27(1):249–261, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p249_s1.

Bayin:1986:FSB

[523] Selçuk Ş. Bayin and J. P. Krisch. Fluid sources for Bianchi I and III space–times. Journal of Mathematical Physics, 27(1):262–264, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p262_s1.

Martin:1986:PTP

[524] J. Martín and J. M. M. Senovilla. Petrov type D perfect-fluid solutions in generalized Kerr–Schild form. Journal of Mathematical Physics, 27(1):265–270, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i1/p265_s1. See erratum [795].

PoncedeLeon:1986:FSU

[525] J. Ponce de León. Fluid spheres of uniform density in general relativity. Journal of Mathematical Physics, 27(1):271–276, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/p271_s1.

SalazarI:1986:MDW

[526] Humberto Salazar I. Magnetized diagonal Weyl metrics. Journal of Mathematical Physics, 27(1):277–280, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p277_s1.

Hojman:1986:SCQ

[527] Sergio Hojman, Luis Nuñez, Alberto Patiño, and Hector Rago. Symmetries and conserved quantities in geodesic motion. Journal of Mathematical Physics, 27(1):281–286, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p281_s1.

Hartle:1986:SMI

[528] James B. Hartle. Simplicial minisuperspace. II. Some classical solutions on simple triangulations. Journal of Mathematical Physics, 27(1):287– 295, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p287_s1.

Brewin:1986:EDR

[529] Leo Brewin. Exterior differentiation in the Regge calculus. Journal of Mathematical Physics, 27(1):296–301, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p296_s1.

Barker:1986:SRC

[530] William A. Barker. Series representations for calculations in quantum statistics. Journal of Mathematical Physics, 27(1):302–304, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p302_s1.

Beretta:1986:TLS

[531] Gian Paolo Beretta. A theorem on Lyapunov stability for dynamical systems and a conjecture on a property of entropy. Journal of Mathematical Physics, 27(1):305–308, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p305_s1.

Saxena:1986:DPB

[532] G. M. Saxena and C. L. Mehta. Dynamics of para-Bose coherent states. Journal of Mathematical Physics, 27(1):309–310, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/p309_s1.

Yngvason:1986:EII

[533] Jakob Yngvason. Euclidean invariant integral representations for Schwinger functionals. Journal of Mathematical Physics, 27(1):311– 320, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p311_s1.

Bishop:1986:AGGa

[534] N. T. Bishop. The action of a general gauge field theory: minimum or stationary. Journal of Mathematical Physics, 27(1):321–324, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p321_s1.

OSe:1986:SIL

[535] D. Ó’Sé, T. N. Sherry, and D. H. Tchrakian. Surface integrals in lower dimensions from higher-order Chern classes and a class of solutions in three dimensions. Journal of Mathematical Physics, 27(1):325–339, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p325_s1.

Chakrabarti:1986:CHM

[536] A. Chakrabarti. Construction of hyperbolic monopoles. Journal of Mathematical Physics, 27(1):340–348, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p340_s1.

Fuchs:1986:PSC

[537] Jürgen Fuchs. Physical state conditions and supersymmetry breaking in quantum mechanics. Journal of Mathematical Physics, 27(1):349– 353, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p349_s1.

Truax:1986:TOP

[538] D. Rodney Truax, V. Alan Kostelecký, and Michael Martin Nieto. On the theory of one-parameter subgroups of supergroups. Journal of Mathematical Physics, 27(1):354–364, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i1/p354_s1.

Kim:1986:NMC

[539] Shoon K. Kim. A new method of constructing the symmetry-adapted linear combinations based on the correspondence theorem and induced representations. Journal of Mathematical Physics, 27(1):365–369, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p365_s1.

Hillion:1986:PEP

[540] P. Hillion and S. Quinnez. Propagation of electromagnetic plane waves in random media. Journal of Mathematical Physics, 27(1):370–376, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i1/p370_s1.

DeSanto:1986:RBC

[541] John A. DeSanto. Relation between the connected diagram and smoothing methods for rough surface scattering. Journal of Mathematical Physics, 27(1):377–379, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i1/p377_s1.

Spencer:1986:HAN

[542] Richard G. Spencer. A Hamiltonian approach to normal mode coupling in a Coulomb plasma. Journal of Mathematical Physics, 27(1):380– 384, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p380_s1.

Hoegh-Krohn:1986:FSP

[543] Raphael Høegh-Krohn, Helge Holden, Steinar Johannesen, and Tore Wentzel-Larsen. The Fermi surface for point interactions. Journal of Mathematical Physics, 27(1):385–405, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p385_s1.

Coley:1986:TFC

[544] A. A. Coley and B. O. J. Tupper. Two-fluid cosmological models. Journal of Mathematical Physics, 27(1):406–416, January 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i1/p406_s1.

Reboucas:1986:IRB

[545] M. J. Rebouças and J. B. S. d’Olival. Imparting rotation to a Bianchi type II space–time. Journal of Mathematical Physics, 27(1):417–418, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/

jmapaq/v27/i1/p417_s1. See comments [870] and erratum [796].

Cronstrom:1986:ETB

[546] C. Cronström and J. Mickelsson. Erratum: “On topological boundary characteristics in nonabelian gauge theory”, [J. Math. Phys. 24 (1983), no. 10, 2528–2531; MR 85e:81083]. Journal of Mathematical Physics, 27 (1):419, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i1/p419_s1. See [4].

Harrell:1986:EHO

[547] E. M. Harrell, II. Erratum: “Hamiltonian operators with maximal eigenvalues”, [J. Math. Phys. 25 (1984), no. 1, 48–51; MR 85e:35089]. Journal of Mathematical Physics, 27(1):419, January 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i1/p419_s2. See [6].

Ramm:1986:ECR

[548] A. G. Ramm. Erratum: “Calculating resonances (natural frequencies) and extracting them from transient fields”, [J. Math. Phys. 26, 1012 (1985)]. Journal of Mathematical Physics, 27(1):419, January 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i1/p419_

s3. See [172].

Ablamowicz:1986:DCC

[549] Rafal Ablamowicz. Deformation and contraction in Clifford algebras. Journal of Mathematical Physics, 27(2):423–427, February 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p423_
s1.

Quesne:1986:CSRa

[550] C. Quesne. Coherent states of the real symplectic group in a complex analytic parametrization. I. Unitary-operator coherent states. Journal of Mathematical Physics, 27(2):428–441, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i2/p428_s1.

Yu:1986:BSI

[551] Z. R. Yu, O. Scholten, and H. Z. Sun. The basis for the symmetric irreducible representations of SO(9). Journal of Mathematical Physics, 27(2):442–446, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i2/p442_s1.

Cicogna:1986:RBS

[552] Giampaolo Cicogna, Giuseppe Gaeta, and Pietro Rossi. Remarks on bifurcation with symmetry, gradient property, and reducible representations. Journal of Mathematical Physics, 27(2):447–450, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/
p447_s1.

Schmidt:1986:UTI

[553] Jeffrey R. Schmidt. A unified treatment of SU(N) inner and outer multiplicities. Journal of Mathematical Physics, 27(2):451–454, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p451_s1.

vanBaal:1986:SCT

[554] Pierre van Baal and Bert van Geemen. A simple construction of twisteating solutions. Journal of Mathematical Physics, 27(2):455–457, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p455_s1.

Cooperstock:1986:DRI

[555] F. I. Cooperstock and P. H. Lim. Differentiation of retarded integrals and the divergence theorem for retarded functions with discontinuities. Journal of Mathematical Physics, 27(2):458–460, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p458_s1.

Dammert:1986:TTS

[556] Örjan Dammert. Transmission through a system of potential barriers. II. Necessary condition for complete transparency. A maximum transmission problem. Journal of Mathematical Physics, 27(2):461–470, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p461_s1.

Sachs:1986:AIA

[557] R. L. Sachs. Another identity among squares of eigenfunctions. Journal of Mathematical Physics, 27(2):471, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i2/p471_s1.

Sen:1986:FST

[558] Diptiman Sen. Fermions in the space–time R×S^{3}. Journal of Mathematical Physics, 27(2):472–482, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i2/p472_s1.

Wang:1986:CLP

[559] Lipo Wang. On the classical limit of phase-space formulation of quantum mechanics: entropy. Journal of Mathematical Physics, 27(2):483– 487, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p483_s1.

Schwartz:1986:NIM

[560] Charles Schwartz. Numerical integration in many dimensions. III. Journal of Mathematical Physics, 27(2):488, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i2/p488_s1.

Brown:1986:PBD

[561] J. David Brown and Marc Henneaux. On the Poisson brackets of differentiable generators in classical field theory. Journal of Mathematical Physics, 27(2):489–491, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i2/p489_s1.

Nash:1986:RYM

[562] C. Nash. Remarks on the Yang–Mills equations in four dimensions: CP^{2}. Journal of Mathematical Physics, 27(2):492–494, February 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p492_
s1.

Larsson:1986:HNH

[563] Jonas Larsson. Hamiltonian and non-Hamiltonian perturbation theory for nearly periodic motion. Journal of Mathematical Physics, 27(2):495– 501, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p495_s1.

Liu:1986:SFN

[564] K. L. Liu and K. Young. Stability of forced nonlinear oscillators via Poincaré map. Journal of Mathematical Physics, 27(2):502–506, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p502_s1.

Goldstein:1986:RB

[565] Harris F. Goldstein and Carl M. Bender. Relativistic brachistochrone. Journal of Mathematical Physics, 27(2):507–511, February 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p507_
s1.

Jaen:1986:FFP

[566] X. Jaen, A. Molina, and V. Iranzo. Front form and point form formulation of predictive relativistic mechanics. Noninteraction theorems. Journal of Mathematical Physics, 27(2):512–518, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p512_s1.

Jaen:1986:RPG

[567] X. Jaen, J. Llosa, F. Marqués, and A. Montoto. Realization of Poincaré group induced by a second-order ordinary differential system. Noninteraction theorem. Journal of Mathematical Physics, 27(2):519–523, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p519_s1.

Iosifescu:1986:BRQ

[568] Mircea Iosifescu and Horia Scutaru. Boson realization from quantum constraints. Journal of Mathematical Physics, 27(2):524–528, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p524_s1.

Kammerer:1986:AMP

[569] J.-B. Kammerer. Analysis of the Moyal product in a flat space. Journal of Mathematical Physics, 27(2):529–535, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i2/p529_s1.

Nomura:1986:PCF

[570] Masao Nomura. Propagation coefficients for fixed-isospin (T, T_{z}) average and related spectroscopic sum rules. Journal of Mathematical Physics, 27(2):536–543, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i2/p536_s1.

Rashid:1986:SFE

[571] M. A. Rashid. Summation-free expression for some special Clebsch– Gordan coefficients. Journal of Mathematical Physics, 27(2):544–548, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p544_s1.

Rashid:1986:SER

[572] M. A. Rashid. Simple expressions for radial functions appearing in the expansions of Y_{l_{1}}^{m_{1}}(∇)F_{l_{2}}^{m_{2}}(r) and ∇^{2n}Y_{l_{1}}^{m_{1}}(∇)F_{l_{2}}^{m_{2}}(r). Journal of Mathematical Physics, 27(2):549–551, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i2/p549_s1.

Gomez:1986:COP

[573] F. J. Gomez, M. F. Rañada, and J. Sesma. Confining optical potentials. Journal of Mathematical Physics, 27(2):552–558, February 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i2/p552_
s1.

Tipler:1986:PDE

[574] Frank J. Tipler. Penrose diagrams for the Einstein, Eddington–Lemaître, Eddington–Lemaître–Bondi, and anti-de Sitter universes. Journal of Mathematical Physics, 27(2):559–561, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i2/p559_s1.

BretonBaez:1986:MGM

[575] Nora Bretón Baez and Alberto García Díaz. The most general magnetized Kerr–Newman metric. Journal of Mathematical Physics, 27(2):562– 563, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p562_s1. See erratum [758].

Letelier:1986:SSV

[576] Patricio S. Letelier. Soliton solutions to the vacuum Einstein equations obtained from a nondiagonal seed solution. Journal of Mathematical Physics, 27(2):564–567, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i2/p564_s1.

Patra:1986:SES

[577] Amar Chandra Patra and Dipankar Ray. Some exact solutions of Einstein–Dirac–Maxwell fields and a massive neutrino. Journal of Mathematical Physics, 27(2):568–572, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i2/p568_s1.

Witt:1986:SGS

[578] Donald M. Witt. Symmetry groups of state vectors in canonical quantum gravity. Journal of Mathematical Physics, 27(2):573–592, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p573_s1.

Lochak:1986:NVS

[579] P. Lochak and J. M. Maillard. Necessary versus sufficient conditions for exact solubility of statistical models on lattices. Journal of Mathematical Physics, 27(2):593–598, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i2/p593_s1.

McQuistan:1986:DON

[580] R. B. McQuistan and J. L. Hock. The density of occupied nearest neighbor pairs on a rectangular 2 × N lattice space. Journal of Mathematical Physics, 27(2):599–602, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i2/p599_s1.

Chen:1986:GGF

[581] M. Chen. On the generalized Gibbs formula and the thermodynamic solution of the Boltzmann equation. Journal of Mathematical Physics, 27(2):603–611, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i2/p603_s1.

Garbaczewski:1986:FEN

[582] Piotr Garbaczewski. Fermion excitations of the nonlinear Schrödinger field in the attractive case. Journal of Mathematical Physics, 27(2):612– 614, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p612_s1.

Letelier:1986:SSS

[583] Patricio S. Letelier. Soliton solutions for self-dual SU(N) gauge fields on Euclidean space. Journal of Mathematical Physics, 27(2):615–619, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p615_s1.

Legare:1986:RYM

[584] M. Légaré and J. Harnad. SO(4) reduction of the SU(N) Yang–Mills– Dirac equations. Journal of Mathematical Physics, 27(2):620–626, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i2/p620_s1.

Leibbrandt:1986:TLF

[585] George Leibbrandt and Su-Long Nyeo. Two-loop Feynman integrals in the physical light-cone gauge. Journal of Mathematical Physics, 27(2):627–632, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i2/p627_s1.

Carchidi:1986:DCW

[586] M. Carchidi and H. T. Fortune. Determination of coexistence wave functions from two-neutron-transfer data. Journal of Mathematical Physics, 27(2):633–644, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i2/p633_s1.

MacBain:1986:EUP

[587] John A. MacBain and J. Bee Bednar. Existence and uniqueness properties for the one-dimensional magnetotellurics inversion problem. Journal of Mathematical Physics, 27(2):645–649, February 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i2/p645_s1.

Ram:1986:PFM

[588] Shri Ram. Perfect fluid models of Bianchi type-VI_{0} in modified Brans– Dicke cosmology. Journal of Mathematical Physics, 27(2):650–652, February 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i2/p650_s1.

Litvin:1986:ETP

[589] D. B. Litvin, J. Fuksa, and V. Kopský. On exomorphic types of phase transitions. Journal of Mathematical Physics, 27(3):661–667, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p661_s1.

Li:1986:VBR

[590] Sai Ping Li, R. V. Moody, M. Nicolescu, and J. Patera. Verma bases for representations of classical simple Lie algebras. Journal of Mathematical Physics, 27(3):668–677, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i3/p668_s1.

Post:1986:PLA

[591] G. Post. Pure Lie algebraic approach to the modified Korteweg–de Vries equation. Journal of Mathematical Physics, 27(3):678–681, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p678_s1.

Gervois:1986:SIIa

[592] A. Gervois and H. Navelet. Some integrals involving three modified Bessel functions. I. Journal of Mathematical Physics, 27(3):682–687, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p682_s1.

Gervois:1986:SIIb

[593] A. Gervois and H. Navelet. Some integrals involving three modified Bessel functions. II. Journal of Mathematical Physics, 27(3):688–695, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p688_s1.

Bassetti:1986:LPU

[594] B. Bassetti, E. Montaldi, and M. Raciti. On the linearization problem for ultraspherical polynomials. Journal of Mathematical Physics, 27(3): 696–698, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i3/p696_s1.

Chalbaud:1986:FAV

[595] Eduardo Chalbaud and Pablo Martín. Fractional approximation to the vacuum-vacuum amplitude of a ϕ^{4}-potential theory in zero dimensions. Journal of Mathematical Physics, 27(3):699–702, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p699_s1.

Leaute:1986:PTS

[596] B. Léauté and G. Marcilhacy. On a particular transcendent solution of the Ernst system generalized on n fields. Journal of Mathematical Physics, 27(3):703–706, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i3/p703_s1.

Chowdhury:1986:CIL

[597] A. Roy Chowdhury and Pranab K. Chanda. To the complete integrability of long-wave-short-wave interaction equations. Journal of Mathematical Physics, 27(3):707–709, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i3/p707_s1.

Rogers:1986:RBI

[598] Alice Rogers. Realizing the Berezin integral as a superspace contour integral. Journal of Mathematical Physics, 27(3):710–717, March 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p710_
s1.

Fischer:1986:RSS

[599] Arthur E. Fischer. Resolving the singularities in the space of Riemannian geometries. Journal of Mathematical Physics, 27(3):718–738, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p718_s1.

Isenberg:1986:EDC

[600] J. Isenberg and G. Horndeski. Effective determinism in a classical field theory with spacelike characteristics. Journal of Mathematical Physics, 27(3):739–745, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i3/p739_s1.

Molelekoa:1986:SSH

[601] Moshe Molelekoa. The structure of the space of homogeneous Yang–Mills fields. Journal of Mathematical Physics, 27(3):746–754, March 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p746_
s1.

Nassar:1986:ENE

[602] Antônio B. Nassar. Ermakov and non-Ermakov systems in quantum dissipative models. Journal of Mathematical Physics, 27(3):755–758, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p755_s1.

Ozawa:1986:IGQ

[603] Masanao Ozawa. On information gain by quantum measurements of continuous observables. Journal of Mathematical Physics, 27(3):759– 763, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i3/p759_s1.

Nash:1986:ECI

[604] Patrick L. Nash. Evaluation of a class of integrals that arise in certain path integrals. Journal of Mathematical Physics, 27(3):764–766, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p764_s1.

Zhang:1986:PIT

[605] Tian Rong Zhang and Bin Kang Cheng. Path integration of the timedependent forced oscillator with a two-time quadratic action. Journal of Mathematical Physics, 27(3):767–771, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i3/p767_s1.

Dattoli:1986:TOT

[606] G. Dattoli, J. Gallardo, and A. Torre. Time-ordering techniques and solution of differential difference equation appearing in quantum optics. Journal of Mathematical Physics, 27(3):772–780, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p772_s1.

Dray:1986:UTW

[607] Tevian Dray. A unified treatment of Wigner D functions, spin-weighted spherical harmonics, and monopole harmonics. Journal of Mathematical Physics, 27(3):781–792, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i3/p781_s1.

Werner:1986:SOR

[608] R. Werner. Screen observables in relativistic and nonrelativistic quantum mechanics. Journal of Mathematical Physics, 27(3):793–803, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p793_s1.

Kristensson:1986:ODI

[609] Gerhard Kristensson. The one-dimensional inverse scattering problem for an increasing potential. Journal of Mathematical Physics, 27(3):804– 815, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i3/p804_s1.

Rawitscher:1986:SEE

[610] G. H. Rawitscher and G. Delic. Sturmian eigenvalue equations with a Bessel function basis. Journal of Mathematical Physics, 27(3):816– 823, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i3/p816_s1.

Ashtekar:1986:NHS

[611] Abhay Ashtekar. A note on helicity and self-duality. Journal of Mathematical Physics, 27(3):824–827, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i3/p824_s1.

Baumann:1986:TPF

[612] Klaus Baumann. On the two-point functions of interacting Wightman fields. Journal of Mathematical Physics, 27(3):828–831, March 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p828_
s1.

Nagamachi:1986:HR

[613] Shigeaki Nagamachi and Nobumichi Mugibayashi. Hyperfunctions and renormalization. Journal of Mathematical Physics, 27(3):832–839, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i3/p832_s1.

Faraut:1986:VAB

[614] J. Faraut and G. A. Viano. Volterra algebra and the Bethe–Salpeter equation. Journal of Mathematical Physics, 27(3):840–848, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/
p840_s1.

Case:1986:VPO

[615] K. M. Case. Variational principles and orthogonal polynomials. Journal of Mathematical Physics, 27(3):849–852, March 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i3/p849_s1.

Bugajska:1986:SST

[616] Krystyna Bugajska. Spinors and space–times. Journal of Mathematical Physics, 27(3):853–858, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i3/p853_s1.

Baake:1986:ERS

[617] M. Baake and P. Reinicke. Explicit reduction of spl(1, 2) to osp(1, 2) as a simplified model for applications of supersymmetry to nuclear physics. Journal of Mathematical Physics, 27(3):859–868, March 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i3/p859_s1.

Quesne:1986:CSRb

[618] C. Quesne. Coherent states of the real symplectic group in a complex analytic parametrization. II. Annihilation-operator coherent states. Journal of Mathematical Physics, 27(3):869–878, March 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i3/p869_s1.

Dobrev:1986:SAE

[619] V. K. Dobrev and O. Ts. Stoytchev. Structural analysis and elementary representations of SL(4, R) and GL(4, R) and their covering groups. Journal of Mathematical Physics, 27(4):883–899, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p883_s1. See

erratum [759].

Haase:1986:SGA

[620] R. W. Haase and R. Dirl. The symmetric group: algebraic formulas for some S_{f} 6-j symbols and S_{f} ⊃ S_{f_{1}} × S_{f_{2}} 3-j symbols. Journal of Mathematical Physics, 27(4):900–913, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i4/p900_s1.

Capps:1986:ERB

[621] Richard H. Capps. Extension and replacement bases for semisimple Lie algebras. Journal of Mathematical Physics, 27(4):914–923, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p914_s1.

Castanos:1986:BRI

[622] O. Castaños, P. Kramer, and M. Moshinsky. Boson realization of sp(4, R). II. The generating kernel formulation. Journal of Mathematical Physics, 27(4):924–935, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p924_s1.

Heiss:1986:FTC

[623] W. D. Heiss and G. M. Welke. The Fourier transform of confining potentials. Journal of Mathematical Physics, 27(4):936–937, April 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p936_
s1.

vanHaeringen:1986:CSG

[624] H. van Haeringen. A class of sums of Gegenbauer functions: twentyfour sums in closed form. Journal of Mathematical Physics, 27(4):938– 952, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i4/p938_s1.

Thomas:1986:PWD

[625] Brian K. Thomas and Victor Franco. Partial wave decomposition of the Glauber amplitude for the elastic scattering of structureless charged particles by atomic hydrogen. Journal of Mathematical Physics, 27(4): 953–965, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i4/p953_s1.

King:1986:SSS

[626] R. B. King. Systematics of strongly self-dominant higher-order differential equations based on the Painlevé analysis of their singularities. Journal of Mathematical Physics, 27(4):966–971, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p966_s1.

Caviglia:1986:STI

[627] Giacomo Caviglia. Symmetry transformations, isovectors, and conservation laws. Journal of Mathematical Physics, 27(4):972–978, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p972_s1.

Fujii:1986:KSS

[628] Yasunori Fujii and Kengo Yamagishi. Killing spinors on spheres and hyperbolic manifolds. Journal of Mathematical Physics, 27(4):979–981, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p979_s1.

Horvathy:1986:PGC

[629] P. A. Horváthy and J. H. Rawnsley. The problem of “global color” in gauge theories. Journal of Mathematical Physics, 27(4):982–990, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p982_s1.

Becker:1986:SAF

[630] W. Becker and M. S. Zubairy. Stochastic averaging by functional differentiation. Journal of Mathematical Physics, 27(4):991–995, April 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p991_
s1.

Lerche:1986:SFT

[631] I. Lerche. Statistical filter theory and the O’Doherty–Anstey effect: dependence on offset. Journal of Mathematical Physics, 27(4):996–1014, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p996_s1.

Weiss:1986:TAD

[632] J. Weiss. Is there action-at-a-distance linear confinement? Journal of Mathematical Physics, 27(4):1015–1022, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i4/p1015_s1.

Weiss:1986:ADL

[633] J. Weiss. Action-at-a-distance linear potentials and conformal conservation laws. Journal of Mathematical Physics, 27(4):1023–1026, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p1023_s1.

Hall:1986:PEL

[634] Richard L. Hall. Potential envelopes and the large-N approximation. Journal of Mathematical Physics, 27(4):1027–1030, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1027_s1.

Seetharaman:1986:EED

[635] M. Seetharaman and S. S. Vasan. Energy eigenvalues of d-dimensional quartic anharmonic oscillator. Journal of Mathematical Physics, 27 (4):1031–1036, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1031_s1.

Castagnino:1986:MFC

[636] Mario A. Castagnino and Jorge B. Sztrajman. Massless fields in curved space–time: the conformal formalism. Journal of Mathematical Physics, 27(4):1037–1047, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1037_s1.

delOlmo:1986:CWR

[637] V. del Olmo and J. Olivert. Compatibility of weak rigidity with some types of elastic schemes. Journal of Mathematical Physics, 27(4):1048– 1055, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i4/p1048_s1.

Krori:1986:KOE

[638] K. D. Krori and R. Bhattacharjee. A Kerr object embedded in a gravitational field. Journal of Mathematical Physics, 27(4):1056–1058, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p1056_s1.

Magnon:1986:GPH

[639] Anne Magnon. A generalization of Penrose’s helicity theorem for space–times with nonzero dual mass. Journal of Mathematical Physics, 27(4):1059–1065, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1059_s1.

Magnon:1986:GTD

[640] Anne Magnon. Global techniques, dual mass, and causality violation. Journal of Mathematical Physics, 27(4):1066–1073, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1066_s1.

Singh:1986:BTT

[641] Tarkeshwar Singh. Birkhoff-type theorem in self-creation cosmology. Journal of Mathematical Physics, 27(4):1074–1075, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1074_s1.

VandenBergh:1986:CRF

[642] N. Van den Bergh. Conformally Ricci-flat perfect fluids. Journal of Mathematical Physics, 27(4):1076–1081, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i4/p1076_s1.

Rieckers:1986:MDF

[643] A. Rieckers and M. Ullrich. On the microscopic derivation of the finitetemperature Josephson relation in operator form. Journal of Mathematical Physics, 27(4):1082–1092, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i4/p1082_s1.

Masson:1986:EBL

[644] D. Masson. Exact Bogoliubov limits for the Bassichis–Foldy model and continued fractions. Journal of Mathematical Physics, 27(4):1093–1098, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i4/p1093_s1.

Phares:1986:TMF

[645] Alain J. Phares and Francis J. Wunderlich. Thermodynamics and molecular freedom of dimers on plane triangular lattices. Journal of Mathematical Physics, 27(4):1099–1109, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i4/p1099_s1.

Willis:1986:SGC

[646] B. L. Willis. Solution of a generalized Chandrasekhar H-equation. Journal of Mathematical Physics, 27(4):1110–1112, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1110_s1.

Volkel:1986:URS

[647] A. H. Völkel. Unbounded representations of symmetry groups in gauge quantum field theory. II. Integration. Journal of Mathematical Physics, 27(4):1113–1127, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1113_s1.

Bandelloni:1986:GAT

[648] Giuseppe Bandelloni. Gravitational anomalies in two dimensions: a cohomological approach. Journal of Mathematical Physics, 27(4):1128– 1138, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i4/p1128_s1.

Kersten:1986:CAL

[649] Paul H. M. Kersten. Creating and annihilating Lie–Bäcklund transformations of the Federbush model. Journal of Mathematical Physics, 27(4):1139–1144, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1139_s1.

Saito:1986:AFI

[650] Yoshimi Saitō. An approximation formula in the inverse scattering problem. Journal of Mathematical Physics, 27(4):1145–1153, April 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1145_
s1.

Ong:1986:CFS

[651] Michael K. Ong. A closed form solution of the s-wave Bethe–Goldstone equation with an infinite repulsive core. Journal of Mathematical Physics, 27(4):1154–1158, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i4/p1154_s1.

Mavroyannis:1986:SAS

[652] Constantine Mavroyannis. Strong amplification of sidebands in a strongly driven three-level atomic system. II. classical description of the laser field. Journal of Mathematical Physics, 27(4):1159–1163, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i4/p1159_s1.

Carrion:1986:ISE

[653] Philip M. Carrion. On inverse scattering in an elastic medium with vertical inhomogeneities. Journal of Mathematical Physics, 27(4):1164– 1168, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i4/p1164_s1.

Conte:1986:ELP

[654] Robert Conte and Jacques Raynal. Erratum: “On a labeling for point group harmonics. I and II”, [J. Math. Phys. 26, 2413, 2441 (1985)]. Journal of Mathematical Physics, 27(4):1169, April 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i4/p1169_s1. See [379, 380].

Li:1986:GMG

[655] De Ming Li, Charles P. Poole, Jr., and Horacio A. Farach. A general method of generating and classifying Clifford algebras. Journal of Mathematical Physics, 27(5):1173–1180, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i5/p1173_s1.

Bremner:1986:DEN

[656] Andrew Bremner. On Diophantine equations and nontrivial Racah coefficients. Journal of Mathematical Physics, 27(5):1181–1184, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/
p1181_s1.

Nash:1986:EEC

[657] Patrick L. Nash. On the exceptional equivalence of complex Dirac spinors and complex space–time vectors. Journal of Mathematical Physics, 27(5): 1185–1190, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1185_s1.

Jantzen:1986:ASS

[658] Robert T. Jantzen and Kjell Rosquist. Adapted slicings of space–times possessing simply transitive similarity groups. Journal of Mathematical Physics, 27(5):1191–1194, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1191_s1.

Syed:1986:MDR

[659] Ansaruddin Syed. The most degenerate representation matrix elements of finite rotations of SO(n−2, 2). Journal of Mathematical Physics, 27(5): 1195–1205, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1195_s1.

Robinson:1986:RPD

[660] P. A. Robinson. Relativistic plasma dispersion functions. Journal of Mathematical Physics, 27(5):1206–1214, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i5/p1206_s1.

Bountis:1986:ISN

[661] T. C. Bountis, V. Papageorgiou, and P. Winternitz. On the integrability of systems of nonlinear ordinary differential equations with superposition principles. Journal of Mathematical Physics, 27(5):1215–1224, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1215_s1.

David:1986:SRK

[662] D. David, N. Kamran, D. Levi, and P. Winternitz. Symmetry reduction for the Kadomtsev–Petviashvili equation using a loop algebra. Journal of Mathematical Physics, 27(5):1225–1237, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i5/p1225_s1.

Leaver:1986:SGS

[663] E. W. Leaver. Solutions to a generalized spheroidal wave equation: Teukolsky’s equations in general relativity, and the two-center problem in molecular quantum mechanics. Journal of Mathematical Physics, 27(5): 1238–1265, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1238_s1.

Radford:1986:PSL

[664] Chris Radford. Prolongation structures and Lie algebra real forms. Journal of Mathematical Physics, 27(5):1266–1270, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1266_s1.

Daubechies:1986:PNE

[665] Ingrid Daubechies, A. Grossmann, and Y. Meyer. Painless nonorthogonal expansions. Journal of Mathematical Physics, 27(5):1271–1283, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1271_s1.

Ben-Menahem:1986:TPV

[666] Shahar Ben-Menahem. The Thomas precession and velocity-space curvature. Journal of Mathematical Physics, 27(5):1284–1286, May 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1284_
s1.

Derrick:1986:SCP

[667] G. H. Derrick. Simulation of classical particle trajectories in a complex two-dimensional space. Journal of Mathematical Physics, 27(5):1287– 1292, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1287_s1.

Weiss:1986:BTP

[668] John Weiss. Bäcklund transformation and the Painlevé property. Journal of Mathematical Physics, 27(5):1293–1305, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1293_s1.

vanWyk:1986:GLT

[669] C. B. van Wyk. General Lorentz transformations and applications. Journal of Mathematical Physics, 27(5):1306–1310, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1306_s1.

vanWyk:1986:LTT

[670] C. B. van Wyk. Lorentz transformations in terms of initial and final vectors. Journal of Mathematical Physics, 27(5):1311–1314, May 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1311_
s1.

Zampino:1986:BST

[671] E. Zampino. A brief study on the transformation of Maxwell equations in Euclidean four-space. Journal of Mathematical Physics, 27(5):1315– 1318, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1315_s1.

Ashtekar:1986:GQC

[672] Abhay Ashtekar and Matthew Stillerman. Geometric quantization and constrained systems. Journal of Mathematical Physics, 27(5):1319–1330, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1319_s1.

Dhara:1986:TDI

[673] A. K. Dhara and S. V. Lawande. Time-dependent invariants and quantum mechanics in two dimensions. Journal of Mathematical Physics, 27(5):1331–1339, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i5/p1331_s1.

Fack:1986:DGA

[674] V. Fack, H. De Meyer, and G. Vanden Berghe. Dynamic-group approach to the x^{2} + λx^{2}/(1 + gx^{2}) potential. Journal of Mathematical Physics, 27(5):1340–1343, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1340_s1.

Mainland:1986:EBM

[675] G. Bruce Mainland. Electromagnetic binding of a minimally interacting, relativistic spin-0 and spin-1/2 constituent: zero four-momentum solutions. Journal of Mathematical Physics, 27(5):1344–1350, May 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1344_
s1.

Solari:1986:GCS

[676] H. G. Solari. Glauber’s coherent states and the semiclassical propagator. Journal of Mathematical Physics, 27(5):1351–1357, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1351_s1.

Kerimov:1986:LQF

[677] G. A. Kerimov. Lobachevskian quantum field theory in 2ν dimensions. Journal of Mathematical Physics, 27(5):1358–1362, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1358_s1.

Hajj-Boutros:1986:SSP

[678] Joseph Hajj-Boutros. Spherically symmetric perfect fluid solutions in isotropic coordinates. Journal of Mathematical Physics, 27(5):1363–1366, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v27/i5/p1363_s1. See erratum [1021].

Papini:1986:CVF

[679] G. Papini and M. Weiss. Charged viscous fluids: Exact solutions. Journal of Mathematical Physics, 27(5):1367–1369, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1367_s1.

Reboucas:1986:NGT

[680] M. J. Rebouças, J. E. Åman, and A. F. F. Teixeira. A note on Gödeltype space–times. Journal of Mathematical Physics, 27(5):1370–1372, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1370_s1.

Cornille:1986:OAM

[681] Henri Cornille. Oscillating and absolute Maxwellians: exact solutions for (d > 1)-dimensional Boltzmann equations. Journal of Mathematical Physics, 27(5):1373–1386, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1373_s1.

Rosen:1986:TPF

[682] Gerald Rosen. Theorems pertaining to Fokker–Planck statistical equilibrium for multidimensional stochastic systems. Journal of Mathematical Physics, 27(5):1387–1390, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1387_s1.

Kitada:1986:PCS

[683] Akihiko Kitada. On a property of a classical solution of the nonlinear mass transport equation u_{t} = u_{xx}/(1 + u_{x}^{2}). Journal of Mathematical Physics, 27(5):1391–1392, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1391_s1.

Muzinich:1986:DGL

[684] I. J. Muzinich. Differential geometry in the large and compactification of higher-dimensional gravity. Journal of Mathematical Physics, 27(5): 1393–1397, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1393_s1.

Panahi:1986:SDK

[685] M. Panahi and R. W. Tucker. Separation of Dirac and Kähler equations in spherically symmetric space–times. Journal of Mathematical Physics, 27(5):1398–1403, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1398_s1.

tenEikelder:1986:SMT

[686] H. M. M. ten Eikelder. Symmetries of the massive Thirring model. Journal of Mathematical Physics, 27(5):1404–1410, May 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i5/p1404_s1.

vanBeveren:1986:DDH

[687] E. van Beveren, T. A. Rijken, and C. Dullemond. Dielectric description of hadrons with anti-de Sitter symmetry. Journal of Mathematical Physics, 27(5):1411–1418, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1411_s1.

Kostelecky:1986:BCH

[688] V. Alan Kostelecký, Michael Martin Nieto, and Rodney Truax. Baker– Campbell–Hausdorff relations for supergroups. Journal of Mathematical Physics, 27(5):1419–1429, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i5/p1419_s1.

Baake:1986:SBS

[689] M. Baake and P. Reinicke. Symmetry breaking of u(6/2j + 1) supersymmetric models. Journal of Mathematical Physics, 27(5):1430–1434, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1430_s1.

Nakayama:1986:SEW

[690] J. Nakayama, K. Mitzutani, and M. Tsuneoka. Scattering of electromagnetic waves from a perfectly conductive slightly random surface: depolarization in backscatter. Journal of Mathematical Physics, 27(5):1435– 1448, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1435_s1.

Navarro-Saad:1986:GTT

[691] Miguel Navarro-Saad and Kurt Bernardo Wolf. The group-theoretical treatment of aberrating systems. I. Aligned lens systems in third aberration order. Journal of Mathematical Physics, 27(5):1449–1457, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1449_s1.

Wolf:1986:GTT

[692] Kurt Bernardo Wolf. The group-theoretical treatment of aberrating systems. II. Axis-symmetric inhomogeneous systems and fiber optics in third aberration order. Journal of Mathematical Physics, 27(5):1458–1465, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1458_s1.

Singh:1986:STM

[693] S. N. Singh, K. R. Choubey, and B. P. Singh. On steady transverse magnetogas flow. Journal of Mathematical Physics, 27(5):1466–1470, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v27/i5/p1466_s1. See erratum [2128].

Kim:1986:MGG

[694] Shoon K. Kim. The 32 minimal general generator sets of 230 double space groups. Journal of Mathematical Physics, 27(5):1471–1483, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1471_s1.

Kim:1986:AGG

[695] Shoon K. Kim. The 38 assemblies of the general generator sets for 1421 magnetic double space groups. Journal of Mathematical Physics, 27(5): 1484–1489, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1484_s1.

Gron:1986:TKS

[696] Ø. Grøn. Transition of a Kantowski–Sachs cosmological model into an inflationary era. Journal of Mathematical Physics, 27(5):1490–1491, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i5/p1490_s1.

Odundun:1986:EOT

[697] Olúwolé Òdúndún. Erratum: “Operators for the two-dimensional harmonic oscillator in an angular momentum basis”, [J. Math. Phys. 24 (1983), no. 9, 2340–2344; MR 84m:81072]. Journal of Mathematical Physics, 27(5):1492, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i5/p1492_s1. See [3].

Papadopoulos:1986:EFI

[698] G. J. Papadopoulos. Erratum: “Functional integrals for spin-Bose systems”, [J. Math. Phys. 27, 221 (1986)]. Journal of Mathematical Physics, 27(5):1492, May 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i5/p1492_s2. See [518].

Hirota:1986:SSD

[699] R. Hirota, B. Grammaticos, and A. Ramani. Soliton structure of the Drinfel’d–Sokolov–Wilson equation. Journal of Mathematical Physics, 27(6):1499–1505, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i6/p1499_s1.

Sachdev:1986:GBE

[700] P. L. Sachdev, K. R. C. Nair, and V. G. Tikekar. Generalized Burgers equations and Euler–Painlevé transcendents. I. Journal of Mathematical Physics, 27(6):1506–1522, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1506_s1.

Iwai:1986:FDC

[701] Toshihiro Iwai and Yoshio Uwano. The four-dimensional conformal Kepler problem reduces to the three-dimensional Kepler problem with a centrifugal potential and Dirac’s monopole field. Classical theory. Journal of Mathematical Physics, 27(6):1523–1529, June 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1523_s1.

Ullah:1986:EAR

[702] Nazakat Ullah. On an expression for the average resolvent using Grassmann integration. Journal of Mathematical Physics, 27(6):1530–1532, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i6/p1530_s1.

Berezin:1986:AMN

[703] Alexander A. Berezin. Asymptotics of the maximum number of repulsive particles on a spherical surface. Journal of Mathematical Physics, 27(6): 1533–1536, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i6/p1533_s1.

Perez:1986:LCU

[704] J. Fernando Perez, F. A. B. Coutinho, and C. P. Malta. Logarithmic corrections to the uncertainty principle and infinitude of the number of bound states of N-particle systems. Journal of Mathematical Physics, 27(6):1537–1540, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1537_s1.

Kibler:1986:HOC

[705] M. Kibler, A. Ronveaux, and T. Négadi. On the hydrogen-oscillator connection: passage formulas between wave functions. Journal of Mathematical Physics, 27(6):1541–1548, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i6/p1541_s1.

Golin:1986:CMS

[706] Simon Golin. Comment on momentum in stochastic mechanics. Journal of Mathematical Physics, 27(6):1549–1555, June 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1549_s1.

Pitowsky:1986:RQP

[707] Itamar Pitowsky. The range of quantum probability. Journal of Mathematical Physics, 27(6):1556–1565, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i6/p1556_s1.

Sundrum:1986:NAA

[708] Raman Sundrum and L. J. Tassie. Non-Abelian Aharonov–Bohm effects, Feynman paths, and topology. Journal of Mathematical Physics, 27(6): 1566–1570, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i6/p1566_s1.

Requardt:1986:RST

[709] Manfred Requardt. Reeh–Schlieder-type density results in one- and nbody Schrödinger theory and the “unique continuation problem”. Journal of Mathematical Physics, 27(6):1571–1577, June 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1571_s1.

Weber:1986:CNM

[710] Edmond Weber. Cosmologies with a noninteracting mixture of dust and radiation. Journal of Mathematical Physics, 27(6):1578–1582, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i6/p1578_s1.

TorresdelCastillo:1986:KSG

[711] G. F. Torres del Castillo. Killing spinors and gravitational perturbations. Journal of Mathematical Physics, 27(6):1583–1585, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1583_s1.

TorresdelCastillo:1986:GPA

[712] G. F. Torres del Castillo. Gravitational perturbations of algebraically special space–times via the HH equation. Journal of Mathematical Physics, 27(6):1586–1588, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1586_s1.

Kamran:1986:PPF

[713] Niky Kamran and Jean-Alain Marck. Parallel-propagated frame along the geodesics of the metrics admitting a Killing–Yano tensor. Journal of Mathematical Physics, 27(6):1589–1591, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i6/p1589_s1.

Hajj-Boutros:1986:MGE

[714] J. Hajj-Boutros. A method for generating exact Bianchi type II cosmological models. Journal of Mathematical Physics, 27(6):1592–1594, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v27/i6/p1592_s1. See comment [1118].

Finkelstein:1986:IOD

[715] R. Finkelstein and M. Villasante. Invariant operators for the ndimensional super-Poincaré algebra and the decomposition of the scalar superfield. Journal of Mathematical Physics, 27(6):1595–1606, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/
p1595_s1.

Villasante:1986:SIS

[716] M. Villasante. Superfield and irreducible superfield structure. Journal of Mathematical Physics, 27(6):1607–1611, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i6/p1607_s1.

Leaf:1986:LPQ

[717] Boris Leaf. Local properties of quantum systems. Journal of Mathematical Physics, 27(6):1612–1617, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i6/p1612_s1.

Sharma:1986:WVD

[718] N. L. Sharma and M. Howard Lee. Wave vector dependent susceptibility of a free electron gas in D dimensions and the singularity at 2k_{F}. Journal of Mathematical Physics, 27(6):1618–1623, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1618_s1.

Bowden:1986:ZDF

[719] R. L. Bowden. On the zeros of the dispersion function in particle transport theory. Journal of Mathematical Physics, 27(6):1624–1632, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i6/p1624_s1.

Willis:1986:MTE

[720] B. L. Willis and C. V. M. van der Mee. Multigroup transport equations with nondiagonal cross-section matrices. Journal of Mathematical Physics, 27(6):1633–1638, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1633_s1.

Doughty:1986:MSM

[721] Noel A. Doughty and Graham P. Collins. Multispinor symmetries for massless arbitrary spin Fierz–Pauli and Rarita–Schwinger wave equations. Journal of Mathematical Physics, 27(6):1639–1645, June 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1639_
s1.

Batrouni:1986:CIF

[722] G. G. Batrouni, H. Kawai, and Pietro Rossi. Coordinate-independent formulation of the Langevin equation. Journal of Mathematical Physics, 27(6):1646–1648, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i6/p1646_s1.

Rumpf:1986:SGP

[723] Helmut Rumpf. A simple Grassmannian path integral representation of the Dirac propagator. Journal of Mathematical Physics, 27(6):1649– 1654, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i6/p1649_s1.

Eastwood:1986:NDT

[724] Michael Eastwood and Richard Jozsa. A notion of duality for twistor functions. Journal of Mathematical Physics, 27(6):1655–1657, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/
p1655_s1.

Auberson:1986:PFQ

[725] G. Auberson, L. Epele, G. Mahoux, and F. R. A. Simão. Partition functions of the quark-gluon black body and density of states. Journal of Mathematical Physics, 27(6):1658–1666, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i6/p1658_s1.

Kristensson:1986:DISa

[726] G. Kristensson and R. J. Krueger. Direct and inverse scattering in the time domain for a dissipative wave equation. I. Scattering operators. Journal of Mathematical Physics, 27(6):1667–1682, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1667_s1.

Kristensson:1986:DISb

[727] G. Kristensson and R. J. Krueger. Direct and inverse scattering in the time domain for a dissipative wave equation. II. Simultaneous reconstruction of dissipation and phase velocity profiles. Journal of Mathematical Physics, 27(6):1683–1693, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1683_s1.

Banik:1986:RRS

[728] N. C. Banik and I. Lerche. Reflection and refraction of a spherical acoustic wave from a thin layer. Journal of Mathematical Physics, 27(6):1694– 1700, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i6/p1694_s1.

Yagle:1986:LSS

[729] Andrew E. Yagle and Bernard C. Levy. Layer-stripping solutions of multidimensional inverse scattering problems. Journal of Mathematical Physics, 27(6):1701–1710, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i6/p1701_s1.

Matsumoto:1986:FRA

[730] H. Matsumoto, H. Umezawa, and J. P. Whitehead. Feynman rules in the Anderson model: generalization to the (2J + 1)-component case. Journal of Mathematical Physics, 27(6):1711–1715, June 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i6/p1711_s1.

Kalnins:1986:SVD

[731] E. G. Kalnins and W. Miller, Jr. Separation of variables on n-dimensional Riemannian manifolds. I. the n-sphere S_{n} and Euclidean n-space R^{n}. Journal of Mathematical Physics, 27(7):1721–1736, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1721_s1.

Fokas:1986:IPM

[732] A. S. Fokas. An inverse problem for multidimensional first-order systems. Journal of Mathematical Physics, 27(7):1737–1746, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1737_s1.

Takayama:1986:CSS

[733] Ken Takayama. A class of solvable second-order ordinary differential equations with variable coefficients. Journal of Mathematical Physics, 27(7):1747–1749, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i7/p1747_s1.

Torrence:1986:WEC

[734] R. J. Torrence. Wave equations with the characteristic propagation property. Journal of Mathematical Physics, 27(7):1750–1753, July 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1750_
s1.

Espindola:1986:LHJ

[735] Maria L. Espindola, Oslim Espindola, and Nelson L. Teixeira. Linearization of the Hamilton–Jacobi equation. Journal of Mathematical Physics, 27(7):1754–1755, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i7/p1754_s1.

Katzin:1986:CFS

[736] Gerald H. Katzin and Jack Levine. Characteristic functional structure of infinitesimal symmetry mappings of classical dynamical systems. III. Systems with cyclic variables. Journal of Mathematical Physics, 27(7):

1756–1759, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i7/p1756_s1.

Spigler:1986:MPR

[737] Renato Spigler. Mean power reflection from a one-dimensional nonlinear random medium. Journal of Mathematical Physics, 27(7):1760–1771, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i7/p1760_s1.

Weston:1986:FIP

[738] V. H. Weston. Formulation of the inverse problem for the reduced wave equation in momentum space. Journal of Mathematical Physics, 27(7): 1772–1781, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i7/p1772_s1.

Gudder:1986:DQM

[739] Stanley P. Gudder. Discrete quantum mechanics. Journal of Mathematical Physics, 27(7):1782–1790, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i7/p1782_s1.

Pulmannova:1986:TPS

[740] Sylvia Pulmannová. Transition probability spaces. Journal of Mathematical Physics, 27(7):1791–1795, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i7/p1791_s1.

Waniewski:1986:MMN

[741] Jacek Waniewski. Mobility and measurements in nonlinear wave mechanics. Journal of Mathematical Physics, 27(7):1796–1799, July 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1796_
s1.

Hajicek:1986:EPN

[742] P. Hájíček. Elementary properties of a new kind of path integral. Journal of Mathematical Physics, 27(7):1800–1805, July 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1800_s1.

Ali:1986:SBG

[743] M. K. Ali, W. Robert Wood, and J. S. Devitt. On the summation of the Birkhoff–Gustavson normal form of an anharmonic oscillator. Journal of Mathematical Physics, 27(7):1806–1812, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i7/p1806_s1.

Brill:1986:JLI

[744] Dieter Brill and C. V. Vishveshwara. Joint linearization instabilities in general relativity. Journal of Mathematical Physics, 27(7):1813–1816, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i7/p1813_s1.

Linet:1986:SCS

[745] B. Linet. The static, cylindrically symmetric strings in general relativity with cosmological constant. Journal of Mathematical Physics, 27(7): 1817–1818, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i7/p1817_s1.

Bilge:1986:CIP

[746] A. H. Bilge and M. Gürses. Cartan ideal, prolongation, and Bäcklund transformations for Einstein’s equations. Journal of Mathematical Physics, 27(7):1819–1833, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i7/p1819_s1.

Khachaturyan:1986:SMA

[747] Armen Khachaturyan. Statistical mechanics approach in minimizing a multivariable function. Journal of Mathematical Physics, 27(7):1834– 1838, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i7/p1834_s1.

Mellein:1986:MFG

[748] B. Mellein. R-mer filling with general range-R cooperative effects. Journal of Mathematical Physics, 27(7):1839–1851, July 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1839_s1.

Chen:1986:PTT

[749] M. Chen. On the phase transition of the three-dimensional Percus–Yevick equation for an arbitrary potential of finite range. Journal of Mathematical Physics, 27(7):1852–1858, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i7/p1852_s1.

Renn:1986:APU

[750] J. Renn and A. Steinmann. An alternate proof of ultraviolet stability of the two-dimensional massive sine-Gordon field theory in all regions of collapse. Journal of Mathematical Physics, 27(7):1859–1866, July 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1859_
s1.

Jack:1986:TLC

[751] I. Jack. Two-loop counterterms for the Wess–Zumino model on anti-de Sitter space. Journal of Mathematical Physics, 27(7):1867–1878, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i7/p1867_s1.

Manoukian:1986:GCD

[752] Edward B. Manoukian. Generalized conditions for the decoupling theorem of quantum field theory in Minkowski space with particles of vanishingly small masses. Journal of Mathematical Physics, 27(7):1879– 1882, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i7/p1879_s1.

Furlan:1986:INS

[753] P. Furlan and R. Raczka. Intrinsic nonlinear spinor wave equations associated with nonlinear spinor representations. Journal of Mathematical Physics, 27(7):1883–1892, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i7/p1883_s1.

Kalnins:1986:MOS

[754] E. G. Kalnins, W. Miller, Jr., and G. C. Williams. Matrix operator symmetries of the Dirac equation and separation of variables. Journal of Mathematical Physics, 27(7):1893–1900, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i7/p1893_s1.

Lucke:1986:PTF

[755] W. Lücke. PCT theorem for fields with arbitrary high-energy behavior. Journal of Mathematical Physics, 27(7):1901–1905, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1901_s1.

Kalinowski:1986:SWE

[756] M. W. Kalinowski and A. Grundland. Simple waves for equation of potential, nonstationary flow of compressible gas. Journal of Mathematical Physics, 27(7):1906–1915, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i7/p1906_s1.

Laurence:1986:SRR

[757] Peter Laurence. Some rigorous results concerning spectral theory for ideal MHD. Journal of Mathematical Physics, 27(7):1916–1926, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i7/p1916_s1.

BretonBaez:1986:EMG

[758] Nora Bretón Baez and Alberto García Díaz. Erratum: “The most general magnetized Kerr–Newman metric”, [J. Math. Phys. 27, 562 (1986)]. Journal of Mathematical Physics, 27(7):1927, July 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i7/p1927_

s1. See [575].

Dobrev:1986:ESA

[759] V. K. Dobrev and O. Ts. Stoytchev. Erratum: “Structural analysis and elementary representations of SL(4, R) and GL(4, R) and their covering groups”, [J. Math. Phys. 27, 883 (1986)]. Journal of Mathematical Physics, 27(7):1927, July 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i7/p1927_s2. See [619].

Watanabe:1986:GTD

[760] Hiroshi Watanabe. Group-theoretic determination of normal coordinates for molecular vibration. Journal of Mathematical Physics, 27(8):1931– 1932, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p1931_s1.

Coleman:1986:SLP

[761] A. J. Coleman. The state labeling problem — a universal solution. Journal of Mathematical Physics, 27(8):1933–1943, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p1933_s1.

Gould:1986:MFWa

[762] M. D. Gould. Multiplicity-free Wigner coefficients for semisimple Lie groups. I. The U(n) pattern calculus. Journal of Mathematical Physics, 27(8):1944–1963, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i8/p1944_s1.

Gould:1986:MFWb

[763] M. D. Gould. Multiplicity-free Wigner coefficients for semisimple Lie groups. II. A pattern calculus for O(n). Journal of Mathematical Physics, 27(8):1964–1979, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i8/p1964_s1.

Gourdin:1986:SRO

[764] Michel Gourdin. Superunitary reduction of the orthosymplectic supertableaux. Journal of Mathematical Physics, 27(8):1980–1986, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p1980_s1.

Klein:1986:CAB

[765] Abraham Klein and Qing Ying Zhang. Closed, analytic, boson realizations for Sp(4). Journal of Mathematical Physics, 27(8):1987–1993, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p1987_s1.

Palev:1986:FDI

[766] T. D. Palev. Finite-dimensional irreducible representations of the Lie superalgebra sl(1, 3) in a Gel’fand–Zetlin basis. Journal of Mathematical Physics, 27(8):1994–2001, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i8/p1994_s1.

Sen:1986:RLG

[767] R. N. Sen. The representation of Lie groups by bundle maps. Journal of Mathematical Physics, 27(8):2002–2008, August 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2002_s1.

Kim:1986:WSS

[768] Chang-Ho Kim, Kee Yong Kim, W. S. l’Yi, Yong Duk Kim, and Young Jai Park. SU(m/n) weight systems and superprojection matrices. Journal of Mathematical Physics, 27(8):2009–2015, August 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2009_
s1.

Newell:1986:HC

[769] Alan C. Newell and Zeng Yunbo. The Hirota conditions. Journal of Mathematical Physics, 27(8):2016–2021, August 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2016_s1.

Dabrowski:1986:SSS

[770] Ludwik Dabrowski and Andrzej Trautman. Spinor structures on spheres and projective spaces. Journal of Mathematical Physics, 27(8):2022– 2028, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2022_s1.

McInnes:1986:SCR

[771] Brett McInnes. Spontaneous compactification and Ricci-flat manifolds with torsion. Journal of Mathematical Physics, 27(8):2029–2038, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p2029_s1.

Fuller:1986:TNN

[772] William Roger Fuller. Transverse nearest neighbor degeneracies on a 2 × N lattice. Journal of Mathematical Physics, 27(8):2039–2042, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p2039_s1.

Hsieh:1986:SSS

[773] S. H. Hsieh, Steven L. Kent, and E. T. Newman. Special solutions of the Sparling equation. Journal of Mathematical Physics, 27(8):2043– 2046, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2043_s1.

Okolowski:1986:NCS

[774] J. A. Okolowski. Nonlinear classical scalar field theory in curved space– time. Journal of Mathematical Physics, 27(8):2047–2050, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/
p2047_s1.

Gotay:1986:CRQ

[775] Mark J. Gotay. Constraints, reduction, and quantization. Journal of Mathematical Physics, 27(8):2051–2066, August 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2051_s1.

Kamupingene:1986:WQS

[776] A. H. Kamupingene, T. D. Palev, and S. P. Tsaneva. Wigner quantum systems. Two particles interacting via a harmonic potential. I. Two-dimensional space. Journal of Mathematical Physics, 27(8):2067– 2075, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2067_s1.

Ohya:1986:CEM

[777] Masanori Ohya and Takashi Matsuoka. Continuity of entropy and mutual entropy in C^{∗}-dynamical systems. Journal of Mathematical Physics, 27(8):2076–2079, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i8/p2076_s1.

Talukdar:1986:GFM

[778] B. Talukdar, U. Laha, and T. Sasakawa. Green’s function for motion in Coulomb-modified separable nonlocal potentials. Journal of Mathematical Physics, 27(8):2080–2086, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i8/p2080_s1.

Esculpi:1986:CSR

[779] M. Esculpi and L. Herrera. Conformally symmetric radiating spheres in general relativity. Journal of Mathematical Physics, 27(8):2087– 2096, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2087_s1.

Herrera:1986:ASG

[780] L. Herrera and J. Jimenez. Axially symmetric gravitational radiation from isolated sources and the deviation from spherical symmetry. Journal of Mathematical Physics, 27(8):2097–2104, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2097_s1.

Magnon:1986:DMS

[781] Anne Magnon. Dual mass, H-spaces, self-dual gauge connections, and nonlinear gravitons with topological origin. Journal of Mathematical Physics, 27(8):2105–2110, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i8/p2105_s1.

OMurchadha:1986:TEM

[782] Niall Ó Murchadha. Total energy momentum in general relativity. Journal of Mathematical Physics, 27(8):2111–2128, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2111_s1.

Xanthopoulos:1986:IVP

[783] Basilis C. Xanthopoulos. The initial value problem for colliding gravitational and hydrodynamic waves. Journal of Mathematical Physics, 27(8):2129–2136, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i8/p2129_s1.

Kamimura:1986:TDG

[784] Kiyoshi Kamimura. Three-dimensional gauge theory in Dirac formalism. Journal of Mathematical Physics, 27(8):2137–2139, August 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2137_
s1.

Kersten:1986:IHI

[785] P. H. M. Kersten and H. M. M. ten Eikelder. Infinite hierarchies of t-independent and t-dependent conserved functionals of the Federbush model. Journal of Mathematical Physics, 27(8):2140–2145, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/
p2140_s1.

Kim:1986:TSL

[786] Jae Kwan Kim, Jeong Yong Kim, and I. G. Koh. “twist” for simply laced groups. Journal of Mathematical Physics, 27(8):2146–2150, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p2146_s1.

tenEikelder:1986:NHS

[787] H. M. M. ten Eikelder and L. J. F. Broer. Non-Hamiltonian symmetries of a Boussinesq equation. Journal of Mathematical Physics, 27(8):2151– 2153, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2151_s1.

Heidenreich:1986:SSM

[788] W. Heidenreich. On solution spaces of massless field equations with arbitrary spin. Journal of Mathematical Physics, 27(8):2154–2159, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p2154_s1.

Nash:1986:GHM

[789] C. Nash. Geometry of hyperbolic monopoles. Journal of Mathematical Physics, 27(8):2160–2164, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v27/i8/p2160_s1.

Piecuch:1986:HOI

[790] Piotr Piecuch. Higher-order interaction energies for a system of Narbitrary molecules in the light of spherical tensor theory. Journal of Mathematical Physics, 27(8):2165–2187, August 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2165_s1.

Paris:1986:ESI

[791] R. B. Paris, N. Auby, and R. Y. Dagazian. The eigenvalues of the simplified ideal MHD ballooning equation. Journal of Mathematical Physics, 27(8):2188–2202, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i8/p2188_s1.

Haubold:1986:RTR

[792] H. J. Haubold and A. M. Mathai. The resonant thermonuclear reaction rate. Journal of Mathematical Physics, 27(8):2203–2207, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i8/p2203_s1.

BenDaniel:1986:ERT

[793] D. J. BenDaniel. Erratum: “The relation of a theory of countable sets to the field equations of physics”, [J. Math. Phys. 26 (1985), no. 4, 585–587; MR 86j:03007]. Journal of Mathematical Physics, 27(8): 2208, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i8/p2208_s1. See [104].

Hostler:1986:ESF

[794] Levere C. Hostler. Erratum: “Scalar formalism for quantum electrodynamics”, [J. Math. Phys. 26 (1985), no. 6, 1348–1354; MR 86j:81190]. Journal of Mathematical Physics, 27(8):2208, August 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/

p2208_s2. See [227].

Martin:1986:EPT

[795] J. Martín and J. M. M. Senovilla. Erratum: “Petrov type D perfectfluid solutions in generalized Kerr–Schild form”, [J. Math. Phys. 27, 265 (1986)]. Journal of Mathematical Physics, 27(8):2209, August 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i8/p2209_

s1. See [524].

Reboucas:1986:EIR

[796] M. J. Rebouças and J. B. S. d’Olival. Erratum: “Imparting rotation to a Bianchi type II space–time”, [J. Math. Phys. 27, 417 (1986)]. Journal of Mathematical Physics, 27(8):2209, August 1986. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i8/p2209_s2. See [545].

Beckers:1986:NES

[797] J. Beckers, V. Hussin, and P. Winternitz. Nonlinear equations with superposition formulas and the exceptional group G_{2}. I. complex and real forms of g_{2} and their maximal subalgebras. Journal of Mathematical Physics, 27(9):2217–2227, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2217_s1.

Han:1986:EPW

[798] D. Han, Y. S. Kim, and D. Son. Eulerian parametrization of Wigner’s little groups and gauge transformations in terms of rotations in twocomponent spinors. Journal of Mathematical Physics, 27(9):2228–2235, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2228_s1.

Aroyo:1986:GRR

[799] M. I. Aroyo, J. N. Kotzev, M. N. Angelova-Tjurkedjieva, R. Dirl, and P. Kasperkovitz. Generating relations for reducing matrices. II. Corepresentations. Journal of Mathematical Physics, 27(9):2236–2246, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i9/p2236_s1.

Kobayashi:1986:AGS

[800] Yuji Kobayashi and Shigeaki Nagamachi. Analysis on generalized superspaces. Journal of Mathematical Physics, 27(9):2247–2256, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i9/p2247_s1.

Tamizhmani:1986:LPP

[801] K. M. Tamizhmani and M. Lakshmanan. Linearization and Painlevé property of Liouville and Cheng equations. Journal of Mathematical Physics, 27(9):2257–2258, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2257_s1.

Bonora:1986:AST

[802] L. Bonora, P. Pasti, and M. Tonin. The anomaly structure of theories with external gravity. Journal of Mathematical Physics, 27(9):2259–2270, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2259_s1.

Ziolkowski:1986:PIR

[803] Richard W. Ziolkowski. A path-integral-Riemann-space approach to the electromagnetic wedge diffraction problem. Journal of Mathematical Physics, 27(9):2271–2281, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2271_s1.

Amann:1986:GRC

[804] Anton Amann. Group-related coherent states. Journal of Mathematical Physics, 27(9):2282–2289, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2282_s1.

Mielnik:1986:EL

[805] Bogdan Mielnik. Evolution loops. Journal of Mathematical Physics, 27 (9):2290–2306, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i9/p2290_s1. See erratum [1789].

Zambrini:1986:VPS

[806] J. C. Zambrini. Variational processes and stochastic versions of mechanics. Journal of Mathematical Physics, 27(9):2307–2330, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i9/p2307_s1.

Chatterjee:1986:LSK

[807] Ashok Chatterjee. Large-N solution of the Klein–Gordon equation. Journal of Mathematical Physics, 27(9):2331–2335, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i9/p2331_s1.

Carot:1986:CTV

[808] J. Carot and Ll. Mas. Conformal transformations and viscous fluids in general relativity. Journal of Mathematical Physics, 27(9):2336–2339, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2336_s1.

Wolf:1986:CPF

[809] Thomas Wolf. A class of perfect fluid metrics with flat three-dimensional hypersurfaces. Journal of Mathematical Physics, 27(9):2340–2353, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2340_s1.

Wolf:1986:AVS

[810] Thomas Wolf. About vacuum solutions of Einstein’s field equations with flat three-dimensional hypersurfaces. Journal of Mathematical Physics, 27(9):2354–2359, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i9/p2354_s1.

Bedran:1986:EAC

[811] M. L. Bedran and B. Lesche. An example of affine collineation in the Robertson–Walker metric. Journal of Mathematical Physics, 27(9):2360– 2361, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2360_s1.

Nackoney:1986:GSI

[812] Raymond W. Nackoney. Gravitational solutions, including radiation, for a perturbed light beam. Journal of Mathematical Physics, 27(9):2362– 2372, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2362_s1.

Baumann:1986:TRP

[813] Klaus Baumann. Three remarks on Powers’ theorem about irreducible fields fulfilling CAR. Journal of Mathematical Physics, 27(9):2373–2378, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2373_s1.

Baker:1986:FDB

[814] George A. Baker, Jr. Four-dimensional boson field theory. II. Existence. Journal of Mathematical Physics, 27(9):2379–2393, September 1986. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i9/p2379_
s1.

Harnad:1986:FPS

[815] J. Harnad and M. Jacques. Formal power series solutions of supersymmetric (N = 3) Yang–Mills equations. Journal of Mathematical Physics, 27(9):2394–2401, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i9/p2394_s1.

Moreschi:1986:FPE

[816] Osvaldo M. Moreschi and George A. J. Sparling. On the formulation of the positive-energy theorem in Kaluza–Klein theories. Journal of Mathematical Physics, 27(9):2402–2408, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i9/p2402_s1.

Haidari:1986:CQY

[817] A. D. Haidari. Conformal quantum Yang–Mills. Journal of Mathematical Physics, 27(9):2409–2414, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2409_s1.

Lord:1986:GTGa

[818] Eric A. Lord and P. Goswami. Gauge theory of a group of diffeomorphisms. I. General principles. Journal of Mathematical Physics, 27 (9):2415–2422, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i9/p2415_s1.

Hostler:1986:SFN

[819] Levere C. Hostler. Scalar formalism for non-Abelian gauge theory. Journal of Mathematical Physics, 27(9):2423–2428, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i9/p2423_s1.

Sathiapalan:1986:SSS

[820] B. Sathiapalan. Spectra of supersymmetric O(N) sigma models in 0 + 1 dimension. Journal of Mathematical Physics, 27(9):2429–2433, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i9/p2429_s1.

Lee:1986:SDP

[821] Taejin Lee. The Schwinger–DeWitt proper-time method for odddimensional gauge theory. Journal of Mathematical Physics, 27(9):2434– 2436, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i9/p2434_s1.

Kupershmidt:1986:MAQa

[822] B. A. Kupershmidt. Mathematical aspects of quantum fluids. II. nonrotating ^{4}He and Clebsch representations of symplectic two-cocycles. Journal of Mathematical Physics, 27(9):2437–2444, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i9/p2437_s1.

Leach:1986:ECA

[823] P. G. L. Leach. Erratum: “Comment on an aspect of a paper by G. Thompson”, [J. Math. Phys. 27, 153 (1986)]. Journal of Mathematical Physics, 27(9):2445–2446, September 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i9/p2445_s1. See [11, 509].

Bacry:1986:CTD

[824] H. Bacry and J. Nuyts. Classification of ten-dimensional kinematical groups with space isotropy. Journal of Mathematical Physics, 27 (10):2455–2457, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i10/p2455_s1.

Steeb:1986:NES

[825] W.-H. Steeb and J. A. Louw. Nahm’s equations, singular point analysis, and integrability. Journal of Mathematical Physics, 27(10):2458– 2460, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2458_s1.

Gaffet:1986:ISS

[826] B. Gaffet. On the integration of the self-similar equations, and the meaning of the Cole–Hopf transformation. Journal of Mathematical Physics, 27(10):2461–2463, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i10/p2461_s1.

Chowdhury:1986:BTH

[827] A. Roy Chowdhury and Swapna Roy. On the Bäcklund transformation and Hamiltonian properties of superevaluation equations. Journal of Mathematical Physics, 27(10):2464–2468, October 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2464_s1.

Abbati:1986:SAG

[828] M. C. Abbati, R. Cirelli, A. Manià, and P. Michor. Smoothness of the action of the gauge transformation group on connections. Journal of Mathematical Physics, 27(10):2469–2474, October 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2469_s1.

Thomas:1986:AER

[829] Lawrence E. Thomas and Zhong Yin. Approach to equilibrium for random walks on graphs and for stochastic infinite particle processes. Journal of Mathematical Physics, 27(10):2475–2477, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2475_s1.

Helfer:1986:SGTa

[830] Adam D. Helfer. Sites and googly twistor spaces. I. Vacuum. Journal of Mathematical Physics, 27(10):2478–2483, October 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2478_s1.

Helfer:1986:SGTb

[831] Adam D. Helfer. Sites and googly twistor spaces. II. Yang–Mills fields. Journal of Mathematical Physics, 27(10):2484–2488, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i10/
p2484_s1.

Hojman:1986:CPG

[832] Sergio Hojman. Comment on a paper: “On the generation of equivalent Hamiltonians”, [J. Math. Phys. 27 (1986), no. 1, 151–152; MR 8XX:xxxxx] by O. Espindola, N. L. Teixeira and M. L. Espindola. Journal of Mathematical Physics, 27(10):2489, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i10/p2489_s1. See [508].

Moreira:1986:CTE

[833] D. T. Moreira and J.-E. Werth. Canonical transformations and the equivalence problem. Journal of Mathematical Physics, 27(10):2490– 2494, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2490_s1.

Olver:1986:DTC

[834] Peter J. Olver. Dirac’s theory of constraints in field theory and the canonical form of Hamiltonian differential operators. Journal of Mathematical Physics, 27(10):2495–2501, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i10/p2495_s1.

Narcowich:1986:DLE

[835] Francis J. Narcowich. A Dyson-like expansion for solutions to the quantum Liouville equation. Journal of Mathematical Physics, 27(10):2502– 2510, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2502_s1.

Duggal:1986:CCA

[836] K. L. Duggal and R. Sharma. Conformal collineations and anisotropic fluids in general relativity. Journal of Mathematical Physics, 27(10):2511– 2513, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2511_s1.

Maharaj:1986:CFG

[837] S. D. Maharaj and R. Maartens. Collision-free gases in static space– times. Journal of Mathematical Physics, 27(10):2514–2519, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i10/p2514_s1.

Dimock:1986:CQS

[838] J. Dimock and Bernard S. Kay. Classical and quantum scattering theory for linear scalar fields on the Schwarzschild metric. II. Journal of Mathematical Physics, 27(10):2520–2525, October 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2520_s1.

Levermore:1986:LTT

[839] C. D. Levermore, G. C. Pomraning, D. L. Sanzo, and J. Wong. Linear transport theory in a random medium. Journal of Mathematical Physics, 27(10):2526–2536, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i10/p2526_s1.

dEmilio:1986:COL

[840] Emilio d’Emilio and Mihail Mintchev. Charge operators without local commutativity. Journal of Mathematical Physics, 27(10):2537–2541, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i10/p2537_s1.

Hislop:1986:SPD

[841] Peter D. Hislop. A simple proof of duality for local algebras in free quantum field theory. Journal of Mathematical Physics, 27(10):2542– 2550, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2542_s1.

Bandelloni:1986:YMC

[842] Giuseppe Bandelloni. Yang–Mills cohomology in four dimensions. Journal of Mathematical Physics, 27(10):2551–2557, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2551_s1.

Guha-Roy:1986:TWS

[843] C. Guha-Roy, B. Bagchi, and D. K. Sinha. Traveling-wave solutions and the coupled Korteweg–de Vries equation. Journal of Mathematical Physics, 27(10):2558–2560, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i10/p2558_s1.

Stubbe:1986:ELS

[844] Joachim Stubbe. Exact localized solutions of a family of two-dimensional nonlinear spinor fields. Journal of Mathematical Physics, 27(10):2561– 2567, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2561_s1.

Bishop:1986:AGGb

[845] N. T. Bishop. The action of a general gauge field theory: minimum or stationary. II. Journal of Mathematical Physics, 27(10):2568–2569, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i10/p2568_s1.

Jaskolski:1986:GIF

[846] Z. Jaskólski. The geometrical interpretation of the Faddeev–Popov determinant in Polyakov’s theory of random surfaces. Journal of Mathematical Physics, 27(10):2570–2575, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i10/p2570_s1.

Garcilazo:1986:BSR

[847] Humberto Garcilazo. Basis states for relativistic three-body calculations of particles with spin. Journal of Mathematical Physics, 27(10):2576– 2583, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i10/p2576_s1.

Yagle:1986:DIM

[848] Andrew E. Yagle. Differential and integral methods for multidimensional inverse scattering problems. Journal of Mathematical Physics, 27(10):2584–2591, October 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i10/p2584_s1.

Sannan:1986:GME

[849] Sigurd Sannan. Gravitational mass in an expanding universe. Journal of Mathematical Physics, 27(10):2592–2597, October 1986. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v27/i10/p2592_s1.

Knop:1986:IEG

[850] Jan V. Knop, Wolfgang R. Müller, Klaus Szymanski, Nenad Trinajstić, Alexander F. Kleiner, and Milan Randić. On irreducible endospectral graphs. Journal of Mathematical Physics, 27(11):2601–2612, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2601_s1.

Bremner:1986:CDZ

[851] Andrew Bremner and Simcha Brudno. A complete determination of the zeros of weight-1 6-j coefficients. Journal of Mathematical Physics, 27 (11):2613–2615, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2613_s1.

Judd:1986:AES

[852] B. R. Judd. Algebraic expressions for some multiplicity-free 6j symbols and isoscalar factors for G_{2}. Journal of Mathematical Physics, 27 (11):2616–2622, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2616_s1.

Leo:1986:GAT

[853] R. A. Leo, L. Martina, and G. Soliani. Group analysis of the three-wave resonant system in (2 + 1) dimensions. Journal of Mathematical Physics, 27(11):2623–2628, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2623_s1.

Engdahl:1986:GGF

[854] Erik Engdahl, Erkki Brändas, Magnus Rittby, and Nils Elander. Generalized Green’s functions and spectral densities in the complex energy plane. Journal of Mathematical Physics, 27(11):2629–2639, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2629_s1.

Nirmala:1986:ABT

[855] N. Nirmala, M. J. Vedan, and B. V. Baby. Auto-Bäcklund transformation, Lax pairs, and Painlevé property of a variable coefficient Korteweg– de Vries equation. I. Journal of Mathematical Physics, 27(11):2640–2643, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2640_s1.

Nirmala:1986:VCK

[856] N. Nirmala, M. J. Vedan, and B. V. Baby. A variable coefficient Korteweg–de Vries equation: similarity analysis and exact solution. II. Journal of Mathematical Physics, 27(11):2644–2646, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i11/
p2644_s1.

Weiss:1986:PFP

[857] John Weiss. Periodic fixed points of Bäcklund transformations and the Korteweg–de Vries equation. Journal of Mathematical Physics, 27 (11):2647–2656, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2647_s1.

Stefanescu:1986:SAC

[858] I. Sabba Stefanescu. On the stable analytic continuation with a condition of uniform boundedness. Journal of Mathematical Physics, 27(11):2657– 2686, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2657_s1.

Ramm:1986:ISG

[859] A. G. Ramm. Inverse scattering for geophysical problems when the background is variable. Journal of Mathematical Physics, 27(11):2687–2689, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2687_s1.

Matsumoto:1986:NTS

[860] Shuichi Matsumoto and Kiyotaka Kakazu. A note on topology of supermanifolds. Journal of Mathematical Physics, 27(11):2690–2692, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2690_s1.

Thompson:1986:KTS

[861] G. Thompson. Killing tensors in spaces of constant curvature. Journal of Mathematical Physics, 27(11):2693–2699, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i11/p2693_s1.

Bugajska:1986:SPS

[862] Krystyna Bugajska. Some problems of spinor and algebraic spinor structures. Journal of Mathematical Physics, 27(11):2700–2705, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2700_s1.

Harriott:1986:HTN

[863] Tina A. Harriott and J. G. Williams. Hydrodynamics of a topologically nontrivial metric. Journal of Mathematical Physics, 27(11):2706– 2708, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2706_s1.

Szabo:1986:QCS

[864] L. Szabó. Quantum causal structures. Journal of Mathematical Physics, 27(11):2709–2710, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2709_s1.

Brull:1986:SSS

[865] L. Brüll and H.-J. Kapellen. Special solutions of singular nonlinear Schrödinger equations with polynomial nonlinearities. Journal of Mathematical Physics, 27(11):2711–2713, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i11/p2711_s1.

Kroger:1986:RNI

[866] Helmut Kröger, Anais Smailagic, and Ralph Girard. On renormalization in nonrelativistic interaction models with infinities. Journal of Mathematical Physics, 27(11):2714–2719, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i11/p2714_s1.

Newton:1986:LES

[867] Roger G. Newton. Low-energy scattering for medium-range potentials. Journal of Mathematical Physics, 27(11):2720–2730, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i11/
p2720_s1.

Emch:1986:NCP

[868] Gérard G. Emch and Gerhard C. Hegerfeldt. New classical properties of quantum coherent states. Journal of Mathematical Physics, 27(11):2731– 2737, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2731_s1.

Froman:1986:PIF

[869] Per Olof Fröman, Finn Karlsson, and Staffan Yngve. Phase-integral formulas for Bessel functions and their relation to already existing asymptotic formulas. Journal of Mathematical Physics, 27(11):2738–2747, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2738_s1.

Jantzen:1986:RCC

[870] Robert T. Jantzen. Rotation in cosmology. Comments on: “Imparting rotation to a Bianchi type II space–time”, [J. Math. Phys. 27 (1986), no. 1, 417–418; MR 88a:83024a] by M. J. Rebouças and J. B. S. d’Olival, and similar papers. Journal of Mathematical Physics, 27(11):2748–2750, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2748_s1. See [545].

Lee:1986:SSD

[871] S.-C. Lee and S. L. Lou. Spherically symmetric dyonic solutions for vacuum Einstein equations in 4+K dimensions. Journal of Mathematical Physics, 27(11):2751–2755, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i11/p2751_s1.

TorresdelCastillo:1986:WSE

[872] G. F. Torres del Castillo. Wavelike solutions to the Einstein equations coupled to neutrino and gauge fields. Journal of Mathematical Physics, 27(11):2756–2758, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2756_s1.

Habisohn:1986:CRG

[873] Chris X. Habisohn. Calculation of radiated gravitational energy using the second-order Einstein tensor. Journal of Mathematical Physics, 27 (11):2759–2769, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2759_s1.

Chatzidimitriou-Dreismann:1986:STN

[874] C. A. Chatzidimitriou-Dreismann. Short-time nonstationary character of correlation functions in classical equilibrium ensembles. Journal of Mathematical Physics, 27(11):2770–2773, November 1986. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i11/p2770_s1.

Glasser:1986:FET

[875] M. L. Glasser. The free energy of the three-dimensional Zamolodchikov model. Journal of Mathematical Physics, 27(11):2774–2775, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2774_s1.

Maillard:1986:AAV

[876] Jean-Marie Maillard. Automorphisms of algebraic varieties and Yang– Baxter equations. Journal of Mathematical Physics, 27(11):2776–2781, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i11/p2776_s1. See erratum [1095].

Wyss:1986:FDE

[877] Walter Wyss. The fractional diffusion equation. Journal of Mathematical Physics, 27(11):2782–2785, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i11/p2782_s1.

Janssen:1986:SPA

[878] H. Janssen and C. Dullemond. Spinor propagators in anti-de Sitter space–time. Journal of Mathematical Physics, 27(11):2786–2790, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2786_s1.

Kersten:1986:INI

[879] P. H. M. Kersten and H. M. M. ten Eikelder. An infinite number of infinite hierarchies of conserved quantities of the Federbush model. Journal of Mathematical Physics, 27(11):2791–2796, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i11/p2791_s1.

Newman:1986:GTH

[880] E. T. Newman. Gauge theories, the holonomy operator, and the Riemann–Hilbert problem. Journal of Mathematical Physics, 27(11): 2797–2802, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i11/p2797_s1.

Reifler:1986:GSA

[881] Frank Reifler and Randall Morris. A gauge symmetric approach to Fierz identities. Journal of Mathematical Physics, 27(11):2803–2806, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i11/p2803_s1.

Middya:1986:GFA

[882] T. R. Middya, A. N. Basu, and S. Sengupta. A Green’s function approach to third-order elastic constants of disordered solids. Journal of Mathematical Physics, 27(11):2807–2812, November 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i11/p2807_s1.

Blank:1986:IRL

[883] J. Blank and M. Havlíček. Irreducible ∗-representations of Lie superalgebras B(0, n) with finite-degenerated vacuum. Journal of Mathematical Physics, 27(12):2823–2831, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i12/p2823_s1.

Gourdin:1986:GAS

[884] Michel Gourdin. The generalized atypical supertableaux of the orthosymplectic groups OSP(2|2p). Journal of Mathematical Physics, 27(12):2832– 2841, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i12/p2832_s1.

VanderJeugt:1986:PFD

[885] Joris Van der Jeugt. Principal five-dimensional subalgebras of Lie superalgebras. Journal of Mathematical Physics, 27(12):2842–2847, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i12/p2842_s1.

Dorizzi:1986:AEK

[886] B. Dorizzi, B. Grammaticos, A. Ramani, and P. Winternitz. Are all the equations of the Kadomtsev–Petviashvili hierarchy integrable. Journal of Mathematical Physics, 27(12):2848–2852, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2848_s1.

Omote:1986:PSN

[887] Minoru Omote. Prolongation structures of nonlinear equations and infinite-dimensional algebras. Journal of Mathematical Physics, 27 (12):2853–2860, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2853_s1.

Pempinelli:1986:NHN

[888] F. Pempinelli and S. Potenza. On a new hierarchy of nonlinear evolution equations containing the Pohlmeyer–Lund–Regge equation. Journal of Mathematical Physics, 27(12):2861–2867, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2861_s1.

Yoshimura:1986:SRN

[889] Tetz Yoshimura. Some remarks on the nonlinear integral equation in Kirkpatrick’s theory of glass transition. Journal of Mathematical Physics, 27(12):2868–2871, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2868_s1.

Klein:1986:CCM

[890] David Klein. A class of continuum models with no phase transitions. Journal of Mathematical Physics, 27(12):2872–2875, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/
p2872_s1.

Mello:1986:CLT

[891] P. A. Mello. Central-limit theorems on groups. Journal of Mathematical Physics, 27(12):2876–2891, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i12/p2876_s1.

Gielerak:1986:DET

[892] R. Gielerak. On the DLR equation for the two-dimensional sine-Gordon model. Journal of Mathematical Physics, 27(12):2892–2902, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i12/p2892_s1.

Mead:1986:ASF

[893] Lawrence R. Mead. Approximate solution of Fredholm integral equations by the maximum-entropy method. Journal of Mathematical Physics, 27 (12):2903–2907, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2903_s1.

Atre:1986:CPI

[894] M. V. Atre and N. Mukunda. Classical particles with internal structure: general formalism and application to first-order internal spaces. Journal of Mathematical Physics, 27(12):2908–2919, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2908_s1.

Cordani:1986:KPM

[895] Bruno Cordani. Kepler problem with a magnetic monopole. Journal of Mathematical Physics, 27(12):2920–2921, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2920_s1.

Krause:1986:QKH

[896] Jorge Krause. Quantum kinematics of the harmonic oscillator. Journal of Mathematical Physics, 27(12):2922–2935, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2922_s1.

Ali:1986:GQM

[897] S. Twareque Ali and Gérard G. Emch. Geometric quantization: modular reduction theory and coherent states. Journal of Mathematical Physics, 27(12):2936–2943, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2936_s1.

Chetouani:1986:CGF

[898] L. Chetouani and T. F. Hammann. Coulomb Green’s functions, in an n-dimensional Euclidean space. Journal of Mathematical Physics, 27 (12):2944–2948, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2944_s1.

Nassar:1986:TDI

[899] Antônio B. Nassar. Time-dependent invariant associated to nonlinear Schrödinger-Langevin equations. Journal of Mathematical Physics, 27 (12):2949–2952, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2949_s1.

Batlle:1986:EBL

[900] C. Batlle, J. Gomis, J. M. Pons, and N. Roman-Roy. Equivalence between the Lagrangian and Hamiltonian formalism for constrained systems. Journal of Mathematical Physics, 27(12):2953–2962, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v27/i12/p2953_s1.

Nash:1986:CPQ

[901] Patrick L. Nash. Chebyshev polynomials and quadratic path integrals. Journal of Mathematical Physics, 27(12):2963, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2963_s1.

Labarthe:1986:PLZ

[902] J. J. Labarthe. Parametrization of the linear zeros of 6-j coefficients. Journal of Mathematical Physics, 27(12):2964–2965, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/
p2964_s1.

Morales:1986:RRT

[903] J. Morales, L. Sandoval, and A. Palma. Recurrence relations for twocenter harmonic oscillator integrals. Journal of Mathematical Physics, 27(12):2966–2972, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2966_s1.

Friedman:1986:RCC

[904] John L. Friedman and Ian Jack. 3+1 Regge calculus with conserved momentum and Hamiltonian constraints. Journal of Mathematical Physics, 27(12):2973–2986, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p2973_s1.

Maartens:1986:KDP

[905] R. Maartens, D. P. Mason, and M. Tsamparlis. Kinematic and dynamic properties of conformal Killing vectors in anisotropic fluids. Journal of Mathematical Physics, 27(12):2987–2994, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2987_s1.

Beesham:1986:PLS

[906] A. Beesham. Power law singularities in the scale covariant theory. Journal of Mathematical Physics, 27(12):2995–2997, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p2995_s1.

Broda:1986:UGP

[907] B. Broda. The unboundedness of the gravitational partition function. Journal of Mathematical Physics, 27(12):2998–3002, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/
p2998_s1.

Kleinert:1986:PIS

[908] H. Kleinert. Path integral for second-derivative Lagrangian L = (κ/2)ẍ^{2} + (m/2)ẋ^{2} + (k/2)x^{2}j(τ)x(τ). Journal of Mathematical Physics, 27(12):3003–3013, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p3003_s1.

Kostin:1986:PDB

[909] M. D. Kostin. Probability density for Bose–Einstein and Fermi–Dirac particles: Slater–Kahn functions. Journal of Mathematical Physics, 27 (12):3014–3015, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p3014_s1.

McQuistan:1986:OSV

[910] R. B. McQuistan and J. L. Hock. The occupation statistics for variously shaped particles on a rectangular 2 × N lattice space. Journal of Mathematical Physics, 27(12):3016–3021, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v27/i12/p3016_s1.

Daher:1986:VPT

[911] N. Daher and G. A. Maugin. Virtual power and thermodynamics for electromagnetic continua with interfaces. Journal of Mathematical Physics, 27(12):3022–3035, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p3022_s1.

Davies:1986:BRT

[912] A. J. Davies and G. C. Joshi. A bimodular representation of tendimensional fermions. Journal of Mathematical Physics, 27(12):3036– 3039, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i12/p3036_s1.

Moriya:1986:ASN

[913] Takeshi Moriya and Yukinori Yasui. Anomalies in supersymmetric nonlinear σ models based on E_{l}-type groups. Journal of Mathematical Physics, 27(12):3040–3050, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v27/i12/p3040_s1.

Lord:1986:GTGb

[914] Eric A. Lord. Gauge theory of a group of diffeomorphisms. II. The conformal and de Sitter groups. Journal of Mathematical Physics, 27 (12):3051–3054, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v27/i12/p3051_s1.

Barut:1986:ESR

[915] A. O. Barut and N. Ünal. An exactly soluble relativistic quantum twofermion problem. Journal of Mathematical Physics, 27(12):3055–3060, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i12/p3055_s1. See comment [1379].

Verosky:1986:FOC

[916] John M. Verosky. First-order conserved densities for gas dynamics. Journal of Mathematical Physics, 27(12):3061–3063, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v27/i12/p3061_s1.

Kupershmidt:1986:MAQb

[917] B. A. Kupershmidt. Mathematical aspects of quantum fluids. III. interior Clebsch representations and transformations of symplectic twococycles for ^{4}He. Journal of Mathematical Physics, 27(12):3064–3073, December 1986. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v27/i12/p3064_s1.

Inoue:1987:CBA

[918] Atsushi Inoue, Hideyuki Ueda, and Toshiyuki Yamauchi. Commutants and bicommutants of algebras of unbounded operators. Journal of Mathematical Physics, 28(1):1–7, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i1/p1_s1.

Kryachko:1987:CTM

[919] Eugene S. Kryachko and Toshikatsu Koga. On classical theory of moments: finite-set-of-moments approach. I. Nonnegative distribution: its even moments and Hankel transform. Journal of Mathematical Physics, 28(1):8–14, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i1/p8_s1.

Brachet:1987:ARO

[920] M. E. Brachet and H. M. Fried. Approximate representations of SU(2) ordered exponentials in the adiabatic and stochastic limits. Journal of Mathematical Physics, 28(1):15–27, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p15_s1.

Garfinkle:1987:AFS

[921] David Garfinkle. Asymptotically flat space–times have no conformal Killing fields. Journal of Mathematical Physics, 28(1):28–32, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i1/p28_s1.

Dowker:1987:QTR

[922] J. S. Dowker. Quantum theory on a regular tetrahedron. Journal of Mathematical Physics, 28(1):33–38, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p33_s1.

Delic:1987:SFM

[923] G. Delic. Spectral function methods for nonlinear diffusion equations. Journal of Mathematical Physics, 28(1):39–59, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i1/p39_s1.

Delbourgo:1987:ISC

[924] R. Delbourgo, P. Hughes, and B. G. Kenny. Islands of stability and complex universality relations. Journal of Mathematical Physics, 28(1): 60–63, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i1/p60_s1.

Derrick:1987:CMR

[925] G. H. Derrick. Classical mechanics with respect to an observer’s past light cone. Journal of Mathematical Physics, 28(1):64–70, January 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i1/p64_s1.

Gordon:1987:MET

[926] William B. Gordon. Maxwell’s equations for a transverse field. Journal of Mathematical Physics, 28(1):71–78, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p71_s1.

Maduemezia:1987:HRQ

[927] Awele Maduemezia. On hyper-relativistic quantum systems. Journal of Mathematical Physics, 28(1):79–84, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p79_s1.

Daubechies:1987:WMP

[928] Ingrid Daubechies, John R. Klauder, and Thierry Paul. Wiener measures for path integrals with affine kinematic variables. Journal of Mathematical Physics, 28(1):85–102, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i1/p85_s1.

Osborn:1987:CRP

[929] T. A. Osborn, L. Papież, and R. Corns. Constructive representations of propagators for quantum systems with electromagnetic fields. Journal of Mathematical Physics, 28(1):103–123, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p103_s1.

Brudno:1987:NZW

[930] Simcha Brudno. Nontrivial zeros of the Wigner (3-j) and Racah (6j) coefficients. II. Some nonlinear solutions. Journal of Mathematical Physics, 28(1):124–127, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i1/p124_s1.

Nomura:1987:SRS

[931] Masao Nomura. The sum rule for spectroscopic factors in the seniority scheme of identical particles. Journal of Mathematical Physics, 28(1):128–133, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i1/p128_s1.

Ahmad:1987:QSD

[932] Siraj Ahmad and A. Roy Chowdhury. The quasiperiodic solutions to the discrete nonlinear Schrödinger equation. Journal of Mathematical Physics, 28(1):134–137, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i1/p134_s1.

Krolak:1987:TPC

[933] Andrzej Królak. Towards the proof of the cosmic censorship hypothesis in cosmological space–times. Journal of Mathematical Physics, 28(1):138– 141, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i1/p138_s1.

Szabados:1987:STC

[934] L. B. Szabados. On singularity theorems and curvature growth. Journal of Mathematical Physics, 28(1):142–145, January 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i1/p142_s1.

Szczyrba:1987:HDH

[935] Victor Szczyrba. Hamiltonian dynamics of higher-order theories of gravity. Journal of Mathematical Physics, 28(1):146–158, January 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i1/p146_
s1.

Knutsen:1987:PPN

[936] Henning Knutsen. On the physical properties of a nonquadratic solution for the McVittie metric. Journal of Mathematical Physics, 28(1):159– 164, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i1/p159_s1.

Letelier:1987:ESG

[937] P. S. Letelier and S. R. Oliveira. Exact self-gravitating disks and rings: a solitonic approach. Journal of Mathematical Physics, 28(1):165–170, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i1/p165_s1.

Demianski:1987:GTC

[938] M. Demiański, Z. Golda, L. M. Sokołowski, M. Szydłowski, and P. Turkowski. The group-theoretical classification of the 11-dimensional classical homogeneous Kaluza–Klein cosmologies. Journal of Mathematical Physics, 28(1):171–173, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i1/p171_s1.

Lassner:1987:CCR

[939] G. Lassner, G. A. Lassner, and C. Trapani. Canonical commutation relations on the interval. Journal of Mathematical Physics, 28(1):174– 177, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i1/p174_s1.

Malvagi:1987:SIE

[940] Fausto Malvagi and G. C. Pomraning. A singular integral equation formulation and solution for transport in semi-infinite ducts. Journal of Mathematical Physics, 28(1):178–183, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p178_s1.

Castagnino:1987:VPC

[941] Mario A. Castagnino, Diego D. Harari, and Carmen A. Núñez. Vacuum polarization in curved backgrounds deduced from Hadamard kernels. Journal of Mathematical Physics, 28(1):184–192, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i1/p184_s1.

Banai:1987:UCQ

[942] Miklós Banai. An unconventional canonical quantization of local scalar fields over quantum space–time. Journal of Mathematical Physics, 28(1):193–214, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i1/p193_s1.

Fronsdal:1987:LSG

[943] C. Fronsdal and W. F. Heidenreich. Linear de Sitter gravity. Journal of Mathematical Physics, 28(1):215–220, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i1/p215_s1.

Anders:1987:NSC

[944] Till B. Anders and Wolfgang Jachmann. On new substitutions connecting different Feynman diagrams. Journal of Mathematical Physics, 28(1):221–233, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i1/p221_s1.

Feher:1987:SPC

[945] L. Gy. Fehér. The O(3, 1) symmetry problem of the charge-monopole interaction. Journal of Mathematical Physics, 28(1):234–239, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i1/p234_s1.

Tafel:1987:RHP

[946] J. Tafel. The Riemann–Hilbert problem for the supersymmetric constraint equations. Journal of Mathematical Physics, 28(1):240–243, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i1/p240_s1.

Berryman:1987:RBS

[947] James G. Berryman. Relationship between specific surface area and spatial correlation functions for anisotropic porous media. Journal of Mathematical Physics, 28(1):244–245, January 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i1/p244_s1.

Furlan:1987:IRC

[948] P. Furlan, V. B. Petkova, and G. M. Sotkov. Indecomposable representations of the conformal group: a nonsingular photon-Weyl graviton system. Journal of Mathematical Physics, 28(2):251–257, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i2/p251_s1.

delOlmo:1987:CGC

[949] M. A. del Olmo, J. Patera, M. A. Rodriguez, and C. Rousseau. Clebsch– Gordan coefficients for SU(5) unification models. Journal of Mathematical Physics, 28(2):258–271, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p258_s1.

Palev:1987:FDRa

[950] Tchavdar D. Palev. Finite-dimensional representations of the Lie superalgebra sl(1, 3) in a Gel’fand–Zetlin basis. II. nontypical representations. Journal of Mathematical Physics, 28(2):272–291, February 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i2/p272_
s1.

VanderJeugt:1987:RSL

[951] Joris Van der Jeugt. Regular subalgebras of Lie superalgebras and extended Dynkin diagrams. Journal of Mathematical Physics, 28(2):292– 301, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i2/p292_s1. See erratum [1232].

Bergvelt:1987:DDA

[952] M. J. Bergvelt and A. P. E. ten Kroode. Differential-difference AKNS equations and homogeneous Heisenberg algebras. Journal of Mathematical Physics, 28(2):302–306, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p302_s1.

Bluman:1987:IPW

[953] George Bluman and Sukeyuki Kumei. On invariance properties of the wave equation. Journal of Mathematical Physics, 28(2):307–318, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i2/p307_s1.

De:1987:IPP

[954] Archan K. De and A. Roy Chowdhury. On the inverse problem and prolongation structure for the modified anisotropic Heisenberg spin chain.

Journal of Mathematical Physics, 28(2):319–322, February 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i2/p319_
s1.

Fernandez:1987:SPT

[955] Francisco M. Fernández, Gustavo A. Arteca, and Eduardo A. Castro. Singular points from Taylor series. Journal of Mathematical Physics, 28(2):323–329, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p323_s1.

Visentin:1987:FAE

[956] Kley Visentin and Pablo Martín. Fractional approximation to elliptic functions. Journal of Mathematical Physics, 28(2):330–333, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i2/p330_s1.

Chakravarty:1987:QSD

[957] S. Chakravarty and E. T. Newman. Quadratures for self-dual GL(2, C) Yang–Mills fields. Journal of Mathematical Physics, 28(2):334–338, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i2/p334_s1.

Demys:1987:CHC

[958] K. Demys. Comment: “On the hyperbolic complex linear symmetry groups and their local gauge transformation actions”, [J. Math. Phys. 26 (1985), no. 3, 404–406; MR 86h:83055 ] by Z. Z. Zhong. Journal of Mathematical Physics, 28(2):339, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p339_s1. See [75] and erratum

[2138].

Arnoldus:1987:MSP

[959] Henk F. Arnoldus and Thomas F. George. Multiplicative stochastic processes involving the time derivative of a Markov process. Journal of Mathematical Physics, 28(2):340–346, February 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i2/p340_s1.

Chen:1987:IMN

[960] H. H. Chen and J. E. Lin. On the integrability of multidimensional nonlinear evolution equations. Journal of Mathematical Physics, 28(2):347– 350, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i2/p347_s1.

Gutkin:1987:CSD

[961] Eugene Gutkin. Continuity of scattering data for particles on the line with directed repulsive interactions. Journal of Mathematical Physics, 28(2):351–359, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p351_s1.

Kristensson:1987:DIS

[962] G. Kristensson and R. J. Krueger. Direct and inverse scattering in the time domain for a dissipative wave equation. III. Scattering operators in the presence of a phase velocity mismatch. Journal of Mathematical Physics, 28(2):360–370, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i2/p360_s1.

Grosse:1987:IGT

[963] H. Grosse and G. Karner. Implementability of gauge transformations and quantization of fermions in external fields. Journal of Mathematical Physics, 28(2):371–375, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i2/p371_s1.

Gudder:1987:TAS

[964] Stanley Gudder and Sylvia Pulmannová. Transition amplitude spaces. Journal of Mathematical Physics, 28(2):376–385, February 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i2/p376_
s1.

Exner:1987:QMH

[965] P. Exner and P. Šeba. Quantum motion on a half-line connected to a plane. Journal of Mathematical Physics, 28(2):386–391, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v28/i2/p386_s1. See erratum [1231].

Ojha:1987:JMM

[966] P. C. Ojha. The Jacobi-matrix method in parabolic coordinates: expansion of Coulomb functions in parabolic Sturmians. Journal of Mathematical Physics, 28(2):392–396, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p392_s1.

Bracken:1987:CAM

[967] A. J. Bracken, D. S. McAnally, and O. A. Òdúndún. Coherent angular momentum states for the two-dimensional oscillator. Journal of Mathematical Physics, 28(2):397–403, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p397_s1.

Dattoli:1987:CES

[968] G. Dattoli, J. C. Gallardo, and A. Torre. Comments on exact solutions for a one-dimensional periodic solid. Journal of Mathematical Physics, 28(2):404–405, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p404_s1.

Mazur:1987:SAK

[969] P. O. Mazur and Luca Bombelli. Stationary axisymmetric Kaluza–Klein black hole. Journal of Mathematical Physics, 28(2):406–409, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i2/p406_s1.

PoncedeLeon:1987:IFR

[970] J. Ponce de León. Imperfect fluids and repulsive gravitation. Journal of Mathematical Physics, 28(2):410–411, February 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i2/p410_s1.

Mazur:1987:GFS

[971] Paul Mazur and Robert H. Barron. The Green’s function of a slab of a diatomic Montroll–Potts crystal. Journal of Mathematical Physics, 28(2):412–424, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p412_s1.

Young:1987:PCE

[972] K. Young. Physical condition for elimination of ambiguity in conditionally convergent lattice sums. Journal of Mathematical Physics, 28(2):425–427, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p425_s1.

McMullan:1987:YMT

[973] David McMullan. Yang–Mills theory and the Batalin–Fradkin– Vilkovisky formalism. Journal of Mathematical Physics, 28(2):428–437, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i2/p428_s1.

Browning:1987:BFV

[974] A. D. Browning and David McMullan. The Batalin, Fradkin, and Vilkovisky formalism for higher-order theories. Journal of Mathematical Physics, 28(2):438–444, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p438_s1.

Kobayashi:1987:GCA

[975] K. Kobayashi. Geometrostatis and the current algebra of nonlinear sigma models on supergroup manifolds. Journal of Mathematical Physics, 28(2):445–447, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p445_s1.

Collins:1987:SAH

[976] Graham P. Collins and Noel A. Doughty. Systematics of arbitrary-helicity Lagrangian wave equations. Journal of Mathematical Physics, 28(2):448– 456, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i2/p448_s1.

Hall:1987:DSO

[977] Richard L. Hall. Dirac spinor orbits. Journal of Mathematical Physics, 28(2):457–462, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p457_s1.

Nash:1987:SSB

[978] C. Nash and G. C. Joshi. Spontaneous symmetry breaking and the Higgs mechanism for quaternion fields. Journal of Mathematical Physics, 28(2):463–467, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i2/p463_s1.

Mitra:1987:GAC

[979] P. Mitra. Geometric approach to classical improvement of lattice actions. Journal of Mathematical Physics, 28(2):468–471, February 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i2/p468_
s1.

Arai:1987:SRS

[980] Asao Arai. Some remarks on scattering theory in supersymmetric quantum systems. Journal of Mathematical Physics, 28(2):472–476, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i2/p472_s1.

Bouquiaux:1987:SFP

[981] L. Bouquiaux, P. Dauby, and V. Hussin. Superspace formulation of N = 2 pseudomechanics and superpotentials. Journal of Mathematical Physics, 28(2):477–484, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i2/p477_s1.

Biyajima:1987:CFG

[982] Minoru Biyajima, Tetsuji Kawabe, and Naomichi Suzuki. On the color factor of n-gluon decay from quarkonium. Journal of Mathematical Physics, 28(2):485–488, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i2/p485_s1.

Andres:1987:DQO

[983] J. Andres. On dissipativity of quantum optical systems. Journal of Mathematical Physics, 28(2):489–491, February 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i2/p489_s1.

Salingaros:1987:ELG

[984] N. A. Salingaros. Erratum: “The Lorentz group and the Thomas precession. II. Exact results for the product of two boosts”, [J. Math. Phys. 27 (1986), no. 1, 157–162; MR 87g:78009]. Journal of Mathematical Physics, 28(2):492–493, February 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i2/p492_s1. See [510, 1513].

Beyer:1987:GTB

[985] W. A. Beyer, J. D. Louck, and P. R. Stein. Group theoretical basis of some identities for the generalized hypergeometric series. Journal of Mathematical Physics, 28(3):497–508, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i3/p497_s1.

Bender:1987:CHP

[986] Carl M. Bender, Lawrence R. Mead, and Stephen S. Pinsky. Continuous Hahn polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 28(3):509–513, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p509_s1.

Basu:1987:GRG

[987] Debabrata Basu. The Gel’fand realization and the generating function of the Clebsch–Gordan coefficients of SL(2, R) in the hyperbolic basis. Journal of Mathematical Physics, 28(3):514–519, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p514_s1.

Beckers:1987:NES

[988] J. Beckers, V. Hussin, and P. Winternitz. Nonlinear equations with superposition formulas and the exceptional group G_{2}. II. classification of the equations. Journal of Mathematical Physics, 28(3):520–529, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i3/p520_s1.

delOlmo:1987:SFR

[989] M. A. del Olmo, M. A. Rodríguez, and P. Winternitz. Superposition formulas for rectangular matrix Riccati equations. Journal of Mathematical Physics, 28(3):530–535, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i3/p530_s1.

Kitada:1987:PCS

[990] Akihiko Kitada and Hiroyuki Umehara. On a property of a classical solution of the nonlinear mass transport equation u_{t} = u_{xx}/(1 + u_{x}^{2}). II. Journal of Mathematical Physics, 28(3):536–537, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p536_s1. See

comment [1339].

Calogero:1987:EPD

[991] F. Calogero. The evolution partial differential equation u_{t} = u_{xxx} + 3(u_{xx}u^{2} + 3u_{x}^{2}u) + 3u_{x}u^{4}. Journal of Mathematical Physics, 28(3):538– 555, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p538_s1.

ElSawi:1987:WA

[992] M. El Sawi. On the WKBJ approximation. Journal of Mathematical Physics, 28(3):556–558, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i3/p556_s1. See comment [1465].

Grundling:1987:ASD

[993] Hendrik Grundling and C. A. Hurst. Algebraic structures of degenerate systems and the indefinite metric. Journal of Mathematical Physics, 28(3):559–572, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p559_s1.

Tuynman:1987:GBK

[994] G. M. Tuynman. Generalized Bergman kernels and geometric quantization. Journal of Mathematical Physics, 28(3):573–583, March 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p573_
s1.

Vourdas:1987:LGT

[995] A. Vourdas. Loop gauge theory and group cohomology. Journal of Mathematical Physics, 28(3):584–591, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i3/p584_s1.

Cheng:1987:FNB

[996] Zheming Cheng and Robert Savit. Fractal and nonfractal behavior in Lévy flights. Journal of Mathematical Physics, 28(3):592–597, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i3/p592_s1.

Chetouani:1987:HAP

[997] L. Chetouani and T. F. Hammann. The hydrogen atom in phase space. Journal of Mathematical Physics, 28(3):598–604, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p598_s1.

Barut:1987:GMO

[998] A. O. Barut, A. Inomata, and R. Wilson. The generalized Morse oscillator in the SO(4, 2) dynamical group scheme. Journal of Mathematical Physics, 28(3):605–611, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p605_s1.

Cerkaski:1987:CCO

[999] Marcin Cerkaski. On a class of 6-j coefficients with one multiplicity index for groups SP(2N), SO(2N), and SO(2N + 1). Journal of Mathematical Physics, 28(3):612–617, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p612_s1.

Dattoli:1987:TOT

[1000] G. Dattoli and A. Torre. SU(2) and SU(1, 1) time-ordering theorems and Bloch-type equations. Journal of Mathematical Physics, 28(3):618– 621, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p618_s1.

Morchio:1987:MSL

[1001] G. Morchio and F. Strocchi. Mathematical structures for long-range dynamics and symmetry breaking. Journal of Mathematical Physics, 28(3): 622–635, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p622_s1.

CaballeroCarretero:1987:TCA

[1002] J. A. Caballero Carretero and A. Martín Sánchez. Transmission coefficients in anharmonic symmetrical potentials. Journal of Mathematical Physics, 28(3):636–642, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i3/p636_s1.

Delfino:1987:ACF

[1003] A. Delfino. Analytical continuation of the Faddeev equation for local potentials. Journal of Mathematical Physics, 28(3):643–646, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i3/p643_s1.

Maia:1987:GSTa

[1004] M. D. Maia. Geometric space–time perturbation. I. Multiparameter perturbations. Journal of Mathematical Physics, 28(3):647–650, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i3/p647_s1.

Maia:1987:GSTb

[1005] M. D. Maia. Geometric space–time perturbation. II. Gauge invariance. Journal of Mathematical Physics, 28(3):651–653, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p651_s1.

Bona:1987:GSS

[1006] C. Bona, J. Stela, and P. Palou. On the generalization of Szafron solutions of Einstein field equations. Journal of Mathematical Physics, 28(3): 654–656, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p654_s1.

Ribeiro:1987:BVF

[1007] Marcelo Byrro Ribeiro and Abhik Kumar Sanyal. Bianchi VI_{0} viscous fluid cosmology with magnetic field. Journal of Mathematical Physics, 28(3):657–660, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i3/p657_s1.

Barker:1987:CPN

[1008] B. M. Barker and R. F. O’Connell. On the completion of the post-Newtonian gravitational two-body problem with spin. Journal of Mathematical Physics, 28(3):661–667, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i3/p661_s1.

Winicour:1987:IVG

[1009] Jeffrey Winicour. An initial value gravitational quadrupole radiation theorem. Journal of Mathematical Physics, 28(3):668–676, March 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p668_
s1.

Burgers:1987:CAD

[1010] G. J. H. Burgers and H. Van Dam. On classical action at a distance theories which contain a cutoff. Journal of Mathematical Physics, 28(3): 677–684, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p677_s1.

Ardalan:1987:QMD

[1011] F. Ardalan and H. Arfaei. Quotient of manifolds by discrete groups. Journal of Mathematical Physics, 28(3):685–689, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p685_s1.

Maslanka:1987:MCL

[1012] Paweł Maślanka. The many-component limit of the Potts lattice gauge model in the external field. Journal of Mathematical Physics, 28(3):690– 696, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p690_s1.

Baumann:1987:RIQ

[1013] Klaus Baumann. On relativistic irreducible quantum fields fulfilling CCR. Journal of Mathematical Physics, 28(3):697–704, March 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p697_
s1.

Moyer:1987:APC

[1014] H. Gardner Moyer. An action principle combining electromagnetism and general relativity. Journal of Mathematical Physics, 28(3):705– 710, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p705_s1.

Bandyopadhyay:1987:SQF

[1015] P. Bandyopadhyay and K. Hajra. Stochastic quantization of a Fermi field: fermions as solitons. Journal of Mathematical Physics, 28(3):711– 716, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p711_s1.

Atkinson:1987:NCQ

[1016] D. Atkinson and M. Koopmans. Nonperturbative confinement in quantum chromodynamics. IV. Improved treatment of Schoenmaker’s equation. Journal of Mathematical Physics, 28(3):717–719, March 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p717_
s1.

Hostler:1987:SEO

[1017] Levere C. Hostler. Self-energy operator for an electron in an external Coulomb potential. Journal of Mathematical Physics, 28(3):720– 727, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p720_s1.

Goedert:1987:RFM

[1018] João Goedert and H. Ralph Lewis. Rational functions of momentum as invariants for one-dimensional, time-dependent potentials: basic theory. Journal of Mathematical Physics, 28(3):728–735, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i3/p728_s1.

Lewis:1987:RFM

[1019] H. Ralph Lewis and João Goedert. Rational functions of momentum as invariants for one-dimensional, time-dependent potentials: two- and three-resonance cases. Journal of Mathematical Physics, 28(3):736– 743, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p736_s1.

Goldin:1987:USQ

[1020] Gerald A. Goldin, Ralph Menikoff, and David H. Sharp. On unusual statistics and quantum point vortices. Journal of Mathematical Physics, 28(3):744–746, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i3/p744_s1.

Hajj-Boutros:1987:ESS

[1021] Joseph Hajj-Boutros. Erratum: “Spherically symmetric perfect fluid solutions in isotropic coordinates”, [J. Math. Phys. 27 (1986), no. 5, 1363–1366; MR 87f:83009]. Journal of Mathematical Physics, 28(3): 747, March 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i3/p747_s1. See [678].

Sheikh:1987:AMA

[1022] J. A. Sheikh, J. Sita, and C. R. Sarma. An angular momentum analysis of symmetric products of paired nucleon states. Journal of Mathematical Physics, 28(4):751–757, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p751_s1.

VanderJeugt:1987:RES

[1023] Joris Van der Jeugt. Representations of N = 2 extended supergravity and unitarity conditions in Osp(N, 4). Journal of Mathematical Physics, 28(4):758–764, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p758_s1.

Moretti:1987:ACW

[1024] Paolo Moretti. The asymptotic calculation of Wiener integrals occurring in quantum and statistical physics: a divergency problem. Journal of Mathematical Physics, 28(4):765–768, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p765_s1.

Suzuki:1987:CCE

[1025] Noboru Suzuki. Comment on “Compact expression for Löwdin’s alpha function”, [J. Math. Phys. 26 (1985), no. 5, 940–942; MR 86j:33014] by A. A. Antone. Journal of Mathematical Physics, 28(4):769–773, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i4/p769_s1. See [159].

Leaute:1987:CSD

[1026] Bernard Léauté and Gilbert Marcilhacy. A class of self-dual solutions for SU(2) gauge fields on Euclidean space. Journal of Mathematical Physics, 28(4):774–776, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p774_s1.

Ragnisco:1987:EPA

[1027] O. Ragnisco, P. M. Santini, S. Chitlaru-Briggs, and M. J. Ablowitz. An example of \overline{∂} problem arising in a finite difference context: direct and inverse problem for the discrete analog of the equation ψ_{xx} + uψ = σψ_{y}. Journal of Mathematical Physics, 28(4):777–780, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p777_s1.

Mehta:1987:EEF

[1028] M. L. Mehta. Eigenvalues and eigenvectors of the finite Fourier transform. Journal of Mathematical Physics, 28(4):781–785, April 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p781_
s1.

Bruzzo:1987:VCS

[1029] Ugo Bruzzo and Roberto Cianci. Variational calculus on supermanifolds and invariance properties of superspace field theories. Journal of Mathematical Physics, 28(4):786–791, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i4/p786_s1.

Atre:1987:CPI

[1030] M. V. Atre and N. Mukunda. Classical particles with internal structure. II. Second-order internal spaces. Journal of Mathematical Physics, 28(4): 792–806, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i4/p792_s1.

Espindola:1987:HSN

[1031] Maria L. Espindola, Nelson L. Teixeira, and Oslim Espindola. Hamiltonization for singular and nonsingular mechanics. Journal of Mathematical Physics, 28(4):807–809, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i4/p807_s1.

Matthews:1987:SME

[1032] N. F. J. Matthews. On the strength of Maxwell’s equations. Journal of Mathematical Physics, 28(4):810–814, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p810_s1.

Noriega:1987:EIP

[1033] R. J. Noriega and C. G. Schifini. The equivariant inverse problem and the Maxwell equations. Journal of Mathematical Physics, 28(4):815– 817, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i4/p815_s1.

Barchielli:1987:MLR

[1034] Alberto Barchielli. Markovian limit for a reduced operation-valued stochastic process. Journal of Mathematical Physics, 28(4):818–826, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i4/p818_s1.

Englefield:1987:ARE

[1035] M. J. Englefield. Algebra representations on eigenfunctions of the Rosen– Morse potential. Journal of Mathematical Physics, 28(4):827–832, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i4/p827_s1.

Redei:1987:PLH

[1036] Miklós Rédei. On the problem of local hidden variables in algebraic quantum mechanics. Journal of Mathematical Physics, 28(4):833–835, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i4/p833_s1.

Felderhof:1987:ATS

[1037] B. U. Felderhof and R. B. Jones. Addition theorems for spherical wave solutions of the vector Helmholtz equation. Journal of Mathematical Physics, 28(4):836–839, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p836_s1.

Guha:1987:ESS

[1038] Asok Guha and S. Mukherjee. Exact solution of the Schrödinger equation with noncentral parabolic potentials. Journal of Mathematical Physics, 28(4):840–843, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p840_s1.

Gutkin:1987:CRC

[1039] Eugene Gutkin. Commutation relations for creation-annihilation operators associated with the quantum nonlinear Schrödinger equation. Journal of Mathematical Physics, 28(4):844–849, April 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i4/p844_s1.

Falkensteiner:1987:UIG

[1040] P. Falkensteiner and H. Grosse. Unitary implementability of gauge transformations for the Dirac operator and the Schwinger term. Journal of Mathematical Physics, 28(4):850–854, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p850_s1.

delOlmo:1987:WRP

[1041] V. del Olmo and J. Olivert. Weak rigidity in the PPN formalism. Journal of Mathematical Physics, 28(4):855–858, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p855_s1.

Punsly:1987:BSK

[1042] Brian Punsly. Bundle structures in the Kerr space–time. Journal of Mathematical Physics, 28(4):859–887, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p859_s1.

Reboucas:1987:CAS

[1043] M. J. Rebouças and J. E. Åman. Computer-aided study of a class of Riemannian space–times. Journal of Mathematical Physics, 28(4):888– 892, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i4/p888_s1.

Lee:1987:STAa

[1044] S.-C. Lee. Soliton transformations for axially symmetric higherdimensional gravity. I. Generalized Neugebauer–Kramer transformations. Journal of Mathematical Physics, 28(4):893–900, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p893_s1.

Lee:1987:STAb

[1045] S.-C. Lee. Soliton transformations for axially symmetric higherdimensional gravity. II. Belinskiĭ-Zakharov N-soliton transformations. Journal of Mathematical Physics, 28(4):901–904, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p901_s1.

Xanthopoulos:1987:PFS

[1046] Basilis C. Xanthopoulos. Perfect fluids satisfying a less than extremely relativistic equation of state. Journal of Mathematical Physics, 28(4): 905–912, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i4/p905_s1.

Pesquera:1987:AOD

[1047] L. Pesquera and R. Blanco. The anharmonic oscillator driven by Gaussian noise. Comparison between direct perturbation and Markov approximation methods. Journal of Mathematical Physics, 28(4):913–920, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i4/p913_s1.

Tanikawa:1987:CPP

[1048] Kiyotaka Tanikawa and Yoshihiro Yamaguchi. Coexistence of periodic points in reversible dynamical systems on a surface. Journal of Mathematical Physics, 28(4):921–928, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i4/p921_s1.

Braden:1987:MCP

[1049] H. W. Braden. Mass corrections to periodic solitons. Journal of Mathematical Physics, 28(4):929–934, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i4/p929_s1.

Legare:1987:SRS

[1050] Martin Légaré. Symmetry reduction and simple supersymmetric models. Journal of Mathematical Physics, 28(4):935–939, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p935_s1.

TorresdelCastillo:1987:GFS

[1051] G. F. Torres del Castillo. On gauge fields with sources. Journal of Mathematical Physics, 28(4):940–944, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p940_s1.

Kramer:1987:ITL

[1052] G. Kramer and B. Lampe. Integrals for two-loop calculations in massless QCD. Journal of Mathematical Physics, 28(4):945–956, April 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p945_
s1.

VanIsacker:1987:GCC

[1053] P. Van Isacker. Generalized coupling coefficients for O(6) ⊃ O(5) and O(5) ⊃ O(3) in Bose–Fermi symmetries for odd-odd nuclei. Journal of Mathematical Physics, 28(4):957–963, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p957_s1.

Iwai:1987:GTQ

[1054] Toshihiro Iwai. A gauge theory for the quantum planar three-body problem. Journal of Mathematical Physics, 28(4):964–974, April 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p964_
s1.

Rosen:1987:AEC

[1055] Gerald Rosen. Analytical estimation of critical parameter values for bound states of screened Coulomb potentials. Journal of Mathematical Physics, 28(4):975–977, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i4/p975_s1.

Brownstein:1987:USU

[1056] K. R. Brownstein. Unique shape of uniformly polarizable dielectrics. Journal of Mathematical Physics, 28(4):978–980, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i4/p978_s1.

Strinati:1987:EPT

[1057] G. Strinati. On the excitonic-polaron theory in angular variables. Journal of Mathematical Physics, 28(4):981–985, April 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i4/p981_s1.

Cerkaski:1987:BRA

[1058] Marcin Cerkaski. Branching rules for sp(2N) algebra reduction on the chain sp(2N−2)×sp(2). Journal of Mathematical Physics, 28(5):989–990, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p989_s1.

Teofilatto:1987:EGL

[1059] Paolo Teofilatto. Enlargeable graded Lie algebras of supersymmetry. Journal of Mathematical Physics, 28(5):991–996, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p991_s1.

Sachdev:1987:GBE

[1060] P. L. Sachdev and K. R. C. Nair. Generalized Burgers equations and Euler–Painlevé transcendents. II. Journal of Mathematical Physics, 28(5):997–1004, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i5/p997_s1.

Kalnins:1987:SVT

[1061] E. G. Kalnins and Willard Miller, Jr. R-separation of variables for the time-dependent Hamilton–Jacobi and Schrödinger equations. Journal of Mathematical Physics, 28(5):1005–1015, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i5/p1005_s1.

Bender:1987:MES

[1062] Carl M. Bender, Lawrence R. Mead, and N. Papanicolaou. Maximum entropy summation of divergent perturbation series. Journal of Mathematical Physics, 28(5):1016–1018, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i5/p1016_s1.

Bokhari:1987:SSS

[1063] Ashfaque H. Bokhari and Asghar Qadir. Symmetries of static, spherically symmetric space–times. Journal of Mathematical Physics, 28 (5):1019–1022, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i5/p1019_s1. See erratum [1416] and correc-

tionBokhari:1990:KVS.

Janssen:1987:PMV

[1064] H. Janssen and C. Dullemond. Propagators for massive vector fields in anti-de Sitter space–time using Stueckelberg’s Lagrangian. Journal of Mathematical Physics, 28(5):1023–1029, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i5/p1023_s1.

Korzeniowski:1987:LTB

[1065] Andrzej Korzeniowski. A limit theorem for basic states of disordered structures. Journal of Mathematical Physics, 28(5):1030–1031, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1030_s1.

Morales:1987:HTP

[1066] J. Morales, J. Lopez-Bonilla, and A. Palma. Hypervirial theorem and parameter differentiation: closed formulation for harmonic oscillator integrals. Journal of Mathematical Physics, 28(5):1032–1035, May 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1032_
s1.

Howard:1987:LSS

[1067] J. E. Howard and R. S. MacKay. Linear stability of symplectic maps. Journal of Mathematical Physics, 28(5):1036–1051, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1036_s1.

Saito:1987:GDS

[1068] S. Saito and N. Saitoh. Gauge and dual symmetries and linearization of Hirota’s bilinear equations. Journal of Mathematical Physics, 28(5): 1052–1055, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i5/p1052_s1.

Caviglia:1987:NTC

[1069] G. Caviglia and A. Morro. Noether-type conservation laws for perfect fluid motions. Journal of Mathematical Physics, 28(5):1056–1060, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1056_s1.

Weston:1987:FWE

[1070] Vaughan H. Weston. Factorization of the wave equation in higher dimensions. Journal of Mathematical Physics, 28(5):1061–1068, May 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1061_
s1.

Arnoldus:1987:CRC

[1071] Henk F. Arnoldus, Daniel Jelski, and Thomas F. George. Confinement and redistribution of charges and currents on a surface by external fields. Journal of Mathematical Physics, 28(5):1069–1074, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1069_s1.

Coll:1987:EFT

[1072] Bartolomé Coll, Francesc Fayos, and Joan Josep Ferrando. On the electromagnetic field and the Teukolsky–Press relations in arbitrary space– times. Journal of Mathematical Physics, 28(5):1075–1079, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/
p1075_s1.

Derezinski:1987:EAM

[1073] Jan Dereziński. Existence and analyticity of many-body scattering amplitudes at low energies. Journal of Mathematical Physics, 28(5):1080– 1088, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i5/p1080_s1.

Low:1987:NCL

[1074] F. E. Low. A new class of lattice identities. Journal of Mathematical Physics, 28(5):1089–1090, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i5/p1089_s1.

Hadeler:1987:TCN

[1075] K. P. Hadeler and H. Lange. Two comments on nonlinear Schrödinger equations. Journal of Mathematical Physics, 28(5):1091–1093, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/
p1091_s1.

Verosky:1987:HSN

[1076] John Verosky. The Hamiltonian structures of the nonlinear Schrödinger equation in the classical limit. Journal of Mathematical Physics, 28(5): 1094–1096, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i5/p1094_s1.

Solari:1987:STS

[1077] Hernán G. Solari. Semiclassical treatment of spin system by means of coherent states. Journal of Mathematical Physics, 28(5):1097–1102, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1097_s1.

Chung:1987:GLT

[1078] S. G. Chung and Thomas F. George. Generalization of Levinson’s theorem to particle-matter interactions. Journal of Mathematical Physics, 28(5):1103–1106, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i5/p1103_s1.

Rozga:1987:DKS

[1079] Krzysztof Rózga. Double Kerr–Schild equivalence and hyperheavens. Journal of Mathematical Physics, 28(5):1107–1113, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1107_s1.

PoncedeLeon:1987:NAM

[1080] J. Ponce de León. New analytical models for anisotropic spheres in general relativity. Journal of Mathematical Physics, 28(5):1114–1117, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1114_s1.

Sussman:1987:SSS

[1081] Roberto A. Sussman. On spherically symmetric shear-free perfect fluid configurations (neutral and charged). I. Journal of Mathematical Physics, 28(5):1118–1136, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i5/p1118_s1.

Jack:1987:LIR

[1082] Ian Jack and Leonard Parker. Linear independence of renormalization counterterms in curved space–times of arbitrary dimensionality. Journal of Mathematical Physics, 28(5):1137–1139, May 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1137_s1.

Toscani:1987:NBE

[1083] G. Toscani and V. Protopopescu. The nonlinear Boltzmann equation with partially absorbing boundary conditions. Global existence and uniqueness results. Journal of Mathematical Physics, 28(5):1140–1145, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1140_s1.

Baker:1987:FDB

[1084] George A. Baker, Jr. and J. D. Johnson. Four-dimensional boson field theory. III. Nontriviality. Journal of Mathematical Physics, 28(5):1146– 1158, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i5/p1146_s1.

Gutkin:1987:FRP

[1085] Eugene Gutkin. A few remarks on the paper: “Necessary versus sufficient conditions for exact solubility of statistical models on lattices”, [J. Math. Phys. 27 (1986), no. 2, 593–598; MR 87h:82083 ] by P. Lochak and J.-M. Maillard. Journal of Mathematical Physics, 28(5):1159–1163, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1159_s1.

Tasaki:1987:MFL

[1086] Hal Tasaki. On the motion of the first Lee–Yang zero. Journal of Mathematical Physics, 28(5):1164–1166, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i5/p1164_s1.

Chau:1987:GRE

[1087] L.-L. Chau and H. C. Yen. Generalized Riccati equations for self-dual Yang–Mills fields. Journal of Mathematical Physics, 28(5):1167–1169, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1167_s1.

Humi:1987:FDL

[1088] Mayer Humi. Finite-dimensional Lorentz covariant bifurcations. Journal of Mathematical Physics, 28(5):1170–1174, May 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1170_s1.

Kim:1987:CGM

[1089] Y. S. Kim and E. P. Wigner. Cylindrical group and massless particles. Journal of Mathematical Physics, 28(5):1175–1179, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i5/p1175_s1.

Scheunert:1987:ISM

[1090] Manfred Scheunert. Invariant supersymmetric multilinear forms and the Casimir elements of P -type Lie superalgebras. Journal of Mathematical Physics, 28(5):1180–1191, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i5/p1180_s1.

Jarvis:1987:GQS

[1091] P. D. Jarvis, Mei Yang, and B. G. Wybourne. Generalized quasispin for supergroups. Journal of Mathematical Physics, 28(5):1192–1197, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1192_s1.

Goldstein:1987:RMT

[1092] Jerome A. Goldstein and Gisèle Ruiz Rieder. A rigorous modified Thomas–Fermi theory for atomic systems. Journal of Mathematical Physics, 28(5):1198–1202, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i5/p1198_s1.

Robinson:1987:RPD

[1093] P. A. Robinson. Relativistic plasma dispersion functions: series, integrals, and approximations. Journal of Mathematical Physics, 28(5):

1203–1205, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i5/p1203_s1.

Kondo:1987:CES

[1094] Kei ichi Kondo and Takashi Hara. Critical exponent of susceptibility for a class of general ferromagnets in d > 4 dimensions. Journal of Mathematical Physics, 28(5):1206–1208, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i5/p1206_s1.

Maillard:1987:EAA

[1095] Jean-Marie Maillard. Erratum: “Automorphisms of algebraic varieties and Yang–Baxter equations”, [J. Math. Phys. 27 (1986), no. 11, 2776– 2781; MR 87k:82146]. Journal of Mathematical Physics, 28(5):1209, May 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i5/p1209_s1. See [876].

Capps:1987:ROA

[1096] Richard H. Capps. Representations of E_{8} and other algebras. Journal of Mathematical Physics, 28(6):1215–1222, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i6/p1215_s1.

Furlan:1987:GEA

[1097] P. Furlan. Generalized Euler angles as intrinsic coordinates for nonlinear spinors. Journal of Mathematical Physics, 28(6):1223–1230, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1223_s1.

LeBlanc:1987:DEA

[1098] R. Le Blanc and D. J. Rowe. Decomposition of the SO^{∗}(8) enveloping algebra under U(4) ⊃ U(3). Journal of Mathematical Physics, 28(6): 1231–1236, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1231_s1.

Lenczewski:1987:IMP

[1099] Romuald Lenczewski. Indecomposable modules of the Poincaré algebra in an energy-cyclic angular momentum basis. Journal of Mathematical Physics, 28(6):1237–1242, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i6/p1237_s1.

Elphick:1987:CES

[1100] Christian Elphick. Classes of exactly solvable nonlinear evolution equations for Grassmann variables: the normal form method. Journal of Mathematical Physics, 28(6):1243–1249, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i6/p1243_s1.

Baumann:1987:LTS

[1101] G. Baumann and T. F. Nonnenmacher. Lie transformations, similarity reduction, and solutions for the nonlinear Madelung fluid equations with external potential. Journal of Mathematical Physics, 28(6):1250– 1260, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1250_s1.

Papachristou:1987:SAI

[1102] C. J. Papachristou and B. Kent Harrison. Some aspects of the isogroup of the self-dual Yang–Mills system. Journal of Mathematical Physics, 28(6): 1261–1264, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1261_s1.

Grosjean:1987:IDC

[1103] C. C. Grosjean. The invariant density for a class of discrete-time maps involving an arbitrary monotonic function operator and an integer parameter. Journal of Mathematical Physics, 28(6):1265–1274, June 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1265_
s1.

Fried:1987:UOE

[1104] H. M. Fried. On unitary SU(N) ordered exponentials in a strong coupling limit. Journal of Mathematical Physics, 28(6):1275–1282, June 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1275_
s1.

Vatsya:1987:EAS

[1105] S. R. Vatsya. Existence and approximation of the solutions of some nonlinear problems. Journal of Mathematical Physics, 28(6):1283–1286, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1283_s1.

Hwang:1987:EIM

[1106] J. S. Hwang. On the existence of infinitely many resonances in the scattering problem based on moment conditions and entire functions. Journal of Mathematical Physics, 28(6):1287–1289, June 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1287_s1.

Bhattacharya:1987:DTC

[1107] Ranjan Bhattacharya and Siddhartha Bhowmick. Do trilinear commutation relations in quantum mechanics admit coordinate space realization in three dimensions. Journal of Mathematical Physics, 28(6):1290– 1292, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1290_s1.

Ziolkowski:1987:ESA

[1108] Richard W. Ziolkowski and William A. Johnson. Electromagnetic scattering of an arbitrary plane wave from a spherical shell with a circular aperture. Journal of Mathematical Physics, 28(6):1293–1314, June 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1293_
s1.

Iwai:1987:GSI

[1109] Toshihiro Iwai. A geometric setting for internal motions of the quantum three-body system. Journal of Mathematical Physics, 28(6):1315– 1326, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1315_s1.

Derrick:1987:QMC

[1110] G. H. Derrick. Quantum mechanics of a charged scalar boson with respect to an observer’s past light cone. Journal of Mathematical Physics, 28(6): 1327–1340, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1327_s1.

Ramm:1987:SCZ

[1111] A. G. Ramm. Sufficient conditions for zero not to be an eigenvalue of the Schrödinger operator. Journal of Mathematical Physics, 28(6):1341– 1343, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1341_s1.

Rosenhouse:1987:KVH

[1112] Avia Rosenhouse and Jacob Katriel. The Korteweg–de Vries hierarchy of isospectral transformations: towards a general explicit expression. Journal of Mathematical Physics, 28(6):1344–1350, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1344_s1.

Goldberg:1987:SLE

[1113] I. B. Goldberg and R. H. Pratt. Schrödinger-like equation for relativistic particles. Journal of Mathematical Physics, 28(6):1351–1359, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1351_s1.

Wu:1987:GAQ

[1114] Yihren Wu. A geometric approach to quantum scattering with group symmetry. Journal of Mathematical Physics, 28(6):1360–1363, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1360_s1.

Magnon:1987:EOS

[1115] Anne M. R. Magnon. Existence and observability of spinor structure. Journal of Mathematical Physics, 28(6):1364–1369, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1364_s1.

Faridi:1987:EME

[1116] Abbas M. Faridi. Einstein–Maxwell equations and the conformal Ricci collineations. Journal of Mathematical Physics, 28(6):1370–1376, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1370_s1.

Iyer:1987:ESS

[1117] B. R. Iyer and C. V. Vishveshwara. Exact solutions for space–times with local rotational symmetry in which the Dirac equation separates. Journal of Mathematical Physics, 28(6):1377–1381, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i6/p1377_s1.

Lorenz-Petzold:1987:CTN

[1118] Dieter Lorenz-Petzold. Comment on the two “new” classes of Bianchi type II solutions: “A method for generating exact Bianchi type II cosmological models”, [J. Math. Phys. 27 (1986), no. 6, 1592–1594; MR 87i:83023 ] by J. Hajj-Boutros. Journal of Mathematical Physics, 28(6): 1382, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1382_s1. See [714].

Singh:1987:NSC

[1119] Tarkeshwar Singh. Nonexistence of static conformally flat solutions in a new scalar-tensor theory. Journal of Mathematical Physics, 28(6):1383– 1384, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i6/p1383_s1.

Barker:1987:SRCa

[1120] William A. Barker. Series representations for calculations in quantum statistics. II. Journal of Mathematical Physics, 28(6):1385–1388, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1385_s1.

Barker:1987:SRCb

[1121] William A. Barker. Series representations for calculations in quantum statistics. III. Journal of Mathematical Physics, 28(6):1389–1393, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1389_s1.

Cohen:1987:STS

[1122] J. S. Cohen and M. Winnink. Stable thermodynamic states. Journal of Mathematical Physics, 28(6):1394–1397, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i6/p1394_s1.

Kandrup:1987:NEC

[1123] Henry E. Kandrup. Nonequilibrium entropy in classical and quantum field theory. Journal of Mathematical Physics, 28(6):1398–1415, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i6/p1398_s1.

Federbush:1987:PCA

[1124] Paul Federbush and Calvin Williamson. A phase cell approach to Yang–Mills theory. II. Analysis of a mode. Journal of Mathematical Physics, 28(6):1416–1419, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i6/p1416_s1.

vanHolten:1987:QNM

[1125] J. W. van Holten. Quantum noncompact σ models. Journal of Mathematical Physics, 28(6):1420–1429, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i6/p1420_s1.

Moreau:1987:DGP

[1126] M. Moreau and B. Gaveau. Definition and general properties of a class of non-Markovian collision processes. Journal of Mathematical Physics, 28(6):1430–1439, June 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i6/p1430_s1.

Barannik:1987:SLA

[1127] L. F. Barannik and W. I. Fushchich. On subalgebras of the Lie algebra of the extended Poincaré group P̃(1, n). Journal of Mathematical Physics, 28(7):1445–1458, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i7/p1445_s1.

Gill:1987:TOO

[1128] Tepper L. Gill and W. W. Zachary. Time-ordered operators and Feynman–Dyson algebras. Journal of Mathematical Physics, 28(7):1459– 1470, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1459_s1.

McLaughlin:1987:UTI

[1129] Joyce R. McLaughlin and William Rundell. A uniqueness theorem for an inverse Sturm–Liouville problem. Journal of Mathematical Physics, 28(7):1471–1472, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i7/p1471_s1.

Sastri:1987:OMC

[1130] C. C. A. Sastri, K. A. Dunn, and D. R. K. S. Rao. Ovsiannikov’s method and the construction of partially invariant solutions. Journal of Mathematical Physics, 28(7):1473–1476, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i7/p1473_s1.

Swart:1987:STE

[1131] J. H. Swart. On the stability of the telegraph equation. Journal of Mathematical Physics, 28(7):1477–1480, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i7/p1477_s1.

Yagle:1987:MIS

[1132] Andrew E. Yagle. Multidimensional inverse scattering: an orthogonalization formulation. Journal of Mathematical Physics, 28(7):1481–1491, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i7/p1481_s1.

Coleman:1987:PME

[1133] Robert Alan Coleman and Herbert Korte. Any physical, monopole equation of motion structure uniquely determines a projective inertial structure and an (n − 1)-force. Journal of Mathematical Physics, 28(7):1492– 1498, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1492_s1.

Neyzi:1987:CSD

[1134] Fahrünisa Neyzi and Yavuz Nutku. Canonical structures for dispersive waves in shallow water. Journal of Mathematical Physics, 28(7):1499– 1504, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1499_s1.

Abounadi:1987:DWE

[1135] Jinane Abounadi and Teymour Darkhosh. The Dirac wave equation in the presence of an external field. Journal of Mathematical Physics, 28(7): 1505–1507, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1505_s1.

Landau:1987:EWD

[1136] Basil V. Landau. Extensions of Wigner’s distribution to particles with spin 1/2. Journal of Mathematical Physics, 28(7):1508–1511, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i7/
p1508_s1.

Bolle:1987:WIA

[1137] D. Bollé, F. Gesztesy, H. Grosse, W. Schweiger, and B. Simon. Witten index, axial anomaly, and Kreĭn’s spectral shift function in supersymmetric quantum mechanics. Journal of Mathematical Physics, 28(7): 1512–1525, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1512_s1.

Solomon:1987:MUS

[1138] Allan I. Solomon and Joseph L. Birman. An SU(8) model for the unification of superconductivity, charge, and spin density waves. Journal of Mathematical Physics, 28(7):1526–1534, July 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i7/p1526_s1.

Carter:1987:SKM

[1139] B. Carter. Separability of the Killing–Maxwell system underlying the generalized angular momentum constant in the Kerr–Newman black hole metrics. Journal of Mathematical Physics, 28(7):1535–1538, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i7/p1535_s1.

Chi:1987:NSS

[1140] L. K. Chi. A new self-similar space–time. Journal of Mathematical Physics, 28(7):1539–1540, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i7/p1539_s1.

Toimela:1987:STO

[1141] Tuomo Toimela. On the solution of the Tolman–Oppenheimer–Volkov equation with the ultrarelativistic equation of state. Journal of Mathematical Physics, 28(7):1541–1543, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i7/p1541_s1.

Krisch:1987:SFS

[1142] J. P. Krisch. Spin fluids in stationary axis-symmetric space–times. Journal of Mathematical Physics, 28(7):1544–1546, July 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i7/p1544_s1.

Accioly:1987:HST

[1143] A. J. Accioly and A. T. Gonçalves. Homogeneous space–times of Gödeltype in higher-derivative gravity. Journal of Mathematical Physics, 28(7): 1547–1552, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1547_s1.

Higuchi:1987:STS

[1144] Atsushi Higuchi. Symmetric tensor spherical harmonics on the N-sphere and their application to the de Sitter group SO(N, 1). Journal of Mathematical Physics, 28(7):1553–1566, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i7/p1553_s1. See erratum [2142].

Cornille:1987:ESD

[1145] Henri Cornille. Exact solutions in 1 + 1 dimensions of the general twovelocity discrete Illner model. Journal of Mathematical Physics, 28(7): 1567–1579, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1567_s1.

Farina:1987:EODa

[1146] J. E. G. Farina. The effect of a one-dimensional potential of finite range on the statistical parameters of an incident ensemble of particles. I. Pure states. Journal of Mathematical Physics, 28(7):1580–1582, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i7/p1580_s1.

Farina:1987:EODb

[1147] J. E. G. Farina. The effect of a one-dimensional potential of finite range on the statistical parameters of an incident ensemble of particles. II. Localized states. Journal of Mathematical Physics, 28(7):1583–1588, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i7/p1583_s1.

Hsieh:1987:PDE

[1148] D. Y. Hsieh. On partial differential equations related to Lorenz system. Journal of Mathematical Physics, 28(7):1589–1597, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i7/p1589_s1.

Furlan:1987:QCF

[1149] P. Furlan, G. M. Sotkov, and I. T. Todorov. Quasiprimary composite fields and null vectors in critical Ising-type models. Journal of Mathematical Physics, 28(7):1598–1607, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i7/p1598_s1.

Panagopoulos:1987:SCV

[1150] Haralambos Panagopoulos and Luc Vinet. Superinvariant chiral and vector fields. Journal of Mathematical Physics, 28(7):1608–1618, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i7/p1608_s1.

Flesch:1987:GFN

[1151] R. J. Flesch and S. E. Trullinger. Green’s functions for nonlinear Klein– Gordon kink perturbation theory. Journal of Mathematical Physics, 28(7):1619–1631, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i7/p1619_s1.

Manias:1987:CAC

[1152] M. V. Manías, M. C. von Reichenbach, F. A. Schaposnik, and M. Trobo. Current algebra for chiral gauge theories. Journal of Mathematical Physics, 28(7):1632–1635, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i7/p1632_s1.

Stark:1987:PCI

[1153] Hans-Georg Stark. Poincaré covariant infraparticle sectors. Journal of Mathematical Physics, 28(7):1636–1643, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i7/p1636_s1.

Kota:1987:GTA

[1154] V. K. B. Kota, J. Van der Jeugt, H. De Meyer, and G. Vanden Berghe. Group theoretical aspects of the extended interacting boson model. Journal of Mathematical Physics, 28(7):1644–1652, July 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i7/p1644_s1.

Sloane:1987:TSM

[1155] N. J. A. Sloane. Theta series and magic numbers for diamond and certain ionic crystal structures. Journal of Mathematical Physics, 28(7):1653– 1657, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1653_s1. See comment [1624].

Weber:1987:GQS

[1156] Edmond Weber. Global qualitative study of Bianchi universes in the presence of a cosmological constant. Journal of Mathematical Physics, 28(7): 1658–1666, July 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i7/p1658_s1.

Giroux:1987:FTR

[1157] Y. Giroux and R. T. Sharp. A folk theorem revisited: Degenerate representations. Journal of Mathematical Physics, 28(8):1671–1672, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1671_s1.

Morel:1987:IRS

[1158] B. Morel, J. Patera, R. T. Sharp, and J. Van der Jeugt. Indices of representations of simple superalgebras. Journal of Mathematical Physics, 28(8):1673–1682, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1673_s1.

Trullinger:1987:PPI

[1159] S. E. Trullinger and R. J. Flesch. Parent potentials for an infinite class of reflectionless kinks. Journal of Mathematical Physics, 28(8):1683– 1690, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i8/p1683_s1.

Post:1987:F

[1160] G. Post. On the τ-functions of A_{2}^{(2)}. Journal of Mathematical Physics, 28(8):1691–1696, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1691_s1.

Chowdhury:1987:PAT

[1161] A. Roy Chowdhury and Siraj Ahmad. On the prolongation approach in three dimensions for the conservation laws and Lax pair of the Benjamin– Ono equation. Journal of Mathematical Physics, 28(8):1697–1699, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1697_s1.

Kobayashi:1987:GCS

[1162] Yuji Kobayashi and Shigeaki Nagamachi. Generalized complex superspace—involutions of superfields. Journal of Mathematical Physics, 28(8):1700–1708, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1700_s1.

Hou:1987:CCS

[1163] Bo Yu Hou and Yao Zhong Zhang. Cohomology in connection space, family index theorem, and abelian gauge structure. Journal of Mathematical Physics, 28(8):1709–1715, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i8/p1709_s1.

Cicuta:1987:LRR

[1164] G. M. Cicuta, L. Molinari, E. Montaldi, and F. Riva. Large rectangular random matrices. Journal of Mathematical Physics, 28(8):1716– 1718, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i8/p1716_s1.

Augusteijn:1987:WBS

[1165] Marijke F. Augusteijn and Ernst Breitenberger. Waves, bifurcations, and solitons in a model with sixfold symmetry. Journal of Mathematical Physics, 28(8):1719–1728, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i8/p1719_s1.

Drijkoningen:1987:HWA

[1166] G. G. Drijkoningen, C. H. Chapman, and C. J. Thompson. On head-wave amplitudes. Journal of Mathematical Physics, 28(8):1729–1731, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1729_s1.

Hietarinta:1987:SBEa

[1167] Jarmo Hietarinta. A search for bilinear equations passing Hirota’s threesoliton condition. I. KdV-type bilinear equations. Journal of Mathematical Physics, 28(8):1732–1742, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i8/p1732_s1.

Hillion:1987:SFW

[1168] Pierre Hillion. Spinor focus wave modes. Journal of Mathematical Physics, 28(8):1743–1748, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i8/p1743_s1.

Blanco:1987:AOF

[1169] R. Blanco, L. Pesquera, and E. Santos. The anharmonic oscillator at a finite temperature. Comparison of quantum and classical stochastic calculations. Journal of Mathematical Physics, 28(8):1749–1758, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1749_s1.

Jordan:1987:DCB

[1170] Thomas F. Jordan. Direct calculation of the Berry phase for spins and helicities. Journal of Mathematical Physics, 28(8):1759–1760, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1759_s1.

Kruszynski:1987:COR

[1171] P. Kruszyński and W. M. de Muynck. Compatibility of observables represented by positive operator-valued measures. Journal of Mathematical Physics, 28(8):1761–1763, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i8/p1761_s1.

Lahti:1987:HIC

[1172] Pekka J. Lahti and Maciej J. Maczyński. Heisenberg inequality and the complex field in quantum mechanics. Journal of Mathematical Physics, 28(8):1764–1769, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1764_s1.

Ashbaugh:1987:MME

[1173] Mark S. Ashbaugh and Evans M. Harrell, II. Maximal and minimal eigenvalues and their associated nonlinear equations. Journal of Mathematical Physics, 28(8):1770–1786, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i8/p1770_s1.

Beumee:1987:AFTa

[1174] J. G. B. Beumee and H. Rabitz. An application of filtering theory to parameter identification using stochastic mechanics. Journal of Mathematical Physics, 28(8):1787–1794, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i8/p1787_s1.

Beumee:1987:AFTb

[1175] J. G. B. Beumee and H. Rabitz. An application of filtering theory to parameter identification in quantum mechanics. Journal of Mathematical Physics, 28(8):1795–1800, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i8/p1795_s1.

Lai:1987:PRE

[1176] C. H. Lai. Perturbative results from the 1/N expansion for screened Coulomb potentials. Journal of Mathematical Physics, 28(8):1801– 1808, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i8/p1801_s1.

Chowdhury:1987:CIS

[1177] A. Roy Chowdhury and Minati Naskar. On the complete integrability of the supersymmetric nonlinear Schrödinger equation. Journal of Mathematical Physics, 28(8):1809–1812, August 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i8/p1809_s1.

Froman:1987:ELC

[1178] Per Olof Fröman, Staffan Yngve, and Nanny Fröman. The energy levels and the corresponding normalized wave functions for a model of a compressed atom. Journal of Mathematical Physics, 28(8):1813–1826, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1813_s1.

Requardt:1987:INS

[1179] Manfred Requardt. An investigation of (nonadditive) scattering invariants in classical mechanics and quantum theory by differential topological methods. Journal of Mathematical Physics, 28(8):1827–1836, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1827_s1.

Hall:1987:UMW

[1180] G. S. Hall and A. D. Rendall. Uniqueness of the metric from the Weyl and energy-momentum tensors. Journal of Mathematical Physics, 28(8):1837–1839, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1837_s1.

McKellar:1987:LIE

[1181] R. J. McKellar. Lorentz invariance and the equivalence of metric, spinor, and frame field concomitants. Journal of Mathematical Physics, 28(8):1840–1843, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i8/p1840_s1.

Brill:1987:LLS

[1182] D. Brill, O. Reula, and B. Schmidt. Local linearization stability. Journal of Mathematical Physics, 28(8):1844–1847, August 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i8/p1844_s1.

Carminati:1987:SFP

[1183] J. Carminati. Shear-free perfect fluids in general relativity. I. Petrov type N Weyl tensor. Journal of Mathematical Physics, 28(8):1848–1853, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1848_s1.

Castagnino:1987:WTT

[1184] M. Castagnino, G. Domenech, R. J. Noriega, and C. G. Schifini. On a Weyl-type theorem for higher-order Lagrangians. Journal of Mathematical Physics, 28(8):1854–1857, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i8/p1854_s1.

Applebaum:1987:SWW

[1185] David Applebaum. The stochastic Wigner–Weisskopf atom. Journal of Mathematical Physics, 28(8):1858–1863, August 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i8/p1858_s1.

Witte:1987:RCB

[1186] Nicholas S. Witte, Ross L. Dawe, and Kenneth C. Hines. Relativistic charged bosons in a magnetic field. I. Wave functions and matrix elements. Journal of Mathematical Physics, 28(8):1864–1874, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1864_s1.

Bandelloni:1987:FDA

[1187] Giuseppe Bandelloni. Field-diffeomorphism anomalies in the nonlinear σ model. Journal of Mathematical Physics, 28(8):1875–1882, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1875_s1.

AlSaad:1987:IPL

[1188] A. Al Saad and I. M. Benn. Ideal-preserving Lorentzian connections. Journal of Mathematical Physics, 28(8):1883–1885, August 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i8/p1883_
s1.

Tsujishita:1987:KIP

[1189] Toru Tsujishita. KdV-invariant polynomial functionals. Journal of Mathematical Physics, 28(8):1886–1900, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i8/p1886_s1.

Brihaye:1987:MLP

[1190] Yves Brihaye and Jean Nuyts. Minimization of Landau potentials invariant under SO(3). Journal of Mathematical Physics, 28(8):1901–1911, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1901_s1.

Morchio:1987:LRD

[1191] G. Morchio and F. Strocchi. Long-range dynamics and broken symmetries in gauge models. The Schwinger model. Journal of Mathematical Physics, 28(8):1912–1919, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i8/p1912_s1.

Liu:1987:CMGa

[1192] Hong Ya Liu. Cosmological models in globally geodesic coordinates. I. Metric. Journal of Mathematical Physics, 28(8):1920–1923, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i8/p1920_s1.

Liu:1987:CMGb

[1193] Hong Ya Liu. Cosmological models in globally geodesic coordinates. II. Near-field approximation. Journal of Mathematical Physics, 28(8):1924– 1927, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i8/p1924_s1.

Schmidt:1987:GLS

[1194] Hans-Jürgen Schmidt. On Geroch’s limit of space–times and its relation to a new topology in the space of Lie groups. Journal of Mathematical Physics, 28(8):1928–1936, August 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v28/i8/p1928_s1. See addendum [1512].

Sharma:1987:STU

[1195] C. S. Sharma and T. J. Coulson. Spectral theory for unitary operators on a quaternionic Hilbert space. Journal of Mathematical Physics, 28 (9):1941–1946, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p1941_s1.

Dirl:1987:GRR

[1196] R. Dirl, P. Kasperkovitz, M. I. Aroyo, J. N. Kotzev, and B. L. Davies. Generating relations for reducing matrices. III. Kronecker products. Journal of Mathematical Physics, 28(9):1947–1971, September 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p1947_
s1.

Bertrand:1987:NRP

[1197] J. Bertrand and G. Rideau. Nonlinear representations of Poincaré group in three dimensions. Journal of Mathematical Physics, 28(9):1972–1977, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p1972_s1.

Bohm:1987:PIC

[1198] Manfred Böhm and Georg Junker. Path integration over compact and noncompact rotation groups. Journal of Mathematical Physics, 28 (9):1978–1994, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p1978_s1.

Alcaras:1987:PTA

[1199] J. A. Castilho Alcarás and V. Vanagas. The polynomial-type analysis of SU(3) group-theoretical quantities. Journal of Mathematical Physics, 28(9):1995–2006, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p1995_s1.

Gourdin:1987:CIS

[1200] Michel Gourdin. Classification and interpretation of the supertableaux of the orthosymplectic groups OSP(m|4). Journal of Mathematical Physics, 28(9):2007–2017, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2007_s1.

Athorne:1987:IED

[1201] Chris Athorne and Allan Fordy. Integrable equations in (2 + 1) dimensions associated with symmetric and homogeneous spaces. Journal of Mathematical Physics, 28(9):2018–2024, September 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2018_s1.

Weiss:1987:PFP

[1202] John Weiss. Periodic fixed points of Bäcklund transformations. Journal of Mathematical Physics, 28(9):2025–2039, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2025_s1.

Montaldi:1987:RSH

[1203] E. Montaldi and G. Zucchelli. Remarks on some hypergeometric orthogonal polynomials of mathematical physics. Journal of Mathematical Physics, 28(9):2040–2045, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i9/p2040_s1.

Shabani:1987:SCU

[1204] J. Shabani. On some classes of unbounded commutants of unbounded operator families. Journal of Mathematical Physics, 28(9):2046–2050, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2046_s1.

Maartens:1987:ACR

[1205] R. Maartens. Affine collineations in Robertson–Walker space–time. Journal of Mathematical Physics, 28(9):2051–2052, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2051_s1.

Luehr:1987:GSU

[1206] C. P. Luehr and M. Rosenbaum. Graded spinors as an underlying geometry for extended supersymmetries. Journal of Mathematical Physics, 28(9):2053–2065, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2053_s1.

Ozsvath:1987:SP

[1207] István Ozsváth and Leland Sapiro. The Schücking problem. Journal of Mathematical Physics, 28(9):2066–2073, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2066_s1.

Bruzzo:1987:GUT

[1208] Ugo Bruzzo. The global Utiyama theorem in Einstein–Cartan theory. Journal of Mathematical Physics, 28(9):2074–2077, September 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2074_
s1.

Celletti:1987:REC

[1209] Alessandra Celletti and Luigi Chierchia. Rigorous estimates for a computer-assisted KAM theory. Journal of Mathematical Physics, 28 (9):2078–2086, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2078_s1.

Guha-Roy:1987:SWS

[1210] C. Guha-Roy. Solitary wave solutions of a system of coupled nonlinear equations. Journal of Mathematical Physics, 28(9):2087–2088, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i9/p2087_s1.

Hogan:1987:NPC

[1211] P. A. Hogan. Newman–Penrose constants for zero-rest-mass fields on Minkowskian space–time. Journal of Mathematical Physics, 28(9):2089– 2093, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2089_s1.

Hietarinta:1987:SBEb

[1212] Jarmo Hietarinta. A search for bilinear equations passing Hirota’s threesoliton condition. II. mKdV-type bilinear equations. Journal of Mathematical Physics, 28(9):2094–2101, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i9/p2094_s1.

Segert:1987:NAB

[1213] Jan Segert. Non-Abelian Berry’s phase, accidental degeneracy, and angular momentum. Journal of Mathematical Physics, 28(9):2102–2114, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2102_s1.

Cieciura:1987:PRW

[1214] Grzegorz Cieciura and Igor Szczyrba. Physical relations and the Weyl group. Journal of Mathematical Physics, 28(9):2115–2127, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i9/p2115_s1.

Cieciura:1987:WGT

[1215] Grzegorz Cieciura and Igor Szczyrba. Weyl group and tensor operators for hadron-type SU(n) representations. Journal of Mathematical Physics, 28(9):2128–2141, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2128_s1.

DeBievre:1987:MOC

[1216] S. De Bièvre. Möller operators in classical relativistic two-particle scattering. Journal of Mathematical Physics, 28(9):2142–2148, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i9/p2142_s1.

Magnon:1987:MDM

[1217] Anne Magnon. Mass, dual mass, and gravitational entropy. Journal of Mathematical Physics, 28(9):2149–2154, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2149_s1.

Ernst:1987:CGPa

[1218] Frederick J. Ernst, Alberto García D., and Isidore Hauser. Colliding gravitational plane waves with noncollinear polarization. I. Journal of Mathematical Physics, 28(9):2155–2161, September 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2155_s1.

Kyriakopoulos:1987:SEE

[1219] E. Kyriakopoulos. Solutions of Einstein’s equations for the interior of a stationary axisymmetric perfect fluid. Journal of Mathematical Physics, 28(9):2162–2166, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2162_s1.

Ross:1987:CQS

[1220] D. K. Ross. Charge quantization without superheavy masses in a Kaluza– Klein description of electromagnetism. Journal of Mathematical Physics, 28(9):2167–2170, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2167_s1.

SalazarI:1987:DRT

[1221] Humberto Salazar I., Alberto García D., and Jerzy Plebański. Duality rotations and type D solutions to Einstein equations with nonlinear electromagnetic sources. Journal of Mathematical Physics, 28(9):2171–2181, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2171_s1.

Mason:1987:KDC

[1222] D. P. Mason and R. Maartens. Kinematics and dynamics of conformal collineations in relativity. Journal of Mathematical Physics, 28(9):2182– 2186, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2182_s1.

Zinoun:1987:AOS

[1223] A. Zinoun and J. Cortois. Algebra of operators for a system of composite and elementary particles. Journal of Mathematical Physics, 28(9):2187– 2191, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2187_s1.

Sen:1987:PBB

[1224] Xiao Yan Sen and Z. R. Yu. The physical basis of an n g-boson system. Journal of Mathematical Physics, 28(9):2192–2193, September 1987. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2192_
s1.

Winternitz:1987:ESM

[1225] P. Winternitz, A. M. Grundland, and J. A. Tuszyński. Exact solutions of the multidimensional classical ϕ^{6}-field equations obtained by symmetry reduction. Journal of Mathematical Physics, 28(9):2194–2212, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2194_s1.

Caprasse:1987:RSR

[1226] H. Caprasse and M. Hans. Renormalization schemes and renormalization group functions in Yang–Mills theories with massive fermions. Journal of Mathematical Physics, 28(9):2213–2219, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i9/p2213_s1.

Szabo:1987:QKS

[1227] László Szabó. Quantum kinematics of strings: quantization on C^{∞}(M^{3}, R^{2})/C^{∞}(M^{3}, GL(2, R)). Journal of Mathematical Physics, 28 (9):2220–2222, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2220_s1.

Chacon:1987:CGV

[1228] E. Chacón, P. O. Hess, and M. Moshinsky. Collectivity and geometry. V. Spectra and shapes in the two-dimensional case. Journal of Mathematical Physics, 28(9):2223–2240, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i9/p2223_s1.

Brevik:1987:PEC

[1229] I. Brevik. Phenomenological electrodynamics in curvilinear space, with application to Rindler space. Journal of Mathematical Physics, 28 (9):2241–2249, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i9/p2241_s1.

Goderis:1987:RDM

[1230] D. Goderis, A. Verbeure, and P. Vets. Relaxation of the Dicke maser model. Journal of Mathematical Physics, 28(9):2250–2253, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i9/p2250_s1.

Exner:1987:EQM

[1231] P. Exner and P. Šeba. Erratum: “Quantum motion on a half-line connected to a plane”, [J. Math. Phys. 28 (1987), no. 2, 386–391; MR 88e:81016]. Journal of Mathematical Physics, 28(9):2254, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i9/p2254_s2. See [965].

VanderJeugt:1987:ERS

[1232] J. Van der Jeugt. Erratum: “Regular subalgebras of Lie superalgebras and extended Dynkin diagrams”, [J. Math. Phys. 28 (1987), no. 2, 292–301; MR 88d:17020]. Journal of Mathematical Physics, 28(9):2254, September 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i9/p2254_s1. See [951].

Bacry:1987:CP

[1233] Henri Bacry. SL(2, C), SU(2), and Chebyshev polynomials. Journal of Mathematical Physics, 28(10):2259–2267, October 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2259_s1.

Krause:1987:PAT

[1234] J. Krause. Pedestrian approach to two-cocycles for unitary ray representations of Lie groups. Journal of Mathematical Physics, 28(10):2268– 2279, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2268_s1.

Palev:1987:FDRb

[1235] Tchavdar D. Palev. Finite-dimensional representations of the special linear Lie superalgebra sl(1, n). I. Typical representations. Journal of Mathematical Physics, 28(10):2280–2303, October 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2280_s1.

Gould:1987:PCA

[1236] M. D. Gould and J. Paldus. Parastatistics and the Clifford algebra unitary group approach to the many-electron correlation problem. Journal of Mathematical Physics, 28(10):2304–2309, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2304_s1.

Santini:1987:IVP

[1237] P. M. Santini, M. J. Ablowitz, and A. S. Fokas. On the initial value problem for a class of nonlinear integral evolution equations including the sine-Hilbert equation. Journal of Mathematical Physics, 28(10):2310– 2316, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2310_s1.

Matsuno:1987:ESN

[1238] Y. Matsuno. Exact solutions for the nonlinear Klein–Gordon and Liouville equations in four-dimensional Euclidean space. Journal of Mathematical Physics, 28(10):2317–2322, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i10/p2317_s1.

Kerber:1987:TIS

[1239] Adalbert Kerber and Thomas Scharf. Tensors with icosahedral symmetry that are invariant under a certain wreath product. Journal of Mathematical Physics, 28(10):2323–2324, October 1987. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2323_s1.

Wang:1987:LGFa

[1240] Ren Chuan Wang. Local Green’s function. I. Journal of Mathematical Physics, 28(10):2325–2328, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i10/p2325_s1.

Wang:1987:LGFb

[1241] Renchuan Wang. Local Green’s function. II. Journal of Mathematical Physics, 28(10):2329–2332, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i10/p2329_s1.

Wang:1987:LGFc

[1242] Renchuan Wang. Local Green’s function. III. Journal of Mathematical Physics, 28(10):2333–2337, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i10/p2333_s1.

Jaulent:1987:SWT

[1243] M. Jaulent and M. Manna. The “spectral Wronskian” tool and the \overline{∂} investigation of the KdV hierarchy. Journal of Mathematical Physics, 28(10):2338–2342, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2338_s1.

Noguchi:1987:YMH

[1244] Mitsunori Noguchi. Yang–Mills–Higgs theory on a compact Riemann surface. Journal of Mathematical Physics, 28(10):2343–2346, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2343_s1.

Betounes:1987:DGA

[1245] David Betounes. Differential geometric aspects of the Cartan form: symmetry theory. Journal of Mathematical Physics, 28(10):2347–2353, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2347_s1.

Nikolov:1987:DRI

[1246] P. A. Nikolov. On the dimensional reduction of invariant fields and differential operators. Journal of Mathematical Physics, 28(10):2354– 2362, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2354_s1.

Bartocci:1987:CS

[1247] Claudio Bartocci and Ugo Bruzzo. Cohomology of supermanifolds. Journal of Mathematical Physics, 28(10):2363–2368, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2363_s1.

Negri:1987:CTC

[1248] Luiz J. Negri, Luimar C. Oliveira, and Julio M. Teixeira. Canonoid transformations and constants of motion. Journal of Mathematical Physics, 28(10):2369–2372, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2369_s1.

Han:1987:DLT

[1249] D. Han, Y. S. Kim, and D. Son. Decomposition of Lorentz transformations. Journal of Mathematical Physics, 28(10):2373–2378, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2373_s1.

Nelson:1987:GLT

[1250] Robert A. Nelson. Generalized Lorentz transformation for an accelerated, rotating frame of reference. Journal of Mathematical Physics, 28(10):2379–2383, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2379_s1. See erratum [2141].

Amann:1987:JPS

[1251] Anton Amann. Jauch–Piron states in W^{∗}-algebraic quantum mechanics. Journal of Mathematical Physics, 28(10):2384–2389, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/
p2384_s1.

Varilly:1987:WTF

[1252] Joseph C. Várilly and José M. Gracia-Bondía. The Wigner transformation is of finite order. Journal of Mathematical Physics, 28(10):2390– 2392, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2390_s1.

Pulmannova:1987:GPT

[1253] Sylvia Pulmannová and Stanley Gudder. Geometric properties of transition amplitude spaces. Journal of Mathematical Physics, 28(10):2393– 2399, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2393_s1.

Gingold:1987:GLD

[1254] H. Gingold. In general, the less degeneracy the less transition. A principle for time-dependent Hamiltonian systems in quantum mechanics. Journal of Mathematical Physics, 28(10):2400–2406, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2400_s1.

Kowalski:1987:TOD

[1255] Jacek M. Kowalski and John L. Fry. Tunneling in one-dimensional ideal barriers. Journal of Mathematical Physics, 28(10):2407–2415, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2407_s1.

Somersalo:1987:ISS

[1256] Erkki Somersalo. Inverse scattering for standing wave solutions of the Schrödinger equation. Journal of Mathematical Physics, 28(10):2416– 2419, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2416_s1.

Galvan:1987:ADE

[1257] J. B. Galvan and J. Sesma. Ambiguities in the Debye expansion of the elastic S matrix. Journal of Mathematical Physics, 28(10):2420– 2426, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2420_s1.

Malament:1987:NAC

[1258] David B. Malament. A note about closed timelike curves in Gödel space– time. Journal of Mathematical Physics, 28(10):2427–2430, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/
p2427_s1.

Letelier:1987:AFT

[1259] Patricio S. Letelier and Enric Verdaguer. Anisotropic fluid with SU(2)- type structure in general relativity: a model of localized matter. Journal of Mathematical Physics, 28(10):2431–2436, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i10/p2431_s1.

Ranganathan:1987:GID

[1260] D. Ranganathan. A geometric interpretation for the Dirac field in curved space. Journal of Mathematical Physics, 28(10):2437–2439, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2437_s1.

Summers:1987:BIQa

[1261] Stephen J. Summers and Reinhard Werner. Bell’s inequalities and quantum field theory. I. General setting. Journal of Mathematical Physics, 28(10):2440–2447, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2440_s1.

Summers:1987:BIQb

[1262] Stephen J. Summers and Reinhard Werner. Bell’s inequalities and quantum field theory. II. Bell’s inequalities are maximally violated in the vacuum. Journal of Mathematical Physics, 28(10):2448–2456, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2448_s1.

ElHassouni:1987:GSE

[1263] A. El Hassouni, E. G. Oudrhiri-Safiani, and E. H. Saidi. Generalization of the N = 1 supersymmetric effective Lagrangian to arbitrary N in the absence of central charges. Journal of Mathematical Physics, 28(10):2457–2462, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2457_s1.

Jacques:1987:IDL

[1264] Michel Jacques and Yvan Saint-Aubin. Infinite-dimensional Lie algebras acting on the solution space of various σ models. Journal of Mathematical Physics, 28(10):2463–2479, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i10/p2463_s1.

Kostelecky:1987:BCH

[1265] V. Alan Kostelecký and D. Rodney Truax. Baker–Campbell–Hausdorff relations for the super-Poincaré group. Journal of Mathematical Physics, 28(10):2480–2487, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2480_s1.

Atkinson:1987:CSBa

[1266] D. Atkinson and P. W. Johnson. Chiral symmetry breakdown. I. Gauge dependence in constant vertex approximation. Journal of Mathematical Physics, 28(10):2488–2493, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i10/p2488_s1.

Atkinson:1987:CSBb

[1267] D. Atkinson. Chiral symmetry breakdown. II. Ansatz for vertex in ultraviolet. Journal of Mathematical Physics, 28(10):2494–2497, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i10/p2494_s1.

Wolf:1987:GTT

[1268] Kurt Bernardo Wolf. The group-theoretical treatment of aberrating systems. III. The classification of asymmetric aberrations. Journal of Mathematical Physics, 28(10):2498–2507, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i10/p2498_s1.

Lewis:1987:SBR

[1269] D. Lewis, J. Marsden, and T. Ratiu. Stability and bifurcation of a rotating planar liquid drop. Journal of Mathematical Physics, 28(10):2508– 2515, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i10/p2508_s1.

Keller:1987:ECP

[1270] Joseph B. Keller. Effective conductivity of periodic composites composed of two very unequal conductors. Journal of Mathematical Physics, 28(10):2516–2520, October 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i10/p2516_s1.

Vanthournout:1987:TSI

[1271] J. Vanthournout, J. Van der Jeugt, H. De Meyer, and G. Vanden Berghe. Totally symmetric irreducible representations of the group SO(6) in the principal SO(3) subgroup basis. Journal of Mathematical Physics, 28 (11):2529–2539, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2529_s1.

Kim:1987:TJS

[1272] Shoon K. Kim. Theorems on the Jordan–Schwinger representations of Lie algebras. Journal of Mathematical Physics, 28(11):2540–2545, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2540_s1.

Saenz:1987:CMA

[1273] A. W. Sáenz. Canonical map approach to channeling stability in crystals. II. Journal of Mathematical Physics, 28(11):2546–2558, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2546_s1.

Nakayama:1987:FWH

[1274] J. Nakayama. A formula on the Wiener–Hermite expansion. Journal of Mathematical Physics, 28(11):2559–2563, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/p2559_s1.

McInnes:1987:SSI

[1275] Brett McInnes. Spontaneous splitting and internal isometries of superstring vacua. Journal of Mathematical Physics, 28(11):2564–2568, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2564_s1.

Misawa:1987:CSD

[1276] Tetsuya Misawa and Kunio Yasue. Canonical stochastic dynamical systems. Journal of Mathematical Physics, 28(11):2569–2573, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2569_s1.

Leznov:1987:DAS

[1277] A. N. Leznov and M. A. Mukhtarov. Deformation of algebras and solution of self-duality equation. Journal of Mathematical Physics, 28(11):2574– 2578, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2574_s1.

Nutku:1987:NCC

[1278] Y. Nutku. On a new class of completely integrable nonlinear wave equations. II. Multi-Hamiltonian structure. Journal of Mathematical Physics, 28(11):2579–2585, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2579_s1.

Hietarinta:1987:SBEc

[1279] Jarmo Hietarinta. A search for bilinear equations passing Hirota’s threesoliton condition. III. Sine-Gordon-type bilinear equations. Journal of Mathematical Physics, 28(11):2586–2592, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/p2586_s1.

Liu:1987:NRQ

[1280] Tai-Ping Liu. Nonlinear resonance for quasilinear hyperbolic equation. Journal of Mathematical Physics, 28(11):2593–2602, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/
p2593_s1.

Melnikov:1987:RWN

[1281] V. K. Mel’nikov. Reflection of waves in nonlinear integrable systems. Journal of Mathematical Physics, 28(11):2603–2609, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/
p2603_s1.

Villarroel:1987:YMS

[1282] Javier Villarroel. Yang–Mills solutions in S^{3} × S^{1}. Journal of Mathematical Physics, 28(11):2610–2613, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i11/p2610_s1.

Lahti:1987:TNQ

[1283] Pekka J. Lahti and Kari Ylinen. On total noncommutativity in quantum mechanics. Journal of Mathematical Physics, 28(11):2614–2617, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2614_s1.

Sazdjian:1987:CTP

[1284] H. Sazdjian. The connection of two-particle relativistic quantum mechanics with the Bethe–Salpeter equation. Journal of Mathematical Physics, 28(11):2618–2638, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2618_s1.

Bacry:1987:BAS

[1285] H. Bacry and M. Boon. Boson algebra as a symplectic Clifford algebra. Journal of Mathematical Physics, 28(11):2639–2649, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2639_s1.

deLange:1987:FVO

[1286] O. L. de Lange. Factorizations of vector operators for the isotropic harmonic oscillator in an angular momentum basis. Journal of Mathematical Physics, 28(11):2650–2654, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2650_s1.

Caliceti:1987:PCP

[1287] E. Caliceti, L. Goldoni, and S. Graffi. Probability of convergence of perturbation theory for hydrogen photoionization. Journal of Mathematical Physics, 28(11):2655–2661, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2655_s1.

Pedrosa:1987:CTE

[1288] I. A. Pedrosa. Canonical transformations and exact invariants for dissipative systems. Journal of Mathematical Physics, 28(11):2662–2664, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2662_s1.

Bejarano:1987:ENR

[1289] J. Diaz Bejarano and C. Miro Rodriguez. Evaporation of nonzero rest mass particles from a black hole. Journal of Mathematical Physics, 28 (11):2665–2668, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2665_s1.

DiazBejarano:1987:ENR

[1290] J. Díaz Bejarano and C. Miró Rodríguez. Evaporation of nonzero rest mass particles from a black hole. Journal of Mathematical Physics, 28 (11):2665–2668, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

Rodrigues:1987:NOL

[1291] João P. Rodrigues. Nonorientable one-loop amplitudes for the bosonic open string: electrostatics on a Möbius strip. Journal of Mathematical Physics, 28(11):2669–2676, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2669_s1.

Matsui:1987:PTD

[1292] Yoshiko Matsui. Pole term decomposition of the resolvent kernel in Fredholm’s form. Journal of Mathematical Physics, 28(11):2677–2682, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2677_s1.

Borde:1987:ESG

[1293] Arvind Borde. The existence of singularities in general relativity despite isolated failures of geodesic focusing. Journal of Mathematical Physics, 28(11):2683–2684, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2683_s1.

Krolak:1987:SCC

[1294] Andrzej Królak. Strong cosmic censorship and the strong curvature singularities. Journal of Mathematical Physics, 28(11):2685–2687, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2685_s1.

Szabados:1987:CPC

[1295] L. B. Szabados. Commutation properties of cyclic and null Killing symmetries. Journal of Mathematical Physics, 28(11):2688–2691, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2688_s1.

DiPrisco:1987:BMC

[1296] A. Di Prisco, L. Herrera, J. Jimenez, V. Galina, and J. Ibáñez. The Bondi metric and conformal motions. Journal of Mathematical Physics, 28(11):2692–2696, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2692_s1.

Wang:1987:CCF

[1297] Xing Xiang Wang. A class of conformally flat solutions for a charged sphere in general relativity. Journal of Mathematical Physics, 28(11): 2697–2699, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2697_s1.

Duggal:1987:RFS

[1298] K. L. Duggal. Relativistic fluids with shear and timelike conformal collineations. Journal of Mathematical Physics, 28(11):2700–2704, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2700_s1.

Mason:1987:RRM

[1299] D. P. Mason and C. A. Pooe. Rotating rigid motion in general relativity. Journal of Mathematical Physics, 28(11):2705–2713, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2705_s1.

Buchert:1987:CSS

[1300] Thomas Buchert and Günter Götz. A class of solutions for selfgravitating dust in Newtonian gravity. Journal of Mathematical Physics, 28(11):2714–2719, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2714_s1.

Tomimatsu:1987:NWP

[1301] A. Tomimatsu and H. Ishihara. Nonlinear wave propagation in the Gauss–Bonnet extended Einstein theory. Journal of Mathematical Physics, 28(11):2720–2725, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2720_s1.

Barker:1987:CSD

[1302] William A. Barker, Daniel Raney, and John Sy. A case study of degeneracy in quantum statistics. I. Journal of Mathematical Physics, 28 (11):2726–2730, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2726_s1.

Arnoldus:1987:CFF

[1303] Henk F. Arnoldus and Thomas F. George. Correlation functions in finite memory-time reservoir theory. Journal of Mathematical Physics, 28 (11):2731–2738, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2731_s1.

Wheeler:1987:CPT

[1304] John C. Wheeler. Comment on a paper: “Thermodynamics of dimers on a rectangular L × M × N lattice”, [J. Math. Phys. 26 (1985), no. 10, 2491–2499; MR 86m:82050] by A. J. Phares and F. J. Wunderlich. Journal of Mathematical Physics, 28(11):2739–2740, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i11/p2739_s1. See [387].

Arimitsu:1987:CFD

[1305] T. Arimitsu, H. Umezawa, and Y. Yamanaka. Canonical formalism of dissipative fields in thermo field dynamics. Journal of Mathematical Physics, 28(11):2741–2752, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i11/p2741_s1. See erratum [1544].

Antoine:1987:CNM

[1306] J.-P. Antoine and B. Piette. Classical nonlinear σ models on Grassmann manifolds of compact or noncompact type. Journal of Mathematical Physics, 28(11):2753–2762, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2753_s1.

Challifour:1987:PSF

[1307] J. L. Challifour and W. A. Wood. Path-space formulas for covariant gauges in Kreĭn spaces. Journal of Mathematical Physics, 28(11):2763– 2774, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2763_s1.

Bandelloni:1987:NYM

[1308] Giuseppe Bandelloni. Nonpolynomial Yang–Mills local cohomology. Journal of Mathematical Physics, 28(11):2775–2780, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/
p2775_s1.

Gaigg:1987:LMP

[1309] P. Gaigg, M. Kreuzer, M. Schweda, and O. Piguet. The Leibbrandt– Mandelstam prescription for general axial gauge one-loop integrals. Journal of Mathematical Physics, 28(11):2781–2785, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i11/
p2781_s1.

Muller:1987:SPV

[1310] Eberhard E. Müller. Scalar potentials for vector fields in quantum electrodynamics. Journal of Mathematical Physics, 28(11):2786–2790, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2786_s1.

El-Dabe:1987:EST

[1311] Nabile T. M. El-Dabe. Electrohydrodynamic stability of two superposed elasticoviscous liquids in plane Couette flow. Journal of Mathematical Physics, 28(11):2791–2800, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i11/p2791_s1.

Ghosh:1987:IAS

[1312] Kamal Kumar Ghosh and Dipankar Ray. Ion acoustic solitons in a warm magnetoplasma. Journal of Mathematical Physics, 28(11):2801–2803, November 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i11/p2801_s1.

Humi:1987:RIE

[1313] Mayer Humi. Representations and invariant equations of E(3). Journal of Mathematical Physics, 28(12):2807–2811, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2807_s1.

Louck:1987:WZC

[1314] J. D. Louck and P. R. Stein. Weight-2 zeros of 3-j coefficients and the Pell equation. Journal of Mathematical Physics, 28(12):2812–2823, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2812_s1.

Smrz:1987:CST

[1315] P. K. Smrz. Construction of space–time by gauge translations. Journal of Mathematical Physics, 28(12):2824–2828, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2824_s1.

Tuynman:1987:QTC

[1316] G. M. Tuynman. Quantization: Towards a comparison between methods. Journal of Mathematical Physics, 28(12):2829–2840, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2829_s1.

Sen:1987:CIP

[1317] Tanaji Sen. A class of integrable potentials. Journal of Mathematical Physics, 28(12):2841–2850, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i12/p2841_s1.

Thompson:1987:SOE

[1318] G. Thompson. Second-order equation fields and the inverse problem of Lagrangian dynamics. Journal of Mathematical Physics, 28(12):2851– 2857, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i12/p2851_s1.

Dassios:1987:ISP

[1319] George Dassios. The inverse scattering problem for the soft ellipsoid. Journal of Mathematical Physics, 28(12):2858–2862, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/
p2858_s1.

Verheest:1987:IRM

[1320] Frank Verheest. Integrability of restricted multiple three-wave interactions. Journal of Mathematical Physics, 28(12):2863–2865, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2863_s1.

Busch:1987:QOC

[1321] Paul Busch, Thomas P. Schonbek, and Franklin E. Schroeck, Jr. Quantum observables: compatibility versus commutativity and maximal information. Journal of Mathematical Physics, 28(12):2866–2872, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2866_s1.

Narcowich:1987:PMP

[1322] Francis J. Narcowich. The problem of moments in the phase-space formulation of quantum mechanics. Journal of Mathematical Physics, 28 (12):2873–2882, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i12/p2873_s1.

Nash:1987:CSQ

[1323] C. G. Nash and G. C. Joshi. Composite systems in quaternionic quantum mechanics. Journal of Mathematical Physics, 28(12):2883–2885, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2883_s1.

Nash:1987:CSC

[1324] C. G. Nash and G. C. Joshi. Component states of a composite quaternionic system. Journal of Mathematical Physics, 28(12):2886–2890, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i12/p2886_s1.

Chierchia:1987:ACS

[1325] Luigi Chierchia. Absolutely continuous spectra of quasiperiodic Schrödinger operators. Journal of Mathematical Physics, 28(12):2891– 2898, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i12/p2891_s1.

Berceanu:1987:CPM

[1326] S. Berceanu and A. Gheorghe. On the construction of perfect Morse functions on compact manifolds of coherent states. Journal of Mathematical Physics, 28(12):2899–2907, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v28/i12/p2899_s1.

Fernandez:1987:ETE

[1327] Francisco M. Fernández. Exponential time-evolution operator for the time-dependent harmonic oscillator. Journal of Mathematical Physics, 28(12):2908, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i12/p2908_s1.

Labarthe:1987:QZR

[1328] J. J. Labarthe. Quadratic zeros of Racah 6-j coefficients: a geometrical approach. Journal of Mathematical Physics, 28(12):2909–2913, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2909_s1.

Bruckstein:1987:DSE

[1329] Alfred M. Bruckstein and Thomas Kailath. On discrete Schrödinger equations and their two-component wave equation equivalents. Journal of Mathematical Physics, 28(12):2914–2924, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2914_s1.

Hardman:1987:TDC

[1330] I. Hardman, H. Umezawa, and Y. Yamanaka. Time-dependent canonical formalism of thermally dissipative fields and renormalization scheme. Journal of Mathematical Physics, 28(12):2925–2938, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/
p2925_s1.

Yip:1987:TNI

[1331] P. F. Yip. There is no isolated p–p wave. Journal of Mathematical Physics, 28(12):2939–2941, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i12/p2939_s1.

Anile:1987:SSE

[1332] A. M. Anile, G. Moschetti, and O. I. Bogoyavlenski. Some solutions of Einstein’s equations with shock waves. Journal of Mathematical Physics, 28(12):2942–2948, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i12/p2942_s1.

Berger:1987:GES

[1333] Sonia Berger, Roberto Hojman, and Jorge Santamarina. General exact solutions of Einstein equations for static perfect fluids with spherical symmetry. Journal of Mathematical Physics, 28(12):2949–2950, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2949_s1.

Ernst:1987:CGPb

[1334] Frederick J. Ernst, Alberto García D., and Isidore Hauser. Colliding gravitational plane waves with noncollinear polarization. II. Journal of Mathematical Physics, 28(12):2951–2960, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2951_s1.

Gadella:1987:GWC

[1335] M. Gadella, L. M. Nieto, J. M. Noriega, and E. Santos. The generalized Weyl correspondence and time-dependent stochastic processes. Journal of Mathematical Physics, 28(12):2961–2972, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2961_s1.

March:1987:EPD

[1336] N. H. March. Effective potential determining one-dimensional Slater sum in independent-particle theory. Journal of Mathematical Physics, 28(12):2973–2974, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v28/i12/p2973_s1.

March:1987:EDS

[1337] N. H. March. Electron density and Slater sum of a semi-infinite electron gas in d dimensions. Journal of Mathematical Physics, 28(12):2975– 2976, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v28/i12/p2975_s1.

Elizalde:1987:TPD

[1338] E. Elizalde. Topologies for point distributions. Journal of Mathematical Physics, 28(12):2977–2981, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i12/p2977_s1.

Kitada:1987:CPC

[1339] Akihiko Kitada. Comment on: “On a property of a classical solution of the nonlinear mass transport equation u_{t} = u_{xx}/1 + u_{x}^{2}. II”, [J. Math. Phys. 28 (1987), no. 3, 536–537; MR 88e:82035] by Kitada and H. Umehara. Journal of Mathematical Physics, 28(12):2982–2983, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2982_s1. See [990].

Hostler:1987:RCS

[1340] Levere Hostler. Relativistic Coulomb Sturmian matrix elements and the Coulomb Green’s function of the second-order Dirac equation. Journal of Mathematical Physics, 28(12):2984–2989, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v28/i12/p2984_s1.

Abbasabadi:1987:NES

[1341] Alireza Abbasabadi and Wayne W. Repko. A note on the electron selfenergy. Journal of Mathematical Physics, 28(12):2990–2993, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v28/i12/p2990_s1.

Atkinson:1987:CSBc

[1342] D. Atkinson, A. Hulsebos, and P. W. Johnson. Chiral symmetry breakdown. III. Delbourgo’s gauge technique. Journal of Mathematical Physics, 28(12):2994–2998, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i12/p2994_s1.

McIlroy:1987:CSW

[1343] Douglas K. McIlroy and David P. Mason. The concept of streamfunctions in the Wien effect on a weak electrolyte. Journal of Mathematical Physics, 28(12):2999–3005, December 1987. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v28/i12/p2999_s1.

Champagne:1988:IDS

[1344] B. Champagne and P. Winternitz. On the infinite-dimensional symmetry group of the Davey–Stewartson equations. Journal of Mathematical Physics, 29(1):1–8, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p1_s1.

Aguirre:1988:GSTa

[1345] M. Aguirre and J. Krause. SL(3, R) as the group of symmetry transformations for all one-dimensional linear systems. Journal of Mathematical Physics, 29(1):9–15, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p9_s1.

Biswas:1988:SPO

[1346] S. N. Biswas and S. K. Soni. Supersymmetry, parastatistics, and operator realizations of a Lie algebra. Journal of Mathematical Physics, 29(1): 16–20, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p16_s1.

Holm:1988:ABS

[1347] Darryl D. Holm and Boris A. Kupershmidt. The analogy between spin glasses and Yang–Mills fluids. Journal of Mathematical Physics, 29(1): 21–30, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p21_s1.

Matori:1988:GHS

[1348] Yoshihiro Matori. Global hyperbolicity of a spatially closed space–time. Journal of Mathematical Physics, 29(1):31, January 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i1/p31_s1.

Handy:1988:SOD

[1349] Carlos R. Handy. A solution to the one-dimensional missing moment problem. Journal of Mathematical Physics, 29(1):32–35, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i1/
p32_s1.

Weston:1988:FWE

[1350] Vaughan H. Weston. Factorization of the wave equation in a nonplanar stratified medium. Journal of Mathematical Physics, 29(1):36–45, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v29/i1/p36_s1. See erratum [1691].

Kentwell:1988:HSM

[1351] G. W. Kentwell. The Hamiltonian structure of the modified and fifthorder Korteweg–de Vries equation. Journal of Mathematical Physics, 29(1):46–48, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p46_s1.

Matsuno:1988:NIN

[1352] Y. Matsuno. New integrable nonlinear integro-differential equations and related solvable finite-dimensional dynamical systems. Journal of Mathematical Physics, 29(1):49–56, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i1/p49_s1. See erratum [1747].

Baylis:1988:SRC

[1353] W. E. Baylis and George Jones. Special relativity with Clifford algebras and 2 × 2 matrices, and the exact product of two boosts. Journal of Mathematical Physics, 29(1):57–62, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i1/p57_s1.

Adachihara:1988:SWF

[1354] H. Adachihara, D. W. McLaughlin, J. V. Moloney, and A. C. Newell. Solitary waves as fixed points of infinite-dimensional maps for an optical bistable ring cavity: analysis. Journal of Mathematical Physics, 29(1): 63–85, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p63_s1.

Bluman:1988:ESW

[1355] George Bluman and Sukeyuki Kumei. Exact solutions for wave equations of two-layered media with smooth transition. Journal of Mathematical Physics, 29(1):86–96, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p86_s1.

Evans:1988:SMI

[1356] J. W. Evans. Splitting methods and invariant imbedding for timeindependent wave propagation in focusing media and wave guides. Journal of Mathematical Physics, 29(1):97–102, January 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i1/p97_s1.

Ichinose:1988:ZDP

[1357] Takashi Ichinose and Hiroshi Tamura. The Zitterbewegung of a Dirac particle in two-dimensional space–time. Journal of Mathematical Physics, 29(1):103–109, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p103_s1.

Grosse:1988:SGN

[1358] H. Grosse and L. Pittner. A supersymmetric generalization of von Neumann’s theorem. Journal of Mathematical Physics, 29(1):110–118, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p110_s1.

Arteca:1988:AEP

[1359] Gustavo A. Arteca and Paul G. Mezey. Approximate eigenvalues of parameter-dependent systems from boundaries of level sets. Journal of Mathematical Physics, 29(1):119–127, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i1/p119_s1.

Mainland:1988:HER

[1360] G. Bruce Mainland. Helicity eigenstates of a relativistic spin-0 and spin-1/2 constituent bound by minimal electrodynamics: zero orbital angular momentum, zero four-momentum solutions. Journal of Mathematical Physics, 29(1):128–138, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i1/p128_s1.

Znojil:1988:VCF

[1361] M. Znojil. Vectorial continued fractions and an algebraic construction of effective Hamiltonians. Journal of Mathematical Physics, 29(1):139– 147, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p139_s1.

Klaus:1988:EBJ

[1362] Martin Klaus. Exact behavior of Jost functions at low energy. Journal of Mathematical Physics, 29(1):148–154, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i1/p148_s1.

McGuire:1988:TIP

[1363] J. B. McGuire and C. A. Hurst. Three interacting particles in one dimension: an algebraic approach. Journal of Mathematical Physics, 29(1):155– 168, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p155_s1.

Diaz:1988:SSK

[1364] Mario C. Díaz, Reinaldo J. Gleiser, and Jorge A. Pullin. Solitonic solutions in the Kaluza–Klein–Jordan formalism as cosmological models in general relativity. Journal of Mathematical Physics, 29(1):169–174, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p169_s1.

GarciaD:1988:TRR

[1365] Alberto García D. and Isidore Hauser. Type-D rigidly rotating perfect fluid solutions. Journal of Mathematical Physics, 29(1):175–178, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p175_s1.

Harriott:1988:ESE

[1366] Tina A. Harriott and J. G. Williams. Exact solutions of the Einstein field equations for a topologically nontrivial metric. Journal of Mathematical Physics, 29(1):179–181, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i1/p179_s1.

Koppar:1988:KNM

[1367] Sharda S. Koppar and L. K. Patel. On the Kerr–Newman metric in cosmological background. Journal of Mathematical Physics, 29(1):182– 185, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p182_s1.

Saraykar:1988:SSS

[1368] R. V. Saraykar. Structure of the space of solutions of Einstein’s equations: a new variables approach. Journal of Mathematical Physics, 29(1):186– 196, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p186_s1.

PoncedeLeon:1988:GRS

[1369] J. Ponce de León. Gravitational repulsion in sources of the Reissner– Nordström field. Journal of Mathematical Physics, 29(1):197–206, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p197_s1.

Pantic:1988:SCD

[1370] Dražen Pantić. Stochastic calculus on distorted Brownian motion. Journal of Mathematical Physics, 29(1):207–209, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i1/p207_s1.

Oevel:1988:DCF

[1371] Walter Oevel. Dirac constraints in field theory: lifts of Hamiltonian systems to the cotangent bundle. Journal of Mathematical Physics, 29(1):210–219, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i1/p210_s1.

Oh:1988:ASY

[1372] C. H. Oh and Rajesh R. Parwani. Analytic solutions of the SU(3) Yang– Mills field equations with external sources. Journal of Mathematical Physics, 29(1):220–223, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i1/p220_s1.

Carter:1988:DOT

[1373] Brandon Carter. Duality and orthogonal transitivity in dimensional reduction. Journal of Mathematical Physics, 29(1):224–234, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p224_s1.

Malec:1988:ASS

[1374] Edward Malec. The absence of small solutions of nonlinear field equations. Journal of Mathematical Physics, 29(1):235–237, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i1/
p235_s1.

Papachristou:1988:NSB

[1375] C. J. Papachristou and B. Kent Harrison. Nonlocal symmetries and Bäcklund transformations for the self-dual Yang–Mills system. Journal of Mathematical Physics, 29(1):238–243, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i1/p238_s1.

Nakamura:1988:TGA

[1376] Yoshimasa Nakamura. Transformation group acting on a self-dual Yang– Mills hierarchy. Journal of Mathematical Physics, 29(1):244–248, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i1/p244_s1.

Thaller:1988:NFA

[1377] Bernd Thaller. Normal forms of an abstract Dirac operator and applications to scattering theory. Journal of Mathematical Physics, 29(1):249– 257, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i1/p249_s1.

Lord:1988:GTG

[1378] Eric A. Lord and P. Goswami. Gauge theory of a group of diffeomorphisms. III. The fiber bundle description. Journal of Mathematical Physics, 29(1):258–267, January 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i1/p258_s1.

Barut:1988:CPE

[1379] A. O. Barut, R. A. Moore, and T. C. Scott. Comment on the paper: “An exactly soluble relativistic quantum two-fermion problem”, [J. Math. Phys. 27 (1986), no. 12, 3055–3060; MR 88a:81238] by Barut and N. Ünal. Journal of Mathematical Physics, 29(1):268, January 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i1/p268_

s1. See [915].

Cougo-Pinto:1988:CTV

[1380] M. V. Cougo-Pinto. The construction of B_{l}^{(1)} in terms of vertex operators and a Lie-admissible not associative algebra. Journal of Mathematical Physics, 29(2):275–280, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p275_s1.

Englert:1988:LAS

[1381] F. Englert, A. Sevrin, W. Troost, A. Van Proeyen, and Ph. Spindel. Loop algebras and superalgebras based on S^{7}. Journal of Mathematical Physics, 29(2):281–286, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p281_s1.

Rowe:1988:VCS

[1382] D. J. Rowe, R. Le Blanc, and K. T. Hecht. Vector coherent state theory and its application to the orthogonal groups. Journal of Mathematical Physics, 29(2):287–304, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p287_s1.

Wolf:1988:LAS

[1383] Fritz Wolf. Lie algebraic solutions of linear Fokker–Planck equations. Journal of Mathematical Physics, 29(2):305–307, February 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/p305_
s1.

Zheng:1988:GTH

[1384] Yu Kun Zheng and W. L. Chan. Gauge transformation and the higher order Korteweg–de Vries equation. Journal of Mathematical Physics, 29(2):308–314, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p308_s1.

Colombeau:1988:MDE

[1385] J. F. Colombeau and A. Y. Le Roux. Multiplications of distributions in elasticity and hydrodynamics. Journal of Mathematical Physics, 29(2):315–319, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p315_s1.

Halilsoy:1988:SST

[1386] M. Halilsoy. Studies in space–times admitting two spacelike Killing vectors. Journal of Mathematical Physics, 29(2):320–326, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/
p320_s1.

Bona:1988:CST

[1387] C. Bona and B. Coll. Classification of the space–times admitting a constant curvature synchronization. Journal of Mathematical Physics, 29(2):327–331, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p327_s1.

DeAngelis:1988:ADF

[1388] Gian Fabrizio De Angelis, Diego de Falco, and Glauco Di Genova. Asymptotic decay of the free field covariance on a Riemannian manifold via heat kernel techniques. Journal of Mathematical Physics, 29(2):332– 335, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i2/p332_s1.

Szmigielski:1988:IDH

[1389] J. Szmigielski. Infinite-dimensional homogeneous manifolds with translational symmetry and nonlinear partial differential equations. Journal of Mathematical Physics, 29(2):336–346, February 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i2/p336_s1.

Xu:1988:LCA

[1390] Bang Qing Xu, Yong Jiu Wang, and Chang Gui Shao. Lax conjecture about the eigenspeed of the KdV equation. Journal of Mathematical Physics, 29(2):347–352, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p347_s1.

Cacuci:1988:DPC

[1391] D. G. Cacuci, R. B. Perez, and V. Protopopescu. Duals and propagators: a canonical formalism for nonlinear equations. Journal of Mathematical Physics, 29(2):353–361, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p353_s1.

Ruggeri:1988:DBR

[1392] G. J. Ruggeri. Dirac bracket revisited. Journal of Mathematical Physics, 29(2):362–364, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p362_s1.

Tateno:1988:BDS

[1393] H. Tateno. Behavior of dynamic soliton solutions in nonintegrable extended Klein–Gordon systems by means of a state plane technique. Journal of Mathematical Physics, 29(2):365–385, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/p365_s1.

Ward:1988:SSI

[1394] R. S. Ward. Soliton solutions in an integrable chiral model in 2 + 1 dimensions. Journal of Mathematical Physics, 29(2):386–389, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p386_s1.

Easwaran:1988:EPR

[1395] C. V. Easwaran and S. R. Majumdar. The evolution of perturbations of the renormalized long wave equation. Journal of Mathematical Physics, 29(2):390–392, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p390_s1.

Krause:1988:GQK

[1396] Jorge Krause. Galilean quantum kinematics. Journal of Mathematical Physics, 29(2):393–397, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i2/p393_s1.

Tikochinsky:1988:GMA

[1397] Yoel Tikochinsky. On the generalized multiplication and addition of complex numbers. Journal of Mathematical Physics, 29(2):398–399, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p398_s1.

Wang:1988:ODQ

[1398] Yu Peng Wang and Fu-Cho Pu. One-dimensional quantum three-wave interaction model with two fermionic fields and one bosonic field. Journal of Mathematical Physics, 29(2):400–402, February 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i2/p400_s1.

Kamran:1988:CCS

[1399] N. Kamran, M. Légaré, R. G. McLenaghan, and P. Winternitz. The classification of complete sets of operators commuting with the Dirac operator in Minkowski space–time. Journal of Mathematical Physics, 29(2):403–411, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p403_s1.

Robinson:1988:SLQ

[1400] Sam L. Robinson. The semiclassical limit of quantum dynamics. I. Time evolution. Journal of Mathematical Physics, 29(2):412–419, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p412_s1.

Hall:1988:CSG

[1401] G. S. Hall and W. Kay. Curvature structure in general relativity. Journal of Mathematical Physics, 29(2):420–427, February 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i2/p420_s1.

Hall:1988:HGG

[1402] G. S. Hall and W. Kay. Holonomy groups in general relativity. Journal of Mathematical Physics, 29(2):428–432, February 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i2/p428_s1.

Kuang:1988:GBC

[1403] Zhi Quan Kuang and Can Bin Liang. On the GKP and BS constructions of the c-boundary. Journal of Mathematical Physics, 29(2):433– 435, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i2/p433_s1.

Lindblom:1988:SUD

[1404] Lee Lindblom. Static uniform-density stars must be spherical in general relativity. Journal of Mathematical Physics, 29(2):436–439, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p436_s1.

Kupeli:1988:VSA

[1405] A. H. Kupeli. Vacuum solutions admitting a geodesic null congruence with shear proportional to expansion. Journal of Mathematical Physics, 29(2):440–442, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p440_s1.

Hiscock:1988:GRF

[1406] William A. Hiscock. General relativistic fluid spheres with nonzero vacuum energy density. Journal of Mathematical Physics, 29(2):443–445, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i2/p443_s1.

Krisch:1988:CVZ

[1407] J. P. Krisch. On the classification of vacuum zero Simon tensor solutions in relativity. Journal of Mathematical Physics, 29(2):446–448, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p446_s1.

Ram:1988:BTS

[1408] Shri Ram. Bianchi type VI_{0} space–times with perfect fluid source. Journal of Mathematical Physics, 29(2):449–450, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/p449_s1.

Kondo:1988:CIM

[1409] Kei ichi Kondo and Yūki Sugiyama. Correlation inequalities for multicomponent ferromagnets. Journal of Mathematical Physics, 29(2):451– 459, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i2/p451_s1.

Castagnino:1988:PMS

[1410] Mario Castagnino and Carlos Laciana. On a parametrization method for the semiclassical vacuum problem. I. Journal of Mathematical Physics, 29(2):460–475, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i2/p460_s1.

Inoue:1988:INC

[1411] Kazuhiko Inoue, Kazuhiko Odaka, and Minoru Omote. Infinite number of conserved quantities and extended conformal algebra in the Thirring model. Journal of Mathematical Physics, 29(2):476–480, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/
p476_s1.

Chakrabarti:1988:SPS

[1412] A. Chakrabarti. Spinors in periodic self-dual gauge field backgrounds. Journal of Mathematical Physics, 29(2):481–486, February 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/p481_
s1.

Carroll:1988:DTS

[1413] C. E. Carroll and F. T. Hioe. Driven three-state model and its analytic solutions. Journal of Mathematical Physics, 29(2):487–509, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i2/p487_s1.

Rai:1988:CSCa

[1414] Jagdish Rai, C. L. Mehta, and N. Mukunda. Correlated states and collective transition operators for multilevel atomic systems. I. Journal of Mathematical Physics, 29(2):510–515, February 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i2/p510_s1.

Kramer:1988:SGT

[1415] Peter Kramer. Space-group theory for a nonperiodic icosahedral quasilattice. Journal of Mathematical Physics, 29(2):516–524, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/
p516_s1.

Bokhari:1988:ESS

[1416] Ashfaque H. Bokhari and Asghar Qadir. Erratum: “Symmetries of static, spherically symmetric space–times”, [J. Math. Phys. 28, 1019 (1987)]. Journal of Mathematical Physics, 29(2):525, February 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i2/p525_s1. See

[1063, 2132].

Plebanski:1988:GQA

[1417] J. F. Plebański and M. Przanowski. Generalizations of the quaternion algebra and Lie algebras. Journal of Mathematical Physics, 29(3):529– 535, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i3/p529_s1.

Epifanio:1988:PAM

[1418] Giuseppina Epifanio and Camillo Trapani. Partial ^{∗}-algebras of matrices and operators. Journal of Mathematical Physics, 29(3):536–540, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i3/p536_s1.

Carinena:1988:NGT

[1419] J. F. Cariñena and Luis A. Ibort. Noncanonical groups of transformations, anomalies, and cohomology. Journal of Mathematical Physics, 29(3):541–545, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i3/p541_s1.

Blank:1988:IRL

[1420] J. Blank and M. Havlíček. Irreducible ∗-representations of the Lie superalgebras B(0, n) with finite-degenerated vacuum. II. Journal of Mathematical Physics, 29(3):546–559, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p546_s1.

Kwon:1988:COM

[1421] P. Kwon and M. Villasante. Casimir operators for massless representations of the super-Poincaré algebra and the reduction of the tendimensional massless scalar superfield. Journal of Mathematical Physics, 29(3):560–571, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i3/p560_s1.

Rowe:1988:DSE

[1422] D. J. Rowe, P. Rochford, and J. Repka. Dynamic structure and embedded representation in physics: the group theory of the adiabatic approximation. Journal of Mathematical Physics, 29(3):572–577, March 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/p572_
s1.

Dieckmann:1988:CSG

[1423] Jan Dieckmann. Cauchy surfaces in a globally hyperbolic space–time. Journal of Mathematical Physics, 29(3):578–579, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/p578_s1.

Dabrowski:1988:DOP

[1424] L. Dabrowski and R. Percacci. Diffeomorphisms, orientation, and Pin structures in two dimensions. Journal of Mathematical Physics, 29(3): 580–593, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i3/p580_s1.

Arik:1988:GMS

[1425] M. Arik and F. Neyzi. Grassmannian σ models and strings. Journal of Mathematical Physics, 29(3):594–598, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i3/p594_s1.

Wille:1988:NCF

[1426] L. T. Wille. A novel class of G function integrals. Journal of Mathematical Physics, 29(3):599–603, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p599_s1.

Fokas:1988:BHF

[1427] A. S. Fokas and P. M. Santini. Bi-Hamiltonian formulation of the Kadomtsev–Petviashvili and Benjamin–Ono equations. Journal of Mathematical Physics, 29(3):604–617, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p604_s1.

Gonzalez-Gascon:1988:IPC

[1428] F. González-Gascón and A. González-López. The inverse problem concerning symmetries of ordinary differential equations. Journal of Mathematical Physics, 29(3):618–621, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p618_s1.

Broadbridge:1988:IFO

[1429] P. Broadbridge. Integrable forms of the one-dimensional flow equation for unsaturated heterogeneous porous media. Journal of Mathematical Physics, 29(3):622–627, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i3/p622_s1.

Hietarinta:1988:SBE

[1430] Jarmo Hietarinta. A search for bilinear equations passing Hirota’s threesoliton condition. IV. Complex bilinear equations. Journal of Mathematical Physics, 29(3):628–635, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p628_s1.

Derrick:1988:COS

[1431] G. H. Derrick. Complete orthonormal sets on the past light cone. I. Functions belonging to mass zero and helicity s. Journal of Mathematical Physics, 29(3):636–641, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i3/p636_s1.

Papiez:1988:QSE

[1432] L. Papież, T. A. Osborn, and F. H. Molzahn. Quantum systems with external electromagnetic fields: the large mass asymptotics. Journal of Mathematical Physics, 29(3):642–659, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i3/p642_s1.

Shabani:1988:FMI

[1433] J. Shabani. Finitely many δ interactions with supports on concentric spheres. Journal of Mathematical Physics, 29(3):660–664, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/
p660_s1.

Patera:1988:PMD

[1434] J. Patera and H. Zassenhaus. The Pauli matrices in n dimensions and finest gradings of simple Lie algebras of type A_{n−1}. Journal of Mathematical Physics, 29(3):665–673, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i3/p665_s1.

Persides:1988:SNS

[1435] Sotirios Persides and Basilis C. Xanthopoulos. Some new stationary axisymmetric asymptotically flat space–times obtained from Painlevé transcendents. Journal of Mathematical Physics, 29(3):674–680, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i3/p674_s1.

Ernst:1988:CGP

[1436] Frederick J. Ernst, Alberto García D., and Isidore Hauser. Colliding gravitational plane waves with noncollinear polarization. III. Journal of Mathematical Physics, 29(3):681–689, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i3/p681_s1.

Taub:1988:CPI

[1437] A. H. Taub. On the collision of planar impulsive gravitational waves. Journal of Mathematical Physics, 29(3):690–695, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/p690_s1.

Ito:1988:MFO

[1438] H. M. Ito. Multiplication formulas of orthogonal polynomials of boson field operators: derivation based on the generalized phase-space method. Journal of Mathematical Physics, 29(3):696–701, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/p696_s1.

Ostrovsky:1988:CEA

[1439] A. A. Ostrovsky and G. A. Vilkovisky. The covariant effective action in QED. One-loop magnetic moment. Journal of Mathematical Physics, 29(3):702–708, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i3/p702_s1.

Goldstein:1988:SPT

[1440] Jerome A. Goldstein and Gisèle Ruiz Rieder. Spin-polarized Thomas– Fermi theory. Journal of Mathematical Physics, 29(3):709–716, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i3/p709_s1.

Handy:1988:MPF

[1441] Carlos R. Handy, Daniel Bessis, and Robert M. Williams. Moment problem formulation of the simplified ideal magnetohydrodynamics ballooning equation. Journal of Mathematical Physics, 29(3):717–720, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i3/p717_s1.

Yeh:1988:EDW

[1442] C. Yeh and C. S. Kim. Erratum: “The diffraction of waves by a penetrable ribbon”, [J. Math. Phys. 4 (1963), 65–71; MR 26 #1023] by Yeh. Journal of Mathematical Physics, 29(3):721–722, March 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i3/p721_s1. See

[1].

Braden:1988:ACT

[1443] H. W. Braden. Anomaly coefficients: their calculation and congruences. Journal of Mathematical Physics, 29(4):727–741, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p727_s1.

Iosifescu:1988:SDK

[1444] Mircea Iosifescu and Horia Scutaru. Second-degree kinematical constraints associated with dynamical symmetries. Journal of Mathematical Physics, 29(4):742–757, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p742_s1.

LeBlanc:1988:MRR

[1445] R. Le Blanc and D. J. Rowe. The matrix representations of g_{2}. I. Representations in an so(4) basis. Journal of Mathematical Physics, 29(4): 758–766, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i4/p758_s1.

LeBlanc:1988:MRI

[1446] R. Le Blanc and D. J. Rowe. The matrix representations of g_{2}. II. Representations in an su(3) basis. Journal of Mathematical Physics, 29(4): 767–776, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i4/p767_s1.

Michel:1988:DTS

[1447] L. Michel, J. Patera, and R. T. Sharp. The Demazure–Tits subgroup of a simple Lie group. Journal of Mathematical Physics, 29(4):777– 796, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i4/p777_s1.

deAzcarraga:1988:SGA

[1448] José A. de Azcárraga and Jerzy Lukierski. Superalgebras with Grassmann-algebra-valued structure constants from superfields. Journal of Mathematical Physics, 29(4):797–801, April 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i4/p797_s1.

Vanthournout:1988:STO

[1449] J. Vanthournout, J. Van der Jeugt, H. De Meyer, and G. Vanden Berghe. On the spectrum of a third-order SO(3) scalar in the enveloping algebra of SO(6). Journal of Mathematical Physics, 29(4):802–805, April 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p802_
s1.

Bluman:1988:NCS

[1450] George W. Bluman, Gregory J. Reid, and Sukeyuki Kumei. New classes of symmetries for partial differential equations. Journal of Mathematical Physics, 29(4):806–811, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p806_s1. See erratum [1660].

Caviglia:1988:CVP

[1451] G. Caviglia. Composite variational principles and the determination of conservation laws. Journal of Mathematical Physics, 29(4):812–816, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p812_s1.

Hofstee:1988:FTD

[1452] P. Hofstee and A. Tip. Fourier transforms and the dilatation group. Journal of Mathematical Physics, 29(4):817–822, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p817_s1.

Matori:1988:TVS

[1453] Yoshihiro Matori. Totally vicious space–times and reflectivity. Journal of Mathematical Physics, 29(4):823–824, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i4/p823_s1.

Ozsvath:1988:EP

[1454] István Ozsváth. An embedding problem. Journal of Mathematical Physics, 29(4):825–834, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p825_s1.

Dassios:1988:HRC

[1455] George Dassios. On the harmonic radius and the capacity of an inverse ellipsoid. Journal of Mathematical Physics, 29(4):835–836, April 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p835_
s1.

Soderberg:1988:CMR

[1456] B. Söderberg. Circle maps and reciprocal winding numbers. Journal of Mathematical Physics, 29(4):837–842, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i4/p837_s1.

Chanda:1988:PTC

[1457] Pranab K. Chanda and A. Roy Chowdhury. On a Painlevé test of a coupled system of Boussinesq and Schrödinger equations. Journal of Mathematical Physics, 29(4):843–850, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i4/p843_s1.

Ogura:1988:SWR

[1458] Hisanao Ogura and Hiroshi Nakayama. Scattering of waves from a random cylindrical surface. Journal of Mathematical Physics, 29(4):851– 860, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i4/p851_s1.

Sabatier:1988:MDM

[1459] P. C. Sabatier and B. Dolveck-Guilpard. On modeling discontinuous media. One-dimensional approximations. Journal of Mathematical Physics, 29(4):861–868, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p861_s1.

Gracia-Bondia:1988:ADS

[1460] José M. Gracia-Bondía and Joseph C. Várilly. Algebras of distributions suitable for phase-space quantum mechanics. I. Journal of Mathematical Physics, 29(4):869–879, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p869_s1.

Varilly:1988:ADS

[1461] Joseph C. Várilly and José M. Gracia-Bondía. Algebras of distributions suitable for phase-space quantum mechanics. II. Topologies on the Moyal algebra. Journal of Mathematical Physics, 29(4):880–887, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p880_s1.

Zagrebnov:1988:TLP

[1462] V. A. Zagrebnov. The Trotter–Lie product formula for Gibbs semigroups. Journal of Mathematical Physics, 29(4):888–891, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p888_s1.

Dutta:1988:EAS

[1463] A. K. Dutta and R. S. Willey. Exact analytic solutions for the quantum mechanical sextic anharmonic oscillator. Journal of Mathematical Physics, 29(4):892–900, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p892_s1.

Vrscay:1988:NSI

[1464] Edward R. Vrscay. Nonlinear “self-interaction” Hamiltonians of the form H^{(0)} + λ⟨r^{p}⟩r^{q} and their Rayleigh–Schrödinger perturbation expansions. Journal of Mathematical Physics, 29(4):901–911, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p901_s1.

Froman:1988:CPW

[1465] Nanny Fröman and Per Olof Fröman. Comments on the paper: “On the WKBJ approximation”, [J. Math. Phys. 28 (1987), no. 3, 556–558; MR 88e:81020 ] by M. El Sawi. Journal of Mathematical Physics, 29 (4):912, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i4/p912_s1. See [992].

Shiekh:1988:CTQ

[1466] A. Y. Shiekh. Canonical transformations in quantum mechanics: a canonically invariant path integral. Journal of Mathematical Physics, 29(4):913–930, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p913_s1.

Yngve:1988:ELC

[1467] Staffan Yngve. The energy levels and the corresponding normalized wave functions for a model of a compressed atom. II. Journal of Mathematical Physics, 29(4):931–936, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p931_s1.

Martin-Pascual:1988:PTI

[1468] F. Martín-Pascual and J. M. M. Senovilla. Petrov types D and II perfectfluid solutions in generalized Kerr–Schild form. Journal of Mathematical Physics, 29(4):937–944, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p937_s1.

Sussman:1988:SSSa

[1469] Roberto A. Sussman. On spherically symmetric shear-free perfect fluid configurations (neutral and charged). II. Equation of state and singularities. Journal of Mathematical Physics, 29(4):945–970, April 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p945_
s1.

TorresdelCastillo:1988:SME

[1470] G. F. Torres del Castillo. The separability of Maxwell’s equations in type-D backgrounds. Journal of Mathematical Physics, 29(4):971–977, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p971_s1.

Holm:1988:HSU

[1471] Christian Holm. The hyperspin structure of unitary groups. Journal of Mathematical Physics, 29(4):978–986, April 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i4/p978_s1. See erra-

tum [2053].

Majorana:1988:FME

[1472] A. Majorana. Finite moment equations for a relativistic simple gas. Journal of Mathematical Physics, 29(4):987–989, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p987_s1.

Hall:1988:GSE

[1473] Richard L. Hall. The ground-state energy of a system of identical bosons. Journal of Mathematical Physics, 29(4):990–994, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p990_s1.

Levermore:1988:RTT

[1474] C. D. Levermore, J. Wong, and G. C. Pomraning. Renewal theory for transport processes in binary statistical mixtures. Journal of Mathematical Physics, 29(4):995–1004, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i4/p995_s1.

Mason:1988:BTA

[1475] L. Mason, S. Chakravarty, and E. T. Newman. Bäcklund transformations for the anti-self-dual Yang–Mills equations. Journal of Mathematical Physics, 29(4):1005–1013, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i4/p1005_s1.

Thomi:1988:RBQ

[1476] P. Thomi. A remark on BRST quantization. Journal of Mathematical Physics, 29(4):1014–1021, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i4/p1014_s1.

Cahen:1988:YME

[1477] M. Cahen, S. Gutt, C. Kozameh, and E. T. Newman. Yang–Mills equations and solvable groups. II. Journal of Mathematical Physics, 29(4):1022–1025, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i4/p1022_s1.

Lhallabi:1988:SWI

[1478] T. Lhallabi. N = 2 supersymmetric Ward identities on the harmonic superspace. Journal of Mathematical Physics, 29(4):1026–1031, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p1026_s1.

Suzuki:1988:DLC

[1479] A. T. Suzuki. The 2ω-dimensional light-cone integrals with momentum shift. Journal of Mathematical Physics, 29(4):1032–1033, April 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i4/p1032_
s1.

Jaskolski:1988:GAP

[1480] Z. Jaskólski. On the Gribov ambiguity in the Polyakov string. Journal of Mathematical Physics, 29(4):1034–1043, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i4/p1034_s1.

Cavalcante:1988:CLN

[1481] J. A. Cavalcante and K. Tenenblat. Conservation laws for nonlinear evolution equations. Journal of Mathematical Physics, 29(4):1044–1049, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p1044_s1.

Kersten:1988:NSL

[1482] Paul H. M. Kersten. Nonlocal symmetries and the linearization of the massless Thirring and the Federbush models. Journal of Mathematical Physics, 29(4):1050–1053, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i4/p1050_s1.

Amigo:1988:APL

[1483] J. M. Amigó. All possible local charges in a local quantum field theory: massive case. Journal of Mathematical Physics, 29(4):1054–1060, April 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i4/p1054_s1.

Sharma:1988:CSR

[1484] C. S. Sharma. Complex structure on a real Hilbert space and symplectic structure on a complex Hilbert space. Journal of Mathematical Physics, 29(5):1069–1078, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i5/p1069_s1.

Fecko:1988:R

[1485] Marián Fecko. A representation of SO(p+q, p+q) on SO(p, q). Journal of Mathematical Physics, 29(5):1079–1082, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i5/p1079_s1.

Garcia-Prada:1988:LOR

[1486] Oscar García-Prada, Mariano A. del Olmo, and Mariano Santander. Locally operating realizations of nonconnected transformation Lie groups. Journal of Mathematical Physics, 29(5):1083–1090, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1083_s1.

Siddall:1988:SSE

[1487] T. H. Siddall and J. J. Sullivan. Substitutional symmetries and extra degeneracies of real symmetric matrices. Journal of Mathematical Physics, 29(5):1091–1096, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i5/p1091_s1.

GonzalezLopez:1988:SLS

[1488] A. González López. Symmetries of linear systems of second-order ordinary differential equations. Journal of Mathematical Physics, 29(5): 1097–1105, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1097_s1.

Louck:1988:DFC

[1489] J. D. Louck, L. C. Biedenharn, and M. A. Lohe. On the denominator function for canonical SU(3) tensor operators. II. Explicit polynomial form. Journal of Mathematical Physics, 29(5):1106–1117, May 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1106_
s1.

Boenkost:1988:IBS

[1490] W. Boenkost, F. Constantinescu, and U. Scharffenberger. The inverse of a Borel summable function. Journal of Mathematical Physics, 29(5): 1118–1121, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1118_s1.

Buchdahl:1988:CFH

[1491] H. A. Buchdahl. Characteristic functions and higher order Lagrangians. Journal of Mathematical Physics, 29(5):1122–1126, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1122_s1.

Calvao:1988:NCH

[1492] M. O. Calvão, M. J. Rebouças, A. F. F. Teixeira, and W. M. Silva, Jr. Notes on a class of homogeneous space–times. Journal of Mathematical Physics, 29(5):1127–1129, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i5/p1127_s1.

Klein:1988:APT

[1493] David Klein and Wei-Shih Yang. Absence of phase transitions for continuum models of dimension d > 1. Journal of Mathematical Physics, 29(5): 1130–1133, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1130_s1.

Carinena:1988:GLA

[1494] José F. Cariñena, Carlos López, and Manuel F. Rañada. Geometric Lagrangian approach to first-order systems and applications. Journal of Mathematical Physics, 29(5):1134–1142, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i5/p1134_s1.

Carinena:1988:OLC

[1495] José F. Cariñena, Carlos López, and Narciso Román-Roy. Origin of the Lagrangian constraints and their relation with the Hamiltonian formulation. Journal of Mathematical Physics, 29(5):1143–1149, May 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1143_
s1.

Francoise:1988:SGI

[1496] J.-P. Françoise. Symplectic geometry and integrable m-body problems on the line. Journal of Mathematical Physics, 29(5):1150–1153, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i5/p1150_s1.

Oh:1988:PGA

[1497] C. H. Oh, C. P. Soo, and C. H. Lai. The propagator in the generalized Aharonov–Bohm effect. Journal of Mathematical Physics, 29(5):1154– 1157, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1154_s1.

Alicki:1988:SPP

[1498] Robert Alicki, Sławomir Rudnicki, and Sławomir Sadowski. Symmetry properties of product states for the system of N n-level atoms. Journal of Mathematical Physics, 29(5):1158–1162, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i5/p1158_s1.

Bohm:1988:RTS

[1499] A. Bohm, M. Kmiecik, and L. J. Boya. Representation theory of superconformal quantum mechanics. Journal of Mathematical Physics, 29(5): 1163–1170, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1163_s1.

Derezinski:1988:TST

[1500] Jan Dereziński. Threshold singularities of two-cluster-two-cluster scattering amplitudes for dilation analytic potentials. Journal of Mathematical Physics, 29(5):1171–1173, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i5/p1171_s1.

Yoon:1988:SCT

[1501] Sung Hyun Yoon. Scattering-into-cones theorem for a certain impurity scattering. Journal of Mathematical Physics, 29(5):1174–1176, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i5/p1174_s1.

Sussman:1988:SSSb

[1502] Roberto A. Sussman. On spherically symmetric shear-free perfect fluid configurations (neutral and charged). III. Global view. Journal of Mathematical Physics, 29(5):1177–1211, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i5/p1177_s1.

OBrien:1988:SCG

[1503] G. M. O’Brien and D. H. Tchrakian. Spin-connection generalized Yang– Mills fields on double-dual generalized Einstein–Cartan backgrounds. Journal of Mathematical Physics, 29(5):1212–1219, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1212_s1.

vanKampen:1988:ECM

[1504] N. G. van Kampen. Explicit calculation of a model for diffusion in nonconstant temperature. Journal of Mathematical Physics, 29(5):1220– 1224, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1220_s1.

Baumann:1988:CIQ

[1505] Klaus Baumann. On canonical irreducible quantum field theories describing bosons and fermions. Journal of Mathematical Physics, 29(5): 1225–1230, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i5/p1225_s1.

Jakobczyk:1988:KSW

[1506] L. Jakóbczyk and F. Strocchi. Kreĭn structures for Wightman and Schwinger functions. Journal of Mathematical Physics, 29(5):1231–1235, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i5/p1231_s1.

Ghosh:1988:SFT

[1507] Pradip Ghosh and Pratul Bandyopadhyay. Stochastic field theory and finite-temperature supersymmetry. Journal of Mathematical Physics, 29(5):1236–1240, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i5/p1236_s1.

Zhang:1988:VAG

[1508] Yao Zhong Zhang. The Virasoro algebra and group in four dimensions. Journal of Mathematical Physics, 29(5):1241–1243, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1241_s1.

Ketov:1988:ISS

[1509] S. V. Ketov and B. B. Lokhvitsky. N = 4 irreducible scalar superfields. Journal of Mathematical Physics, 29(5):1244–1252, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i5/p1244_s1.

Lopuszanski:1988:EWW

[1510] Jan Łopuszański. Extended Weinberg–Witten theorem. Journal of Mathematical Physics, 29(5):1253–1257, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i5/p1253_s1.

Kiessling:1988:EUP

[1511] Michael Kiessling. Existence and uniqueness of the pressure profile p(A) of a current-carrying quasineutral equilibrium plasma. Journal of Mathematical Physics, 29(5):1258–1263, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i5/p1258_s1.

Schmidt:1988:AGL

[1512] Hans-Jürgen Schmidt. Addendum: “On Geroch’s limit of space–times and its relation to a new topology in the space of Lie groups”, [J. Math. Phys. 28 (1987), no. 8, 1928–1936; MR 88i:22042]. Journal of Mathematical Physics, 29(5):1264, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i5/p1264_s1. See [1194].

Salingaros:1988:ELG

[1513] N. A. Salingaros. Erratum: “The Lorentz group and the Thomas precession. II. Exact results for the product of two boosts”, [J. Math. Phys. 27 (1986), no. 1, 157–162; MR 87g:78009]. Journal of Mathematical Physics, 29(5):1265, May 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i5/p1265_s1. See [510, 984].

Gunaydin:1988:UHW

[1514] Murat Günaydin. Unitary highest weight representations of noncompact supergroups. Journal of Mathematical Physics, 29(6):1275–1282, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i6/p1275_s1.

Palmer:1988:GCI

[1515] John Palmer. A Grassmann calculus for infinite spin groups. Journal of Mathematical Physics, 29(6):1283–1299, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i6/p1283_s1.

Stovicek:1988:GMF

[1516] P. Štovíček. Gell-Mann formula for simple complex Lie groups and geometric quantization. Journal of Mathematical Physics, 29(6):1300–1307, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i6/p1300_s1.

Bergvelt:1988:FZC

[1517] M. J. Bergvelt and A. P. E. ten Kroode. τ functions and zero curvature equations of Toda–AKNS type. Journal of Mathematical Physics, 29(6): 1308–1320, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1308_s1.

Ma:1988:IES

[1518] Zhong Qi Ma and An Ying Dai. Integral equations with symmetrical kernel applied to a system with a Dirac-type spectrum. Journal of Mathematical Physics, 29(6):1321–1324, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i6/p1321_s1.

Iwai:1988:GSQ

[1519] Toshihiro Iwai. A geometric setting for the quantum planar n-body system, and a U(n − 1) basis for the internal states. Journal of Mathematical Physics, 29(6):1325–1337, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i6/p1325_s1.

Marshall:1988:WHS

[1520] Ian Marshall and Stefan Wojciechowski. When is a Hamiltonian system separable? Journal of Mathematical Physics, 29(6):1338–1346, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i6/p1338_s1.

Rose:1988:GEE

[1521] James H. Rose and Margaret Cheney. Generalized eigenfunction expansions for scattering in inhomogeneous three-dimensional media. Journal of Mathematical Physics, 29(6):1347–1355, June 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i6/p1347_s1.

Hojman:1988:IRT

[1522] R. Hojman, G. Palma, L. Peña, and R. Tabensky. Interference in the radiation of two relativistic charges. Journal of Mathematical Physics, 29(6):

1356–1360, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1356_s1.

Lozada:1988:FMA

[1523] Abraham Lozada and P. L. Torres. Four-momentum and angular momentum in classical electrodynamics. Journal of Mathematical Physics, 29(6): 1361–1367, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1361_s1.

Novoselsky:1988:CFP

[1524] Akiva Novoselsky, Jacob Katriel, and Robert Gilmore. Coefficients of fractional parentage in the L-S coupling scheme. Journal of Mathematical Physics, 29(6):1368–1388, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i6/p1368_s1.

Blood:1988:RQM

[1525] F. A. Blood. A relativistic quantum mechanical harmonic oscillator without space–time variables. Journal of Mathematical Physics, 29(6):1389– 1395, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1389_s1.

Kondo:1988:TDE

[1526] Anne E. Kondo and D. Rodney Truax. On a time-dependent extension of the Morse potential. Journal of Mathematical Physics, 29(6):1396– 1404, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1396_s1.

Kummer:1988:NCO

[1527] Martin Kummer and Raghuramaiah Gompa. Nonlinearly coupled oscillators in quantum mechanics: a normal form approach. Journal of Mathematical Physics, 29(6):1405–1416, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i6/p1405_s1.

Meyer:1988:TDR

[1528] H.-D. Meyer, J. Kučar, and L. S. Cederbaum. Time-dependent rotated Hartree: formal development. Journal of Mathematical Physics, 29(6):

1417–1430, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1417_s1.

Ramm:1988:CZE

[1529] A. G. Ramm. Conditions for zero not to be an eigenvalue of the Schrödinger operator. II. Journal of Mathematical Physics, 29(6):1431– 1432, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1431_s1.

Znojil:1988:PSC

[1530] M. Znojil. On the power-series construction of bound states. I. the energies as zeros of the infinite Hill determinants. Journal of Mathematical Physics, 29(6):1433–1439, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i6/p1433_s1.

Bona:1988:NPG

[1531] C. Bona. A new proof of the generalized Birkhoff theorem. Journal of Mathematical Physics, 29(6):1440–1441, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i6/p1440_s1.

GarciaD:1988:MIC

[1532] Alberto García D. The \overline{S}(a, b, c/m) metrics interpreted as colliding wave solutions. Journal of Mathematical Physics, 29(6):1442, June 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i6/p1442_
s1.

Maharaj:1988:SST

[1533] S. D. Maharaj. On self-similar Tolman models. Journal of Mathematical Physics, 29(6):1443–1444, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i6/p1443_s1.

Papacostas:1988:HMS

[1534] Taxiarchis Papacostas. Hauser–Malhiot spaces admitting a perfect fluid energy-momentum tensor. Journal of Mathematical Physics, 29(6):1445– 1450, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1445_s1.

VandenBergh:1988:CRF

[1535] N. Van den Bergh. Conformally Ricci-flat perfect fluids. II. Journal of Mathematical Physics, 29(6):1451–1454, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i6/p1451_s1.

Lott:1988:RA

[1536] J. Lott. Real anomalies. Journal of Mathematical Physics, 29(6):1455– 1464, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1455_s1.

Arsenault:1988:CEI

[1537] Guy Arsenault, Michel Jacques, and Yvan Saint-Aubin. Collapse and exponentiation of infinite symmetry algebras of Euclidean projective and Grassmannian σ models. Journal of Mathematical Physics, 29(6):1465– 1471, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1465_s1.

Aldrovandi:1988:QPS

[1538] R. Aldrovandi and J. G. Pereira. On the quantization of Poincaré and de Sitter gauge models. Journal of Mathematical Physics, 29(6):1472– 1476, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1472_s1.

Ong:1988:SWB

[1539] Michael K. Ong. The solution of the s-wave Bethe–Goldstone equation with a standard hard core potential. Journal of Mathematical Physics, 29(6):1477–1485, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i6/p1477_s1.

Schwarzmeier:1988:VND

[1540] J. L. Schwarzmeier and H. R. Lewis. Variational nature of dispersion equations revisited. Journal of Mathematical Physics, 29(6):1486–1497, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i6/p1486_s1.

Lowen:1988:SPO

[1541] H. Löwen. Spectral properties of an optical polaron in a magnetic field. Journal of Mathematical Physics, 29(6):1498–1504, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i6/p1498_s1.

Lowen:1988:ABG

[1542] H. Löwen. Analytical behavior of the ground-state energy and pinning transitions for a bound polaron. Journal of Mathematical Physics, 29(6): 1505–1513, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i6/p1505_s1.

Gron:1988:VPS

[1543] Öyvind Grön and Harald H. Soleng. The Vaidya–Patel solution with Robertson–Walker metric as a rotating inflationary scenario. Journal of Mathematical Physics, 29(6):1514–1517, June 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i6/p1514_s1. See er-

ratum [1661].

Arimitsu:1988:ECF

[1544] T. Arimitsu, H. Umezawa, and Y. Yamanaka. Erratum: “Canonical formalism of dissipative fields in the thermo field dynamics”, [J. Math. Phys. 28, 2741 (1987)]. Journal of Mathematical Physics, 29(6):1518, June 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i6/p1518_s1. See [1305].

Fawcett:1988:SOU

[1545] R. J. B. Fawcett and A. J. Bracken. Simple orthogonal and unitary compact quantum systems and the inönü-wigner contraction. Journal of Mathematical Physics, 29(7):1521–1528, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i7/p1521_s1.

Tratnik:1988:MCH

[1546] M. V. Tratnik. Multivariable continuous Hahn polynomials. Journal of Mathematical Physics, 29(7):1529–1534, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i7/p1529_s1.

Reeh:1988:RCC

[1547] Helmut Reeh. A remark concerning canonical commutation relations. Journal of Mathematical Physics, 29(7):1535–1536, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/p1535_s1.

Chang:1988:EBP

[1548] Shau-Jin Chang and Richard Friedberg. Elliptical billiards and Poncelet’s theorem. Journal of Mathematical Physics, 29(7):1537–1550, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i7/p1537_s1.

Crampin:1988:ALS

[1549] M. Crampin and G. E. Prince. Alternative Lagrangians for spherically symmetric potentials. Journal of Mathematical Physics, 29(7):1551– 1554, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1551_s1.

Melzer:1988:STR

[1550] Ezer Melzer. N = 2 supertori and their representation as algebraic curves. Journal of Mathematical Physics, 29(7):1555–1568, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/
p1555_s1.

Rendall:1988:CGS

[1551] A. D. Rendall. Curvature of generic space–times in general relativity. Journal of Mathematical Physics, 29(7):1569–1574, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/p1569_s1.

Bryant:1988:S

[1552] Peter Bryant. GH^{∞} supermanifolds. Journal of Mathematical Physics, 29(7):1575–1579, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i7/p1575_s1.

Colegrave:1988:ITD

[1553] R. K. Colegrave and M. A. Mannan. Invariants for the time-dependent harmonic oscillator. Journal of Mathematical Physics, 29(7):1580–1587, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i7/p1580_s1.

Galvao:1988:QCG

[1554] Carlos A. P. Galvão and Nivaldo A. Lemos. On the quantization of constrained generalized dynamics. Journal of Mathematical Physics, 29(7): 1588–1592, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1588_s1.

Ragnisco:1988:ROB

[1555] O. Ragnisco and P. M. Santini. Recursion operator and bi-Hamiltonian structure for integrable multidimensional lattices. Journal of Mathematical Physics, 29(7):1593–1603, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i7/p1593_s1.

Vujanovic:1988:AHJ

[1556] B. Vujanović and A. M. Strauss. Application of the Hamilton–Jacobi method to linear nonconservative vibration theory. Journal of Mathematical Physics, 29(7):1604–1609, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i7/p1604_s1.

Olver:1988:HSS

[1557] Peter J. Olver and Yavuz Nutku. Hamiltonian structures for systems of hyperbolic conservation laws. Journal of Mathematical Physics, 29(7): 1610–1619, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1610_s1.

Sazdjian:1988:SPT

[1558] H. Sazdjian. The scalar product in two-particle relativistic quantum mechanics. Journal of Mathematical Physics, 29(7):1620–1633, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/
p1620_s1.

Balantekin:1988:CSH

[1559] A. B. Balantekin, H. A. Schmitt, and B. R. Barrett. Coherent states for the harmonic oscillator representations of the orthosymplectic supergroup Osp(1/2N, R). Journal of Mathematical Physics, 29(7):1634– 1639, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1634_s1.

Krisch:1988:CRS

[1560] J. P. Krisch and L. L. Smalley. Cartan and relativistic spin fluids in a rotating cylinder. Journal of Mathematical Physics, 29(7):1640–1644, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i7/p1640_s1.

Krori:1988:ESS

[1561] K. D. Krori, P. Borgohain, P. K. Kar, and Dipali Das (Kar). Exact scalar and spinor solutions in some rotating universes. Journal of Mathematical Physics, 29(7):1645–1649, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i7/p1645_s1.

Stephani:1988:SEF

[1562] H. Stephani. Symmetries of Einstein’s field equations with a perfect fluid source as examples of Lie–Bäcklund symmetries. Journal of Mathematical Physics, 29(7):1650–1654, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v29/i7/p1650_s1.

Moffat:1988:SFG

[1563] J. W. Moffat. Spinor fields and the GL(4, R) gauge structure in the nonsymmetric theory of gravitation. Journal of Mathematical Physics, 29(7): 1655–1660, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1655_s1.

Zardecki:1988:GGT

[1564] Andrew Zardecki. Gravity as a gauge theory with Cartan connection. Journal of Mathematical Physics, 29(7):1661–1666, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/p1661_s1.

Cornille:1988:ESD

[1565] Henri Cornille. Exact solutions for discrete kinetic models with ternary collisions. Journal of Mathematical Physics, 29(7):1667–1677, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/
p1667_s1.

Sinha:1988:TCF

[1566] Suresh K. Sinha. Theory for correlation functions of fermion quantum fluids. Journal of Mathematical Physics, 29(7):1678–1686, July 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i7/p1678_
s1.

Antoine:1988:SEM

[1567] J.-P. Antoine and B. Piette. Solutions of Euclidean σ models on noncompact Grassmann manifolds. Journal of Mathematical Physics, 29(7): 1687–1697, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1687_s1.

Heidenreich:1988:QCI

[1568] W. F. Heidenreich and M. Lorente. Quantization of conformally invariant Bargmann–Wigner equations with gauge freedom. Journal of Mathematical Physics, 29(7):1698–1704, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i7/p1698_s1.

Beckers:1988:HOS

[1569] J. Beckers, D. Dehin, and V. Hussin. On the Heisenberg and orthosymplectic superalgebras of the harmonic oscillator. Journal of Mathematical Physics, 29(7):1705–1711, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i7/p1705_s1.

Kooij:1988:TDS

[1570] Bert Jan Kooij and Dirk Quak. Three-dimensional scattering of impulsive acoustic waves by a semi-infinite crack in the plane interface of a half-space and a layer. Journal of Mathematical Physics, 29(7):1712– 1721, July 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i7/p1712_s1.

Bender:1988:POO

[1571] Carl M. Bender and Gerald V. Dunne. Polynomials and operator orderings. Journal of Mathematical Physics, 29(8):1727–1731, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/
p1727_s1.

Capps:1988:GCS

[1572] Richard H. Capps. Geometrical classification scheme for weights of Lie algebras. Journal of Mathematical Physics, 29(8):1732–1735, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1732_s1.

Cummins:1988:PII

[1573] C. J. Cummins and J. Patera. Polynomial icosahedral invariants. Journal of Mathematical Physics, 29(8):1736–1745, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1736_s1.

Aguirre:1988:GSTb

[1574] M. Aguirre and J. Krause. SL(3, R) as the group of symmetry transformations for all one-dimensional linear systems. II. Realizations of the Lie algebra. Journal of Mathematical Physics, 29(8):1746–1752, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1746_s1.

Baake:1988:BRU

[1575] Michael Baake. Branching rules for unitary representations of Virasoro and super-Virasoro algebras at c = 1. Journal of Mathematical Physics, 29(8):1753–1757, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1753_s1.

Braunss:1988:RST

[1576] G. Braunss and E. Teubner. On the representation of Slater-type densities by Gaussian densities. Journal of Mathematical Physics, 29(8):1758– 1760, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i8/p1758_s1.

Ditto:1988:NSN

[1577] William L. Ditto and Thomas J. Pickett. Nonperturbative solutions of nonlinear differential equations using continued fractions. Journal of Mathematical Physics, 29(8):1761–1770, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1761_s1.

Hillion:1988:SWE

[1578] P. Hillion. Solutions of the wave equations with boundary conditions on the hyperplane z − ct = 0. Journal of Mathematical Physics, 29(8):1771– 1775, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i8/p1771_s1.

Wilhelmsson:1988:SRD

[1579] Hans Wilhelmsson. Solutions of reaction-diffusion integral equations describing explosive evolution of densities for localized structures. Journal of Mathematical Physics, 29(8):1776–1779, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1776_s1.

Quan:1988:IDN

[1580] Xiang Lin Quan, Lian Ji Wang, and He Shan Song. Introducing division by an “a” number and a new “b” number in particle physics. Journal of Mathematical Physics, 29(8):1780–1785, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1780_s1.

Krolak:1988:ERP

[1581] Andrzej Królak. The existence of regular partially future asymptotically predictable space–times. Journal of Mathematical Physics, 29(8):1786– 1788, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i8/p1786_s1.

Bartocci:1988:CSS

[1582] Claudio Bartocci and Ugo Bruzzo. Cohomology of the structure sheaf of real and complex supermanifolds. Journal of Mathematical Physics, 29(8):1789–1794, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1789_s1. See erratum [1983].

Lundell:1988:GAA

[1583] Albert T. Lundell and Yasunari Tosa. Global anomalies and algebraic topology. Journal of Mathematical Physics, 29(8):1795–1806, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1795_s1.

Leach:1988:MSA

[1584] P. G. L. Leach and F. M. Mahomed. Maximal subalgebra associated with a first integral of a system possessing sl(3, R) algebra. Journal of Mathematical Physics, 29(8):1807–1813, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1807_s1.

Skorupski:1988:DBP

[1585] Andrzej A. Skorupski. Detailed behavior of the phase-integral approximations at zeros and singularities of the square of the base function. Journal of Mathematical Physics, 29(8):1814–1823, August 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1814_
s1.

Skorupski:1988:EIO

[1586] Andrzej A. Skorupski. Efficient integration of the one-dimensional time independent wave equation for bound states and for wave propagation. Journal of Mathematical Physics, 29(8):1824–1831, August 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1824_
s1.

Aguirregabiria:1988:EAM

[1587] J. M. Aguirregabiria and J. R. Etxebarria. Electromagnetic angular momentum radiated by two-point charges. Journal of Mathematical Physics, 29(8):1832–1836, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1832_s1.

Okolowski:1988:PSN

[1588] J. A. Okolowski and M. Slomiana. Particlelike solutions to nonlinear classical real scalar field theories. Journal of Mathematical Physics, 29(8):1837–1839, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1837_s1.

Bel:1988:RSE

[1589] L. Bel and E. Ruiz. Relativistic Schrödinger equations. Journal of Mathematical Physics, 29(8):1840–1846, August 1988. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i8/p1840_s1.

DiazBejarano:1988:GEC

[1590] J. Díaz Bejarano and A. Martín Sánchez. Generalized exponential, circular, and hyperbolic functions for nonlinear wave equations. Journal of Mathematical Physics, 29(8):1847–1853, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1847_s1.

Basu:1988:HWG

[1591] Debabrata Basu and T. K. Kar. The Heisenberg–Weyl group in the coherent state basis and the Bargmann transform. Journal of Mathematical Physics, 29(8):1854–1859, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i8/p1854_s1.

Nishioka:1988:TRR

[1592] Michio Nishioka. A theorem on the rate of regeneration in decay processes. Journal of Mathematical Physics, 29(8):1860–1861, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/
p1860_s1.

Ajanapon:1988:FQP

[1593] Pimon Ajanapon. Field quantization for pure and mixed states and its classical limit. Journal of Mathematical Physics, 29(8):1862–1865, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1862_s1.

Kundu:1988:MSS

[1594] Prasun K. Kundu. Multipole structure of stationary gravitational fields at null infinity. Journal of Mathematical Physics, 29(8):1866–1878, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1866_s1.

Nahmad-Achar:1988:GKS

[1595] E. Nahmad-Achar. On generalized Kerr–Schild transformations. Journal of Mathematical Physics, 29(8):1879–1884, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1879_s1.

Trapani:1988:SDQ

[1596] Camillo Trapani. States and derivations on quasi-∗-algebras. Journal of Mathematical Physics, 29(8):1885–1890, August 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1885_s1.

Choquet-Bruhat:1988:CPS

[1597] Yvonne Choquet-Bruhat. The Cauchy problem for stringy gravity. Journal of Mathematical Physics, 29(8):1891–1895, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i8/p1891_s1.

Chen:1988:NDP

[1598] Y. Y. Chen. The nonlinear desingularization phenomenon in the abelian Higgs model for isotropic solutions. Journal of Mathematical Physics, 29(8):1896–1900, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1896_s1.

Guadagnini:1988:LEM

[1599] E. Guadagnini. The loop effective measure for sigma models coupled with gravity. Journal of Mathematical Physics, 29(8):1901–1908, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1901_s1.

Bakas:1988:GGC

[1600] I. Bakas and D. McMullan. Gauge group cohomology in the monopole sector of Yang–Mills theories. Journal of Mathematical Physics, 29 (8):1909–1915, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1909_s1.

Embacher:1988:FBS

[1601] Franz Embacher. Free bosonic string field theory without supplementary fields. Journal of Mathematical Physics, 29(8):1916–1929, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i8/p1916_s1.

Hostler:1988:SEO

[1602] Levere Hostler. Self-energy operator for an electron in an external Coulomb potential. II. Journal of Mathematical Physics, 29(8):1930– 1936, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i8/p1930_s1.

Hogreve:1988:MDA

[1603] H. Hogreve. On the monotonic dependence of atomic excitation energies on the nuclear charge parameter Z. Journal of Mathematical Physics, 29(8):1937–1942, August 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i8/p1937_s1.

Saidi:1988:CCS

[1604] E. H. Saidi. SU(2) classification of N = 2 complex supersymmetric representations. Journal of Mathematical Physics, 29(9):1949–1957, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i9/p1949_s1.

Vanthournout:1988:SRM

[1605] J. Vanthournout, H. De Meyer, and G. Vanden Berghe. On the symmetric representations of SU(5). Matrix elements of the generators in the subgroup bases SU(2) × SU(2) × U(1) and SU(2) × SU(2). Journal of Mathematical Physics, 29(9):1958–1963, September 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i9/p1958_s1.

Manoranjan:1988:SAI

[1606] V. S. Manoranjan, T. Ortega, and J. M. Sanz-Serna. Soliton and antisoliton interactions in the “good” Boussinesq equation. Journal of Mathematical Physics, 29(9):1964–1968, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i9/p1964_s1.

Sachs:1988:CEI

[1607] Robert L. Sachs. On a class of eigenfunction identities for Hill operators. Journal of Mathematical Physics, 29(9):1969–1971, September 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p1969_
s1.

Collinson:1988:PAC

[1608] C. D. Collinson. Proper affine collineations in Robertson–Walker space– times. Journal of Mathematical Physics, 29(9):1972–1973, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i9/p1972_s1.

Helminck:1988:MHV

[1609] G. F. Helminck. Matrix hierarchies and vector bundles over Riemann surfaces. Journal of Mathematical Physics, 29(9):1974–1978, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i9/p1974_s1.

Bowick:1988:RC

[1610] M. J. Bowick and A. Lahiri. The Ricci curvature of Diff S^{1}/SL(2, R). Journal of Mathematical Physics, 29(9):1979–1981, September 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p1979_
s1.

Nakayama:1988:WAB

[1611] J. Nakayama and E. Omori. Wiener analysis of a binary hysteresis system. Journal of Mathematical Physics, 29(9):1982–1989, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i9/p1982_s1.

Kupershmidt:1988:LWE

[1612] B. A. Kupershmidt. Long wave equations and metrizable Lie algebras. Journal of Mathematical Physics, 29(9):1990–2000, September 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p1990_
s1.

Lim:1988:TFH

[1613] Paul H. Lim. Tensor formulation of Hamilton’s equations. Journal of Mathematical Physics, 29(9):2001–2009, September 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i9/p2001_s1.

Schmid:1988:QHT

[1614] Rudolf Schmid. The quadratic-Hamiltonian theorem in infinite dimensions. Journal of Mathematical Physics, 29(9):2010–2011, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i9/p2010_s1.

Leon:1988:STS

[1615] Jérôme J.-P. Leon. Spectral transform and solitons for generalized coupled Bloch systems. Journal of Mathematical Physics, 29(9):2012–2019, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i9/p2012_s1.

Gudder:1988:TA

[1616] Stanley Gudder. A theory of amplitudes. Journal of Mathematical Physics, 29(9):2020–2035, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i9/p2020_s1.

Narcowich:1988:CCT

[1617] Francis J. Narcowich. Conditions for the convolution of two Wigner distributions to be itself a Wigner distribution. Journal of Mathematical Physics, 29(9):2036–2041, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i9/p2036_s1.

Jordan:1988:BPU

[1618] Thomas F. Jordan. Berry phases and unitary transformations. Journal of Mathematical Physics, 29(9):2042–2052, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p2042_s1.

Turbiner:1988:AOC

[1619] A. V. Turbiner and A. G. Ushveridze. Anharmonic oscillator: constructing the strong coupling expansions. Journal of Mathematical Physics, 29(9):2053–2063, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i9/p2053_s1.

Hasse:1988:GKC

[1620] W. Hasse and V. Perlick. Geometrical and kinematical characterization of parallax-free world models. Journal of Mathematical Physics, 29 (9):2064–2068, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i9/p2064_s1.

Mathai:1988:GIM

[1621] A. M. Mathai, H. J. Haubold, J. P. Mücket, S. Gottlöber, and V. Müller. Gravitational instability in a multicomponent cosmological medium.

Journal of Mathematical Physics, 29(9):2069–2077, September 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p2069_
s1.

TorresdelCastillo:1988:TSI

[1622] G. F. Torres del Castillo. The Teukolsky–Starobinsky identities in type D vacuum backgrounds with cosmological constant. Journal of Mathematical Physics, 29(9):2078–2083, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i9/p2078_s1.

Giraud:1988:RSM

[1623] B. G. Giraud, S. Kessal, and A. Weiguny. A 29-root soluble model for the nonlinear calculation of a four-body propagator. Journal of Mathematical Physics, 29(9):2084–2089, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i9/p2084_s1.

Hall:1988:CPT

[1624] George L. Hall. Comment on the paper: “Theta series and magic numbers for diamond and certain ionic crystal structures”, [J. Math. Phys. 28 (1987), no. 7, 1653–1657; MR 88g:82058 ] by N. J. A. Sloane. Journal of Mathematical Physics, 29(9):2090–2091, September 1988. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i9/p2090_

s1. See [1155].

Manolessou:1988:EMN

[1625] Marietta M. Manolessou. The Φ^{4} equations of motion. I. A new constructive method: the Φ iteration. Journal of Mathematical Physics, 29 (9):2092–2105, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i9/p2092_s1.

Wollenberg:1988:EQF

[1626] Manfred Wollenberg. The existence of quantum fields for local nets of algebras of observables. Journal of Mathematical Physics, 29(9):2106– 2111, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i9/p2106_s1.

Hurd:1988:PCF

[1627] T. R. Hurd. A power counting formula for short distance singularities in quantum field theory. Journal of Mathematical Physics, 29(9):2112–2116, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i9/p2112_s1.

Winicour:1988:MFN

[1628] J. Winicour. Massive fields at null infinity. Journal of Mathematical Physics, 29(9):2117–2121, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i9/p2117_s1.

Chmielowski:1988:CMU

[1629] Piotr Chmielowski. A class of models of the universe filled with multidimensional objects. Journal of Mathematical Physics, 29(9):2122–2126, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i9/p2122_s1.

Marques:1988:DEN

[1630] S. Marques. The Dirac equation in a non-Riemannian manifold. I. An analysis using the complex algebra. Journal of Mathematical Physics, 29(9):2127–2131, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i9/p2127_s1.

Villarroel:1988:SYM

[1631] Javier Villarroel. Stochasticity in Yang–Mills theory. Journal of Mathematical Physics, 29(9):2132–2136, September 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i9/p2132_s1.

Turkowski:1988:LDR

[1632] P. Turkowski. Low-dimensional real Lie algebras. Journal of Mathematical Physics, 29(10):2139–2144, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2139_s1.

Gagnon:1988:NES

[1633] L. Gagnon, V. Hussin, and P. Winternitz. Nonlinear equations with superposition formulas and exceptional group G_{2}. III. the superposition formulas. Journal of Mathematical Physics, 29(10):2145–2155, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2145_s1.

Cianci:1988:SFO

[1634] Roberto Cianci. Superspace first-order partial differential equations through the Cartan–Kähler integration theorem. Journal of Mathematical Physics, 29(10):2156–2161, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2156_s1.

Plebanski:1988:WGS

[1635] Jerzy F. Plebański. Wronskians, geometry, and some general solutions to the nonlinear Liouville-like PDEs. Journal of Mathematical Physics, 29(10):2162–2174, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i10/p2162_s1.

Farrell:1988:ISP

[1636] R. A. Farrell, E. P. Gray, and R. W. Hart. An infinite sum of products of Jacobi polynomials. Journal of Mathematical Physics, 29(10):2175– 2177, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i10/p2175_s1.

Misawa:1988:NTS

[1637] Tetsuya Misawa. Noether’s theorem in symmetric stochastic calculus of variations. Journal of Mathematical Physics, 29(10):2178–2180, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2178_s1.

Carinena:1988:CTG

[1638] José F. Cariñena and Manuel F. Rañada. Canonoid transformations from a geometric perspective. Journal of Mathematical Physics, 29(10):2181– 2186, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i10/p2181_s1.

Kersten:1988:SSM

[1639] P. H. M. Kersten. Symmetries for the super modified KdV equation. Journal of Mathematical Physics, 29(10):2187–2189, October 1988.

CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i10/
p2187_s1.

Piette:1988:ESG

[1640] B. Piette. Explicit solutions of Grassmannian σ models. Journal of Mathematical Physics, 29(10):2190–2196, October 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i10/p2190_s1.

Verheest:1988:IRM

[1641] Frank Verheest. Integrability of restricted multiple three-wave interactions. II. Coupling constants with ratios 1 and 2. Journal of Mathematical Physics, 29(10):2197–2201, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2197_s1.

Dendy:1988:RIA

[1642] R. O. Dendy. Resonant interval action transfer between coupled harmonic oscillators. Journal of Mathematical Physics, 29(10):2202–2204, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2202_s1.

Weston:1988:FDW

[1643] V. H. Weston. Factorization of the dissipative wave equation and inverse scattering. Journal of Mathematical Physics, 29(10):2205–2218, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2205_s1.

Hillion:1988:SME

[1644] Pierre Hillion. Solutions of Maxwell’s equations with boundary conditions on the hyperplane z − ct = 0. Journal of Mathematical Physics, 29(10):2219–2222, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i10/p2219_s1.

Bona:1988:DCQ

[1645] Pavel Bóna. The dynamics of a class of quantum mean-field theories. Journal of Mathematical Physics, 29(10):2223–2235, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i10/
p2223_s1.

Beattie:1988:LES

[1646] Christopher Beattie and Mary Beth Ruskai. Location of essential spectrum of intermediate Hamiltonians restricted to symmetry subspaces. Journal of Mathematical Physics, 29(10):2236–2240, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2236_s1.

Dabrowski:1988:FMS

[1647] L. Dabrowski and J. Shabani. Finitely many sphere interactions in quantum mechanics: nonseparated boundary conditions. Journal of Mathematical Physics, 29(10):2241–2244, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i10/p2241_s1.

Newton:1988:PCS

[1648] Paul K. Newton. The perturbed cubic Schrödinger equation: selection mechanism, resonant limits, and spatial chaos. Journal of Mathematical Physics, 29(10):2245–2249, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2245_s1.

Huestis:1988:DSS

[1649] David L. Huestis. Degenerate Siegert states. Journal of Mathematical Physics, 29(10):2250–2253, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2250_s1.

Lin:1988:ISM

[1650] Jian Cheng Lin. Inverse scattering for the mixed spectrum of δ potentials. Journal of Mathematical Physics, 29(10):2254–2255, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2254_s1.

Molzahn:1988:QWA

[1651] F. H. Molzahn. A quantum WKB approximation without classical trajectories. Journal of Mathematical Physics, 29(10):2256–2267, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2256_s1.

Kamath:1988:BAF

[1652] S. G. Kamath. The brachistochrone in almost flat space. Journal of Mathematical Physics, 29(10):2268–2272, October 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i10/p2268_s1.

Holm:1988:NSE

[1653] Christian Holm. Neutrino spectrum of Einstein universes. Journal of Mathematical Physics, 29(10):2273–2279, October 1988. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v29/i10/p2273_s1.

Novotny:1988:ETM

[1654] M. A. Novotny. Equivalence of transfer matrices. Journal of Mathematical Physics, 29(10):2280–2287, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i10/p2280_s1.

Fannes:1988:TSB

[1655] M. Fannes and B. Nachtergaele. Translating the spin-boson model into a classical system. Journal of Mathematical Physics, 29(10):2288–2293, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i10/p2288_s1.

Wang:1988:DRC

[1656] Rong Tai Wang. The derivation of the regularized chiral Jacobian using the zeta function method. Journal of Mathematical Physics, 29(10):2294– 2299, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i10/p2294_s1.

Wickerhauser:1988:HFK

[1657] Mladen Victor Wickerhauser. Hamilton’s form for the Kadomtsev– Petviashvili equation. Journal of Mathematical Physics, 29(10):2300– 2302, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i10/p2300_s1.

Majid:1988:LTY

[1658] Shahn Majid. Liouville theorem for the Yang–Mills self-duality equations. Journal of Mathematical Physics, 29(10):2303–2310, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i10/

p2303_s1. See erratum [1947].

Furlan:1988:MMC

[1659] P. Furlan, G. M. Sotkov, and Ya. Stanev. Minimal models of U(1) conformal current algebra. Journal of Mathematical Physics, 29(10):2311– 2319, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i10/p2311_s1.

Bluman:1988:ENC

[1660] G. W. Bluman, G. J. Reid, and Sukeyuki Kumei. Erratum: “New classes of symmetries for partial differential equations”, [J. Math. Phys. 29, 806 (1988)]. Journal of Mathematical Physics, 29(10):2320, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i10/

p2320_s1. See [1450].

Gron:1988:EVP

[1661] Öyvind Grön and Harald H. Soleng. Erratum: “The Vaidya–Patel solution with Robertson–Walker metric as a rotating inflationary scenario”, [J. Math. Phys. 29, 1514 (1988)]. Journal of Mathematical Physics, 29(10):2320, October 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i10/p2320_s2. See [1543].

Shaw:1988:TCA

[1662] Ronald Shaw. Ternary composition algebras and Hurwitz’s theorem. Journal of Mathematical Physics, 29(11):2329–2333, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i11/
p2329_s1.

Plebanski:1988:NCP

[1663] J. F. Plebański and M. Przanowski. Notes on a cross product of vectors. Journal of Mathematical Physics, 29(11):2334–2337, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2334_s1.

Wilmot:1988:SCA

[1664] G. P. Wilmot. The structure of the Clifford algebra. Journal of Mathematical Physics, 29(11):2338–2345, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i11/p2338_s1.

Wilmot:1988:CAP

[1665] G. P. Wilmot. Clifford associativity and the Pfaffian expansion. Journal of Mathematical Physics, 29(11):2346–2350, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i11/p2346_s1.

Alisauskas:1988:BOC

[1666] Sigitas Ališauskas. On biorthogonal and orthonormal Clebsch–Gordan coefficients of SU(3): analytical and algebraic approaches. Journal of Mathematical Physics, 29(11):2351–2366, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i11/p2351_s1.

Gunaydin:1988:ULW

[1667] M. Günaydin and S. J. Hyun. Unitary lowest weight representations of the noncompact supergroup OSp(2n/2m, R). Journal of Mathematical Physics, 29(11):2367–2376, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2367_s1.

Klimyk:1988:IOG

[1668] A. U. Klimyk and I. I. Kachurik. Infinitesimal operators of group representations in noncanonical bases. Journal of Mathematical Physics, 29 (11):2377–2383, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i11/p2377_s1.

Pan:1988:SIF

[1669] Feng Pan and Yu Fang Cao. Some isoscalar factors for SO_{N} ⊃ SO_{N−1} and state expansion coefficients for SO_{N} ⊃ SO_{N−1} in terms of SU_{N} ⊃ SU_{N−1} ⊃ SO_{N−1}. Journal of Mathematical Physics, 29(11):2384–2388, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2384_s1.

Teofilatto:1988:DSS

[1670] Paolo Teofilatto. Discrete supergroups and super Riemann surfaces. Journal of Mathematical Physics, 29(11):2389–2396, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i11/
p2389_s1.

Sachdev:1988:GBE

[1671] P. L. Sachdev, K. R. C. Nair, and V. G. Tikekar. Generalized Burgers equations and Euler–Painlevé transcendents. III. Journal of Mathematical Physics, 29(11):2397–2404, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2397_s1.

Shabani:1988:SUC

[1672] J. Shabani. Some unbounded commutants of a set of operators on a partial inner product space. Journal of Mathematical Physics, 29(11):2405– 2410, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2405_s1.

Sharma:1988:SO

[1673] C. S. Sharma and D. F. Almeida. Semilinear operators. Journal of Mathematical Physics, 29(11):2411–2420, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i11/p2411_s1.

Sharma:1988:PCS

[1674] Ramesh Sharma. Proper conformal symmetries of conformal symmetric space–times. Journal of Mathematical Physics, 29(11):2421–2422, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2421_s1.

Dodd:1988:FZC

[1675] R. K. Dodd and S. K. Paul. Factorization of zero curvature representations. Journal of Mathematical Physics, 29(11):2423–2429, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2423_s1.

Lucey:1988:CH

[1676] C. A. Lucey and E. T. Newman. On the construction of Hamiltonians. Journal of Mathematical Physics, 29(11):2430–2433, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2430_s1.

Gottlieb:1988:WFC

[1677] H. P. W. Gottlieb. Wake-free cylindrical and spherical waves in inhomogeneous media. Journal of Mathematical Physics, 29(11):2434–2437, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2434_s1.

Benoit:1988:SPC

[1678] Louis Benoit and Luc Vinet. Supermultiplet of particles in a Coulomb potential: states and wave functions from the representation theory of OSp(2, 1). Journal of Mathematical Physics, 29(11):2438–2442, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2438_s1.

Rai:1988:CSCb

[1679] Jagdish Rai, C. L. Mehta, and N. Mukunda. Correlated states and collective transition operators for multilevel atomic systems. II. Symmetry properties. Journal of Mathematical Physics, 29(11):2443–2449, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2443_s1.

Matsui:1988:URF

[1680] Yoshiko Matsui. The unitarity relations for the four-body scattering amplitude. Journal of Mathematical Physics, 29(11):2450–2461, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2450_s1.

Bona:1988:ICV

[1681] C. Bona. Invariant conformal vectors in space–times admitting a group of G_{3} of motions acting on spacelike orbits S_{2}. Journal of Mathematical Physics, 29(11):2462–2464, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2462_s1.

Hall:1988:ACS

[1682] G. S. Hall and J. da Costa. Affine collineations in space–time. Journal of Mathematical Physics, 29(11):2465–2472, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i11/p2465_s1.

Qadir:1988:SSS

[1683] Asghar Qadir and M. Ziad. Static spherically symmetric space–times with six Killing vectors. Journal of Mathematical Physics, 29(11):2473– 2474, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2473_s1. See addendum [2127].

Kuang:1988:BTT

[1684] Zhi Quan Kuang and Can Bin Liang. Birkhoff and Taub theorems generalized to metrics with conformal symmetries. Journal of Mathematical Physics, 29(11):2475–2478, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2475_s1.

PoncedeLeon:1988:SEE

[1685] J. Ponce de León. Solutions of Einstein’s equations relevant to the description of “bubbles” in the early universe. Journal of Mathematical Physics, 29(11):2479–2485, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2479_s1.

Uehara:1988:SST

[1686] Shozo Uehara and Yukinori Yasui. Super-Selberg trace formula from the chaotic model. Journal of Mathematical Physics, 29(11):2486–2490, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2486_s1.

Caticha:1988:LTS

[1687] Nestor Caticha. Limit theorems for spin systems with an abelian discrete symmetry. Journal of Mathematical Physics, 29(11):2491–2493, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i11/p2491_s1.

Braden:1988:AC

[1688] H. W. Braden. Anomalous congruences. Journal of Mathematical Physics, 29(11):2494–2498, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i11/p2494_s1.

Mathieu:1988:SEK

[1689] Pierre Mathieu. Supersymmetric extension of the Korteweg–de Vries equation. Journal of Mathematical Physics, 29(11):2499–2506, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2499_s1.

Hammer:1988:CSP

[1690] C. L. Hammer and J. E. Shrauner. Coherent-state path-integral S-matrix formalism applied to the Lee model. Journal of Mathematical Physics, 29(11):2507–2512, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i11/p2507_s1.

Weston:1988:EFW

[1691] Vaughan H. Weston. Erratum: “Factorization of the wave equation in a nonplanar stratified medium”, [J. Math. Phys. 29 (1988), no. 1, 36–45; MR 88k:35113]. Journal of Mathematical Physics, 29(11):2513, November 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i11/p2513_s1. See [1350].

Bergdolt:1988:PBA

[1692] G. Bergdolt. Projector bases and algebraic spinors. Journal of Mathematical Physics, 29(12):2519–2522, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v29/i12/p2519_s1.

dePrunele:1988:CVM

[1693] E. de Prunelé. SU(1, 1), its connections with SU(2), and the vector model. Journal of Mathematical Physics, 29(12):2523–2532, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i12/p2523_s1.

Gazeau:1988:ISF

[1694] Jean-Pierre Gazeau and Michel Hans. Integral-spin fields on (3 + 2)- de Sitter space. Journal of Mathematical Physics, 29(12):2533–2552, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i12/p2533_s1.

Zeng:1988:WOR

[1695] Gao Jian Zeng and Xiou Hua Yuan. Wigner operator and Racah operator of the Lie superalgebra OSp(1, 2). Journal of Mathematical Physics, 29 (12):2553–2562, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2553_s1.

Leach:1988:ASN

[1696] P. G. L. Leach, M. R. Feix, and S. Bouquet. Analysis and solution of a nonlinear second-order differential equation through rescaling and through a dynamical point of view. Journal of Mathematical Physics, 29(12):2563–2569, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2563_s1.

Zheng:1988:GBT

[1697] Yu Kun Zheng and W. L. Chan. Gauge and Bäcklund transformations for the variable coefficient higher-order modified Korteweg–de Vries equation. Journal of Mathematical Physics, 29(12):2570–2575, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i12/p2570_s1.

Gruszczak:1988:GMS

[1698] J. Gruszczak, M. Heller, and P. Multarzyński. A generalization of manifolds as space–time models. Journal of Mathematical Physics, 29 (12):2576–2580, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2576_s1.

Goderis:1988:TFS

[1699] D. Goderis, A. Verbeure, and P. Vets. Theory of fluctuations and small oscillations for quantum lattice systems. Journal of Mathematical Physics, 29(12):2581–2585, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i12/p2581_s1.

Dattoli:1988:ESF

[1700] G. Dattoli, M. Richetta, and A. Torre. Evolution of SU(2) and SU(1, 1) states: a further mathematical analysis. Journal of Mathematical Physics, 29(12):2586–2588, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i12/p2586_s1.

Palev:1988:FDR

[1701] Tchavdar D. Palev. Finite-dimensional representations of the special linear Lie superalgebra sl(1, n). II. Nontypical representations. Journal of Mathematical Physics, 29(12):2589–2598, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i12/p2589_s1.

Omote:1988:IDS

[1702] Minoru Omote. Infinite-dimensional symmetry algebras and an infinite number of conserved quantities of the (2 + 1)-dimensional Davey– Stewartson equation. Journal of Mathematical Physics, 29(12):2599– 2603, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i12/p2599_s1.

Scialom:1988:AST

[1703] Márcia A. G. Scialom and Rafael J. Iório, Jr. Adiabatic switching for time-dependent electric fields. Journal of Mathematical Physics, 29 (12):2604–2610, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2604_s1.

Znojil:1988:PON

[1704] M. Znojil. Padé oscillators and a new formulation of perturbation theory. Journal of Mathematical Physics, 29(12):2611–2617, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i12/
p2611_s1.

Zhou:1988:EMA

[1705] Qi Huang Zhou. The exterior metric approach to a charged axially symmetric celestial body—the fourth-order approximate solutions of Einstein–Maxwell equations. Journal of Mathematical Physics, 29 (12):2618–2621, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2618_s1.

Taub:1988:CIG

[1706] A. H. Taub. Collision of impulsive gravitational waves followed by dust clouds. Journal of Mathematical Physics, 29(12):2622–2627, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i12/p2622_s1.

Barouch:1988:BHF

[1707] E. Barouch, A. S. Fokas, and V. G. Papageorgiou. The bi-Hamiltonian formulation of the Landau–Lifshiftz equation. Journal of Mathematical Physics, 29(12):2628–2633, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i12/p2628_s1.

Suranyi:1988:ACP

[1708] P. Suranyi. Algebraic construction of partition functions for c < 1 minimal conformal field theories. Journal of Mathematical Physics, 29 (12):2634–2640, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2634_s1.

Oh:1988:SGP

[1709] C. H. Oh, K. Singh, and C. H. Lai. Some global properties and invariance of bundle metrics in the Kaluza–Klein scheme. Journal of Mathematical Physics, 29(12):2641–2652, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i12/p2641_s1.

Zablocki:1988:SMO

[1710] Krzysztof Zabłocki. A simplified model for orbifold compactification. Journal of Mathematical Physics, 29(12):2653–2658, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v29/i12/
p2653_s1.

Bender:1988:EFD

[1711] Carl M. Bender and H. F. Jones. Evaluation of Feynman diagrams in the logarithmic approach to quantum field theory. Journal of Mathematical Physics, 29(12):2659–2665, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v29/i12/p2659_s1.

Leo:1988:PPN

[1712] R. A. Leo, G. Mancarella, G. Soliani, and L. Solombrino. On the Painlevé property of nonlinear field equations in 2 + 1 dimensions: the Davey– Stewartson system. Journal of Mathematical Physics, 29(12):2666–2671, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i12/p2666_s1.

Nakamura:1988:PRS

[1713] Yoshimasa Nakamura. Periodic reduction of self-dual Yang–Mills equations. Journal of Mathematical Physics, 29(12):2672–2674, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v29/i12/p2672_s1.

Roberts:1988:ESC

[1714] Thomas M. Roberts, Harold A. Sabbagh, and L. David Sabbagh. Electromagnetic scattering for a class of anisotropic layered media. Journal of Mathematical Physics, 29(12):2675–2681, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v29/i12/p2675_s1.

Sen:1988:NEV

[1715] Diptiman Sen. A new eight-vertex model with an infinite number of commensurate phases. Journal of Mathematical Physics, 29(12):2682– 2687, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i12/p2682_s1.

DellAntonio:1988:GRE

[1716] G. F. Dell’Antonio and V. Nesi. A general representation for the effective dielectric constant of a composite. Journal of Mathematical Physics, 29 (12):2688–2694, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v29/i12/p2688_s1.

McIlroy:1988:SMA

[1717] Douglas K. McIlroy. A simple model of asymmetry in Wien dissociation of a weak electrolyte. Journal of Mathematical Physics, 29(12):2695– 2699, December 1988. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v29/i12/p2695_s1.

Basu:1989:RHS

[1718] Debabrata Basu. Representations of SL(2, R) in a Hilbert space of analytic functions and a class of associated integral transforms. Journal of Mathematical Physics, 30(1):1–8, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i1/p1_s1.

Kotzev:1989:GRR

[1719] J. N. Kotzev, M. K. Peev, M. I. Aroyo, R. Dirl, and P. Kasperkovitz. Generating relations for reducing matrices. IV. Subduced representations. Journal of Mathematical Physics, 30(1):9–17, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p9_s1.

Vretare:1989:CCS

[1720] L. Vretare. Character calculations for semisimple Lie algebras. Journal of Mathematical Physics, 30(1):18–22, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i1/p18_s1.

Znojil:1989:SAA

[1721] M. Znojil. Singular anharmonicities and the analytic continued fractions. Journal of Mathematical Physics, 30(1):23–27, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p23_s1.

Calogero:1989:UIN

[1722] F. Calogero. Universality and integrability of the nonlinear evolution PDEs describing N-wave interactions. Journal of Mathematical Physics, 30(1):28–40, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i1/p28_s1.

Inayat-Hussain:1989:NGH

[1723] A. A. Inayat-Hussain. A new generalization of the Hankel integral transform. Journal of Mathematical Physics, 30(1):41–44, January 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p41_s1.

Ivancovich:1989:GFE

[1724] J. Ivancovich, C. Kozameh, and E. T. Newman. Green’s functions of the edh operators. Journal of Mathematical Physics, 30(1):45–52, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i1/p45_s1.

Kreider:1989:TDI

[1725] K. L. Kreider. Time-dependent inverse scattering from gradient-type interfaces using an exact solution. Journal of Mathematical Physics, 30(1): 53–58, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p53_s1.

Razon:1989:BTQ

[1726] Aharon Razon, L. P. Horwitz, and L. C. Biedenharn. On a basic theorem of quaternion modules. Journal of Mathematical Physics, 30(1): 59, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p59_s1.

Abdalla:1989:QMD

[1727] M. Sebawe Abdalla and Ursula A. T. Ramjit. Quantum mechanics of the damped pulsating oscillator. Journal of Mathematical Physics, 30(1): 60–65, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p60_s1.

Arshansky:1989:QRTa

[1728] R. Arshansky and L. P. Horwitz. The quantum relativistic two-body bound state. I. The spectrum. Journal of Mathematical Physics, 30(1): 66–80, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p66_s1.

Baseia:1989:ESE

[1729] B. Baseia. Extended Schrödinger equations leading to critical phenomena. Journal of Mathematical Physics, 30(1):81–82, January 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p81_s1.

Chen:1989:DRF

[1730] Robert L. W. Chen. Derivation of the real form of Schrödinger’s equation for a nonconservative system and the unique relation between ℜ(ψ) and ℑ(ψ). Journal of Mathematical Physics, 30(1):83–86, January 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p83_s1.

Dunn:1989:SK

[1731] K. A. Dunn and J. G. Williams. De Sitter kinks. Journal of Mathematical Physics, 30(1):87–89, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i1/p87_s1.

Burnett:1989:HFL

[1732] Gregory A. Burnett. The high-frequency limit in general relativity. Journal of Mathematical Physics, 30(1):90–96, January 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i1/p90_s1.

Papacostas:1989:CGE

[1733] Taxiarchis Papacostas and Basilis C. Xanthopoulos. Collisions of gravitational and electromagnetic waves that do not develop curvature singularities. Journal of Mathematical Physics, 30(1):97–103, January 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p97_s1.

Ferrara:1989:CGB

[1734] S. Ferrara and M. Villasante. Curvatures, Gauss–Bonnet and Chern– Simons multiplets in old-minimal N = 1 supergravity. Journal of Mathematical Physics, 30(1):104–114, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i1/p104_s1.

Gielerak:1989:FED

[1735] Roman Gielerak. Free energy density for continuous systems with and without superstability assumptions. Journal of Mathematical Physics, 30(1):115–133, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i1/p115_s1.

Schneider:1989:FDW

[1736] W. R. Schneider and W. Wyss. Fractional diffusion and wave equations. Journal of Mathematical Physics, 30(1):134–144, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/p134_s1.

Lawler:1989:LDE

[1737] Gregory F. Lawler. Low-density expansion for a two-state random walk in a random environment. Journal of Mathematical Physics, 30(1):145– 157, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p145_s1.

vanderMee:1989:WPS

[1738] C. V. M. van der Mee. Well-posedness of stationary and time-dependent Spencer–Lewis equations modeling electron slowing down. Journal of Mathematical Physics, 30(1):158–165, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i1/p158_s1.

Chau:1989:KMA

[1739] Ling-Lie Chau and Mo Lin Ge. Kac–Moody algebra from infinitesimal Riemann–Hilbert transform. Journal of Mathematical Physics, 30(1):166–171, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i1/p166_s1.

Tanaka:1989:CFS

[1740] K. Tanaka. Curvature form and solutions of nonlinear models. Journal of Mathematical Physics, 30(1):172–174, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i1/p172_s1.

Manolessou:1989:EMIa

[1741] Marietta Manolessou. The Φ^{4} equations of motion. II. The zero-, one- , and two-dimensional solutions. Journal of Mathematical Physics, 30(1):175–196, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i1/p175_s1.

Willis:1989:CMS

[1742] Barton L. Willis. Conservation of momentum for systems of charged particles. Journal of Mathematical Physics, 30(1):197–200, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i1/
p197_s1.

Kwon:1989:DSS

[1743] P. Kwon and M. Villasante. Decomposition of the scalar superfield in ten dimensions. Journal of Mathematical Physics, 30(1):201–212, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i1/p201_s1.

Arshansky:1989:RPS

[1744] R. Arshansky and L. P. Horwitz. Relativistic potential scattering and phase shift analysis. Journal of Mathematical Physics, 30(1):213–218, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i1/p213_s1.

Klein:1989:BQR

[1745] Abraham Klein and Eugene R. Marshalek. On the boson-quasifermion realization of the particle-hole SO(2Ω + 1) algebra. Journal of Mathematical Physics, 30(1):219–232, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i1/p219_s1.

Hogan:1989:EFE

[1746] P. A. Hogan and G. F. R. Ellis. Electromagnetic fields in an expanding universe. Journal of Mathematical Physics, 30(1):233–240, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i1/p233_s1.

Matsuno:1989:ENI

[1747] Y. Matsuno. Erratum: “New integrable nonlinear integro-differential equations and related solvable finite-dimensional dynamical systems”, [J. Math. Phys. 29 (1988), no. 1, 49–56; MR 89a:35192]. Journal of Mathematical Physics, 30(1):241, January 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i1/p241_s1. See [1352].

Kopczynski:1989:SIC

[1748] Wojciech Kopczyński. Spinorial idempotents in the Clifford algebras of Minkowski spaces. Journal of Mathematical Physics, 30(2):243–248, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p243_s1.

Wene:1989:CAI

[1749] G. P. Wene. The Clifford algebra of an infinite-dimensional space. Journal of Mathematical Physics, 30(2):249–251, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i2/p249_s1.

Eastaugh:1989:SOR

[1750] A. Eastaugh and P. van Nieuwenhuizen. Small oscillations of regular polytopes in d ≥ 4 space dimensions. Journal of Mathematical Physics, 30(2):252–260, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p252_s1.

Naito:1989:FIT

[1751] Seichi Naito, Masahiro Nishimoto, and Tohru Hirokane. Fierz identities in ten-dimensional superspace. Journal of Mathematical Physics, 30(2):261–273, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p261_s1.

Balantekin:1989:CSN

[1752] A. B. Balantekin, H. A. Schmitt, and P. Halse. Coherent states for the noncompact supergroups OSp(2/2N, R). Journal of Mathematical Physics, 30(2):274–279, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i2/p274_s1.

Barannik:1989:CSG

[1753] L. F. Barannik and W. I. Fushchich. Continuous subgroups of the generalized Schrödinger groups. Journal of Mathematical Physics, 30(2):280– 290, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p280_s1.

Fatyga:1989:BCH

[1754] B. W. Fatyga, V. Alan Kostelecký, and D. Rodney Truax. Baker– Campbell–Hausdorff relations for the orthosymplectic group OSP(1/2). Journal of Mathematical Physics, 30(2):291–294, February 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i2/p291_
s1.

Critchfield:1989:CEF

[1755] Charles L. Critchfield. Computation of elliptic functions. Journal of Mathematical Physics, 30(2):295–297, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p295_s1.

Brenner:1989:SHO

[1756] Terence Brenner. On the spectra of hypoelliptic operators in R^{2}. Journal of Mathematical Physics, 30(2):298–301, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p298_s1.

Baker-Jarvis:1989:SBV

[1757] James Baker-Jarvis. Solution to boundary value problems using the method of maximum entropy. Journal of Mathematical Physics, 30 (2):302–306, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p302_s1.

Spies:1989:ECB

[1758] Günther Otto Spies. Eigenvalue curves and boundary conditions. Journal of Mathematical Physics, 30(2):307–312, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i2/p307_s1.

Gagnon:1989:EMW

[1759] L. Gagnon. Equations of motion and wave functions in spherical coordinates for an integrable Hamiltonian system. Journal of Mathematical Physics, 30(2):313–317, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i2/p313_s1.

Rivas:1989:CRS

[1760] Martín Rivas. Classical relativistic spinning particles. Journal of Mathematical Physics, 30(2):318–329, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i2/p318_s1.

Tu:1989:TIP

[1761] Gui Zhang Tu. The trace identity, a powerful tool for constructing the Hamiltonian structure of integrable systems. Journal of Mathematical Physics, 30(2):330–338, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i2/p330_s1.

Felderhof:1989:DTS

[1762] B. U. Felderhof and R. B. Jones. Displacement theorems for spherical solutions of the linear Navier–Stokes equations. Journal of Mathematical Physics, 30(2):339–342, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i2/p339_s1.

AAkerstedt:1989:ASF

[1763] Hans O. Åkerstedt. Analytic solutions of a fourth-order differential equation with a second-order turning point. Journal of Mathematical Physics, 30(2):343–344, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p343_s1.

Lozada:1989:PIC

[1764] Abraham Lozada and P. L. Torres. Poincaré invariance—classical electrodynamics. Journal of Mathematical Physics, 30(2):345–353, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p345_s1.

Loffredo:1989:LVP

[1765] Maria I. Loffredo and Laura M. Morato. Lagrangian variational principle in stochastic mechanics: gauge structure and stability. Journal of Mathematical Physics, 30(2):354–360, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p354_s1.

Pottinger:1989:GQA

[1766] Jürgen Pöttinger. Global quantities in algebraic quantum mechanics of infinite systems: classical observables or parameters. Journal of Mathematical Physics, 30(2):361–368, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i2/p361_s1.

Sharma:1989:AFO

[1767] C. S. Sharma and D. F. Almeida. Additive functionals and operators on a quaternionic Hilbert space. Journal of Mathematical Physics, 30(2):369– 375, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p369_s1.

Watanabe:1989:HOS

[1768] Shuji Watanabe. A harmonic oscillator system with the generalized commutation relations. Journal of Mathematical Physics, 30(2):376– 379, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p376_s1.

Arshansky:1989:QRTb

[1769] R. Arshansky and L. P. Horwitz. The quantum relativistic two-body bound state. II. The induced representation of SL(2, C). Journal of Mathematical Physics, 30(2):380–392, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p380_s1.

Morales:1989:GBT

[1770] J. Morales and A. Flores-Riveros. The generalization of the binomial theorem. Journal of Mathematical Physics, 30(2):393–397, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p393_s1.

Rashid:1989:RRT

[1771] Muneer A. Rashid. Recurrence relations for two-center harmonic oscillator integrals and their solutions using generating functions. Journal of Mathematical Physics, 30(2):398–405, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p398_s1.

Leach:1989:RFA

[1772] P. G. L. Leach, G. P. Flessas, and V. M. Gorringe. A Rodrigues formula approach to determining closed-form solutions to the Schrödinger equation for symmetric anharmonic oscillators. Journal of Mathematical Physics, 30(2):406–412, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i2/p406_s1.

Znojil:1989:PSC

[1773] M. Znojil. On the power-series construction of Schrödinger bound states. II. the effective hill determinants. Journal of Mathematical Physics, 30(2):413–419, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p413_s1.

Bolle:1989:LGS

[1774] D. Bollé, C. Danneels, and T. A. Osborn. Local and global spectral shift functions in R^{2}. Journal of Mathematical Physics, 30(2):420–432, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p420_s1.

Krasinski:1989:SFN

[1775] Andrzej Krasiński. Shear-free normal cosmological models. Journal of Mathematical Physics, 30(2):433–441, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p433_s1.

Oliver:1989:FIC

[1776] Gabriel Oliver and Enric Verdaguer. A family of inhomogeneous cosmological Einstein–Rosen metrics. Journal of Mathematical Physics, 30(2):442–445, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p442_s1.

TorresdelCastillo:1989:RSF

[1777] G. F. Torres del Castillo. Rarita–Schwinger fields in algebraically special vacuum space–times. Journal of Mathematical Physics, 30(2):446– 451, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p446_s1.

Hartle:1989:SMI

[1778] James B. Hartle. Simplicial minisuperspace. III. integration contours in a five-simplex model. Journal of Mathematical Physics, 30(2):452– 460, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p452_s1.

Redei:1989:HVP

[1779] Miklós Rédei. The hidden variable problem in algebraic relativistic quantum field theory. Journal of Mathematical Physics, 30(2):461–463, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p461_s1.

Spehler:1989:CSN

[1780] D. Spehler and G. C. Marques. Classical solutions of a nonrelativistic model exhibiting spontaneous symmetry breakdown. Journal of Mathematical Physics, 30(2):464–469, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i2/p464_s1.

Thomi:1989:BQS

[1781] P. Thomi. BRST quantization of the spinning particle. Journal of Mathematical Physics, 30(2):470–476, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i2/p470_s1.

McMullan:1989:CFOa

[1782] D. McMullan and J. Paterson. Covariant factor ordering of gauge systems using ghost variables. I. constraint rescaling. Journal of Mathematical Physics, 30(2):477–486, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i2/p477_s1.

McMullan:1989:CFOb

[1783] D. McMullan and J. Paterson. Covariant factor ordering of gauge systems using ghost variables. II. States and observables. Journal of Mathematical Physics, 30(2):487–497, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i2/p487_s1.

McInnes:1989:GGS

[1784] Brett McInnes. The geometry of gauge symmetry breaking in the superstring context. Journal of Mathematical Physics, 30(2):498–505, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p498_s1.

Grossman:1989:MST

[1785] Bernard Grossman. Membranes in string theory, trees, the Weil conjectures, and the Ramanujan numbers. Journal of Mathematical Physics, 30(2):506–511, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p506_s1.

Arai:1989:SEM

[1786] Asao Arai. Supersymmetric embedding of a model of a quantum harmonic oscillator interacting with infinitely many bosons. Journal of Mathematical Physics, 30(2):512–520, February 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i2/p512_s1.

Rylov:1989:ASE

[1787] Yu. A. Rylov. The algebraical structure of the electromagnetic tensor and description of charged particles moving in the strong electromagnetic field. Journal of Mathematical Physics, 30(2):521–536, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i2/p521_s1.

Mielnik:1989:ETR

[1788] Bogdan Mielnik and David J. Fernández C. An electron trapped in a rotating magnetic field. Journal of Mathematical Physics, 30(2):537– 549, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i2/p537_s1.

Mielnik:1989:EEL

[1789] Bogdan Mielnik. Erratum: “Evolution loops”, [J. Math. Phys. 27 (1986), no. 9, 2290–2306; MR 87j:81030]. Journal of Mathematical Physics, 30(2):550, February 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i2/p550_s1. See [805].

Kamupingene:1989:FDR

[1790] Alfred H. Kamupingene, Nguyen Anh Ky, and Tchavdar D. Palev. Finite-dimensional representations of the Lie superalgebra gl(2/2) in a gl(2) ⊕ gl(2) basis. I. Typical representations. Journal of Mathematical Physics, 30(3):553–570, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i3/p553_s1.

Bouwknegt:1989:LAA

[1791] Peter Bouwknegt. Lie algebra automorphisms, the Weyl group, and tables of shift vectors. Journal of Mathematical Physics, 30(3):571– 584, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p571_s1.

Buccella:1989:OSE

[1792] F. Buccella, A. Della Selva, and A. Sciarrino. Octonions and subalgebras of the exceptional algebras. Journal of Mathematical Physics, 30(3):585– 593, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p585_s1.

Doebner:1989:BOR

[1793] H. D. Doebner, Bruno Gruber, and Miguel Lorente. Boson operator realizations of su(2) and su(1, 1) and unitarization. Journal of Mathematical Physics, 30(3):594–600, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i3/p594_s1.

Common:1989:CTK

[1794] A. K. Common and A. Martin. Constraints on the total and kinetic energy of ground states in a class of potential models. Journal of Mathematical Physics, 30(3):601–607, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i3/p601_s1.

Tanikawa:1989:CSP

[1795] Kiyotaka Tanikawa and Yoshihiro Yamaguchi. Coexistence of symmetric periodic points in the standard map. Journal of Mathematical Physics, 30(3):608–616, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i3/p608_s1.

Noriega:1989:SCM

[1796] R. J. Noriega and C. G. Schifini. Scalar concomitants of a metric and a curvature form. II. Journal of Mathematical Physics, 30(3):617– 618, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p617_s1.

Bugajska:1989:LSO

[1797] Krystyna Bugajska. Lorentzian structures of open, spin manifolds. Journal of Mathematical Physics, 30(3):619–623, March 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i3/p619_s1.

Nester:1989:GCO

[1798] James M. Nester. A gauge condition for orthonormal three-frames. Journal of Mathematical Physics, 30(3):624–626, March 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i3/p624_s1.

Tratnik:1989:MBH

[1799] M. V. Tratnik. Multivariable biorthogonal Hahn polynomials. Journal of Mathematical Physics, 30(3):627–634, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i3/p627_s1.

Schutz:1989:ASI

[1800] John W. Schutz. Affine structure and isotropy imply Minkowski space– time and the orthochronous Poincaré group. Journal of Mathematical Physics, 30(3):635–638, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i3/p635_s1.

Calogero:1989:SCI

[1801] F. Calogero. Solutions of certain integrable nonlinear PDEs describing nonresonant N-wave interactions. Journal of Mathematical Physics, 30(3):639–654, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i3/p639_s1.

Chetouani:1989:EPI

[1802] L. Chetouani, L. Guechi, and T. F. Hammann. Exact path integral solution of the Coulomb plus Aharonov–Bohm potential. Journal of Mathematical Physics, 30(3):655–658, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i3/p655_s1.

Giavarini:1989:GCS

[1803] G. Giavarini and E. Onofri. Generalized coherent states and Berry’s phase. Journal of Mathematical Physics, 30(3):659–663, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i3/
p659_s1.

Kruszynski:1989:RCS

[1804] Paweł Kruszyński. Ring characterization of scalar product spaces and Gel’fand triples. Journal of Mathematical Physics, 30(3):664–667, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i3/p664_s1.

Zhou:1989:QIS

[1805] Huan Qiang Zhou and Lin Jie Jiang. Quantum inverse scattering method for a nonlinear N-wave resonance interaction system. Journal of Mathematical Physics, 30(3):668–672, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i3/p668_s1.

Sidharth:1989:PSC

[1806] B. G. Sidharth. Phase shifts in the collision of massive particles. Journal of Mathematical Physics, 30(3):673–677, March 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i3/p673_s1.

Li:1989:FEC

[1807] Wei Li and Frederick J. Ernst. A family of electrovac colliding wave solutions of Einstein’s equations. Journal of Mathematical Physics, 30(3): 678–682, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p678_s1.

Levy:1989:DNR

[1808] Laurent P. Lévy and Andrew T. Ogielski. Dissipation in nonlinear response. Journal of Mathematical Physics, 30(3):683–688, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i3/
p683_s1.

Cleary:1989:LEM

[1809] Paul W. Cleary. Lyapunov exponents as a measure of the size of chaotic regions. Journal of Mathematical Physics, 30(3):689–695, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i3/
p689_s1.

Atag:1989:LIS

[1810] Satilmis Atag. Large-N iterative solution of the Dirac equation. Journal of Mathematical Physics, 30(3):696–699, March 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i3/p696_s1.

Buchdahl:1989:CRW

[1811] H. A. Buchdahl. On the compatibility of relativistic wave equations in Riemann–Cartan spaces. Journal of Mathematical Physics, 30(3):700– 705, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p700_s1.

Tafel:1989:CSG

[1812] Jacek Tafel. A comparison of solution generating techniques for the self-dual Yang–Mills equations. Journal of Mathematical Physics, 30(3): 706–710, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p706_s1.

Lazarides:1989:TGS

[1813] G. Lazarides and Q. Shafi. Three generation superstring models with maximal discrete symmetries. Journal of Mathematical Physics, 30(3): 711–718, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p711_s1.

Mitra:1989:FDI

[1814] Bani Mitra. Finite-dimensional irreducible representations of the SU(3/1) superalgebra. Journal of Mathematical Physics, 30(3):719– 731, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p719_s1. See erratum [2137].

Willis:1989:CAM

[1815] Barton L. Willis. Conservation of angular momentum for systems of charged particles. Journal of Mathematical Physics, 30(3):732– 734, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i3/p732_s1.

Huang:1989:SIP

[1816] Jau-Inn Huang. Solution to the inverse problem of the 1-D wave equation. Journal of Mathematical Physics, 30(3):735–739, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i3/p735_s1.

Litvin:1989:ACO

[1817] Daniel B. Litvin. Algebraic construction of one-dimensional quasiperiodic tilings. Journal of Mathematical Physics, 30(3):740–743, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i3/p740_s1.

Adam:1989:NEP

[1818] John A. Adam. A nonlinear eigenvalue problem in astrophysical magnetohydrodynamics: some properties of the spectrum. Journal of Mathematical Physics, 30(3):744–756, March 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i3/p744_s1.

Ram:1989:SHC

[1819] Shri Ram. Spatially homogeneous cosmological models of Bianchi type III. Journal of Mathematical Physics, 30(3):757–759, March 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i3/p757_
s1.

Koornwinder:1989:MPP

[1820] Tom H. Koornwinder. Meixner–Pollaczek polynomials and the Heisenberg algebra. Journal of Mathematical Physics, 30(4):767–769, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p767_s1.

Dowker:1989:HKE

[1821] J. S. Dowker. Heat kernel expansion on a generalized cone. Journal of Mathematical Physics, 30(4):770–773, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p770_s1.

Weniger:1989:ATF

[1822] E. Joachim Weniger and E. Otto Steinborn. Addition theorems for B functions and other exponentially declining functions. Journal of Mathematical Physics, 30(4):774–784, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i4/p774_s1.

Guler:1989:DSC

[1823] Y. Güler. On the dynamics of singular, continuous systems. Journal of Mathematical Physics, 30(4):785–788, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p785_s1.

Cornille:1989:PDE

[1824] Henri Cornille. Positive (2 + 1)-dimensional exact shock waves solutions to the Broadwell model. Journal of Mathematical Physics, 30(4):789– 797, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i4/p789_s1.

Chang:1989:BSQ

[1825] Shau-Jin Chang and Kang Jie Shi. Billiard systems on quadric surfaces and the Poncelet theorem. Journal of Mathematical Physics, 30(4):798– 804, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i4/p798_s1.

Herrmann:1989:FUM

[1826] Robert A. Herrmann. Fractals and ultrasmooth microeffects. Journal of Mathematical Physics, 30(4):805–808, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p805_s1.

Low:1989:GSN

[1827] R. J. Low. The geometry of the space of null geodesics. Journal of Mathematical Physics, 30(4):809–811, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p809_s1.

Bergvelt:1989:NSF

[1828] M. J. Bergvelt. A note on super Fock space. Journal of Mathematical Physics, 30(4):812–815, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i4/p812_s1.

Cabra:1989:RAF

[1829] D. Cabra and F. A. Schaposnik. On regularization ambiguities in fermionic models. Journal of Mathematical Physics, 30(4):816–823, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p816_s1.

Yang:1989:GSM

[1830] Yi Song Yang. Generalized Skyrme model on higher-dimensional Riemannian manifolds. Journal of Mathematical Physics, 30(4):824–828, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p824_s1.

Singh:1989:CIB

[1831] G. S. Singh and S. N. Singh. Configurational interference in boundaryvalue problems governed by the Helmholtz equation. Journal of Mathematical Physics, 30(4):829–833, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i4/p829_s1.

Okubo:1989:IDSa

[1832] Susumu Okubo. Integrable dynamical systems with hierarchy. I. Formulation. Journal of Mathematical Physics, 30(4):834–843, April 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i4/p834_
s1.

Lai:1989:RRS

[1833] Shan Tao Lai and Ying Nan Chiu. On the recurrence relations, summations, and integrations of finite rotation matrix elements. Journal of Mathematical Physics, 30(4):844–853, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p844_s1.

Kanter:1989:OVE

[1834] Marek Kanter. Observation versus evolution: a short time theorem for positive definite functions, unitary groups, and quantum processes. Journal of Mathematical Physics, 30(4):854–857, April 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i4/p854_s1.

deLange:1989:OAS

[1835] O. L. de Lange. An operator analysis for the Schrödinger, Klein–Gordon, and Dirac equations with a Coulomb potential. Journal of Mathematical Physics, 30(4):858–866, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i4/p858_s1.

Banerjee:1989:VFS

[1836] A. Banerjee and S. B. Dutta Choudhury. Viscous fluid spheres in general relativity. Journal of Mathematical Physics, 30(4):867–871, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p867_s1.

Hauser:1989:IVPa

[1837] Isidore Hauser and Frederick J. Ernst. Initial value problem for colliding gravitational plane waves. I. Journal of Mathematical Physics, 30(4): 872–887, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i4/p872_s1.

Costa:1989:SCP

[1838] Isaias Costa. Separable coordinates and particle creation. II. Two new vacua related to accelerating observers. Journal of Mathematical Physics, 30(4):888–891, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i4/p888_s1.

Mansouri:1989:PCF

[1839] Freydoon Mansouri and Xi Zeng Wu. Parastatistics and conformal field theories in two dimensions. Journal of Mathematical Physics, 30(4): 892–901, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i4/p892_s1.

Hsu:1989:EHO

[1840] Lucas Hsu, Niky Kamran, and Peter J. Olver. Equivalence of higherorder Lagrangians. II. The Cartan form for particle Lagrangians. Journal of Mathematical Physics, 30(4):902–906, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p902_s1.

Manolessou:1989:EMIb

[1841] Marietta Manolessou. The Φ^{4} equations of motion. III. The sweeping factors. Journal of Mathematical Physics, 30(4):907–929, April 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i4/p907_
s1.

Newboult:1989:CNS

[1842] G. K. Newboult, D. F. Parker, and T. R. Faulkner. Coupled nonlinear Schrödinger equations arising in the study of monomode step-index optical fibers. Journal of Mathematical Physics, 30(4):930–936, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p930_s1.

Gunaydin:1989:VOC

[1843] M. Günaydin. Vertex operator construction of nonassociative algebras and their affinizations. Journal of Mathematical Physics, 30(4):937–942, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i4/p937_s1.

Vanthournout:1989:SRM

[1844] J. Vanthournout, H. De Meyer, and G. Vanden Berghe. On the symmetric representations of SU(5). Matrix elements of the generators in the SO(3) basis. Journal of Mathematical Physics, 30(4):943–948, April 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i4/p943_
s1.

Mathew:1989:VPC

[1845] George Mathew and M. J. Vedan. Variational principle and conservation laws for nonbarotropic flows. Journal of Mathematical Physics, 30(4): 949–952, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i4/p949_s1.

Lehnigk:1989:LND

[1846] Siegfried H. Lehnigk. On the log-normal diffusion process. Journal of Mathematical Physics, 30(4):953–955, April 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i4/p953_s1.

Fairlie:1989:EEK

[1847] D. B. Fairlie, P. Fletcher, and J. Nuyts. Eigenvalue equations for Krichever–Novikov algebras. Journal of Mathematical Physics, 30(5):

957–964, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p957_s1.

Cantoni:1989:CIR

[1848] Vittorio Cantoni. Construction of indecomposable representations of inhomogeneous Lie groups. Journal of Mathematical Physics, 30(5): 965–969, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p965_s1.

Chacon:1989:CGV

[1849] E. Chacón, P. Hess, and M. Moshinsky. Collectivity and geometry. VI. Spectra and shapes in the three-dimensional case. Journal of Mathematical Physics, 30(5):970–980, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i5/p970_s1.

Constantinescu:1989:ITS

[1850] F. Constantinescu and H. F. de Groote. The integral theorem for supersymmetric invariants. Journal of Mathematical Physics, 30(5):981–992, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p981_s1.

Plebanski:1989:LUT

[1851] Jerzy F. Plebański and J. D. Finley, III. On linear unitary transformations of two canonical variables. Journal of Mathematical Physics, 30(5): 993–999, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p993_s1.

Simon:1989:TSN

[1852] R. Simon, N. Mukunda, and E. C. G. Sudarshan. The theory of screws: a new geometric representation for the group SU(1, 1). Journal of Mathematical Physics, 30(5):1000–1006, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i5/p1000_s1.

daCosta:1989:SFR

[1853] G. A. T. F. da Costa and M. Gomes. The Schwinger functions of a rational interaction: local existence of the Borel transform. Journal of Mathematical Physics, 30(5):1007–1008, May 1989. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i5/p1007_s1.

Nahmad-Achar:1989:NDQ

[1854] E. Nahmad-Achar. A new derivation of the quadrupole formula for angular momentum. Journal of Mathematical Physics, 30(5):1009–1012, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p1009_s1.

Brugarino:1989:PPA

[1855] T. Brugarino. Painlevé property, auto-Bäcklund transformation, Lax pairs, and reduction to the standard form for the Korteweg–de Vries equation with nonuniformities. Journal of Mathematical Physics, 30(5): 1013–1015, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1013_s1.

Rao:1989:NTS

[1856] K. Srinivasa Rao and V. Rajeswari. A note on the triple sum series for the 9-j coefficient. Journal of Mathematical Physics, 30(5):1016– 1017, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1016_s1.

Bazanski:1989:HJF

[1857] Stanisław Bażański. Hamilton–Jacobi formalism for geodesics and geodesic deviations. Journal of Mathematical Physics, 30(5):1018–1029, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p1018_s1.

Coleman:1989:ADF

[1858] Robert Alan Coleman and Herbert Korte. All directing fields that are polynomial in the (n − 1) velocity are geodesic. Journal of Mathematical Physics, 30(5):1030–1033, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i5/p1030_s1.

Wene:1989:MSQ

[1859] G. P. Wene. The metric of a space and quadratic algebras. Journal of Mathematical Physics, 30(5):1034–1038, May 1989. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i5/p1034_s1.

McHugh:1989:NNT

[1860] Andrew McHugh. A Newlander–Nirenberg theorem for supermanifolds. Journal of Mathematical Physics, 30(5):1039–1042, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/p1039_s1.

Okubo:1989:IDSb

[1861] Susumu Okubo. Integrable dynamical systems with hierarchy. II. Solutions. Journal of Mathematical Physics, 30(5):1043–1052, May 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/p1043_
s1.

Doebner:1989:TEQ

[1862] H. D. Doebner, H. J. Elmers, and W. F. Heidenreich. On topological effects in quantum mechanics: the harmonic oscillator in the pointed plane. Journal of Mathematical Physics, 30(5):1053–1059, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/p1053_s1.

Boukraa:1989:TMS

[1863] S. Boukraa and J.-L. Basdevant. Technical methods for solving boundstate equations in momentum space. Journal of Mathematical Physics, 30(5):1060–1072, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i5/p1060_s1.

Jakubiec:1989:ULL

[1864] Andrzej Jakubiec and Jerzy Kijowski. On the universality of linear Lagrangians for gravitational field. Journal of Mathematical Physics, 30(5): 1073–1076, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1073_s1.

Jakubiec:1989:TGN

[1865] Andrzej Jakubiec and Jerzy Kijowski. On theories of gravitation with nonsymmetric connection. Journal of Mathematical Physics, 30(5):1077– 1080, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1077_s1.

Boyer:1989:SSD

[1866] C. P. Boyer and P. Winternitz. Symmetries of the self-dual Einstein equations. I. The infinite-dimensional symmetry group and its lowdimensional subgroups. Journal of Mathematical Physics, 30(5):1081– 1094, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1081_s1.

Mitskievic:1989:GFS

[1867] Nikolai Vsevolodovich Mitskievic and Krishnadeva K. Kumaradtya. The gravitational field of a spinning pencil of light. Journal of Mathematical Physics, 30(5):1095–1099, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i5/p1095_s1.

Nagatomo:1989:EDA

[1868] Kiyokazu Nagatomo. Explicit description of ansatz E_{n} for the Ernst equation in general relativity. Journal of Mathematical Physics, 30(5): 1100–1102, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1100_s1.

Habisohn:1989:MSF

[1869] Chris X. Habisohn. Massless scalar fields at null and spatial infinity in the Schwarzschild space–time. Journal of Mathematical Physics, 30(5): 1103–1114, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1103_s1.

Hammond:1989:SOE

[1870] Richard T. Hammond. Second-order equations from a second-order formalism. Journal of Mathematical Physics, 30(5):1115–1116, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/
p1115_s1.

Lavenda:1989:CAI

[1871] B. H. Lavenda and J. Dunning-Davies. The case against intermediate statistics. Journal of Mathematical Physics, 30(5):1117–1121, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p1117_s1.

Saito:1989:DFS

[1872] Yoshihiko Saito, Reiji Sugano, Tadayuki Ohta, and Toshiei Kimura. A dynamical formalism of singular Lagrangian system with higher derivatives. Journal of Mathematical Physics, 30(5):1122–1132, May 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/p1122_

s1. See addendum [2140].

Elizalde:1989:EZF

[1873] E. Elizalde and A. Romeo. Expressions for the zeta-function regularized Casimir energy. Journal of Mathematical Physics, 30(5):1133–1139, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/

v30/i5/p1133_s1. See erratum [2129].

Oevel:1989:SYB

[1874] Walter Oevel. R structures, Yang–Baxter equations, and related involution theorems. Journal of Mathematical Physics, 30(5):1140–1149, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p1140_s1.

Majid:1989:PWL

[1875] Shahn Majid. Partial waves for the linearized Yang–Mills equation about a monopole background. Journal of Mathematical Physics, 30(5):1150– 1157, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1150_s1.

Zhong:1989:DSF

[1876] Zai Zhe Zhong. Double-solution family of self-dual SU(2) gauge fields. Journal of Mathematical Physics, 30(5):1158–1160, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i5/p1158_s1.

Fried:1989:OED

[1877] H. M. Fried. Ordered exponentials and differential equations. Journal of Mathematical Physics, 30(5):1161–1163, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i5/p1161_s1.

Arai:1989:EIM

[1878] Asao Arai. Existence of infinitely many zero-energy states in a model of supersymmetric quantum mechanics. Journal of Mathematical Physics, 30(5):1164–1170, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i5/p1164_s1.

Schmid:1989:IDV

[1879] Rudolf Schmid and Alberto Simoni. On infinite-dimensional variational principles with constraints. Journal of Mathematical Physics, 30(5): 1171–1176, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i5/p1171_s1.

Frosali:1989:CRP

[1880] Giovanni Frosali, Cornelis V. M. van der Mee, and Stefano L. Paveri-Fontana. Conditions for runaway phenomena in the kinetic theory of particle swarms. Journal of Mathematical Physics, 30(5):1177–1186, May 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i5/p1177_s1.

Bohm:1989:PID

[1881] M. Böhm and G. Junker. Path integration over the n-dimensional Euclidean group. Journal of Mathematical Physics, 30(6):1195–1197, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i6/p1195_s1.

Li:1989:RHT

[1882] Wei Li and Bo Yu Hou. The Riemann–Hilbert transformation for an approach to a representation of the Virasoro group. Journal of Mathematical Physics, 30(6):1198–1204, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i6/p1198_s1.

Gould:1989:RTS

[1883] M. D. Gould. Representation theory of the symplectic groups. I. Journal of Mathematical Physics, 30(6):1205–1218, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1205_s1.

Klimyk:1989:RBS

[1884] A. U. Klimyk and N. Ja. Vilenkin. Relations between spherical functions of compact groups. Journal of Mathematical Physics, 30(6):1219– 1225, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1219_s1.

Malbouisson:1989:SRM

[1885] A. P. C. Malbouisson, F. R. A. Simão, and A. Ferraz de Camargo Filho. Some results from a Mellin transform expansion for the heat kernel. Journal of Mathematical Physics, 30(6):1226–1229, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1226_s1. See er-

ratum [2126] and comment [2131].

Kriele:1989:CMS

[1886] Marcus Kriele. Comment on: “A metric space construction for the boundary of space–time”, [J. Math. Phys. 27 (1986), no. 1, 124–131; MR 87c:83053 ] by D. A. Meyer. Journal of Mathematical Physics, 30(6): 1230–1232, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1230_s1. See [505].

Bruzzo:1989:CCS

[1887] Ugo Bruzzo and Daniel Hernández Ruipérez. Characteristic classes of super vector bundles. Journal of Mathematical Physics, 30(6):1233– 1237, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1233_s1.

Ali:1989:BGN

[1888] M. K. Ali and W. Robert Wood. The Birkhoff–Gustavson normal form of one-dimensional double-well Hamiltonians. Journal of Mathematical Physics, 30(6):1238–1244, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i6/p1238_s1.

Gutkin:1989:AVC

[1889] Eugene Gutkin. Asymptotic velocities in classical mechanics. Journal of Mathematical Physics, 30(6):1245–1249, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i6/p1245_s1.

Case:1989:ETD

[1890] K. M. Case and A. Monge. Explicitly time-dependent constants/ symmetries of the higher-order KP equations. Journal of Mathematical Physics, 30(6):1250–1253, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i6/p1250_s1.

Besieris:1989:BTP

[1891] Ioannis M. Besieris, Amr M. Shaarawi, and Richard W. Ziolkowski. A bidirectional traveling plane wave representation of exact solutions of the scalar wave equation. Journal of Mathematical Physics, 30(6):1254– 1269, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1254_s1.

Skarke:1989:UPS

[1892] Harald Skarke. On the uniqueness of prescriptions for spurious poles in Feynman integrals. Journal of Mathematical Physics, 30(6):1270– 1272, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1270_s1.

Fan:1989:NUO

[1893] Hong Yi Fan. A new unitary operator for a pair of coupled oscillators. Journal of Mathematical Physics, 30(6):1273–1276, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1273_s1.

Arai:1989:LTB

[1894] Asao Arai. Long-time behavior of two-point functions of a quantum harmonic oscillator interacting with bosons. Journal of Mathematical Physics, 30(6):1277–1288, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i6/p1277_s1.

Jurco:1989:CYB

[1895] Branislav Jurčo. Classical Yang–Baxter equations and quantum integrable systems. Journal of Mathematical Physics, 30(6):1289–1293, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i6/p1289_s1.

AresdeParga:1989:LP

[1896] Gonzálo Ares de Parga, Oscar Chavoya A., and José L. López Bonilla. Lanczos potential. Journal of Mathematical Physics, 30(6):1294–1295, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i6/p1294_s1.

Lester:1989:ZLR

[1897] J. A. Lester. Zeeman’s lemma on Robertson–Walker space–times. Journal of Mathematical Physics, 30(6):1296–1300, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1296_s1.

Whiting:1989:MSK

[1898] Bernard F. Whiting. Mode stability of the Kerr black hole. Journal of Mathematical Physics, 30(6):1301–1305, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i6/p1301_s1.

Dray:1989:DCS

[1899] Tevian Dray, Ravi Kulkarni, and Joseph Samuel. Duality and conformal structure. Journal of Mathematical Physics, 30(6):1306–1309, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i6/p1306_s1.

GarciaDiaz:1989:MGC

[1900] Alberto García Díaz and Nora Bretón Baez. Magnetic generalizations of the Carter metrics. Journal of Mathematical Physics, 30(6):1310– 1315, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1310_s1.

Duggal:1989:RFM

[1901] K. L. Duggal. Relativistic fluids and metric symmetries. Journal of Mathematical Physics, 30(6):1316–1322, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i6/p1316_s1.

TorresdelCastillo:1989:DPR

[1902] G. F. Torres del Castillo. Debye potentials for Rarita–Schwinger fields in curved space–times. Journal of Mathematical Physics, 30(6):1323– 1328, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1323_s1.

Chen:1989:TSB

[1903] M. Chen. On the thermodynamic solution of the Boltzmann equation and nonlinear irreversible thermodynamics. Journal of Mathematical Physics, 30(6):1329–1337, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i6/p1329_s1.

Arik:1989:MHS

[1904] Metin Arik, Fahrünisa Neyzi, Yavuz Nutku, Peter J. Olver, and John M. Verosky. Multi-Hamiltonian structure of the Born–Infeld equation. Journal of Mathematical Physics, 30(6):1338–1344, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1338_s1.

Batlle:1989:NTG

[1905] C. Batlle, J. Gomis, X. Gràcia, and J. M. Pons. Noether’s theorem and gauge transformations: Application to the bosonic string and CP_{2}^{n−1} model. Journal of Mathematical Physics, 30(6):1345–1350, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i6/
p1345_s1.

deLeon:1989:ATS

[1906] Manuel de León and Paulo R. Rodrigues. n^{k} almost-tangent structures and the Hamiltonization of higher-order field theories. Journal of Mathematical Physics, 30(6):1351–1353, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i6/p1351_s1.

Zastawniak:1989:NPS

[1907] Tomasz Zastawniak. The nonexistence of the path-space measure for the Dirac equation in four space–time dimensions. Journal of Mathematical Physics, 30(6):1354–1358, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i6/p1354_s1.

Bandyopadhyay:1989:SFT

[1908] Pratul Bandyopadhyay, Brahmananda Bhattacharya, and Pradip Ghosh. Stochastic field theory at finite temperature and equilibrium states in quantum statistical mechanics. Journal of Mathematical Physics, 30(6): 1359–1364, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1359_s1.

Zhang:1989:ITC

[1909] Yu Shun Zhang. Inelastic two-composite-particle system scattering theory at high energy. Journal of Mathematical Physics, 30(6):1365–1372, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i6/p1365_s1.

ONeil:1989:HDV

[1910] Kevin A. O’Neil. On the Hamiltonian dynamics of vortex lattices. Journal of Mathematical Physics, 30(6):1373–1379, June 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1373_s1.

Maeder:1989:NNL

[1911] J. M. Maeder and R. B. McQuistan. Nearest neighbor lattice statistics on semi-infinite two-dimensional rectangular lattices of various widths. Journal of Mathematical Physics, 30(6):1380–1384, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i6/p1380_s1.

Fassari:1989:SPK

[1912] Silvestro Fassari. Spectral properties of the Kronig–Penney Hamiltonian with a localized impurity. Journal of Mathematical Physics, 30(6):1385– 1392, June 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i6/p1385_s1.

DeBievre:1989:CSS

[1913] S. De Bièvre. Coherent states over symplectic homogeneous spaces. Journal of Mathematical Physics, 30(7):1401–1407, July 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1401_s1.

Walton:1989:CBR

[1914] Mark A. Walton. Conformal branching rules from Kac–Moody automorphisms. Journal of Mathematical Physics, 30(7):1408–1414, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1408_s1.

LeBlanc:1989:HWR

[1915] R. Le Blanc and D. J. Rowe. Highest weight representations for gl(m/n) and gl(m + n). Journal of Mathematical Physics, 30(7):1415–1432, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1415_s1.

Palev:1989:IFD

[1916] Tchavdar D. Palev. Irreducible finite-dimensional representations of the Lie superalgebra gl(n/1) in a Gel’fand–Zetlin basis. Journal of Mathematical Physics, 30(7):1433–1442, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i7/p1433_s1.

Cooperstock:1989:RDC

[1917] F. I. Cooperstock, P. H. Lim, and P. C. Peters. The role of discontinuities in conservation laws. Journal of Mathematical Physics, 30(7): 1443–1446, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1443_s1.

Bender:1989:NPA

[1918] C. M. Bender, Kimball A. Milton, Stephen S. Pinsky, and L. M. Simmons, Jr. A new perturbative approach to nonlinear problems. Journal of Mathematical Physics, 30(7):1447–1455, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i7/p1447_s1.

Milev:1989:DLP

[1919] N. V. Milev and D. D. Baĭnov. Dichotomies for linear periodic differential equations with impulses. Journal of Mathematical Physics, 30(7):1456– 1458, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1456_s1.

Baker-Jarvis:1989:SDE

[1920] James Baker-Jarvis, Michael Racine, and Jihad Alameddine. Solving differential equations by a maximum entropy-minimum norm method with applications to Fokker–Planck equations. Journal of Mathematical Physics, 30(7):1459–1463, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i7/p1459_s1.

Fatyga:1989:GVF

[1921] B. W. Fatyga, V. Alan Kostelecký, and D. Rodney Truax. Grassmannvalued fluid dynamics. Journal of Mathematical Physics, 30(7):1464– 1472, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1464_s1.

Cardin:1989:CMS

[1922] Franco Cardin and Giovanni Zanzotto. On constrained mechanical systems: d’Alembert’s and Gauss’ principles. Journal of Mathematical Physics, 30(7):1473–1479, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i7/p1473_s1.

Ewen:1989:GFF

[1923] Holger Ewen and Heinz-Jürgen Schmidt. Geometry of free fall and simultaneity. Journal of Mathematical Physics, 30(7):1480–1486, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1480_s1.

Schilling:1989:SAS

[1924] Randolph James Schilling. A systematic approach to the soliton equations of a discrete eigenvalue problem. Journal of Mathematical Physics, 30(7):1487–1501, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i7/p1487_s1.

Jaen:1989:LFR

[1925] X. Jaén, J. Llosa, and A. Molina. Lagrangian formalism and retarded classical electrodynamics. Journal of Mathematical Physics, 30(7):1502– 1504, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1502_s1.

Jaworski:1989:CTS

[1926] W. Jaworski. The concept of a time-of-sojourn operator and spreading of wave packets. Journal of Mathematical Physics, 30(7):1505–1514, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1505_s1.

Takasaki:1989:INH

[1927] K. Takasaki. An infinite number of hidden variables in hyper-Kähler metrics. Journal of Mathematical Physics, 30(7):1515–1521, July 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1515_
s1.

Fernandez:1989:TEO

[1928] Francisco M. Fernández. On the time evolution operator for timedependent quadratic Hamiltonians. Journal of Mathematical Physics, 30(7):1522–1524, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i7/p1522_s1.

Taylor:1989:ESS

[1929] D. R. Taylor and P. G. L. Leach. Exact solutions of the Schrödinger equation for nonseparable anharmonic oscillator potentials in two dimensions. Journal of Mathematical Physics, 30(7):1525–1532, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1525_s1.

Chandler:1989:BQS

[1930] Colston Chandler and A. G. Gibson. N-body quantum scattering theory in two Hilbert spaces. V. Computation strategy. Journal of Mathematical Physics, 30(7):1533–1544, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i7/p1533_s1. See erratum [2139].

Doughty:1989:GBQ

[1931] Noel A. Doughty and Richard A. Arnold. Gupta–Bleuler quantization of free massless Lagrangian gauge fields of arbitrary helicity: the bosonic case. Journal of Mathematical Physics, 30(7):1545–1553, July 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1545_
s1.

Sahni:1989:EGE

[1932] D. C. Sahni. Equivalence of generic equation method and the phenomenological model for linear transport problems in a two-state random scattering medium. Journal of Mathematical Physics, 30(7):1554– 1559, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1554_s1.

Davidson:1989:SCS

[1933] W. Davidson. A static cylindrically symmetric solution for perfect fluid in general relativity. Journal of Mathematical Physics, 30(7):1560–1561, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1560_s1.

Economou:1989:MSS

[1934] A. Economou and D. Tsoubelis. Multiple-soliton solutions of Einstein’s equations. Journal of Mathematical Physics, 30(7):1562–1569, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/
p1562_s1.

Cho:1989:HDU

[1935] Y. M. Cho and D. S. Kimm. Higher-dimensional unification by isometry. Journal of Mathematical Physics, 30(7):1570–1578, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1570_s1.

Rosenbaum:1989:GMG

[1936] M. Rosenbaum, J. C. D’Olivo, E. Nahmad-Achar, R. Bautista, and J. Muciño. Geometric model for gravitational and electroweak interactions. Journal of Mathematical Physics, 30(7):1579–1590, July 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1579_
s1.

Zak:1989:NLM

[1937] J. Zak. von Neumann lattices and magnetic orbitals in finite phase space. Journal of Mathematical Physics, 30(7):1591–1596, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1591_s1.

Semmler:1989:SRE

[1938] U. Semmler. Schrödinger representation in Euclidean quantum field theory. Journal of Mathematical Physics, 30(7):1597–1605, July 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1597_
s1.

Carillo:1989:ASS

[1939] Sandra Carillo and Benno Fuchssteiner. The abundant symmetry structure of hierarchies of nonlinear equations obtained by reciprocal links. Journal of Mathematical Physics, 30(7):1606–1613, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1606_s1.

Lou:1989:RAS

[1940] Sen Yue Lou and Guang Jiong Ni. The relations among a special type of solutions in some (D + 1)-dimensional nonlinear equations. Journal of Mathematical Physics, 30(7):1614–1620, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i7/p1614_s1.

Hostler:1989:FWT

[1941] Levere Hostler. Foldy–Wouthuysen-type reduction of the second-order Dirac equation. Journal of Mathematical Physics, 30(7):1621–1630, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i7/p1621_s1.

Atakishiyev:1989:GCS

[1942] N. M. Atakishiyev and R. M. Mir-Kasimov. Generalized coherent states for a three-dimensional relativistic model of the oscillator. Journal of Mathematical Physics, 30(7):1631–1634, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i7/p1631_s1.

Nachbagauer:1989:GDC

[1943] Herbert Nachbagauer, Ulrike Kraemmer, and Anton Rebhan. Gauge dependences of the covariant effective action in QED. Journal of Mathematical Physics, 30(7):1635–1639, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i7/p1635_s1.

SalazarI:1989:SRE

[1944] Humberto Salazar I. Some remarks on Einstein–Maxwell solitons. Journal of Mathematical Physics, 30(7):1640–1642, July 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1640_s1.

Merle:1989:NES

[1945] Frank Merle. A note on the existence of a solution for a model of McLaughlin, Moloney, and Newell for optical ring cavities. Journal of Mathematical Physics, 30(7):1643–1647, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i7/p1643_s1.

Broadbridge:1989:ESM

[1946] P. Broadbridge. Exact solvability of the Mullins nonlinear diffusion model of groove development. Journal of Mathematical Physics, 30(7):1648– 1651, July 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i7/p1648_s1.

Majid:1989:ELT

[1947] Shahn Majid. Erratum: “Liouville theorem for the Yang–Mills selfduality equations”, [J. Math. Phys. 29 (1988), no. 10, 2303–2310; MR 89i:58153]. Journal of Mathematical Physics, 30(7):1652, July 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i7/p1652_

s1. See [1658].

Beckers:1989:COL

[1948] J. Beckers and J. F. Cornwell. On chains of orthosymplectic Lie superalgebras and the n-dimensional quantum harmonic oscillator. Journal of Mathematical Physics, 30(8):1655–1661, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1655_s1.

Jaulent:1989:MLG

[1949] Marcel Jaulent, Miguel A. Manna, and Luis Martínez Alonso. Multiseries Lie groups and asymptotic modules for characterizing and solving integrable models. Journal of Mathematical Physics, 30(8):1662– 1673, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1662_s1.

Sciarrino:1989:ELA

[1950] A. Sciarrino. Exceptional Lie algebras in terms of fermionic creationannihilation operators. Journal of Mathematical Physics, 30(8):1674– 1678, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1674_s1.

Zeng:1989:CPF

[1951] Yun Bo Zeng and Yi Shen Li. The constraints of potentials and the finite-dimensional integrable systems. Journal of Mathematical Physics, 30(8):1679–1689, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1679_s1.

Li:1989:PFI

[1952] Miao Li. Proof of the family index theorem via a path integral. Journal of Mathematical Physics, 30(8):1690–1694, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1690_s1.

Neyzi:1989:DHS

[1953] Fahrünisa Neyzi. Diagonalization and Hamiltonian structures of hyperbolic systems. Journal of Mathematical Physics, 30(8):1695–1698, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1695_s1.

Swierczynski:1989:CPY

[1954] Z. Świerczyński. On the Cauchy problem for Yang–Mills equations with external current. Journal of Mathematical Physics, 30(8):1699–1706, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1699_s1.

Wanelik:1989:ISS

[1955] Kazimierz Wanelik. On the iterative solutions of some nonlinear eigenvalue problems. Journal of Mathematical Physics, 30(8):1707–1712, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1707_s1.

Lozada:1989:GFE

[1956] Abraham Lozada. General form of the equation of motion for a point charge. Journal of Mathematical Physics, 30(8):1713–1720, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i8/
p1713_s1.

Sutherland:1989:NPM

[1957] Roderick I. Sutherland. Null path model for classical interactions. Journal of Mathematical Physics, 30(8):1721–1726, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1721_s1.

Benedict:1989:MIB

[1958] M. G. Benedict, L. Gy. Fehér, and Z. Horváth. Monopoles and instantons from Berry’s phase. Journal of Mathematical Physics, 30(8):1727– 1731, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1727_s1.

Beckers:1989:GCS

[1959] J. Beckers and N. Debergh. On generalized coherent states with maximal symmetry for the harmonic oscillator. Journal of Mathematical Physics, 30(8):1732–1738, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1732_s1.

Jurco:1989:QIM

[1960] Branislav Jurčo. On quantum integrable models related to nonlinear quantum optics. An algebraic Bethe ansatz approach. Journal of Mathematical Physics, 30(8):1739–1743, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i8/p1739_s1.

Vaklev:1989:GTQ

[1961] Y. S. Vaklev. Gauge transformations for the quadratic bundle. Journal of Mathematical Physics, 30(8):1744–1755, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1744_s1. See er-

ratum [2130].

Zhou:1989:CEQ

[1962] Yu Kui Zhou and Guo-Hun Yun. Construction of the eigenstates for the quantum nonlinear Schrödinger model with the most general supermatrices. Journal of Mathematical Physics, 30(8):1756–1759, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1756_s1.

Moncrief:1989:ABN

[1963] Vincent Moncrief. The asymptotic behavior of nonlinear waves near a cosmological Cauchy horizon. Journal of Mathematical Physics, 30 (8):1760–1768, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1760_s1.

Kuchar:1989:WSD

[1964] K. V. Kuchař and C. G. Torre. World sheet diffeomorphisms and the canonical string. Journal of Mathematical Physics, 30(8):1769–1793, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1769_s1.

Bazanski:1989:GDS

[1965] Stanisław L. Bażański and Piotr Jaranowski. Geodesic deviation in the Schwarzschild space–time. Journal of Mathematical Physics, 30(8):1794– 1803, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1794_s1.

Banach:1989:ITDa

[1966] Zbigniew Banach and Slawomir Piekarski. Irreducible tensor description. I. A classical gas. Journal of Mathematical Physics, 30(8):1804– 1815, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1804_s1.

Banach:1989:ITDb

[1967] Zbigniew Banach and Slawomir Piekarski. Irreducible tensor description. II. A quasiparticle gas. Journal of Mathematical Physics, 30(8):1816– 1825, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1816_s1.

Banach:1989:ITDc

[1968] Zbigniew Banach and Slawomir Piekarski. Irreducible tensor description. III. Thermodynamics of a low-temperature phonon gas. Journal of Mathematical Physics, 30(8):1826–1836, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1826_s1.

Lee:1989:CPD

[1969] M. Howard Lee. Chemical potential of a D-dimensional free Fermi gas at finite temperatures. Journal of Mathematical Physics, 30(8):1837– 1839, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1837_s1.

Courbage:1989:ASE

[1970] M. Courbage. On the approach to statistical equilibrium in an infiniteparticle lattice dynamic model. Journal of Mathematical Physics, 30 (8):1840–1850, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1840_s1.

Chimento:1989:MRP

[1971] Luis P. Chimento and Alejandro S. Jakubi. A modified renormalization procedure which may avoid use of bare parameters. Journal of Mathematical Physics, 30(8):1851–1860, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i8/p1851_s1.

Xu:1989:NHD

[1972] Jian Jun Xu and Xin Zhou Li. New higher-dimensional instanton and monopole solutions. Journal of Mathematical Physics, 30(8):1861– 1865, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1861_s1.

Aldrovandi:1989:PCF

[1973] R. Aldrovandi and R. A. Kraenkel. Perturbative coherence in field theory. Journal of Mathematical Physics, 30(8):1866–1870, August 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1866_
s1.

DiazBejarano:1989:GFT

[1974] J. Díaz Bejarano and A. Martín Sánchez. Generalized Fourier transforms for nonlinear systems. Journal of Mathematical Physics, 30(8):1871– 1876, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1871_s1.

Marzban:1989:HRM

[1975] C. Marzban, B. F. Whiting, and H. Van Dam. Hamiltonian reduction for massive fields coupled to sources. Journal of Mathematical Physics, 30(8):1877–1892, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1877_s1.

Longhi:1989:MTF

[1976] G. Longhi, L. Lusanna, and J. M. Pons. On the many-time formulation of classical particle dynamics. Journal of Mathematical Physics, 30(8):1893–1912, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i8/p1893_s1.

Brihaye:1989:SCI

[1977] Yves Brihaye and Jutta Kunz. Summable chains of instantons and their symmetries. Journal of Mathematical Physics, 30(8):1913–1917, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i8/p1913_s1.

Ardalan:1989:TFR

[1978] F. Ardalan, H. Arfaei, and M. Zarkesh. Theta function representation of modular group for SU(n). Journal of Mathematical Physics, 30(8):1918– 1920, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1918_s1.

Gorman:1989:CPA

[1979] N. Gorman, L. O’Raifeartaigh, and W. McGlinn. Cartan-preserving automorphisms of untwisted and twisted Kac–Moody algebras. Journal of Mathematical Physics, 30(8):1921–1932, August 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i8/p1921_s1.

Buendia:1989:PDT

[1980] G. M. Buendía, U. Percoco, and A. Trias. Path-dependent techniques for electromagnetism with magnetic charges. Journal of Mathematical Physics, 30(8):1933–1936, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i8/p1933_s1.

Long:1989:AWL

[1981] Gui Lu Long and Hong Zhou Sun. The analytical wavefunctions of the SU(3) limit in the interacting boson model. Journal of Mathematical Physics, 30(8):1937–1941, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i8/p1937_s1.

Chen:1989:NVG

[1982] Y. Y. Chen. Nonsymmetric vortices for the Ginzberg–Landau equations on the bounded domain. Journal of Mathematical Physics, 30(8):1942– 1950, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1942_s1.

Bartocci:1989:ECS

[1983] C. Bartocci and U. Bruzzo. Erratum: “Cohomology of the structure sheaf of real and complex supermanifolds”, [J. Math. Phys. 29 (1988), no. 8, 1789–1794; MR 89k:58023]. Journal of Mathematical Physics, 30(8): 1951, August 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i8/p1951_s1. See [1582].

Ekhaguere:1989:UPH

[1984] G. O. S. Ekhaguere. Unbounded partial Hilbert algebras. Journal of Mathematical Physics, 30(9):1957–1964, September 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i9/p1957_s1.

dePrunele:1989:VMC

[1985] E. de Prunelé. SU(1, 1) and the vector model for continuous bases. Journal of Mathematical Physics, 30(9):1965–1970, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p1965_s1.

Shaw:1989:FGC

[1986] R. Shaw. Finite geometries and Clifford algebras. Journal of Mathematical Physics, 30(9):1971–1984, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p1971_s1.

Lev:1989:MRP

[1987] Felix Lev. Modular representations as a possible basis of finite physics. Journal of Mathematical Physics, 30(9):1985–1998, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p1985_
s1.

Rashid:1989:SPR

[1988] M. A. Rashid. A simple proof of the result that the Wigner transformation is of finite order. Journal of Mathematical Physics, 30(9):1999–2000, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p1999_s1.

Tratnik:1989:MWP

[1989] M. V. Tratnik. Multivariable Wilson polynomials. Journal of Mathematical Physics, 30(9):2001–2011, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2001_s1.

Montaldi:1989:ADP

[1990] Emilio Montaldi and Giuseppe Zucchelli. Additive decomposition for the product of two θ_{3} functions and modular equations. Journal of Mathematical Physics, 30(9):2012–2015, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2012_s1.

Malbouisson:1989:MTT

[1991] A. P. C. Malbouisson, M. A. R. Monteiro, and F. R. A. Simão. A Mellin transform technique for the heat kernel expansion. Journal of Mathematical Physics, 30(9):2016–2018, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2016_s1.

Gaveau:1989:STP

[1992] B. Gaveau and L. S. Schulman. Sensitive terms in the path integral: ordering and stochastic options. Journal of Mathematical Physics, 30 (9):2019–2022, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i9/p2019_s1.

MenciaBravo:1989:CPH

[1993] J. Mencia Bravo, R. M. Velasco, and J. M. Sancho. A classical particle in heat bath under the influence of external noise. Journal of Mathematical Physics, 30(9):2023–2027, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i9/p2023_s1.

Thomas:1989:BMG

[1994] Lawrence E. Thomas. Bound on the mass gap for a stochastic contour model at low temperature. Journal of Mathematical Physics, 30(9):2028– 2034, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p2028_s1.

Bose:1989:DMC

[1995] Asok Bose. Dynkin’s method of computing the terms of the Baker– Campbell–Hausdorff series. Journal of Mathematical Physics, 30(9): 2035–2037, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i9/p2035_s1.

Katzin:1989:CFS

[1996] Gerald H. Katzin and Jack Levine. Characteristic functional structure of infinitesimal symmetry mappings of classical dynamical systems. IV. Classical (velocity-independent) mappings of second-order differential equations. Journal of Mathematical Physics, 30(9):2038–2053, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2038_s1.

Pardo:1989:HCT

[1997] Francisco Pardo. The Helmholtz conditions in terms of constants of motion in classical mechanics. Journal of Mathematical Physics, 30 (9):2054–2061, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i9/p2054_s1.

Teixeira:1989:CTG

[1998] Julio M. Teixeira, Luiz J. Negri, and Luimar C. De Oliveira. Canonoid transformations in generalized mechanics. Journal of Mathematical Physics, 30(9):2062–2064, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i9/p2062_s1.

Chan:1989:BTC

[1999] W. L. Chan and Yu Kun Zheng. Bäcklund transformations for the Caudrey–Dodd–Gibbon–Sawada–Kotera equation and its λ-modified equation. Journal of Mathematical Physics, 30(9):2065–2068, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2065_s1.

Stephas:1989:FWF

[2000] Paul Stephas. Fokker–Wheeler–Feynman interactions without integrals. Journal of Mathematical Physics, 30(9):2069–2071, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p2069_
s1.

Leese:1989:EWS

[2001] Robert Leese. Extended wave solutions in an integrable chiral model in (2 + 1) dimensions. Journal of Mathematical Physics, 30(9):2072–2077, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p2072_s1.

Schroeck:1989:EFU

[2002] Franklin E. Schroeck, Jr. On the entropic formulation of uncertainty for quantum measurements. Journal of Mathematical Physics, 30(9):2078– 2082, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p2078_s1.

Alberti:1989:BAT

[2003] Peter M. Alberti and Volker Heinemann. Bounds for the C^{∗}-algebraic transition probability yield best lower and upper bounds to the overlap. Journal of Mathematical Physics, 30(9):2083–2089, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p2083_
s1.

Chrusciel:1989:SSIa

[2004] Piotr T. Chruściel. On the structure of spatial infinity. I. The Geroch structure. Journal of Mathematical Physics, 30(9):2090–2093, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2090_s1.

Chrusciel:1989:SSIb

[2005] Piotr T. Chruściel. On the structure of spatial infinity. II. Geodesically regular Ashtekar–Hansen structures. Journal of Mathematical Physics, 30(9):2094–2100, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i9/p2094_s1.

Plebanski:1989:TSC

[2006] J. F. Plebański and M. Przanowski. Two-sided conformally recurrent four-dimensional Riemannian manifolds. Journal of Mathematical Physics, 30(9):2101–2109, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i9/p2101_s1.

Rago:1989:NCS

[2007] H. Rago. Nonstatic charged spheres admitting a conformal Killing vector. Journal of Mathematical Physics, 30(9):2110–2113, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p2110_
s1.

TorresdelCastillo:1989:SPA

[2008] G. F. Torres del Castillo. Spin-3/2 perturbations of algebraically special solutions of the Einstein–Maxwell equations. Journal of Mathematical Physics, 30(9):2114–2119, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i9/p2114_s1.

Lopes:1989:ACD

[2009] Artur O. Lopes. An analogy of the charge distribution on Julia sets with the Brownian motion. Journal of Mathematical Physics, 30(9):2120– 2124, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p2120_s1.

Budinich:1989:FSD

[2010] Paolo Budinich and Andrzej Trautman. Fock space description of simple spinors. Journal of Mathematical Physics, 30(9):2125–2131, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2125_s1.

Shishkin:1989:DEEa

[2011] German V. Shishkin and Víctor M. Villalba. Dirac equation in external vector fields: separation of variables. Journal of Mathematical Physics, 30(9):2132–2142, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i9/p2132_s1.

Cho:1989:DEZ

[2012] Hing Tong Cho, Kimball A. Milton, Stephen S. Pinsky, and L. M. Simmons, Jr. The delta expansion in zero dimensions. Journal of Mathematical Physics, 30(9):2143–2150, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2143_s1.

Ferreira:1989:HES

[2013] Sabino José Ferreira, Ricardo S. Schor, and Michael L. O’Carroll. Highly excited stable bound states in strongly coupled lattice gauge theories. Journal of Mathematical Physics, 30(9):2151–2155, September 1989. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i9/p2151_
s1.

Szlachanyi:1989:CET

[2014] K. Szlachànyi and P. Vecsernyès. Cluster expansion in terms of knots in gauge theories with finite nonabelian gauge groups. Journal of Mathematical Physics, 30(9):2156–2159, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2156_s1.

Botelho:1989:SWE

[2015] Luiz C. L. Botelho. String wave equations in Polyakov’s path integral framework. Journal of Mathematical Physics, 30(9):2160–2167, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2160_s1.

Ho:1989:FTG

[2016] Choon Lin Ho. Field theory of geometric p-branes. Journal of Mathematical Physics, 30(9):2168–2176, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i9/p2168_s1.

Vassiliadis:1989:WLG

[2017] Dimitris V. Vassiliadis. Wigner’s little group and decomposition of Lorentz transformations. Journal of Mathematical Physics, 30(9):2177– 2180, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i9/p2177_s1.

Carey:1989:GAG

[2018] Alan Carey and John Palmer. Gauge anomalies on S^{2} and group extensions. Journal of Mathematical Physics, 30(9):2181–2191, September 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i9/p2181_s1.

Battle:1989:PSL

[2019] Guy Battle. Phase space localization theorem for ondelettes. Journal of Mathematical Physics, 30(10):2195–2196, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2195_s1.

Perjes:1989:FST

[2020] Zoltán Perjés. Factor structure of the Tomimatsu–Sato metrics. Journal of Mathematical Physics, 30(10):2197–2200, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2197_s1.

Clarkson:1989:NSR

[2021] Peter A. Clarkson and Martin D. Kruskal. New similarity reductions of the Boussinesq equation. Journal of Mathematical Physics, 30(10):2201– 2213, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2201_s1.

Cleary:1989:IOQ

[2022] Paul W. Cleary. Integrability and orbits in quartic polynomial potentials. Journal of Mathematical Physics, 30(10):2214–2225, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i10/
p2214_s1.

Blau:1989:GCL

[2023] Matthias Blau. Group cocycles, line bundles, and anomalies. Journal of Mathematical Physics, 30(10):2226–2232, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2226_s1.

Piette:1989:SCG

[2024] Bernard Piette and Wojciech J. Zakrzewski. Some classes of general solutions of the U(N) chiral σ models in two dimensions. Journal of Mathematical Physics, 30(10):2233–2237, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2233_s1.

Urbantke:1989:LDG

[2025] H. Urbantke. Local differential geometry of null curves in conformally flat space–time. Journal of Mathematical Physics, 30(10):2238–2245, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i10/p2238_s1.

Ward:1989:TD

[2026] R. S. Ward. Twistors in 2 + 1 dimensions. Journal of Mathematical Physics, 30(10):2246–2251, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i10/p2246_s1.

Fodor:1989:MMA

[2027] G. Fodor, C. Hoenselaers, and Z. Perjés. Multipole moments of axisymmetric systems in relativity. Journal of Mathematical Physics, 30 (10):2252–2257, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2252_s1.

Carinena:1989:PMC

[2028] José F. Cariñena and Manuel F. Rañada. Poisson maps and canonoid transformations for time-dependent Hamiltonian systems. Journal of Mathematical Physics, 30(10):2258–2266, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2258_s1.

Harnett:1989:TUA

[2029] Gerald Harnett. Twistors for uniform acceleration. Journal of Mathematical Physics, 30(10):2267–2272, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i10/p2267_s1.

Colombeau:1989:ESP

[2030] Jean-François Colombeau. The elastoplastic shock problem as an example of the resolution of ambiguities in the multiplication of distributions. Journal of Mathematical Physics, 30(10):2273–2279, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i10/p2273_s1.

Das:1989:CEP

[2031] A. Das and D. Kay. A class of exact plane wave solutions of the Maxwell– Dirac equations. Journal of Mathematical Physics, 30(10):2280–2284, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2280_s1.

Blinder:1989:GRA

[2032] S. M. Blinder and E. L. Pollock. Generalized relations among Ndimensional Coulomb Green’s functions using fractional derivatives. Journal of Mathematical Physics, 30(10):2285–2287, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i10/
p2285_s1.

Silvestre-Brac:1989:FGF

[2033] B. Silvestre-Brac, C. Gignoux, and Y. Ayant. Free Green’s function in a harmonic oscillator basis. Journal of Mathematical Physics, 30(10):2288– 2290, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2288_s1.

ArredondoR:1989:ATP

[2034] J. H. Arredondo R. Asymptotic transition probabilities. Journal of Mathematical Physics, 30(10):2291–2296, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2291_s1.

Moncrief:1989:SPD

[2035] Vincent Moncrief. Symmetry preserving deformations of generalized Taub–NUT space–times to all orders in perturbation theory. Journal of Mathematical Physics, 30(10):2297–2314, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2297_s1.

Krori:1989:SSS

[2036] K. D. Krori, P. Borgohain, and Kanika Das. Spherically symmetric solutions in higher dimensions. Journal of Mathematical Physics, 30 (10):2315–2318, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2315_s1.

Rosquist:1989:KTC

[2037] Kjell Rosquist. Killing tensor conservation laws and their generators. Journal of Mathematical Physics, 30(10):2319–2321, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i10/
p2319_s1.

Hauser:1989:IVPb

[2038] Isidore Hauser and Frederick J. Ernst. Initial value problem for colliding gravitational plane waves. II. Journal of Mathematical Physics, 30(10):2322–2336, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2322_s1.

Baumann:1989:TGR

[2039] Klaus Baumann. The theorem of Greenberg and Robinson for twodimensional quantum field theories. Journal of Mathematical Physics, 30(10):2337–2339, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2337_s1.

Bruning:1989:HQF

[2040] Erwin Brüning and Shigeaki Nagamachi. Hyperfunction quantum field theory: basic structural results. Journal of Mathematical Physics, 30 (10):2340–2359, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2340_s1.

Kalnins:1989:KYT

[2041] E. G. Kalnins, W. Miller, Jr., and G. C. Williams. Killing–Yano tensors and variable separation in Kerr geometry. Journal of Mathematical Physics, 30(10):2360–2365, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i10/p2360_s1.

Roy:1989:TAF

[2042] Ashim Roy and Pratul Bandyopadhyay. Topological aspects of a fermion and the chiral anomaly. Journal of Mathematical Physics, 30(10):2366– 2372, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2366_s1.

Shishkin:1989:DEEb

[2043] German V. Shishkin and Víctor M. Villalba. Dirac equation in external vector fields: new exact solutions. Journal of Mathematical Physics, 30(10):2373–2381, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2373_s1.

Lopez:1989:EIP

[2044] M. C. López, R. J. Noriega, and C. G. Schifini. The equivariant inverse problem and the uniqueness of the Yang–Mills equations. Journal of Mathematical Physics, 30(10):2382–2387, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2382_s1. See ad-

dendum [2133].

Zemba:1989:TGC

[2045] Guillermo Zemba and Barton Zwiebach. Tadpole graph in covariant closed string field theory. Journal of Mathematical Physics, 30(10):2388– 2396, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2388_s1.

Nomura:1989:RCG

[2046] Masao Nomura. Relations for Clebsch–Gordan and Racah coefficients in su_{q}(2) and Yang–Baxter equations. Journal of Mathematical Physics, 30(10):2397–2405, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2397_s1.

Papaloucas:1989:VCS

[2047] L. C. Papaloucas, J. Rembieliński, and W. Tybor. Vectorlike coherent states with noncompact stability group. Journal of Mathematical Physics, 30(10):2406–2410, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i10/p2406_s1.

Rideau:1989:UPF

[2048] G. Rideau. Uniqueness problems in formal nonlinear representations of the Poincaré group in 2+1 dimensions. Journal of Mathematical Physics, 30(10):2411–2418, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i10/p2411_s1.

Filippov:1989:MCN

[2049] G. F. Filippov and A. L. Blokhin. Microscopic collective nuclear models with horizontal mixture. Journal of Mathematical Physics, 30(10):2419– 2431, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i10/p2419_s1.

Manning:1989:RTT

[2050] Robert M. Manning. Radiative transfer theory for inhomogeneous media with random extinction and scattering coefficients. Journal of Mathematical Physics, 30(10):2432–2440, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i10/p2432_s1.

Elshehawey:1989:IER

[2051] Elsayed F. Elshehawey. Intervals of electrohydrodynamic Rayleigh– Taylor instability. I. Effect of a tangential periodic field. Journal of Mathematical Physics, 30(10):2441–2444, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2441_s1.

Palczewski:1989:GSB

[2052] A. Palczewski and G. Toscani. Global solution of the Boltzmann equation for rigid spheres and initial data close to a local Maxwellian. Journal of Mathematical Physics, 30(10):2445–2450, October 1989. CODEN JMA-PAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL

http://jmp.aip.org/resource/1/jmapaq/v30/i10/p2445_s1.

Holm:1989:EHS

[2053] Christian Holm. Erratum: “The hyperspin structure of unitary groups”, [J. Math. Phys. 29 (1988), no. 4, 978–986; MR 90b:83081]. Journal of Mathematical Physics, 30(10):2451, October 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i10/p2451_s1. See [1471].

Atakishiyev:1989:RCAa

[2054] Natig M. Atakishiyev, Wolfgang Lassner, and Kurt Bernardo Wolf. The relativistic coma aberration. I. Geometrical optics. Journal of Mathematical Physics, 30(11):2457–2462, November 1989. CODEN JMAPAQ.

ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i11/p2457_s1.

Atakishiyev:1989:RCAb

[2055] Natig M. Atakishiyev, Wolfgang Lassner, and Kurt Bernardo Wolf. The relativistic coma aberration. II. Helmholtz wave optics. Journal of Mathematical Physics, 30(11):2463–2468, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i11/p2463_s1.

Doeraene:1989:BRH

[2056] J. P. Doeraene and G. Iommi Amunátegui. Branching rules for the hyperoctahedral group. Journal of Mathematical Physics, 30(11):2469–2475, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2469_s1.

Grigore:1989:CCG

[2057] D. R. Grigore and O. T. Popp. The complete classification of generalized homogeneous symplectic manifolds. Journal of Mathematical Physics, 30 (11):2476–2483, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2476_s1. See erratum [2135].

Robinson:1989:RNP

[2058] P. A. Robinson. Relativistic and nonrelativistic plasma dispersion functions. Journal of Mathematical Physics, 30(11):2484–2487, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i11/p2484_s1.

Hajicek:1989:TPS

[2059] Peter Hajicek. Topology of parametrized systems. Journal of Mathematical Physics, 30(11):2488–2497, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i11/p2488_s1.

Sanchez:1989:LKT

[2060] Richard Sanchez. Linear kinetic theory in stochastic media. Journal of Mathematical Physics, 30(11):2498–2511, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/p2498_s1.

Leyvraz:1989:SPT

[2061] F. Leyvraz and T. H. Seligman. Sequences of point transformations and linear canonical transformations in classical and quantum mechanics. Journal of Mathematical Physics, 30(11):2512–2515, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2512_s1.

Rylov:1989:EIM

[2062] Yu. A. Rylov. The equations for isentropic motion of inviscid fluid in terms of wave function. Journal of Mathematical Physics, 30(11):2516– 2520, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2516_s1.

Chan:1989:NSV

[2063] W. L. Chan and Kam-Shun Li. Nonpropagating solitons of the variable coefficient and nonisospectral Korteweg–de Vries equation. Journal of Mathematical Physics, 30(11):2521–2526, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/p2521_s1.

Kull:1989:WTW

[2064] H. J. Kull, R. J. Kashuba, and H. L. Berk. WKB theory of wave tunneling for Hermitian and nearly Hermitian vector systems of integral equations. Journal of Mathematical Physics, 30(11):2527–2544, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i11/p2527_s1.

Weston:1989:WSR

[2065] Vaughan H. Weston. Wave splitting and the reflection operator for the wave equation in R^{3}. Journal of Mathematical Physics, 30(11):2545– 2562, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2545_s1.

Exner:1989:OMT

[2066] Pavel Exner. One more theorem on the short-time regeneration rate. Journal of Mathematical Physics, 30(11):2563–2564, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2563_s1.

Narcowich:1989:DPT

[2067] Francis J. Narcowich. Distributions of h-positive type and applications. Journal of Mathematical Physics, 30(11):2565–2573, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2565_s1.

Exner:1989:BSC

[2068] P. Exner and P. Šeba. Bound states in curved quantum waveguides. Journal of Mathematical Physics, 30(11):2574–2580, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2574_s1.

Barrachina:1989:SCP

[2069] R. O. Barrachina and J. H. Macek. Scattering of charged particles. Journal of Mathematical Physics, 30(11):2581–2584, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2581_s1.

Sabatier:1989:MDM

[2070] P. C. Sabatier. On modeling discontinuous media. Three-dimensional scattering. Journal of Mathematical Physics, 30(11):2585–2598, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i11/p2585_s1.

Nardini:1989:ASE

[2071] Franco Nardini. Approximation of Schrödinger eigenvalues and eigenfunctions by canonical perturbation theory: the periodically driven quantum rotator. Journal of Mathematical Physics, 30(11):2599–2606, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2599_s1.

BretonB:1989:SAF

[2072] Nora Bretón B. Stationary axisymmetric fluid solutions of the Einstein– Born–Infeld equations. Journal of Mathematical Physics, 30(11):2607– 2610, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2607_s1.

Bona:1989:NAB

[2073] Carles Bona and Pere Palou. A new approach to Bianchi V models. Journal of Mathematical Physics, 30(11):2611–2613, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2611_s1.

TorresdelCastillo:1989:KSS

[2074] G. F. Torres del Castillo. Killing spinors and separability of Maxwell’s equations. Journal of Mathematical Physics, 30(11):2614–2615, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i11/p2614_s1.

Coley:1989:SCK

[2075] A. A. Coley and B. O. J. Tupper. Special conformal Killing vector space– times and symmetry inheritance. Journal of Mathematical Physics, 30 (11):2616–2625, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2616_s1.

Xanthopoulos:1989:LPP

[2076] Basilis C. Xanthopoulos. Linear perturbations of plane polarized plane waves. I. The absence of purely incoming perturbations. Journal of Mathematical Physics, 30(11):2626–2637, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i11/p2626_s1.

Grigore:1989:PGC

[2077] D. R. Grigore. A possible generalization of the concept of symmetry in analytical mechanics. Journal of Mathematical Physics, 30(11):2638– 2645, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2638_s1.

Grigore:1989:LCA

[2078] D. R. Grigore. Localizability and covariance in analytical mechanics. Journal of Mathematical Physics, 30(11):2646–2652, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2646_s1.

Grigore:1989:SLF

[2079] D. R. Grigore. Symmetries in the Lagrangian formalism. Journal of Mathematical Physics, 30(11):2653–2663, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/p2653_s1.

Oevel:1989:MAV

[2080] Walter Oevel, Benno Fuchssteiner, Hongwei Zhang, and Orlando Ragnisco. Mastersymmetries, angle variables, and recursion operator of the relativistic Toda lattice. Journal of Mathematical Physics, 30(11):2664– 2670, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i11/p2664_s1.

Noriega:1989:UEM

[2081] R. J. Noriega and C. G. Schifini. The uniqueness of the energy momentum tensor in nonabelian gauge field theories. Journal of Mathematical Physics, 30(11):2671–2673, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i11/p2671_s1.

Cadavid:1989:DAF

[2082] A. C. Cadavid and R. J. Finkelstein. Dyons in affine field theories. Journal of Mathematical Physics, 30(11):2674–2682, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2674_s1.

Hsu:1989:QES

[2083] J. P. Hsu. Quantum equations for simple interacting strings: their exact solutions and physical implications. Journal of Mathematical Physics, 30(11):2683–2688, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2683_s1.

SalazarIbarguen:1989:SNE

[2084] Humberto Salazar Ibarguen, A. García, and J. Plebański. Signals in nonlinear electrodynamics invariant under duality rotations. Journal of Mathematical Physics, 30(11):2689–2691, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/p2689_s1.

Cassinelli:1989:SNR

[2085] G. Cassinelli, G. Olivieri, P. Truini, and V. S. Varadarajan. On some nonunitary representations of the Poincaré group and their use for the construction of free quantum fields. Journal of Mathematical Physics, 30(11):2692–2707, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2692_s1.

Kouadik:1989:LRS

[2086] Smain Kouadik and Pierre Minnaert. Linear realizations of the superrotation and super-Lorentz symmetries. I. Journal of Mathematical Physics, 30(11):2708–2713, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2708_s1.

Schmitt:1989:PDS

[2087] H. A. Schmitt, P. Halse, B. R. Barrett, and A. B. Balantekin. Positive discrete series representations of the noncompact superalgebra Osp(4/2, R). Journal of Mathematical Physics, 30(11):2714–2720, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic),

1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/v30/i11/
p2714_s1.

Shen:1989:UCD

[2088] Hao-Ming Shen and Tai T. Wu. The universal current distribution near the end of a tubular antenna. Journal of Mathematical Physics, 30 (11):2721–2729, November 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i11/p2721_s1.

Figueroa-OFarrill:1989:DGA

[2089] José M. Figueroa-O’Farrill. Deformations of the Galilean algebra. Journal of Mathematical Physics, 30(12):2735–2739, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2735_s1.

Tratnik:1989:MMK

[2090] M. V. Tratnik. Multivariable Meixner, Krawtchouk, and Meixner– Pollaczek polynomials. Journal of Mathematical Physics, 30(12):2740– 2749, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2740_s1.

Hou:1989:ACE

[2091] Bo Yu Hou and Hua Wei. Algebras connected with the Z_{n} elliptic solution of the Yang–Baxter equation. Journal of Mathematical Physics, 30(12):2750–2755, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2750_s1.

Patera:1989:FSA

[2092] J. Patera. The four sets of additive quantum numbers of SU(3). Journal of Mathematical Physics, 30(12):2756–2762, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2756_s1.

Patera:1989:NGR

[2093] J. Patera, R. T. Sharp, and J. Van Der Jeugt. New gradings of sl(3, C) representations. Journal of Mathematical Physics, 30(12):2763–2769, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2763_s1.

Mahomed:1989:LAA

[2094] F. M. Mahomed and P. G. L. Leach. Lie algebras associated with scalar second-order ordinary differential equations. Journal of Mathematical Physics, 30(12):2770–2777, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2770_s1.

Penna:1989:GAQ

[2095] Vittorio Penna and Mauro Spera. A geometric approach to quantum vortices. Journal of Mathematical Physics, 30(12):2778–2784, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2778_s1.

Castrigiano:1989:NAP

[2096] D. P. L. Castrigiano and F. Stärk. New aspects of the path integrational treatment of the Coulomb potential. Journal of Mathematical Physics, 30(12):2785–2788, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2785_s1.

Estrada:1989:AET

[2097] Ricardo Estrada, José M. Gracia-Bondía, and Joseph C. Várilly. On asymptotic expansions of twisted products. Journal of Mathematical Physics, 30(12):2789–2796, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2789_s1.

Lehman:1989:ARM

[2098] D. R. Lehman and W. C. Parke. Angular reduction in multiparticle matrix elements. Journal of Mathematical Physics, 30(12):2797–2806, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2797_s1.

Jaen:1989:CFP

[2099] X. Jaén, R. Jáuregui, J. Llosa, and A. Molina. Canonical formalism for path-dependent Lagrangians. Coupling constant expansions. Journal of Mathematical Physics, 30(12):2807–2814, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2807_s1.

Srivastava:1989:APT

[2100] Sunita Srivastava and Vishwamittar. Application of perturbation theory to the damped sextic oscillator. Journal of Mathematical Physics, 30 (12):2815–2818, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2815_s1. See erratum [2136].

Hislop:1989:ASD

[2101] Peter D. Hislop and John David Crawford. Application of spectral deformation to the Vlasov–Poisson system. II. Mathematical results. Journal of Mathematical Physics, 30(12):2819–2837, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2819_s1.

Srivastava:1989:SDE

[2102] Sushil K. Srivastava. Solution of the Dirac equation in Kasner’s space– time. Journal of Mathematical Physics, 30(12):2838–2844, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2838_s1.

Ruiz:1989:ARQ

[2103] E. Ruiz. Approximate relativistic quantum Hamiltonians for N interacting particle systems. Journal of Mathematical Physics, 30(12):2845– 2853, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2845_s1.

Ruelle:1989:QMA

[2104] Ph. Ruelle, E. Thiran, D. Verstegen, and J. Weyers. Quantum mechanics on p-adic fields. Journal of Mathematical Physics, 30(12):2854–2874, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2854_s1.

Dittrich:1989:DOS

[2105] J. Dittrich, P. Exner, and P. Šeba. Dirac operators with a spherically symmetric δ-shell interaction. Journal of Mathematical Physics, 30(12):2875–2882, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2875_s1.

Mann:1989:CTC

[2106] A. Mann, M. Revzen, K. Nakamura, H. Umezawa, and Y. Yamanaka. Coherent and thermal coherent state. Journal of Mathematical Physics, 30(12):2883–2890, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2883_s1.

dePrunele:1989:GTE

[2107] E. de Prunelé. o(4, 2) × o(4, 2) group theoretical expression of the interelectronic Coulomb potential. Journal of Mathematical Physics, 30 (12):2891–2894, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2891_s1.

Bezerra:1989:GAA

[2108] V. B. Bezerra. Gravitational analogs of the Aharonov–Bohm effect. Journal of Mathematical Physics, 30(12):2895–2899, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2895_s1.

Faris:1989:SWP

[2109] William G. Faris and Woody J. Tsay. Scattering of a wave packet by an interval of random medium. Journal of Mathematical Physics, 30 (12):2900–2903, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2900_s1.

Gavazzi:1989:WOF

[2110] Giuliano M. Gavazzi. Weyl-ordered fermions and path integrals. Journal of Mathematical Physics, 30(12):2904–2906, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2904_s1.

Moncrief:1989:REE

[2111] Vincent Moncrief. Reduction of the Einstein equations in 2+1 dimensions to a Hamiltonian system over Teichmüller space. Journal of Mathematical Physics, 30(12):2907–2914, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2907_s1.

Li:1989:SSP

[2112] Jian Zeng Li and Can Bin Liang. Static “semi-plane-symmetric” metrics yielded by plane-symmetric electromagnetic fields. Journal of Mathematical Physics, 30(12):2915–2917, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http:/

/jmp.aip.org/resource/1/jmapaq/v30/i12/p2915_s1.

Coll:1989:TPF

[2113] Bartolomé Coll and Joan Josep Ferrando. Thermodynamic perfect fluid. Its Rainich theory. Journal of Mathematical Physics, 30(12):2918–2922, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v30/i12/p2918_s1.

Jakubiec:1989:CPT

[2114] Andrzej Jakubiec and Jerzy Kijowski. On the Cauchy problem for the theory of gravitation with nonlinear Lagrangian. Journal of Mathematical Physics, 30(12):2923–2924, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2923_s1. See erratum [2134].

Kalnins:1989:TSI

[2115] E. G. Kalnins, W. Miller, Jr., and G. C. Williams. Teukolsky– Starobinsky identities for arbitrary spin. Journal of Mathematical Physics, 30(12):2925–2929, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2925_s1.

Meister:1989:CMM

[2116] A. Meister. Complex manifold methods in quantum field theory in curved space–time. Journal of Mathematical Physics, 30(12):2930–2942, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2930_s1.

Wakolbinger:1989:SVC

[2117] A. Wakolbinger. A simplified variational characterization of Schrödinger processes. Journal of Mathematical Physics, 30(12):2943–2946, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2943_s1.

Book:1989:DCI

[2118] David L. Book and Todd A. Brun. Disproof of a conjectured inequality arising in the theory of magnetic flux diffusion. Journal of Mathematical Physics, 30(12):2947–2950, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2947_s1.

Pak:1989:HMN

[2119] N. K. Pak and R. Percacci. Hamiltonian methods for nonlinear sigma models. Journal of Mathematical Physics, 30(12):2951–2962, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2951_s1.

Bracken:1989:SIR

[2120] A. J. Bracken and A. Cant. Singular indecomposable representations of sl(2, C) and relativistic wave equations. Journal of Mathematical Physics, 30(12):2963–2972, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2963_s1.

Arai:1989:DGS

[2121] Asao Arai. On the degeneracy in the ground state of the N = 2 Wess– Zumino supersymmetric quantum mechanics. Journal of Mathematical Physics, 30(12):2973–2977, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2973_s1.

Devchand:1989:ESG

[2122] C. Devchand. N = 3 extended supersymmetric gauge theories and an explicit construction of higher conservation laws. Journal of Mathematical Physics, 30(12):2978–2983, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v30/i12/p2978_s1.

Frappat:1989:VOR

[2123] L. Frappat, A. Sciarrino, and P. Sorba. Vertex operator representations of basic affine superalgebras. Journal of Mathematical Physics, 30 (12):2984–2990, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v30/i12/p2984_s1.

Kleinert:1989:SDE

[2124] H. Kleinert. Solution of Dyson equation for elastic membranes. Journal of Mathematical Physics, 30(12):2991–2993, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427.

URL http://jmp.aip.org/resource/1/jmapaq/v30/i12/p2991_s1.

Bona:1989:ESC

[2125] Pavel Bóna. Equilibrium states of a class of quantum mean-field theories. Journal of Mathematical Physics, 30(12):2994–3007, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v30/i12/p2994_s1.

Malbouisson:1989:ESR

[2126] A. P. C. Malbouisson, F. R. A. Simão, and A. F. de Camargo Filho. Erratum: “Some results from a Mellin transform expansion for the heat kernel”, [J. Math. Phys. 30, 1226 (1989)]. Journal of Mathematical Physics, 30(12):3008, December 1989. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.

org/resource/1/jmapaq/v30/i12/p3008_s1. See [1885, 2131].

Qadir:1990:ASS

[2127] Asghar Qadir and M. Ziad. Addendum: “Static spherically symmetric space–times with six Killing vectors”, [J. Math. Phys. 29 (1988), no. 11, 2473–2474; MR 90a:83015]. Journal of Mathematical Physics, 31(1):254, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1683].

Biedenharn:1990:EST

[2128] L. C. Biedenharn and H. van Dam. Erratum: “On steady transverse magnetogas flow”, [J. Math. Phys. 27 (1986), no. 5, 1466–1470; MR 87f:76128 ] by S. N. Singh, K. R. Choubey and B. P. Singh. Journal of Mathematical Physics, 31(1):v, January 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [693].

Elizalde:1990:EEZ

[2129] E. Elizalde and A. Romeo. Erratum: “Expressions for the zeta-function regularized Casimir energy”. Journal of Mathematical Physics, 31(3): 771, March 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v31/i3/p771_s1. See [1873].

Vaklev:1990:EGT

[2130] Y. Vaklev. Erratum: “Gauge transformations for the quadratic bundle”. Journal of Mathematical Physics, 31(4):1026, April 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v31/i4/p1026_

s1. See [1961].

Malbouisson:1990:CSR

[2131] A. P. C. Malbouisson, F. R. A. Simão, and A. Ferraz de Camargo Filho. Comment on: “Some results from a Mellin transform expansion for the heat kernel”, [J. Math. Phys. 30 (1989), no. 6, 1226–1229; MR 90m:58204a]. Journal of Mathematical Physics, 31(5):1114–1115, May 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1885, 2126].

Bokhari:1990:KVS

[2132] Ashfaque Bokhari and Asghar Qadir. Killing vectors of static spherically symmetric metrics. Correction to: “Symmetries of static, spherically symmetric space–times”, [J. Math. Phys. 28 (1987), no. 5, 1019– 1022; erratum; MR 90e:83046a b]. Journal of Mathematical Physics, 31 (6):1463, June 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1063, 1416].

Noriega:1990:AEI

[2133] R. J. Noriega and C. G. Schifini. Addendum to: “The equivariant inverse problem and the uniqueness of the Yang–Mills equations”, [J. Math. Phys. 30 (1989), no. 10, 2382–2387] by M. C. López, R. J. Noriega and C. G. Schifini. Journal of Mathematical Physics, 31(6):1503, June 1990. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [2044].

Swierczynski:1990:ECP

[2134] Z. Świerczyński. Erratum: “On the Cauchy problem for Yang–Mills equations with external current”, [J. Math. Phys. 30 (1989), no. 8, 1699–1706; MR 90j:58023]. Journal of Mathematical Physics, 31(9):2334, September 1990. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v31/i9/p2334_s1. See [2114].

Grigore:1991:ECC

[2135] D. R. Grigore and O. T. Popp. Erratum: “The complete classification of generalized homogeneous symplectic manifolds”, [J. Math. Phys. 30 (1989), no. 11, 2476–2483; MR 90i:58052]. Journal of Mathematical Physics, 32(2):564, February 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v32/i2/p564_s2. See [2057].

Srivastava:1991:EAP

[2136] Sunita Srivastava and Vishwamittar. Erratum: “Application of perturbation theory to the damped sextic oscillator”, [J. Math. Phys. 30, 2815 (1989)]. Journal of Mathematical Physics, 32(2):564, February 1991. CO-DEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-

2427. URL http://jmp.aip.org/resource/1/jmapaq/v32/i2/p564_

s1. See [2100].

Mitra:1991:EFD

[2137] Bani Mitra. Erratum: “Finite-dimensional irreducible representations of the SU(3/1) superalgebra”, [J. Math. Phys. 30 (1989), no. 3, 719– 731; MR 90a:17018]. Journal of Mathematical Physics, 32(4):1109, April 1991. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (elec-

tronic), 1527-2427. URL http://jmp.aip.org/resource/1/jmapaq/
v32/i4/p1109_s1. See [1814].

Demys:1992:ECH

[2138] K. Demys. Erratum: “Comment: ‘On the hyperbolic complex linear symmetry groups and their local gauge transformation actions’ by Z. Z. Zhong”, [J. Math. Phys. 26, 404 (1985)]; [J. Math. Phys. 28 (1987), no. 2, 339; MR 88a:83062]. Journal of Mathematical Physics, 33(10): 3602, October 1992. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-

7658 (electronic), 1527-2427. URL http://jmp.aip.org/resource/1/
jmapaq/v33/i10/p3602_s1. See [958].

Chandler:1993:EBQ

[2139] Colston Chandler and A. G. Gibson. Erratum: “N-body quantum scattering theory in two Hilbert spaces. V. Computation strategy”, [J. Math. Phys. 30 (1989), no. 7, 1533–1544; MR 90h:81198]. Journal of Mathematical Physics, 34(2):886, February 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.

aip.org/resource/1/jmapaq/v34/i2/p886_s1. See [1930].

Saito:1993:ADF

[2140] Yoshihiko Saito, Reiji Sugano, Tadayuki Ohta, and Toshiei Kimura. Addendum to: “A dynamical formalism of singular Lagrangian system with higher derivatives”, [J. Math. Phys. 30 (1989), no. 5, 1122–1132; MR 90d:70047]. Journal of Mathematical Physics, 34(8):3775–3779, August 1993. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1872].

Nelson:1994:EGL

[2141] Robert A. Nelson. Erratum: “Generalized Lorentz transformation for an accelerated, rotating frame of reference”, [J. Math. Phys. 28 (1987), no. 10, 2379–2383; MR 89c:83006]. Journal of Mathematical Physics, 35 (11):6224–6225, November 1994. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. URL http://jmp.aip.org/

resource/1/jmapaq/v35/i11/p6224_s1. See [1250].

Higuchi:2002:EST

[2142] Atsushi Higuchi. Erratum: “Symmetric tensor spherical harmonics on the N-sphere and their application to the de Sitter group SO(N, 1)”, [J. Math. Phys. 28, 1553 (1987)]. Journal of Mathematical Physics, 43 (12):6385, December 2002. CODEN JMAPAQ. ISSN 0022-2488 (print), 1089-7658 (electronic), 1527-2427. See [1144].